A Bibliography of Publications about Benford’s Law, Heaps’ Law, and Zipf’s Law

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

14 August 2024 Version 1.156

Title word cross-reference

(0, 1) [BNS10a, BNS10b]. (3x + 1) [Sin03]. (m, k) [AI99]. 1 [Ant91, BB76, Kem75, Kui69, MN08a, Sch91a, Sri13, Sze10]. 1/f [BTW87, Che12b, KA84, MS83, Nic88, NT89]. 2 [Bha12, LP15, PGLL10]. $29.95 [Ale05]. 3 [CNS15]. 3n + 1 [Wir98]. 3x + 1 [KM05, Lag85, LW92, Lag03, LS05, LS06, Lag10, MK05, Sim07]. 80/20 [Lip09]. [0, 1] [CI96]. A [Fle66]. A(n, α)B [LZNR11]. b [Ber15b]. β [NWR09]. mod1 [Cig64, Dia77, Sch83b, Sch84, Sch88a, Tsu52]. ∪ [LS11]. d = 0 [LS11]. H [KPK05, KPK06]. H_{∞} [Sch83a, Tic87]. j [PU86, UP85]. k [LWY10, MAE06, XLW14]. k = 2 [AI99]. L [KM05, MK05]. log p [Sur93]. m = 6 [AI99]. H_{∞} [Sch87b]. N! [Kun87]. N [Rou15, Cha11, CSV96, KR13, KR15, Rou76a, VA99]. P [GBA04, SBB01]. P (X_{2} < X_{1}) [Jee97]. π [She12, Zor11]. ψ [GSR12]. Q [Yua84, Abe02, MML01]. R = Pr(X > Y) [Gun15]. R^{d} [Ber15a]. ρ [TW11]. XY [RMB^{+}14]. ζ [GSR12].

-ary [Ber15b]. -Decay [NWR09]. -digit [KR13, KR15]. -échantillon [Rou76a]. -exponential [Abe02, MML01]. -function [KPK05, KPK06]. -functions [KM05, MK05]. -gram [Cha11]. -Mathematics [LS11]. -order [Yua84]. -sample [Rou76a]. -summability [Sch83a]. -tuple [CSV96]. -uniform [Sch87b]. -Zipf [AI99, VA99].

0 [Ale05]. 0-691-09983-9 [Ale05]. 06-10-1826R [HYYH08].

1 [Bro98]. 11th [NKZ12]. ’12 [NKZ12]. 14th [NKZ12]. 16th [GCB^{+}05]. 1800s [Mil15c, Mil15d]. 1826R [HYYH08]. 19 [GBF21, KO20, LHJ20, MM21].

2 [GBF21]. 2-cycles [Sim07]. 2000s [MK13]. 2005 [NSGP05]. 2009 [Meb10c, PS09, Rou14]. 2010 [And11]. 2012 [NKZ12]. 2020 [CLT20]. 27th [CLT20]. 2nd [SM11].

41 [Ano93].

500 [CEZ10]. 573 [Ano50].

9 [Ale05].

abgegriffenen [Sch06]. Absaroka [Bar11]. Abuse [San87a, San87b, San88, San92]. academic [TV21]. accelerated [AS15]. acceptance [KMN09]. access [PQ00]. according [AYS16]. Accounting [AA10, Bha02, BK08, Cal13, DHP04, Iud12, Nig93, Nig12, OW12, Wan11b, BL95, FGZZ11a, KB07, LWFW14, Moc02, Nig11e, QW03, Sav04, SS10]. Accuracy [Ben43, LP15, dMH06, CFH^{+}20, HSR^{+}08a, HSR^{+}08b, Has03, ZIWU14]. accurate [ZHS09]. Achieving [MYP14, Est16]. Achilles [Moc17]. Achillesferse [Moc17]. ACM [PS09]. ACMOS [NKZ12]. Acquisition [NKRS15, Est16]. across [Per10, Sch06]. act [AG08, Bec14]. activity [EKJ^{+}20, Mir14]. actually [BKS05]. adaptation [Sjo14]. adapted [LCC^{+}05]. Adaptive [LEC^{+}15, LBC06]. add [BOT06]. Adding [Nig99a]. Addison [Ano50]. Addison-Wesley [Ano50]. addition [FT86]. additive [BE69, Jec92, MN14, Ree06]. adjusted [Rou92]. Advance [Sta12]. adventures [GM97b]. Adversarial [MWX^{+}20, SJP^{+}15]. AdWords [GML15]. aerobiological [DdA^{+}09]. aeruginosa} [CLRTFM08]. aesthetics [MRM^{+}05, NSGP05]. affiliation [MG15]. after [eSHK05]. Again [Rai85]. against [HNY20, MPS11]. agenda [Nor13]. agent [MG07a]. agent-based [MG07a]. agglomeration [MML01]. agglomerations [GB13]. aggregated [MAC14b]. Aggregation [RRZR09, EKJ^{+}20]. ago [Sch06]. Agrarprodukte [GTM12]. Agree [SW10]. agricultural [GTM12]. Aid [Ban00, NM97]. Air [Pér07, Bro05, Sto16]. al [Bur04, PT04]. al. [MBG^{+}96]. aléatoire [Rou76a].

Alfred [Har09]. algebraic [Var99]. algébriques [Var99]. algebro [ZF15]. algebro-geometric [ZF15]. algorithm [FC14, Gar11, KSH11, PGLL10, WYTD93, WDY93]. Algorithmic [McA03, McA05]. algorithmically [TGD05]. Algorithms [Haf13, Knu69, Knu98, BFS03, Yua84]. allocation [HG14, KSH11]. Almost [Dur74, Sch88c]. along [JD11, JD12]. alpha [BMP93]. alternative [BP06, HSR^{+}08a, Da 22, HSR^{+}08b, Uze10]. Altmetric [GSB24]. Always [Da 20]. ambient [Bro05]. ambiguities [HN05]. Ambiguity [Wil72]. America [Car15, Meb08a]. American [Meb06b, Meb08b, Meb08c]. AMEX [Bar11]. among [AA11, LWFW14]. amount [LAL15]. AMR [LYH15]. Amsterdam [AW89]. Analiz [AA10]. Analogous [Can03a, Can03b]. Analyse [Bah15]. analyses [Bat09]. Analysing [PK06]. Analysis [AA10, Ans09, BR15, CS11, CL87b, CT05, Cod99, FWW85, HR10, LDJ89, Lin08, Liu15, MK10, MSA^{+}11, Möl09, Nig96a, Nig97, Nig01, OB13, Ped94, QBP^{+}04, RVAN15, RGE13, RGL14, Sen73, Sim98, SB15, SPGM12, Van15a, Van15b, WG68, WSY11, ŽD11, AYS16, Ano05, AA11, AI99, Bah15, BSM15, CSC14, CMHS95, CSV96, DMF09, FDA^{+}04, GBF21, GSB24, Has03, Jon02, Lan00, LOR06, Man59, MBG^{+}95, MWS99, Mil15b, Mir16b, MS08, Mül03, Nig92, Nig99a, Nig00b, Nig00c, Nil10, Par75, Pos04c, Pos10, RR03, SF11, SH10, Tay05, VA99, Wan11a, dGGd21]. Analytical [KZ15, NM97, Bro05, Bro07, DN00]. analytics [Nig11e]. Analyzing [AR13]. Anastasi [Lim06]. Andersen [Nig05]. Anesthesia [HKKS06, Mar12, HZKS12]. angewendet [GTM12]. anharmonic [Mas06a]. animal [Ano05, DMJH11, MDJH05]. animals [BGP07]. ANN [BXK11]. ANN-based [BXK11]. annotated [Lag03]. Anomalies [Car88, SGW04, Rou09]. Anomalous [Ben38, San02a, San02b, San02c, CI96, Ste93]. Anomaly [CT22, PT11, Rou14]. anomaux [San02c]. answers [Zen11]. AntiBenford [CT22]. Antiforensics [MTT13]. Anwendbarkeit [BBJ11]. Anwendungsmöglichkeiten [Mül03]. aplicaciones [PT04]. Apparatus [SF11]. Apparent [Kub77, BH74]. Appearance [NKRS15]. applicability [BSWB16, BBJ11, STJ09, WSB15, WSB16, YLX14]. Application [AdSPG^{+}76, Bak07, BCCP18, Dun67, FGZZ11a, Fed82a, Fed82b, Gut16, Hsü48, HSM17, Izs06, Jin09a, KŽ08a, ŽK09, pLCjW04, MRR^{+}10, NNF21, Nig96b, NM07b, OB13, PGQA^{+}15, Shi09, Wor75, ZSB^{+}12, Boy94, BMM84, CSC14, GD14, Hür03, Hür06b, Hür09, Jin09b, KŽ08b, KB03, Mas05d, MF21, PGHA07, Sch88a, TRB^{+}12]. Applications [Egg90b, EG12a, EG12b, Fel68, Ism04, Iud12, Mül03, Nig12, SG16, Tor06, Ber15b, Dia73, Fer13, FSS07, Giu11, Hil02, Hor09b, KCCT23, Kos15, MM04, MM14, Mil15a, PT04, PT11, Rai03, Krä15]. Applied [NM07a, DdA^{+}09, GTM12, MPS11, Mas06e, NM06, SH10]. Applying [BG11b, CT05, GM12, HSD12]. Apprendre [Rou10]. Approach [MZLH15a, MZLH15b, Möl09, Asa81, ACC16, Bec14, HSR^{+}08a, CL87b, CL87a, Che91, GSR12, HSR^{+}08b, LCJ10, LB14, LSS00, MG07a, Mei09, Nil10].

Approaches [RZS^{+}15]. appropriate [MDJH05]. Approximate [You13, ANRV07]. approximation [BK07, DL07, DL08]. Approximations [BX18, BX19]. April [NKZ12, Rai03]. arcane [Mül11]. area [Sud10]. Areas [WDB95]. Argentina [CS11]. arguments [AS68, Egg07]. ARITH [CLT20]. Arithmetic [BB85, CLT20, JL08, Kre73, Kre75, CHL19]. Arithmetik [Kre75, Kre73]. arithmétiques [JL08]. Arnold [Gre86, Sha89]. articles [AYS14, CC11, Mir16b]. artifact [LYY09]. Artificial [Sun28]. ary [Ber15b]. Asia [BCM02]. Asian [Jin09a, Jin09b]. Aspects [FM95, JP01, MdSSZ09, Bro80a, Bro80b, Bro80c, Hea78, MNW^{+}05, Pha13, PL86, SCD01]. Aspectual [Nov10]. Assessing [Has02, Mac14a, Nig11a, Wal02, dMH06]. Assessment [Moh79, Nig05]. assignment [XS09]. assimilation [Sea02a, Sea02b]. Assist [DHP04]. assisted [DN00]. Association [XP89]. associations [FMS10]. Associativity [LB09]. associés [BMM10]. Assumption [Ami07]. assumptions [XS09]. Assurance [DCH^{+}17]. Astonishing [Liv02]. Astronom [Sch06]. astronomer [Sch06]. Astronomy [AL14b]. Astrophysical [MdSPZ06]. Asymmetric [BFLB10, KP05, CL04a]. asymmetries [BFLB10]. Asymptotic [GBD^{+}15, GBTB^{+}16, Rou76a, Sta05, Loy73, Mas06b, Pet03, Sch83b, Sch84, Sch86, Sch87a]. asymptotics [OJM15, SM06b]. asymptotiques [Rou76a]. Atherton [Gro67]. atmospheric [LF16]. Atomic [MG22, Pai08, Pai13]. Attacks [ALY19, MTT13]. Attempted [Meb06a]. attractors [Nic88, NT89]. attribution [FNP^{+}15, Rou92]. Auctions [Gil05, GML15, Gil07, LG10]. audio [LYH15]. Audit [GC13, Lan03, Sta12, WSU08, BBE10, KMN09, Mül03]. Auditing [KŽ08a, ŽK09, Lu07, Möl09, Nig00a, Nig12, Yor00, BL95, SS10]. Auditor [Nig96a, BXK11]. Auditors [Nig01, Rob97, Nig00c]. audits [GD14, KCCT23, She94, TS94]. Auerbach [RC23]. Aufdeckung [Moc02, HB18]. August [AW89]. authenticating [PGLL10]. Authentication [ZHS10]. authenticity [Has02, Has03]. Author [KR01, Egg94, Lag03]. Authors [WDB95, MG07b]. Authorship [SDK14, Aus14, OM11, Row05]. Auto [LCY^{+}15]. Auto-Generated [LCY^{+}15]. autocatalytic [HS01, MBS99]. Autocratic [Sjo14]. Automated [ABBM10]. Automatic [Fed82b, Kle17, NKRS15, SSMW05, YSC13, LFCR04, NKZ12, Pet03]. Automatically [LP15, Par75]. autonomous [HJ15, Liu15]. aux [Jan12]. Availability [Kub77]. Avalanches [NS96]. Averages [Dia74]. Averaging [JD11, JD12]. Avoidance [Man08]. Aware [LHK^{+}15, IG08, SLS^{+}16]. axial [RDFM10]. Azerbaijan [Her10]. aziende [Bra76].

B [Sha89]. B. [Gre86]. back [BH11b]. background [Sea02a, Sea02b]. Badly [MAdV05]. ball [CM11, Wap12]. ball-bin [CM11]. bank [YD15]. Banking [GP13, OB13]. bankruptcy [Fuj04]. Bargaining [KM11]. barriers [DDS98, DK09, LV94, LG10]. Base [Hil95a, Bha12, Sch98]. Base-invariance [Hil95a]. Baseball [Reg12]. Based [AA10, BS10a, CS09b, Kle17, LHK^{+}15, RBXM15, SG16, Van15a, Van15b, YSC13, Yul25, ANRV07, BXK11, CW97, CZ04, DS01, Egg94, ER02, FGZZ11a, GJS07, LZL^{+}12, MG07a, MZLH15a, MZLH15b, MF21, OJM15, PGLL10, RAV10, SZVHA10, SLS^{+}16, Sud10, SH10, TA12, TH13, WC03, Wu08, XWLD11, ZDCW12, dGGd21]. bases [AW89, Sch81]. basic [BH11a]. Basics [Nig11b]. Basis [Kre73, Kre75, Moo36, Mor06, Abe02, Bah15, HSD12, Los01, Nag05, RRR04]. Basiswahl [Kre73, Kre75]. Batched [LDJ89]. Bayes [Jee97, McG11, Sol00]. Bayesian [AP89, JKK^{+}08, JKK^{+}09]. Be [HKKS06, LCC^{+}05, Mac14a, Nig93, Sav04]. Beale [Was21]. Beautiful [MPM02]. beauty [MA18]. BEC [LS11]. Beckmann [Shi89]. been [OV93]. Before [WHN13]. begin [TFGS07]. Begründung [Hum08]. Behaving [MAdV05]. Behavior [BDE^{+}14, BDE^{+}17, Car88, Gor12a, Ano50, AJ14, BMT96, Che95, CCT95, CSV96, DM19, MNW^{+}05, Mir14, OV95, RV98]. behavioral [DZ03]. Behaviour [RBXM15, DD06, HKL^{+}06, Sch87a, Zip49]. Beijing [Sto16]. Being [CEZ17, CEZ18]. beizukommen [SHC^{+}12]. Belies [Lee18]. benchmarking [Wal02]. Benchmarks [RGE13]. Benford [Bah15, BCCP18, BP20, Die04a, For10, Haf13, Hum06, Hum18, Moc02, Moc17, MdSSZ09, Mül03, Nig00b, Ole10, Ole14, PT04, Pos04c, Pos10, Tor06, AS05, AA10, AR13, AP10, ALY19, Ale09, AL14b, AL14a, All97, AR09, AR10, AR11, AYS14, AYS16, ARS11, AOT13, Ano11, Ano13, AG05, AA11, AJ14, AHI14, AHI15, ACC16, Bah15, BSO10, Bal12, Bal15, BNS10a, BNS10b, Ban13, BCP21, BCCP22, BCD23, BG09, BG10, BG11a, BGTM^{+}11, BBC^{+}18, Ben19, Ber03, Ber04, Ber05b, Ber05a, Ber05c, BBH05, BH07, BH09, BH10a, BH10b, Ber10, BH11b, BH11a, BHKR11, Ber11, BE14a, BE14b, BH15b, Ber15a, BH15a, BE16, BH17, BX18, BX19, BH21, BK07, BDE^{+}14, BDE^{+}17, Bha12, Bha02, BKS05, BK08, BXK11, BSM14, BSM15, BS10b]. Benford [BJS06, BMM10, BSWB16, BF09, Bra78, BBJ11, BG11b, Bro05, Bro07, Bro98, BMP93, Bul11, BK91, Bur10a, Bur09, BS92, BW98, CHL19, CFH^{+}20, CC11, Cam24, CS22, CBC^{+}19, CM21, CP23, CL23, CS09a, CLPR23, CPS16, CG07, CKT^{+}09, Cio05, CT05, CA12, Con10, CEZ10, CEZ17, CEZ18, CLRTFM08, CSC14, Cra10, CLM15, D’A20, Da 20, Da 22, DD06, DCH^{+}17, DDS98, Dea93, DMO10, DMO11, DS05, DJ10a, DJ13, Dep08, Die04a, Die04b, Die07, DJ10b, Die12, DdA^{+}09, DN00, DL07, DL08, DM19, DHL^{+}20, DHP04, EHM^{+}14, Egg11, EG12a, EG12b, Eli13, EL03, FC20, FMSV24, FGZZ11b, FG10, FLM^{+}21, Fel17, Few09, FGPM12, FGPM16, Fru16, FSS07, GBD^{+}15, GBTB^{+}16, GD08, GD09b, GD09a, GD11, GR07, GW01, GW04, GM12, Gil05, Gil07, GJ10, GC13]. Benford [GGP09, GJARPGAC10, Goo16, Gor12a, Got92, GN02, GP13, GGA03, GJS07, GM22, GTM12, GT09, GSB24, HCS09, HB18, Has02, Hay09, HSD12, HR10, Hil95a, Hil97, Hil02, Hil11, HF16, HV01, Hop16, Hor09a, HKKS06, HNY20, HJJYM11, HSM17, Hum06, Hum08, Hum18, Hün07, Hür03, Hür04, Hür06a, Hür06b, Hür09, Hür15a, IMS15, IHPS14, IHWZ17, Iud12, Jam01, JTY14, JTY16, JKK^{+}08, JKK^{+}09, JdlR04, JdlR09, Jas10, JHTS10, JSZ11, JJ21, Jol01, Jol05, Jol09, JS06, JS09, Kaf09b, KNRS88, KSG^{+}16, KBHR10, KK20, Kaz23, KR12, KR13, KR15, Kim12, KCCT23, Kle17, KO20, KM05, Kos15, KŽ08a, KŽ08b, ŽK09, Krä15, KV14, Kre03, Kry09, KP04, KP05, KPK05, KPK06, KB03, KB07, Lac18, LS05, LS06, LCJ15, LHJ20, LSE00, Lee15]. Benford [pLCjW04, LFY15, LF16, LK12, LZNR11, Lol08, Low00a, Low00b, LB05, LBC06, LB06b, LG10, LL08, LS11, MWX^{+}20, MRR^{+}10, MM21, ME22, Meb06a, Meb06b, Meb06c, Meb08b, Meb08c, Meb11, Meb15, MTT13, Mil04, MK05, MN06a, MN06b, MN08a, MN08b, Mil15a, Mil15b, Mil15e, Mir14, MAC14b, Mir16b, MA18, MG22, Moc02, Moc17, Moc21, MW04, Möl09, MdSPZ06, Mor06, Mör01, Mül03, MF21, Mur23, Nag84, NS87, NKS90, Nag05, NWNG22, NHD16, NNF21, NR08, NWR09, Nig93, Nig96b, NM97, Nig98, Nig00c, NM06, NM07a, NM09, Nig11a, Nig11b, Nig11c, Nig11d, Nig12, Nig15, Nov10, NM07b, Ole10, Ole14, OMNO10, OB13, Pai08, PSM10, Pav81, Pav82, PGHA07, PGQA^{+}15, PT04, PT11, Pet81a, Pet81b, Pha13, PTTV01, PR22, Pos04a, Pos04b, Pos04c, Pos10, QZH10, QZHC11, QW03, RMB^{+}14]. Benford [RGE13, RGBK14, Rav08, Reg12, Reg82, RC18, Rob11, RM11, RR09, Ros11, Rou09, Rou15, RS19, STJ09, STJ10, STA11, SM06a, Sch88c, Sch88a, Sch88b, Sch88d, Sch89, Sch90, SN91, Sch91a, Sch98, Sch10, Sch08, Sch15, SF01, Sea02a, Sea02b, SS11, Sen73, SLS^{+}16, SG16, SW10, SF11, SM11, SPP17, SB15, Sin74, SID13, SID15, Smi97, Smi13, SCD01, SPGM12, Sta10, Str10, SH10, SB93, SCE03, TRB^{+}12, Tao09, Töd09, TBL00, TV21, cTnHwH17, Val12, Van15a, Van15b, Var72, VW22, Vog00, Vol20, Vuk08, Wag10, WCCK09, Wan11b, Wan11a, Was21, Was81, WSU08, Wei11, WCP^{+}07, WSB15, WSB16, WSY11, Wój13a, Wlo71, Wój13b, Wój14, Yor00, ŽD11, ZHS09, ZHS10, Zor11, Zun19, dSFSL20, dGGd21, dVM13, eSHK05]. Benford-Gesetz [Hum06, Hum18]. Benford-Gesetzes [Hum08]. Benford-like [GJS07]. Benford-Tests [HB18]. Benfordov [Jas10]. Benfords [Pos05, ARS10, Nig01, Sav06, GTM12, Hün07]. Benford’sche [Bul11, BBJ11]. Benoît [And11, Baw11, Val11]. BEQS [LS11]. Berichtigung [Kre75]. Bernie [Hen11]. Best [BX18, BX19, Lan04]. Beta [ZSB^{+}12, RS19, UP85]. Betrayal [Jud04]. Betriebsprüfung [BBE10]. Better [Haf13, Kos06, RMB^{+}14]. Between [VB08, CL86a, CL86b, Che12a, Egg05b, Egg13, GBA04, Hav95, Hub78, Irm97, LAL15, Nic88, NT89, Orl76]. Bevölkerungskonzentration [Aue13]. bevorzugt [Hum18]. Beyond [LFY15, Mon01, BBL07, Bro07, DM19, EC12, FNP^{+}15, HHC^{+}14, Knu01, SMS10]. biased [VA99]. biases [Nig00b]. Bible [Sal97]. Bibliographic [Fed82a, Fed82b]. bibliographies [Egg90b]. Bibliography [BH09, Vla79, Hür06a, Lag03, Li11, MS10]. Bibliometric [Fai69, Fai05, Nic89, AYS16, OV97, NGJ03]. bibliometrics [Per90]. Big [Mil15c, Mil15d]. bin [CM11]. Binary [HM65, CMHS95, Den97]. binomial [Sar73]. Bioinformatics [Tor06]. Biological [GM08]. Biology [SG00, Mel06, Zip35, Zip65]. Biometric [IHWZ17, IHPS14]. birth [AHI14, AHI15, KPH10]. Bivariate [CT07, Jee97, GN07]. Black [Sch08]. Bladder [ZSB^{+}12]. blind [MS09, MS10]. Block [JHTS10, LYY09]. Blocs [Lee18]. blogosphere [Dep08]. blood [D’A20]. blooper [Her98]. Blunders [Hil11]. body [DG12]. Bolsa [dGGd21]. Book [Ale05, Ano50, Bal66, Gre86, Har09, Lim06, Mel06, Sha89, BBM11b, Gei48]. Boolean [VK90]. bootstrap [Rou15, SH10, VAGE09]. Bose [Mas05a, Mas06a]. Bosnian [Jas10]. bots [ME22]. bound [Hür15c]. Bounds [Kem75, BCIS09, DL07, DL08]. Bradford [Asa81, Bas92, Bas94, Boo94, Bro77, Bro79, Bur91, CL86a, CL86b, CL87a, Che95, CCT95, CZ98, DG78, Egg85, Egg86, Egg90b, Egg90a, Fai69, Fai05, Gro67, Has76, HN05, HHC^{+}14, Hub76, Hub78, KLC84, Lei80, MM84, MK79, Mor81, NH07, OV93, OV95, OV97, Per90, PL86, Pop75, PK78, Sud10, Wil72, Wor75]. Bradfordizing [May08]. Braid [dLKHS04a, dLKHS06]. Brain [QBP^{+}04]. Brazil [Soo13]. Brazilian [GBF21, GR07, MR06]. BRD [Bah15]. Breakdown [AHI14, AHI15, Rou92, BSWB16, KR01]. Breaking [CG07]. Bribery [Bec14]. Brief [Egg10b, EG12b, Gro67, Hub78, KH95, KG72a, Lou92, MK79, Mit03, Mur73a, Nic87a, Sch75c, Voo74a]. broken [Low00b]. Brown [KsS73]. Bucket [KNT89]. budget [Vog00]. Budgetsündern [Vog00]. Bug [CL04a]. Builder [Lu07, KA15]. Building [LB06b]. Bundestag [BG11b]. Burr [CGJ10, KP09]. Burridge [Has07]. Bursting [Zom14]. Business [NCT10, NCT13]. busters [Ano12]. busy [PQ00].

C [Gre86, Sha89, Yul25]. cache [DS01, JK08, VM04]. caching [HS08, SM06b]. CAD [PGLL10]. Calcul [Poi12]. Calculation [Lon92, Egg90b, Lou92, Poi12]. calculations [FG10]. Calculus [ABBM10, Bro79, Uze10]. Calibration [SBB01]. Cambridge [Ano50, Mel06]. came [Sch06]. Camera [LEC^{+}15]. campaign [CG07]. Can [AP10, HKKS06, IKC96, LCC^{+}05, Meb15, Mil15c, Mil15d, MWE95, Nig93, Nig99b, Ano11, RR03, Sav04, Sta10, Meb10b]. Cancer [LY02, ZSB^{+}12]. Cape [NNF21]. Capital [Bar11]. capture [MG22]. cards [Che12c]. cards-shuffling [Che12c]. Carolina [Meb10d]. Carving [SG16]. Case [GR07, Lu07, Vuk08, Bro05, CP23, CEZ10, Kim12, Knu01, LB06b, NNF21, Rou92]. cases [FGZZ11a, GBF21]. Cauchy [EHM^{+}14]. cautionary [Bur04]. Cayley [SS09]. Cayley-tree-like [SS09]. cells [CLRTFM08]. censored [KS10, RAV10, TH13, Wu08]. censoring [AS15, SZVHA10, WC03]. Cent [HKWE14]. Central [MN06a, MN08a, Per96, TbG98, ŽD11]. centuries [McG11]. Certain [BD70, Gil05, KsS73, AS68, Gil07, LG10, Sch88d]. CFE [Sta10]. Chain [KV14, CS09a, HCS09]. Chains [BHKR11, JKK^{+}08, JKK^{+}09, KBHR10, RS19]. challenge [Bro80b, Lag10]. chance [Tij04, Tij07, Tij12]. change [Nig05, STA11, Zip29]. changes [FMS10]. changing [Bro81]. channel [SPGC10]. channels [MWS99]. Chaos [LFY15, Sch91b, KA84]. Chaotic [KN90, KN91, NNN89, Nic88, NT89]. Chapter [NEM09]. Characteristic [Egg10a, XS09]. Characterization [All97, BE14b, BE16, OJM15, Tag70, TA12, Wój14]. Characterizations [BCCP22, TH13]. Characterizing [CNS15, JK08]. Chart [Wil06]. cheats [Sch06]. check [LSS00]. Checking [CEZ10]. Chemical [YSC13]. Chess [BT09]. chiffre [FL84, FL94, FL96, Hil99b]. chiffres [Var99]. China [Bar11, BM08, Hol14, HNY20, Soo13]. Chinese [GB13, Pen10, SW10, ZCW09]. Choice [Kre73, Kre75, SID13, SID15, Hop16]. Choices [Kub77]. Chosen [Sim98]. cifre [Gin57, Her57]. ciphers [Was21]. Circular [Haf69]. citation [HHC^{+}14]. Citations [Mir16b, CC11]. Cities [Bat08, BS00, Gab99, IO03, Knu01, MPS11, MG07a, MAC14b, MR06, Pen10, Sem08, Soo05]. citing [Mir16b]. City [DKM07, ZM97, BBL07, CZ04, GLS06, RC23, VB08]. city-size [CZ04, GLS06, RC23]. clarifying [WSB15, WSB16]. Class [Sim55, CHL19, Man59]. classes [Bro80b]. Classic [Gor12b]. Classical [BH10a, BH10b, OV97, RRZR09, Uze10]. Classification [IHWZ17, IG08, LB06a, LY02, CL05, MRM^{+}05]. Classification-aware [IG08]. classifiers [CL05]. Classifying [MPM02]. Clinching [GML15]. Clinical [Mac14a, Ora17, TMS02]. Clintons [She94, TS94]. Cloud [Zom14]. CLT [AP10]. cluster [KSH11]. Clustering [AVV08, BCK12, BCM02, BM08]. Clusters [SPGM12]. co [Aus14]. co-authorship [Aus14]. coagulation [Whi82]. coauthorship [MG07b]. COBIB [Pet08]. code [McG11]. Coding [Lin20, MBG^{+}95, WR03]. coefficient [CGJ10]. Coefficients [ARS10, GS01, JTY14, MWX^{+}20, ARS11, FSS07, HW15, JTY16, Sar73, SN91, SF11]. coexisting [Nic88, NT89]. Cognitive [Kle17, Van02, AVV08]. coherent [NS96]. collaboration [MG07b]. collection [FGZZ11b]. collisions [Kim12]. colonies [CLRTFM08]. Color [BEJ07]. Columbia [SDK14]. Columns [LLD15]. Combat [Wal03]. Combinations [Sim98]. combinatoric [Pet81b]. command [Che91]. Comment [Meb11, Sti87, SS09, Vos96]. Comments [LWY10, Sco13]. commerce [HKWE14]. commercial [GHP09]. Common [Ban13, For10, BCC05, FC20, RH02, Urz11]. Communication [Egg10b, EG12b, MM86, Ano05, CMFS11, DMJH11, KA15, MDJH05]. Communications [Gro67, Hub78, KH95, KG72a, Lou92, MK79, Mur73a, Nic87a, Pet81b, Sch75c, Voo74a, BOT06]. Communities [Vla72]. Companies [Vuk08, ŽD11, OTT99, Sav06, YD15]. comparative [SH10]. Compare [AP10]. Comparing [Lan04]. Comparison [LW14, Pra88, AYS14, KG72a, KG72b, NG04, SF08]. competing [HSM11]. competition [KZ15]. Complementing [Rou15]. Completely [KP05]. completion [Egg90b]. Complex [Ada11, MK09, NBH^{+}98, GM97b, Pai08, SAG^{+}00, Uze10]. Complexity [BB76, SG00, GSW92, KMN09, WZS12]. Compliance [Nig96b, dVM13]. components [BS00]. composition [Els78, Hil70]. compressed [MTT13]. compressible [TGD05]. Compression [CS09b, Bak07, McA05, QZH10, ZHS09]. Computation [DW15, Gut16, Ste74, Won69, Zen11]. Computational [NSGP05, Hea78, NSGP05]. Computer [CLT20, DN00, Lan03, OM11, RK79, RVAN15, Tur87, Hor09b, Tic88, XWLD11]. computer-oriented [Tic88]. computergerechte [Tic88]. Computers [MPM02, Lin86]. Computing [HM65, Sch88b, LWY10, MAE06, Rai03]. Concentration [Nal03, Aue13, Egg05c]. concentrations [Bro05]. concept [HN05]. Concerning [Kui69, Sco13]. Conclusions [Yul25]. condensate [Mas06b]. condition [FMSV24]. Conditional [BKS05, AG05, GGO15, GGA03]. conditions [BSO10]. Conference [ACM00, And09, AW89, GCB^{+}05, NKZ12]. confidence [Fer13]. confirms [Aro13]. Conform [HR10, San02a, San02b, San02c]. Conformity [Nig11a, CM21, KK20, MAC14b]. Confronting [Kru96a]. congruence [FMSV24]. conjecture [Ben75, CL04b, Hai69, TW71]. connected [Hai69]. connection [MS11, Orl76]. connections [EC12]. connues [Rou76a]. Consensus [Lee18]. consequence [Egg11, TbG98]. Consequences [Egg85, FRB05]. conservatism [LWFW14]. considerations [WSB15, WSB16]. considered [And09]. Constant [Hav03, AS15, Ale05]. constant-partially [AS15]. Constants [ALT16, Ben43, GM08, BK91]. constrained [Izs06]. Constraints [LP15, MZLH15a, MZLH15b]. Constructal [Pér07]. Constructing [Cha11, Kaz23]. construction [ZF15, ER02]. contain [Ora17]. content [PQ00, SLS^{+}16]. content-aware [SLS^{+}16]. Context [NKRS15, Rav08]. Context-Free [Rav08]. continued [JL08, Sch90]. continues [JL08]. Continuous [BGTM^{+}11, Nig00a, BBC^{+}18, Egg05b, Egg05c, Mei09, Wój13a, Wój13b]. contour [RV09]. Contrast [AP10]. contributions [Jan12]. control [DdA^{+}09, NKZ12]. controlling [Ole10, Ole14]. Controllinginstrumente [Ole10, Ole14]. controversy [McG11]. convenience [MTTT08]. converge [Wój13a, Wój13b]. Convergence [FG76, Sch88c, Dur74, Giu11, Yua84]. convergents [JL08]. converges [KR15]. Convexity [RRZR09]. convolution [Egg94, NK07]. cook [FMS10]. cooperative [Liu15]. copula [DM19]. corporate [Ana03, Lim06]. Correction [Kre75, QBP^{+}04, RME98]. correlated [KUY05]. Correlates [BEJ07]. Correlations [CMHS95]. Correspondence [Hay09]. Corrigenda [AG72]. corruption [Sto16]. cortex [BFLB10]. cosmetic [AG08, LW14]. Costs [ZM08]. could [Mül11]. count [Meb15]. Countries [MS13, vZ13]. country [Soo05]. Counts [DCH^{+}17, Meb06a, Meb06c, MK10, Meb10d, Rou92, Tay05, Yeh07]. County [GM22]. County-Level [GM22]. Cours [Lef82, Par97]. Course [Par97]. Court [Lee18]. CoV [GBF21]. COVID [GBF21, KO20, LHJ20, MM21, VW22]. COVID-19 [GBF21, KO20, LHJ20, MM21]. CPAs [Nig99b]. cracked [McG11]. Creativity [Ben84]. credibility [FGZZ11a]. Credit [HH13, ACC16]. Crime [Aro13, HR10]. Crime-fighting [Aro13]. Crisis [HH13]. Critical [LK12, MNW^{+}05, NH07, Rob11, Sta99]. criticality [BTW87, Bak96, SM96]. critique [Man59]. Croatia [Vuk08]. Cross [DM92, LZNR11, MG22, SF08, Soo05]. cross-country [Soo05]. Cross-linguistic [DM92, SF08]. cross-sections [LZNR11]. crude [BCK12]. crystal [Wap12]. CS [RVAN15]. Cubes [Ros12]. Cult [ZM08]. Cultural [NKRS15, BCM02]. Culture [BM08, MSA^{+}11, ZM01]. cumulative [SZVHA10]. curiosities [Wap12]. curiosity [STA11]. Curious [Cou10]. current [Bar11, MWS99]. curve [Egg06, LHJ20]. curves [ANRV07]. Customs [BCCP18, Lac18]. Cut [Gai06]. Cut-out [Gai06]. cyanobacterium [CLRTFM08]. cycles [Sim07].

D [CNS15, LP15, PGLL10]. D. [KPK05]. D.C.L. [Doo10]. damage [HB18]. Damaso [Key36]. Dangers [CT05]. Darmstadt [RVAN15]. Darstellungen [Tic88]. Data [AR13, BHM08, BS10b, CH97, CM21, CSN09b, Con10, DG10, Die04b, Die07, DHP04, EHM^{+}14, GGP09, Gra10, Gut16, Han07, Hil99a, HHG14, JS09, Kle17, Kos12, Kos13, KV14, Kri10, Lan04, LY02, LB05, Lu07, Mac14a, MRR^{+}10, MM21, MS13, Mil15c, Mil15d, MWE95, MDDK02, NM07a, NM09, Pre81, RGBE11, RGL14, VW22, Wag10, Wan11b, WG68, WAT15, YSC13, dMH06, AS15, AYS14, AW89, AP89, AHI14, AHI15, ACC16, Bah15, Bia15, Bro05, Bro07, CP23, HSR^{+}08a, CCC09, CSN07, CSN09a, CA12, DD06, DdA^{+}09, EG12a, EG12b, FGZZ11a, FGZZ11b, GW04, GD14, GJS07, Gro67, GTM12, GSB24, HSR^{+}08b, Has02, Has03, Hil96, HV01, HRJB02a, HRJB02b, HNY20, HDV13, HSM17, JS06, KG72a, KG72b, LCJ10, LFCR04, LWY10, Mar12, MK79, McA03, McA05, MAE06]. data [Mir14, MF21, NM06, Ora17, OMNO10, PT11, QW03, RRRB11, Row05, Sav04, Sav06, Sch10, Sea02a, Sea02b, SS12, Sto16, SCE03, TMS02, TGD05, VA99, Wal02, WYTD93, WDY93, ZIWU14, dVM13, Die04a, Ole10, Ole14]. data-based [GJS07]. Database [Fed82a, Fed84, Fed87, IG08, LD04, PS09, Par75]. Databases [DCH^{+}17]. DATAS(R) [Lan00]. Daten [Bah15]. Datenfälschung. [Die04a]. DCT [SF11]. DDR [Bah15]. Death [DCH^{+}17, Hen11, KPH10]. debugging [CL04a]. Decade [AR13, CS11]. Decay [NWR09, BMP93]. Deception [Han07]. decimal [FL84, FL94, FL96, FL94, FL96, FL84]. Decision [ANRV07, BXK11, CS09a, CL05, HCS09, Wal02]. Decision-making [ANRV07, CS09a, HCS09]. Decisions [BG54]. Decomposition [HHG14, FMSV24]. Decompositions [BDE^{+}14, BDE^{+}17]. decompressed [LYH15]. d’économie [Par97]. deconstruction [LAL15]. Decreasing [Wan11b]. deep [KCCT23]. default [ACC16]. Defense [SJP^{+}15]. Deficit [RGBK14]. defraud [Rou10]. delay [LC15]. delight [War03b]. della [Bra76, Her57]. demand [NNF21]. Democracy [CS11, BG11b]. Democratic [Meb10d]. demography [Tag70]. Denetiminde [AA10]. Denmark [Knu01]. dénombrable [Rou76a]. dénominateurs [JL08]. denominators [JL08]. densities [MS11]. Density [She68, AG05, GGA03]. d’Entretien [Lef82]. denumerable [Rou76a]. Depend [GS01]. Dependence [WDB95, Art94, DB14, Egg06, Los01]. Dependent [BGTM^{+}11, BBC^{+}18, LL08, Yeh07]. Derivation [BH10a, BH10b, Kaf09b, BJS06, Dea93, Hil95c, Hum08, Lip09, RR09, TR91, vLvdW05]. Derivative [HH13]. describing [Mac22]. Description [Fai69, Fai05]. Design [KV14, LHK^{+}15, Tur87, Ben75, CL05, HKL^{+}06]. designed [LACL14]. d’Estoup [Pet73]. Detailed [BBK^{+}15]. Detect [CT05, Die04b, Die07, GP13, Vuk08, Ano11, FGZZ11b, Lac18, Lan00, Nig94, Nig00b, Rou10, Sav06, TB00, TMS02]. Detected [HKKS06]. détecter [Rou10]. Detecting [AHH08, BG09, BG10, BG11a, DMO10, DHP04, GW04, HDV13, JS06, JS09, KB07, Lan03, LFCR04, LB05, Meb06a, Nig15, RGL14, Sta12, SCE03, WSY11, BKS05, HSM17, MF21]. Detection [ALY19, Ban00, BCCP18, BH02, BPS11, CS09b, CT22, Cod99, DMO11, DJ10b, Kle17, KV14, LZL^{+}12, MWX^{+}20, Meb11, Nig12, NEM09, RRRB11, SLS^{+}16, SG16, Smi13, ZHS10, AOT13, BCP21, BM13, BMM10, CBC^{+}19, CG07, DG10, DMF09, Gra10, GHP09, GD14, HYYH08, Kos15, Kri10, KB03, LYY09, LFCR04, Low00a, LBC06, ME22, MTT13, NHD16, Nig92, PT11, Sch10, ZHS09, dGGd21, HB18, BMM10]. Detective [Kos12]. detector [STA11]. detects [SS11]. determinant [Zip29]. Determination [She12]. Deterring [AHH08]. deutschen [QW03]. deuxiè [Poi12]. developed [LW14]. Developing [MVW^{+}03]. Development [Jin09a, LD04, WDB95, ZM97, Che12c, MZ98a, NGJ03]. developments [CH61, eSHK05]. Deviation [LZZ13]. deviations [HSM11]. Devil [BS09]. diagnose [Meb15, Die02]. Diagnosis [Die02]. Diagnostics [BVT96, MK09, NM09]. diagram [AI99]. dictionaries [Mas05c, Mas05b]. Dictionary [LZZ13]. Did [WHN13, VM04]. Difference [BE14a, LCY^{+}15, Meb10b, BS07, WR03]. Different [HKKS06, Afi10, BP84, Egg90a, Nag05, New81]. Differential [CGPT15, GÁVRRL14]. Difficulties [BG09, BG10, BG11a]. Difficulty [Hil99a]. Digit [ABBM10, BS10a, CS09b, Die04b, Die07, Dun69, Gor12a, Hay08a, Hay08b, Hay09, Hil98, KC86, KBRV04, Kun87, Lin08, LB07, LG78, LMH99, ML84, Meb08a, MK10, Meb12, MK13, Pos04c, Rai85, Rod04, Sal97, SM10b, SM10c, Shi09, SM11, Sri13, AS68, Adh69, AP99, Bia15, Boy94, BMP93, BE69, Cio05, Coh76, CK84, DMF09, FGZZ11a, Fle66, FL84, FL94, FL96, Has03, Hil88, Hil95b, Hil95c, Hil99b, Hum06, Hum18, Hür15b, Hür15c, KNRS88, KR13, KR15, Kre02, LZL^{+}12, LL08, LL09, LS11, MM15, MS11, MS12, MS15, Meb06b, Meb08b, Meb08c, Meb10b, Meb10d, Mir12a, MAC14b, Mir16a, Mül03, NKS90, Nig00b, Poc06, Pos04a, Rai76, Rou14, SM09, SM10d, SM10e, Sta05, Sur93, TFGS07, XWLD11, Pos10]. Digit-Based [BS10a]. Digital [AA10, BM13, CNS15, CT05, Cod99, HR10, LOR06, LP15, Nig96a, Nig97, Nig00b, Nig00c, Nig01, OB13, RGL14, Shi09, BBE10, BBJ11, Lan00, Nig92, Nig94, Nig99a, RR03, SB15, SH10]. digitalen [BBE10, BBJ11]. Digitized [MSA^{+}11]. dígito [AP99]. Digits [BS09, Dav76, FG76, FM95, Fri84, HH13, Kos06, Ley96, Meb07, MWE95, MDDK02, Nig11c, PU86, Pin61, Pre81, Ros12, Sli98, Sun28, Tsa74, Web75, Whi72, Won69, BHS11, Ber04, Ber05a, BK91, Dia73, Dia77, GSR12, Gin57, Her57, HS05, HW86, Hün07, Irm97, IRU83, Jec92, KR81, KR04, KB07, LCJ10, Lem86, LS14, MM50, Meb15, Mir11, Mir12b, Mör01, New81, Nig99c, Pha13, PK06, Rai69a, STA11, Sar73, Sti45, TW71, Tur82, UP85, Var99, War03a]. Dijital [AA10]. Dijkstra [FC14]. dimension [Mas06b]. dimensional [Ber05c, BBH05, Has07, MS08, SCD01]. dimensioni [Bra76]. diophantine [Tic85]. Direct [Kaf09b, Lu07]. Dirichlet [PR22]. Dirty [NKL03, NKL04, MA18]. disaster [Mül11]. disciplinary [Boo94]. disclosure [Moc02]. discovering [GÁVRRL14]. discovery [BP06, OMNO10]. discrepancies [ACC16]. discrepancy [Kem75]. Discrete [BE14b, KP09, BE16, Egg05b, KPF15, Mei09, SCD01, Yeh90, Yeh07]. Discrete-Time [BE14b, BE16]. Discussion [Gra10, Kri10, pLCjW04, Möl09, ZIWU14, Val12]. Dishonesty [Han07]. disparate [Par75]. distance [Cam24, Hav95]. Distances [HF16]. Distinguish [IKC96]. Distinguishing [LFY15]. Distortion [BHM08]. distractors [Hop16]. Distributed [Blu70, DW15, LEC^{+}15, MYP14, HG14, KR13, MN14, Sch89]. Distribution [AMP06, AS68, Ami07, Ant91, Axt01, BT09, BGVV99, BD70, CH97, CT07, Coi77, EHM^{+}14, FG76, iCE10, FGPM12, Fri84, FH45, Ham70, HW87, Ism04, Jee97, KC86, Kui69, KPK06, Ley96, Lot26, MSSvK08, MYZ67, NP12, PU86, Pav82, PTTV98, Pin61, Rai69b, Rav08, SDK14, Sch73, Sch88b, SM10b, SPP17, Sim55, Sim98, Tur87, Vla79, Web75, Won69, XP89, YZ16, ZSB^{+}12, AAMAE03, ADAI13, Adh69, Afi10, Asa81, BNS10a, BNS10b, Bel88, BBL07, BS07, Bra76, BK91, Bur91, CH61, Cig64, Cio05, CLM15, Dia73, Dia77, Die04a, DN10, DN11, DS01, DHL^{+}20, EMS11, EMS13, FWW85, FP04, FGPM16, Gar11, Gun15, GTC^{+}07, HZZ14, HW86, Hol69, Hum06, IMS15, Irm97, Jec92, KSG^{+}16, KR81, KR04, KPF15, KN74, KN06, KPK05, LCJ10, Lem86, Li92]. distribution [Loy73, Mac22, MM50, MML01, MPS11, Man59, Man60, MRSCCS11, MK79, Mir12a, MAC14b, Mir16a, Mor81, Mör01, NG04, OJM15, OTT99, OV97, Orl76, OC83, OM11, Par75, Pav81, Per90, Pet03, PTTV01, PL86, Pos04b, PK78, Rai69a, RAV10, RDFM10, SZVHA10, Sch81, Sch83b, Sch83a, Sch84, Sch86, Sch87a, Sch87b, Sch88d, Sco13, Sem08, Sha05, SM09, SS12, SBH^{+}95, Sti45, SM06b, Tao09, TW71, Tic85, Tic86, Tic87, Too92, TKP14, Tsu52, Tur82, UP85, VAGE09, Vol20, WZS12, War03a, Wey15, WC03, Wu08, ZM01, ZDCW12, ZCW09, dSFSL20, dCS15, vZ13]. Distributions [BB85, CSN09b, DKM07, Dun67, Fai69, Fai05, For10, JL08, JKK^{+}08, Kaf09a, Kon65, LSE00, LB05, Mit03, Mor06, Pre81, Rod04, SM10a, Tsa74, Whi83, Abe02, Bas92, BCC05, BSM14, BG78, CZ04, CM11, CGJ10, CSN07, CSN09a, ER02, FC20, GGPS03, GLS06, GJS07, GLSW96, GN07, Hai69, HT05a, HT05b, HT06, Hil70, Hil96, HRJB02a, HRJB02b, HDV13, Hür15c, Izs06, JKK^{+}09, JdlR04, Jon02, KK20, Kaz23, KUY05, KS10, KG72a, KG72b, KP09, KP04, MBS99, MPS11, Mas05a, MM06, MS83, NK07, New05, Nig92, Per10, Ree06, RS19, RC23, Tag70, TA12, TH13, cTnHwH17, Yeh02, You13, ZA95, Gre86, Sha89]. distribuzione [Bra76, Her57]. Divergence [IHWZ17]. Divergent [Har49, Har91]. dividends [AVV08]. dividing [GSR12]. division [Per96]. DNA [Vos96, BHB^{+}96, GÁVRRL14, IKC96, MBG^{+}94, MBG^{+}95]. Do [AVV08, iCE10, FDA^{+}04, IHPS14, LLD15, MS13, San02a, San02b, SPP17, TBL00, WHN13, TFGS07, Vog00, San02c]. Doctoral [Sud10]. doctored [LYY09]. Document [BMLRV10]. Documentation [Fai69, Fai05]. Documents [YLX14, Pet73]. Does [Ber10, GT09, GTC^{+}07, Hay08a, HR10, SW10, DB14, Moc02, SM11]. domain [NHD16]. dominates [Car10]. done [Meb10d]. Double [CS09b, OR07, MTT13, TRB^{+}12]. doubly [Wu08]. Dow [LV94, Nig98]. Dowe [McA05]. down [McG11]. Dr. [Yul25]. Drivers [DKM07]. Droplet [Kim12, FM89]. drug [OMNO10]. DSS# [HYYH08]. dual [Egg86, ER02]. Duality [ER02]. dubious [KB07]. d’un [Rou76a]. Duncan [Kui69, KsS73]. d’une [Rou76a]. Duplication [Nig11c]. During [HH13, AA11, MK13, OM11]. Dynamic [CCT95, KNT89, LEC^{+}15, LP15, BBL07, Che95, OV95, Par75, Zip35, Zip65]. dynamical [Ber03, Ber05b, Ber05c, BBH05, Ber11, MG15, Nil10, OR07, SCD01, TBL00, Wir98]. Dynamics [Ber04, Ber05a, KN91, LC15, MZS99, ACCG13, AFS04, HA99, KN90, dLKHS04a, dLKHS06, NNN89, PQ00].

e-commerce [HKWE14]. Earnings [AA11, SGW04, Van02, Vuk08, YD15, AG08, DZ03, LW14, LWFW14, Nig05, Tho89]. earthquake [BCC05, Has07]. earthquakes [SS11]. Eases [Ban13]. East [ŽD11]. Easy [Ban13, RR03]. easy-to-use [RR03]. eBay [Gil07, LG10, Gil05]. échantillon [Rou76a]. ECML [GCB^{+}05]. ecology [HSM11]. econometrics [Jin09b]. Economic [GT09, KCCT23, MS13, RGBE11, CP23, ZCW09, Sch10]. Economics [Ism04, Jin09a, NCT10, NCT13, Töd09, Jin09b, Wan11b]. economies [SR01]. économiques [Gib31]. economist [Key36]. Economy [Kru96b, NM07b, Art94, Par97]. Econophysics [ACCG13]. ecosystem [HJ15]. ed [Krä15]. eddy [LF16]. edge [Hol18]. Editor [Goo65, Hil11, Bas94, Bro79, Bur04, Che95, Hay08a, Hub77, Kri77, Lon92, OV93, OV95, Var72]. Editorial [BCP21]. eDonkey [HKL^{+}06]. Education [RVAN15]. Effect [KNT89, QBP^{+}04, WCP^{+}07, DB14, EKJ^{+}20, GLSW96, Her10, SF08, Shi89]. Effective [DHP04, MWX^{+}20]. Effectiveness [RGBK14]. Effects [May08, Mil57, RME98, LF16]. Efficient [Bac98, DN10, DN11, RZS^{+}15, LWY10, MAE06, Par75, Urz00]. effort [FS03, LAL15, Zip49, Ano50]. eight [Ora17]. eighth [PS09]. Eins [Hum18, Wey15]. Einsatz [BBE10]. Einstein [Hol18, Mas05a]. Election [AHH08, Ans09, BS10a, DMO11, Meb06a, Meb06b, Meb06c, Meb08a, Meb08b, Meb08c, Meb09, Meb10a, Meb10b, Meb11, Meb12, Mebxx, Bat09, Her10, Rou09, Rou14, SM11, Sjo14, Meb10c]. Elections [DMO10, GM22, Meb06b, Meb08b, BG11b, Meb08c, Meb15]. Electoral [AHH08, KM11, MK09, MK10, Tor06, BG11b, LB14, Nor13, PT11]. electronic [KH95]. electronics [NKZ12]. Electrons [BRS38]. Elegant [Da 20]. elementarmathematische [Hum08]. Elementary [Gou77, Hum08]. Elements [Kle17, LWY10, MBS99, MK79, MAE06, Pet03]. elicitation [XLW14]. elles [San02c]. Elucidating [KSG^{+}16]. elusive [Bas94, OV93]. Embedded [LHK^{+}15]. Emerald [Row05]. emerged [McG11]. Emergence [BA99, CMFS11, RDFM10, SL96, FGPM16, MG07a]. emerging [LW14]. emissions [ZIWU14]. empirica [Bra76]. Empirical [CG77, CSN09b, Fai69, Fai05, GM97a, LSS00, MSSvK08, Möl09, RGE13, RGBK14, SF01, SM10a, TGD05, Van15a, Van15b, ZS11, Bro84, CSN07, CSN09a, DG78, GJS07, HSD12, IO03, LW14, McA03, McA05, MW04, Nic89, Pao86a, Pao86b, Rou05, Tay05, Bra76]. employee [HSR^{+}08a, HSR^{+}08b]. employee-reported [HSR^{+}08a, HSR^{+}08b]. Encoded [MPM02]. ended [Sto16]. Energy [LP15, MZLH15a, MZLH15b]. Engineering [ACM00, Moo36, LSS00, PGLL10]. engineers [Ham62]. English [NG04]. ENIAC [She12]. Enigma [McG11]. Enron [Nig05]. ensembles [BP06]. Enterprise [SJP^{+}15]. entlarven [Sch06]. entries [DG10, Gra10, Kri10]. Entropic [RRR04]. entropies [Abe02]. Entropy [Cio05, HT01, Kaf09a, Kaf09b, Abe02, HT05a, HT05b, HT06, LAL15, Liu08, Yav74]. entry [dGL13]. Entwicklungen [CH61]. Environment [Fed82a, Fed87, Wan11b]. environments [LC15]. episode [Nig05]. Eponymy [WHN13]. EPS [DZ03]. Equations [BE14a, BG15, BS07, Whi82]. équidistribuées [Jol05, Jol09]. equidistributed [Jol05, Jol09]. equilibrium [Kim12]. Equipartitions [IMS15]. Equity [LK12]. Equivalent [Pet81a]. era [AA11]. Ergebnisse [Die04a]. ergodic [Jan12]. ergodicity [Dur74]. ergodique [Jan12]. Erratum [Wat96a]. Error [BB85, Bor20, Lon92, Sti87, ZM08, BCIS09, DL07, DL08, Lou92, Sav06, SCE03]. error-prone [SCE03]. Errors [CT05, Tsa74, Die02, Nig00b, Nig15, OR07, SHC^{+}12, Sea02a, Sea02b]. Erscheinungen [Fur46, Gei48]. ersten [Hum06]. Esperanto [MPPH06]. Essere [CEZ17]. Essex [Rai03]. established [BG11b]. establishment [KLS23]. estimate [Gar11]. Estimates [FG76, Sch87b]. Estimating [AS15, QZH10]. Estimation [AMP06, ADAI13, AG70, AG72, BVT96, CT07, GMV09, HW87, Jee97, Nic87a, Nic87b, Afi10, AP89, DN10, DN11, HZZ14, Izs06, KS10, Sco13, SS12, Wu08]. Estimator [Ism04, GGO15, HDV13]. estimators [TKP14]. Estoup [Pet73]. ethics [Wan11b]. EU [RGBE11]. EU-Governmental [RGBE11]. Euclid [Lin86]. Euclidean [Cam24]. Euler [Ale05, Hav03, LS14]. euro [eSHK05]. Eurographics [NSGP05]. European [GCB^{+}05, DK09, FDA^{+}04, FAD^{+}04, LK12, ŽD11]. evaders [Hün07]. Evaluating [RBXM15]. Evaluation [CZ98, Fed87, KŽ08b, LK12, NHD16, dMH06, CL05, FGZZ11a, Fed84, MM82, May08]. Evasion [CG93, Nig92]. Eveiller [GD10]. Events [CYG07, MZ99]. Everyday [Wan11b, Tij04, Tij07, Tij12]. everything [Cio05]. Everywhere [Mur23]. Evidence [BCM02, BM08, Car88, CG93, Mac14a, MS13, RGBK14, Soo13, YZ16, ZIWU14, LW14, Mar12]. EvoBIO [Rai03]. EvoCOP [Rai03]. EvoIASP [Rai03]. Evolution [GB13, KM11, Yul25, CGPT15, CMFS11, GÁVRRL14, SM96, TR91, TST97]. Evolutionary [MVW^{+}03, OKO06, Rai03]. EvoMUSART [Rai03]. EvoROB [Rai03]. EvoSTIM [Rai03]. EvoWorkshops [Rai03]. Exact [Lei80, KZ15, Mor81]. Examination [MDDK02, Rob97, Sun28, BBJ11, DG78, IO03, NH07, Pao86a, Pao86b, Sav06]. Examinations [Lan04]. examined [MG22]. Examining [Has03]. Example [BS10b, GD10, Moc21, Per10]. Examples [MWX^{+}20, Dur91, Dur05, Got92]. exams [Hop16]. Excel [RR03]. Excessive [Man08]. exchange [Car15, Sav06]. excursions [Hol18]. exemple [GD10]. exercices [Est16]. Exercises [SID13, SID15, Est16]. Exhibit [iCE10, Li92, Par75]. existence [Lol08, Sim07]. existing [LX05]. Exoplanets [SPP17]. expanding [CI96]. expectations [HW15, Wap12]. Expenditure [RGL14, MTTT08]. experiment [CS09a, HCS09]. Experimental [Da 20, Hsü48, dLKHS04b, MNW^{+}05, Da 22]. Experimenten [Die04a]. Experiments [LHK^{+}15, Sun28, Die04a, FM89, LACL14, Wan11b]. Explaining [PTTV01, Smi97, Egg10b, Egg10c, GLS06]. Explanation [AP10, Few09, SF01, BTW87, BH11b, Bha12, Coh76, Egg94, Gab99, OKO06, Ree03, TW11, ZCW09]. Explicit [DL07, DL08, Zör15]. exploratory [DG10, Gra10, Kri10]. Explore [YLX14]. Exploring [Ale05, FNP^{+}15, Hav03, LB07, VW22, Cha11]. exponent [cTnHwH17]. Exponential [HW15, KC86, Abe02, EL03, HS01, KS10, MML01]. Exponentiated [ZSB^{+}12, Afi10]. exponents [RMB^{+}14]. expose [Sch06]. Expression [FK03, RV98]. Expressions [ALT16, JP01]. extended [MRSCCS11]. Extension [Pra88, AG05, BKS05, GGA03, Ree03, SB93]. Extensions [Sch08]. Extinction [SM96, HW15, NP02]. Extraction [YSC13, LYY09]. Extrasolar [SPP17]. extraterrestrial [DMJH11]. Extremal [MN14]. extreme [GMV09, Per82]. eyelids [BBK^{+}15].

F [KPK05]. F. [Pav81]. F.R.S [Doo10]. F.R.S. [Yul25]. fabricated [Hil96, HSM17]. Fabrication [MWE95, Mar12]. fabulous [Har09]. Face [MM87, Sti87]. Fact [RGBE11]. Factor [Vla76, Egg13]. factorials [Sar73]. Factorization [LS11]. Factors [Vla72, BCM02, CC11]. Facts [DJ10b, Lin86, SM11]. Fahrtenbüchern [BBJ11]. failure [Ano93, Tra92]. fair [HG14]. Fairthorne [Rou05]. fait [Hil99b]. fake [Pos04c]. Faked [SSMW05]. Faking [Hil99a]. Fälschungsaufspürung [Pos04c, Pos05]. False [Ioa05, Lin08, Per05]. Falsification [MK09, Die04a]. falsified [HZKS12]. Falsify [MS13]. Familia [dGGd21]. Families [Mor06, RH02]. family [GJS07, Hür15c, ZM01]. far [OV93]. Fast [MS15, Wój13a, Wój13b]. fat [Yeh90]. fat-tailed [Yeh90]. fault [SM06b]. Fear [Lan03, Hün07]. Feature [CCC09, ME22, YSC13, NHD16, QZHC11, RRRB11]. Feature-preserved [CCC09]. Features [MBG^{+}94, Vos96, Egg94, FNP^{+}15, KMN09, LZL^{+}12, Par75]. Federalism [KM11]. Federico [Key36]. feedforward [VK90]. Fehlerquellen [Die02]. Feller [BH10a, BH10b, Val12]. Fibonacci [Har09, BHS11, Dun67, FM95, Kui69, LS14, She68, SB93, Was81, Web75, Wlo71]. Fiction [RGBE11]. field [CS09a, HCS09, Sea02a, Sea02b]. Fields [Da 20, Da 22, Per10]. Fifteenth [AW89]. Fight [Mil15c, Mil15d]. Fighting [AH09, LOR06, Aro13]. fights [Dub01]. Figures [Ans09, FH45, GF44, Rai69b, Wel05]. File [Fed87, HKL^{+}06, RH02, XS09]. Files [LDJ89, SPGC10]. filings [HSD12]. Filtering [LP15]. Final [Pre81, Sun28]. finance [CG07]. Financial [AR13, Ber95, BPS11, CT22, DG12, EHM^{+}14, GR07, HH13, JJ21, RGE13, Van15a, Van15b, AI99, BKS05, BS00, CA12, CEZ10, FGZZ11a, GGPS03, GD14, HSD12, HS01, KCCT23, KB03, Mir12a, Mir16a, RRRB11, RC18, RR03, SHC^{+}12, VA99]. Finanz [SHC^{+}12]. Finanz- [SHC^{+}12]. Find [LCY^{+}15]. Find-the-Difference [LCY^{+}15]. Findings [Ioa05]. finds [Mar12]. fingerprints [SPGC10]. Finite [BHM08, BHKR11, BX18, BX19, KBHR10, NKL04, NKL03]. Finite-state [BHKR11, KBHR10]. Finitely [BE69, Jec92]. FIR [LP15]. Firm [Axt01, KLS23, dGL13, HNY20]. firm-level [HNY20]. Firms [Bal66, Wan11b, AA11, Bra76, FDA^{+}04, FAD^{+}04, Fuj04, SBH^{+}95, ZCW09]. First [CS09b, Die04b, Die07, Fri84, Gor12a, Hil98, HH13, KC86, KBRV04, Kre02, Kun87, LG78, LMH99, PU86, Pin61, Rai69b, Rai85, Rod04, Ros12, SM09, SM10b, SM10c, SM10d, Sri13, Web75, AP99, Bia15, BMP93, BE69, Cio05, Coh76, CK84, FGZZ11a, FL84, FL94, FL96, Hil88, Hil99b, HW86, Hum06, Hür15b, Hür15c, Irm97, KNRS88, KR81, LCJ10, Lem86, LZL^{+}12, LS14, LL08, LL09, MM50, MS11, MS12, MS15, NKS90, Nig99c, Pos04a, Rai69a, Rai76, Rou14, STA11, TFGS07, UP85, Shi09]. First-Digit [KBRV04, SM10c, SM10d, Bia15, Rou14]. Fisher [HPV^{+}09]. fishery [GHP09]. Fishpole [Ano32]. Fit [BCCP18, KG72a, KG72b, OJM15, SZVHA10, SS06, VAGE09, Zör15]. Fitness [MVW^{+}03]. fitted [vZ13]. Fitting [Bur04, CH97]. Fixed [CKT^{+}09, Gar11, Sav04]. Fixed-Odds [CKT^{+}09]. fixed-point [Gar11]. flag [RR03]. flattening [LHJ20]. Flaws [BH10a, BH10b]. Floating [BB85, Kre73, Kre75, Sch88c, Ste74, FT86, Sch73, Sch86]. Floating-Point [Kre73, Kre75, Sch88c, FT86, Sch73, Sch86]. flow [GMV09]. flows [Ber15a, Bia15, CP23, Mir12a, Mir16a]. fluctuations [GGPS03, MBS99, MWS99, Rob11]. focus [Mül03]. focused [LACL14]. Foil [Ber95]. Follow [Bur10a, Bur09, SPP17, GTC^{+}07, IHPS14, TBL00]. Following [Bro98, LL08]. Follows [Haf13]. food [Liu15, Wan11b]. food-limited [Liu15]. Force [Can03a, Can03b]. Forecasting [GT09]. Forensic [Bha02, BK08, Cal13, Iud12, Nig93, Nig11e, Nig12, OW12, SG16, BL95, GBF21, Kos15, KB07, QZHC11, SS10]. Forensics [Meb06b, Meb06c, Meb08a, Meb08b, Mebxx, Piv13, Shi09, ZHS10, Ana03, DG12, FSS07, Lim06, MS09, Meb08c, QZH10, SB15, TRB^{+}12, WCCK09, XWLD11, ZHS09]. forgery [BM13, MS10]. Form [Pet81a, Che80, Hil74, Whi82]. Formal [MM86, vLvdW05]. Formation [ZM97, PK06]. Forms [ARS10, ARS11, HW75b, Hub78]. formula [Dub01, Has76, Mas06e, Mas06d]. Formulation [Lei80, Asa81, PK78]. Forschung [Die02]. Forum [BCP21, Dun80, Moh80]. Foundation [Fed82a]. foundations [Bro80a, Bro80b, Bro80c, Bro81, Dur06]. Founded [Da 20]. Fourier [Boy94, MWX^{+}20, Mil15b]. FPGAs [WAT15]. Fractal [Den97, CZ04, Che12b, Gai06]. Fractals [Sch91b, Egg05a]. Fractional [Pav82, Sur93, ER02, Giu11]. fractions [JL08, Pav81, Sch90]. Fragile [ZHS10, PGLL10]. framework [BCD23, Egg10a, Kaz23]. France [NKZ12]. Fraud [AH09, AHH08, Ban13, Ban00, BCCP18, BG09, BG10, BG11a, BS10a, Ber95, BL95, BH02, BPS11, CT05, Cod99, DMO10, DMO11, DJ10b, DHP04, GR07, Han07, Jud04, Kle17, Kos12, KV14, Lan03, Lan04, LB05, Lu07, MK10, Meb10a, Meb10b, Meb10c, Meb11, Meb12, Mil15c, Mil15d, Nig99b, Nig12, NEM09, Rob97, SS10, Smi13, Sta12, Töd09, Wal03, WSY11, Ana03, Ano11, Bat09, BKS05, CG07, DG10, DMF09, Dub01, GW04, Gra10, GHP09, GD14, HB18, HZKS12, HYYH08, Kos15, Kri10, KB03, Lac18, Lan00, LB14, LOR06, Lim06, LFCR04, Low00a, LBC06, LB06b, MF21, Nig94, Nig00b, Nig15, RRRB11, Rou10, Sav06, Sch10, SHC^{+}12, SM11, Sjo14, TB00, TMS02, dGGd21, BMM10]. frauder [Rou10]. fraudes [BMM10, Rou10]. Frauds [KM11, BCP21, CBC^{+}19]. Fraudster [Sta10]. Fraudulent [CS11, Die04b, Die07, GP13, SSMW05, SCE03]. Free [Rav08, SS09]. French [Par97, Bel88, BMM10, DJ10a, DJ13, Dev57, Est16, Fra17, FL84, FL94, FL96, GD08, GD09b, GD10, Hil99b, JL08, Jan12, Jol05, Jol09, Lef82, Pet73, Poi12, Rou76a, Rou76b, Rou78, Rou10, San02c, Var99]. frequencies [Bia15, CM11, FP04, LL08, LL09, NWNG22, Nig94, Zör15]. Frequency [BEJ07, BG78, CYG07, Coi77, DMO06, EC11, Lot26, Vla79, DM92, ER02, iCV18, Gin57, Has03, Her57, Hil74, Li92, Mac22, Mas05c, Mas05b, NHD16, New81, Par75, SEOV04, Tul85, Zip29, Zip32]. Frequency-rank [BG78]. frequent [LWY10, MAE06]. frequenza [Gin57]. Freund [MM99, MM04, MM14]. FRG [Bah15]. front [LACL14]. führende [Hum18, Hün07]. Fujii [Mar12]. Full [SM10b, SM09]. Fully [NKRS15]. Function [BMLRV10, Egg05b, Egg06, ER12, KPK05, KPK06, LS05, LS06, LS14, SZVHA10, Sur93, Wir98]. functional [Egg13]. Functions [ARS10, MVW^{+}03, ML84, Sim55, AS68, ARS11, BFLB10, BNS10a, BNS10b, CLPR23, GSR12, KM05, Man59, MK05, Pos04b]. Fundamental [BH10a, BH10b, DKM07]. Fundamentals [HT01]. Fürchten [Hün07]. Furlan [Gei48, Reg82]. Further [BGVV99, Sen73, Tur84, Val11]. Fuzzy [AP10, Kle17, NKL03].

G. [Mel06]. Galaxies [HF16]. Game [CKT^{+}09, LCY^{+}15, Mor10, RBXM15]. Games [BG54]. gamma [DHL^{+}20, NK07, Hav03, Ale05]. Gammes [Est16]. Gardner [McA05]. gas [Mas06a, Mas06c]. Gaussian [HS01]. Gaussian-like [HS01]. GDP [Hol14]. GDR [Bah15]. gem [BH11b]. Gene [FK03, RH02]. genera [RH02]. General [Ben43, BS10b, Asa81, Kaz23, Kos15, Lol08, Mac22, Mas05a, OV97, GD09b]. générale [GD09b]. generalisation [Par75]. Generalised [ZHS10, ZHS09]. Generalization [Won69, Orl70, PGHA07]. Generalizations [Lag85]. Generalized [AR09, BDE^{+}17, CH97, Gun15, HW87, CGJ10, CMHS95, EG12a, EG12b, FSS07, Hür03, LAL15, LL08, Orl76, Ris08, Sha05, SF11, SS12, TA12, TH13, VAGE09, Wój14, dCS15, vZ13, BP20]. Generalizing [Hür06b, Hür09, LCJ15]. Generate [MWE95, RDFM10]. Generated [LCY^{+}15, LP15, MN14, Wir98, Zör15]. generates [Hub02]. Generating [KN91, XLW14, KN90]. Generation [LB07, YSC13]. Generative [Mit03]. generators [Bac98, Dia73]. Generic [SL96]. Genes [LY02, NP12, KSG^{+}16]. genetic [FC14]. Genome [FGPM12]. geographical [Moc21]. geography [Soo13]. Geometric [MdSSZ09, ZF15]. Geometrical [AAMAE03]. Geometry [NKRS15, Lee15]. Geophysical [NM07a, NM06]. George [Ano50, VW06]. Geosciences [STJ09]. German [Mör01, Szé90, Zel58, ZB70, Aue13, Bah15, BBE10, BBJ11, Bul11, CH61, Die02, Die04a, Fur46, Gei48, GTM12, HB18, Hum06, Hum08, Hum18, Kre73, Kre75, Moc02, Moc17, Mül03, Ole10, Ole14, Pos04c, Pos10, QW03, SSMW05, Sch73, Sch06, SHC^{+}12, Tic86, Tic87, Tic88, Vog00, Wey15]. Germany [BG11b]. Gesetz [Aue13, BBJ11, GTM12, Hum06, Hum18, Hün07, Mör01, Pos04c, Pos10, Sch06, Vog00, Bul11, Pos05]. Gesetzes [Hum08]. Gibbs [Mas05a]. Gibrat [AFS04, FDA^{+}04, FAD^{+}04, Soo13]. Gini [Bur91]. Girona [NSGP05]. given [Par97]. gives [RMB^{+}14]. glare [QZHC11]. Gleichverteilung [Wey15, CH61, Tic86, Tic87]. Gleitkomma [Kre75, Kre73]. Gleitkomma-Arithmetik [Kre75, Kre73]. Gleitkommadarstellung [Sch73]. Global [DKM07, LD04]. go [VM04]. Goal [Car88]. Goal-Oriented [Car88]. Gödel [Hol18]. Golden [Liv02]. Good [Mac14a]. Goodness [BCCP18, OJM15, KG72a, KG72b, SZVHA10, VAGE09, Zör15]. Goodness-of-Fit [BCCP18, OJM15, Zör15]. Got [Ber95, Nig99b]. Governmental [RGBE11]. GRACE [KNT89]. Grain [DB14]. gram [Cha11]. Graph [BGP07, Asa81, BP06]. graph-oriented [Asa81]. Grapheme [BEJ07]. Grapheme-Color [BEJ07]. Graphical [MZLH15a, MZLH15b]. Graphics [NSGP05, PGLL10, XWLD11]. Great [Jud04]. greater [HHC^{+}14]. Greek [Mül11]. grid [LYY09]. Grouped [AG70, AG72]. groups [MM15]. growing [MS15, ZS11]. Growth [Bal66, HA99, MZS99, Ros11, Ano02, FM89, MHS95, MG07b]. Grundlage [Bah15]. guessing [Hil88]. Guesstimation [ABBM10]. guide [GD14].

H. [KPK05, Man59]. H.F.D. [KPK06]. Hacker [War03b]. Hadron [SM10b, SM09]. Half [NWR09, BMP93, FG10, NR08]. half-life [FG10]. Half-Lives [NWR09, NR08]. Hall [DB14]. Halstead [Pra88]. Handelsdaten [GTM12]. harmful [And09]. harmonic [Per82]. Harmoniegesetz [Fur46, Gei48]. Harmony [Fur46, Gei48]. Hash [KNT89, SD78, WYTD93]. Havil [Ale05]. hazard [SZVHA10]. headings [Pet08]. Health [LB05]. Heap [Egg07, LB09, LZZ13]. Heaps [EC12, GS01, vLvdW05]. Heavy [KLS23, BCIS09]. heel [Moc17]. height [Egg06]. heirs [Pus04]. Help [Möl09, Nig99b, KCCT23]. helps [Lac18]. Herdan [Egg07]. Heritage [NKRS15]. Heterogeneity [LEC^{+}15]. heuristic [BP06]. heuristics [GT99]. heute [Sch06]. Hidden [BHB^{+}96, IG08, SPGC10]. hidden-web [IG08]. Hierarchical [Bas92, JKK^{+}08, Che12a, Che12c, JKK^{+}09]. Hierarchy [Kru96a, Sem08, Che12b, ZF15]. Hilfe [HB18]. Hill [HDV13]. hint [Bat09]. History [Mit03, Sch75b, Pet73]. hitters [BCIS09]. Hold [GT09, FDA^{+}04]. holding [YD15]. homogeneous [CW97]. Hubble [HF16]. Human [LZZ13, NP12, Zip49, Bro80b, FS03, OKO06, Ano50]. Humanities [Mur73b, Hub77, Mur73a]. Humans [Haf13]. hundert [Sch06]. hundred [Sch06]. hunted [McG11]. Hybrid [KNT89, GÁVRRL14]. hybridized [FC14]. Hydrology [NM07a, NM06]. Hyperbolic [Fai69, Fai05, HT05a, HT05b, HT06]. Hypergeometric [BMLRV10]. Hypotheses [SBB01]. hypothesis [BCD23, BCK12, DDS98].

i.e [CEZ17]. Ibry [WHN13]. ICSE [ACM00]. ideal [Mas06c]. Identification [IHWZ17, JHTS10, LYH15, SSMW05]. identifying [MS10]. ideograms [NG04]. ideological [Mir11]. IEEE [Ano93, CLT20]. II [Bro80b, Jol09, MS08, RAV10, SZVHA10, Tic86, WC03, Wu08]. III [Bro80c]. illicit [Mir12a, Mir16a]. Illuminating [Moo36]. Illustrates [Nig98]. Illustration [Sta14, BMP93]. im [Moc02]. Image [NHD16, Piv13, QZHC11, SM06a, SG16, ZHS09, ZHS10, BM13, FSS07, LYY09, MS09, MS10, QZH10, SLS^{+}16, SB15, TRB^{+}12, WCCK09]. Images [AS05, DS05, GJARPGAC10, IHWZ17, Jol01, IHPS14, LZL^{+}12, MTT13, PGHA07, PGQA^{+}15, ZHS09]. Imaging [NSGP05]. immune [BMT96]. Impact [KMN09, AG08, CC11, Egg13, MG22, RRZR09]. Implication [Hay08b]. Implications [Kry09, MK13, Mor81, Tur87, YZ16, HKL^{+}06]. Implies [GD11, HF16, Hil95a]. Imply [Ber10, GD09a]. Importance [LY02, CEZ17, CEZ18]. importanza [CEZ17]. Improved [BR15, TKP14, ZIWU14]. Improving [ZHS10, CS09a, HCS09, QZHC11]. impulsive [LC15]. inaccurate [Ora17]. incidence [Nig05]. Income [Car88, ŽD11, Mac22, Man60, MAC14b, Nig92, OTT99]. incomes [RH02, Ree03]. Incomplete [LB05]. Incongruence [GBA04]. inconsistencies [MS09]. increase [BOT06]. Increasing [Art94, Wój13a, Wój13b]. independence [DM19]. Independent [KC86, VM04, XS09]. Independently [Blu70]. Index [BVT96, Tor06, Bur91, CL87b, CL87a, CM11, GGO15, LV94, PK06]. Indexe [Ley96]. Indexed [Fed87]. India [Soo13]. Indian [GB13, RVAN15]. indication [dVM13]. Indicator [Töd09]. indicators [AYS16, RRZR09]. Indices [SW10, Nal03]. individual [HSD12]. individuality [Bro80b]. Individuals [MZ98b]. Induced [Van02]. inductive [Sim07]. Industrial [ALY19]. Industry [Ros11]. inefficiencies [Nig00b]. Inégalités [Gib31]. Inequalities [BHM08, Tic85, Tic86]. Inequality [Kaf08]. inexpensive [HSR^{+}08a, HSR^{+}08b]. Infamous [CS11]. Inference [WC03, dCS15]. inferences [Gun15]. Infinite [CI96, NKL04, NKL03]. Inflation [GR07, KR01]. influence [BCM02]. influential [HDV13]. Information [Bra34, Bra85, DMJH11, GN07, Hea78, JJ21, KMN09, Voo74b, Baw11, Ben75, Bro80a, Bro80b, Bro80c, Bro81, Egg05d, Moc21, Pop75, Voo74a, Bro81]. Informational [Man53]. informetric [Egg05a, Egg07, EG12a, EG12b]. informetrics [Egg05d]. Ingmar [Har09]. Initial [Dav76, Dun69, Sli98, Whi72, BHS11, BK91, Fle66, Gin57, Her57]. iniziali [Gin57, Her57]. input [LC15]. instabilities [Abe02]. Institutional [Lee18, LWFW14]. institutions [Per10]. Instrument [WSU08, Sch10]. instruments [Ole10, Ole14]. Insurance [LB05, HB18]. Integer [Hür04, ML84, FMSV24, Fle66, Hür06b, Hür09, Hür15b, Yeh90]. Integers [JD11, JD12, Dia73]. Integrated [QBP^{+}04, GGO15, LWY10, MAE06]. Integrating [RBXM15]. integrators [BG15]. Integrity [Mac14a, LSS00, Nor13]. intellectual [Ana03, Lim06]. intelligence [DMJH11]. inter [LAL15]. inter-textual [LAL15]. intérêt [GD10]. Interacting [MZ98b, NBH^{+}98]. interdependence [Fur46, Gei48]. Interdependenz [Fur46, Gei48]. interest [GD10]. Intermittency [MZ98a, ZM97, Poc98, Rob11]. Intermittent [Mil57, KA84, MBS99]. Internal [NBH^{+}98, KMN09]. International [ACM00, AW89, Jin09a, NKZ12, NM07b, CBC^{+}19, Lac18]. Internet [AJ14, BFS03, BS00, GTC^{+}07, HA99, LB09, RH02]. Interoccurrence [Has07]. Interpretation [AdSPG^{+}76, Bra76, Egg05a, GJ10, Hub76, OR07, TR91]. Interpretations [Dun80, Moh80]. interpretativi [Bra76]. Interrogation [Lan04]. Interval [Wu08, Per96]. Intervals [NKL04, NKL03]. intervention [LFCR04]. interventional [CSC14]. Interviews [SSMW05]. intrigued [Olv14]. Introduction [Fel50, Fel57, Fel66, Fel68, Gut16, HW75a, BH15b, BH15a, Hor09b, Mil15e, Zip35, Zip65, eSHK05]. Intuition [Lee18]. Intuitive [WSB15, WSB16]. invariance [Hil95a, KR12, SAG^{+}00, Vol20, Wój14]. Invariant [All97, GM08, Whi83, CI96]. Invariant-Sum [All97]. Inventory [dMH06]. inverse [DB14, DHL^{+}20, Per05]. Inversion [LS11]. Investigating [MPPH06, Ana03, Lim06, RC18]. Investigation [SF01, Fur46, Gei48, HSD12, HYYH08, LSS00, Soo05]. Investigations [Wan11b, Nig11e]. Investigative [Bha02]. Investing [Wal03]. investments [AB03]. investors [Bar11, LWFW14]. Invitation [MTB06]. ion [MG22]. ionic [MWS99]. ionization [MG22]. IR [BSWB16]. Iran [Ans09, Meb09, Meb10c]. Iranian [Rou14, Bat09, Rou09]. Ireland [ACM00]. Irregularities [Nig99b, BG11b, FGZZ11b, Nig00b, RR03, TMS02]. Irrelevance [DMO10]. Irrespective [Hay08b]. Irrungen [SHC^{+}12]. IRS [Her98]. ISBN [Ale05]. Ising [SS09]. Island [PS09, NNF21]. isomorphism [BP06]. ISPs [CL05]. issues [RRZR09]. Italian [Bra76, CEZ17, Gin57, Her57, MAC14b]. iterated [Sch88a, Sch91a]. IV [Bro81]. I’ve [Nig99b].

J [Krä15, Lim06, Yul25]. jaguar [GM97b]. Jahren [Sch06]. Japanese [NG04, SGW04]. JCR [AYS16]. JCR(R) [AYS14]. Jersey [Lim06]. Jobs [ZM08]. Johannesburg [Sav06]. John [Lim06, MM99, MM04, MM14]. Join [KNT89, WYTD93, WDY93]. Jones [LV94]. Josef [Bal66]. Journal [SDK14, DG10, Gra10, Kri10, Mar12, MG07b]. journals [AYS14, CC11, Ora17]. JPEG [AOT13, FSS07, LYY09, LZL^{+}12, SF11, TRB^{+}12, ZHS09]. JPEG2000 [QZH10]. Jules [WHN13]. Julian [Ale05]. July [PS09]. June [ACM00, CLT20, PS09]. Justice [ZM08].

Kanunu [AA10]. Kautilya [Bha02]. Keenan [Gro67]. Keys [SD78]. Kfz [HB18]. Kfz-Schäden [HB18]. kind [Mas07]. Kinematics [AFS04]. Kinetic [GM08]. Kinetics [FM89]. Kingsley [VW06]. klassische [Szé90]. KMUs [Moc17]. Knopoff [Has07]. knowledge [DG12]. known [Rou76a]. Kolmogorov [Lon92, Lou92]. Krylov [TR91].

L [Kui69]. L. [Gei48]. l.s.d [Tur84, Tur87]. lacunarity [Mas06c]. Lady [Wea63, Wea82]. lakes [LD04]. Lamp [Ano32]. landscapes [Bro80c]. Language [BHB^{+}96, BMLRV10, GS01, IKC96, Mel06, Cha11, DMZ05, FS03, KG72a, KG72b, Par75, Zip32, Zip35, Zip65]. Languages [LZZ13, Man53, Rav08, Che91, MPPH06, TST97]. Large [Ber10, DCH^{+}17, ZM97, AW89, BGP07, CI96, CL05, SM96]. Large-Scale [ZM97, SM96]. Largely [Bur10a, Bur09]. Laser [FNP^{+}15]. Last [Nig11c, DMF09]. Last-Two [Nig11c]. latest [Lin86]. Latin [Car15]. Lausanne [Par97]. Lavelette [Pop02]. Law [AL14b, ARS10, Ben38, BH07, Bor20, Bra78, BT37, CSN09b, CBP12, Da 20, GM97a, Hum06, Jol05, Jol09, KN91, Lon92, LMH99, LS11, Mir11, Mit03, Moc02, Moc17, MdSPZ06, Nig01, NBH^{+}98, Ole10, Ole14, Pet08, QW03, San02a, SDK14, Sav06, SM10c, Sta14, Ada11, ARMG03, ARS11, Ano93, Ano02, Ano05, Ano07, AG05, AFS04, AA11, Aro13, Aue13, AB03, ANFF16, ACC16, BBM11a, Bak07, Bal12, BNS10a, BNS10b, BCP21, BBC^{+}18, BBL07, Ben19, Ben75, Ber03, Ber05b, Ber05a, Ber05c, BBH05, BH11b, BH21, BK07, BKS05, BXK11, BCC05, BJS06, BMM10, BP84, BBJ11, Bra76, Bro84, Bro05, Bro07, BMP93, BK91, BS92, CFH^{+}20, CC11, Cam24, CI96, CK05, CS22, CBC^{+}19, CP23, CL23, CLPR23, Cha11, Che78, Che80]. law [CL87b, Che91, CW97, Che12a, Che12b, Che12c, CG07, Cio05, CSN07, CSN09a, CEZ10, CEZ17, CEZ18, CMS10, CMFS11, CLRTFM08, Cou10, CSV96, D’A20, Da 22, DMZ05, DD06, Dea93, DJ13, Den97, Dev57, DdA^{+}09, DL07, DL08, Dur06, DM19, DHL^{+}20, EKJ^{+}20, Egg99a, Egg99b, Egg11, EG12a, EG12b, Eli11, EC12, EL03, FWW85, FC20, FMSV24, FGZZ11a, FGZZ11b, FG10, FLM^{+}21, iCV18, FS02, FRB05, FSS07, Fuj04, Gab99, GGPS03, Gai06, GLS06, GB13, GD08, GD09b, GD09a, GM12, GJ10, Goo16, GGA03, GSW92, GLSW96, GSB24, HT05a, HT05b, HT06, Has02, HPV^{+}09, Hil70, Hil74, HW75b, Hil95c, Hil99b, Hil02, HSM11, HN05, HV01, HHC^{+}14, HNY20, HSM17, Hür03, Hür04, Hür06b, Hür09, IO03, Irm97, JdlR04, JdlR09, JSZ11, Jol01, JS06, KNRS88]. law [KN90, KH02, Kaz23, KUY05, Knu01, KA84, KM05, Kos15, Krä15, Kre02, Kre03, KP04, KPK05, KB03, KB07, Lac18, LAL15, LS05, LS06, LHJ20, LBL01, Li92, LX05, LF16, Lip09, Liu08, LZNR11, Lol08, LCC^{+}05, LBC06, LB06b, LL08, Mac22, MBS99, MRM^{+}05, MPPH06, Man60, Man62, Mas05a, Mas05c, Mas05b, Mas05d, Mas06a, Mas06b, Mas06c, MM06, Mas07, Mas06e, Mas06d, ME22, MDJH05, Mil04, MK05, MN08b, Mir12b, MAC14b, Mir16b, MA18, MTTT08, Moc21, MW04, Mon01, Mör01, MR06, Mül11, MF21, NG04, Nag84, NS87, NKS90, Nag05, Nal03, NWNG22, NNF21, New05, NR08, NH07, Nic88, NNN89, NT89, Nig00b, Nig00c, NM06, OKO06, OTT99, OV97, Orl70, Orl76, PBP07, Pai08, Pao85, Pav81, Pen10, Per10, PT11, Per96, Pet81b, Pet73, Pha13]. law [Poc06, Pop02, Pos04a, Pos04b, Pos10, RRR04, Ree06, Ree03, Reg82, RDFM10, RC18, Rou76a, Rou78, Row05, RS19, RC23, SMS10, STJ10, STA11, Sch88a, Sch88d, Sch89, Sch90, SN91, Sch91a, Sch06, SS06, Sea02a, Sem08, SS11, SM10d, SM10e, SLS^{+}16, Sin74, SCD01, SR01, Ste93, Str10, Tag70, TW11, Tao09, TBL00, TFGS07, TV21, Tra92, TbG98, TR91, Tul85, Urz00, Urz11, Vol20, Wan11a, Wat96a, Wat96b, WH75, XWLD11, Yav74, Yor00, ZS11, Zör15, Zor11, Zun19, dSFSL20, vLvdW05, AS05, AG70, AG72, AA10, AR13, AP10, ALY19, Ale09, AL14a, All97, AdSPG^{+}76, AYS14, AYS16, AOT13, Ano11, Ano13, AJ14, AHI14, AHI15, Bah15, BSO10, Bal15, Ban13, BCCP18, BCCP22, Bas92, Bas94, BG09, BG10, BG11a, BGTM^{+}11, Ben84, Ber04]. Law [BH10a, BH10b, Ber10, BH11a, BHKR11, BE14b, BH15b, BH15a, BE16, BH17, BX18, BX19, Bha12, BK08, BT09, Boo94, BSWB16, BF09, BG11b, BOT06, Bro77, Bro79, Bro98, Bul11, Bur10a, Bur09, Bur04, BW98, CHL19, CM21, CS09a, CL86a, CL86b, CL87a, Che95, CCT95, CZ98, CKT^{+}09, CT05, CA12, Con10, CLM15, DCH^{+}17, DDS98, DMO10, DMO11, DS05, DJ10a, Die04b, Die07, DJ10b, Die12, DN00, DG78, DHP04, EHM^{+}14, EC11, Egg85, Egg86, Egg90b, Egg94, Egg07, Egg10b, Egg10c, Eli13, Fed82a, iCE10, Few09, For10, FGPM16, Fru16, Fur46, FK03, GR07, Gei48, GW01, GW04, Gil05, Gil07, GC13, GGP09, GJARPGAC10, GN02, GP13, GM22, Gro67, GTM12, GT09, Haf13, HCS09, Has76, Hay09, HSD12, HR10, Hil95a, Hil97, Hil11]. Law [HF16, Hop16, Hor09a, HKKS06, HYYH08, Hub02, Hub76, Hub77, Hub78, HJJYM11, Hum08, Hum18, Hün07, Hür06a, Hür15a, IMS15, IHPS14, IHWZ17, Iud12, Jam01, JTY14, JTY16, JKK^{+}08, JKK^{+}09, Jas10, JHTS10, JJ21, Jin09a, Jin09b, JS09, Kaf08, Kaf09b, KK95, KLC84, KBHR10, KK20, Kim12, KCCT23, Kle17, KO20, KH95, Kos15, KŽ08a, KŽ08b, ŽK09, KV14, KR01, Kri77, Kry09, KP05, KPK06, LB09, LCJ15, Lee18, LSE00, Lee15, Lef82, Lei80, LY02, pLCjW04, Li11, LFY15, LK12, Lou92, Low00a, Low00b, LB05, LG10, MM82, MM84, MSSvK08, MPM02, MVW^{+}03, Man08, MRR^{+}10, MG07a, MM21, MZ98b, Meb06a, Meb06b, Meb06c, Meb08b, Meb08c, Meb11, Meb15, MTT13, MN06a, MN06b, MN08a, Mil15a, Mil15b, Mil15e, Mir14, Moc02, Moc17, Möl09, MdSSZ09, MG07b, Mor06, Mül03, Mur73a]. Law [Mur73b, Mur23, NP12, NHD16, NGJ03, NWR09, Nic89, Nig93, Nig96b, NM97, Nig98, NM07a, NM09, Nig11a, Nig11b, Nig11c, Nig11d, Nig12, Nig15, Nov10, NM07b, OW12, Ole10, Ole14, OV93, OV95, OMNO10, OM11, OB13, PSM10, Pao86a, Pao86b, Pav82, PGHA07, PGQA^{+}15, PT04, Pet81a, PMdM^{+}99, Pop75, Pos04c, Pot81, QZH10, QZHC11, QW03, RK79, RMB^{+}14, RGE13, RGBK14, Reg12, RM11, RR09, Ros11, Rou09, Rou15, Rou92, STJ09, SM06a, San02b, San02c, SDK14, Sch88c, Sch88b, Sch74, Sch75c, Sch75b, Sch75a, Sch10, Sch08, Sch15, SF01, Sea02b, Sen73, SG16, SW10, Shi09, SF11, SM11, SB15, SID13, SID15, Smi97, Soo05, Sor12, SPGM12, Sta10, SH10, SB93, SCE03, TRB^{+}12, Töd09, Tor06, cTnHwH17, TST97, Van15a, Van15b, Var72, Vla72]. Law [Vla76, Vla79, Vog00, Vuk08, WDB95, Wag10, WHN13, WCCK09, Wan11b, Was21, Was81, WSU08, Wei11, WCP^{+}07, WSB15, WSB16, Wil72, Wój13a, Wlo71, Wój13b, Wój14, Wor75, YLX14, ŽD11, ZHS09, ZHS10, dGGd21, dVM13, eSHK05, GD11, BP20, MG22, Sch98, Sud10, Bal66]. Law-Based [AA10]. Law-Like [iCE10]. Law-Test [SM11]. Laws [AR09, KP05, Rob11, Sch91b, VW22, AR10, AR11, BS00, Che89, ER02, Egg05d, Egg05a, Eli11, FDA^{+}04, Hür06b, Hür09, LB06a, MM15, Mei09, New05, Per82, Per05, PTTV01, Ree01, RH02, Rou76b, Rou05, SM08, GS01, LZZ13]. LC50 [dVM13]. lead [ZIWU14]. Leading [Kos06, MM15, MZ98b, Dia73, Dia77, Has03, Hum18, Hün07, IRU83, Jec92, Mir11, Mir12a, Mir12b, Mir16a, Pha13, Sti45, Tur82, Var99, War03a, XWLD11]. leads [FWW85, KR01, Liu08, Mas05a]. Learning [CS11, Kle17, LB05, MYP14, RBXM15, VK90, GCB^{+}05, Jas10, KCCT23, LB06b, Rou10]. Least [FS03, Ano50, SM06b, Yua84, Zip49]. least-recently-used [SM06b]. Leaving [SPGC10]. Lecons [Poi12]. Lectures [Poi12]. Legal [Sch75b]. Legends [DJ10b]. legge [Bra76, CEZ17]. Lehmann [Har09]. lehrt [Hün07]. Leimkuhler [Egg90b]. leistet [Moc02]. L’Emploi [Lef82]. length [RDFM10, SEOV04]. Lerch [KPH10]. less [HKL^{+}06]. Letter [Bar11, Goo65, Hay08a, Hil11, Var72]. Letters [Bas94, Bro79, Bur04, Che95, Hub77, Kri77, Lon92, OV93, OV95]. Level [GM22, MM86, HNY20]. levels [BP84]. Levy [Man60, HS01, LC15, MBS99, Sch15]. Lévy-stable [MBS99]. Lexical [EC11]. Ley [PT04]. L’histoire [Pet73]. Libor [AMKMS12, RGE13, AMVBJ11]. Librarianship [Sch75a, Sch75c]. Library [Sch74]. lie [Vog00, MM15]. lies [Hen11]. Life [Ami07, SG00, Wan11b, AS15, Aro14, FG10, Tij04, Tij07, Tij12]. Lifetime [NKL04, Sol00, NKL03]. Light [LK12, MNW^{+}05]. Lightning [MRR^{+}10]. Like [BMLRV10, iCE10, Ber03, Ber05b, DS01, GJS07, GTC^{+}07, HS01, Li92, RV98, SS09, Too92]. Likelihood [dCS15, GJS07, Izs06]. likely [Bur10b, Ora17, Sto16]. Limerick [ACM00]. Limit [Eli11, MN06a, Per96, TbG98, KPH10, MN08a, Sze10]. limitation [Cha11]. Limited [LZZ13, Liu15]. Limitierungsverfahren [Zel58, ZB70]. Limiting [CM11, Zel58, ZB70]. Limits [Dia73, You13]. line [RDFM10]. linéaires [DJ10a, DJ13]. Linear [BE14a, BE14b, RM11, Ber15a, BE16, MM15, NS87, SN91, DJ10a, DJ13]. lines [RV09]. Linguistic [KN91, MBG^{+}94, Vos96, DM92, KN90, SF08]. Linguistics [KK95, MBG^{+}95, MM06, Mon01, Tag70]. Linux [MSSvK08]. Lisp [CG77]. List [CG77, Ped94]. list-update [Ped94]. Listed [Vuk08, Sav06]. Lists [WAT15]. Literatur [SHC^{+}12]. Literature [RK79, DG78, MG07b, Pop75, SHC^{+}12]. Litigation [Nig96a]. Lives [NWR09, ZM08, BMP93, NR08]. LL.D [Doo10]. lobster [GHP09]. local [CHL19, Dur06, Rou15]. locality [BMM84, VM04]. Location [HM65, RBXM15, Shi89]. Location-Based [RBXM15]. Log [LS11, Bra76, Bro84, PGLL10, Per10, SPGC10, Too92]. Log-Law [LS11]. log-like [Too92]. log-normal [Bra76, Per10]. log-polar [PGLL10]. Logarithm [NKL04, NKL03, Sch88a, Sch91a, TW71]. logarithmes [Fra17]. Logarithmic [BB85, Jon02, Jec92, LB06a, MA18, NKS90, Sch81, Sch86]. logarithmically [CPS16]. Logarithms [BD70, Kui69, KsS73, BK07, Fra17]. logbooks [BBJ11]. Logic [Bor20, Kle17]. Logical [GBD^{+}15, GBTB^{+}16]. logistic [OR07]. Lognormal [Mit03, MPS11]. logonormale [Bra76]. Loi [GD09b, Jol05, Jol09, Lef82, BMM10, DJ10a, DJ13, Dev57, GD08, Hil99b, Pet73, San02c]. Lois [Rou78, Rou76a, Rou76b]. long [Ben19, MS83]. Looking [Aro14, Bro98, WHP99]. Lorenz [Egg06]. loss [FT86, HT05a, HT05b, HT06]. Lotka [Lon92, AdSPG^{+}76, BG15, Bur04, CL86a, CL86b, Coi77, Egg85, Egg94, ER02, Egg05b, Egg10b, Egg10c, ER12, HJ15, HW15, Hub02, Hub77, KH95, KR01, Kri77, KZ15, Liu15, LC15, Lou92, MM82, MG07b, Mur73a, Mur73b, NGJ03, Nic89, OM11, Pao85, Pao86a, Pao86b, Pet08, Pot81, RK79, Rou92, Row05, RC23, SDK14, Sch74, Sch75c, Sch75b, Sch75a, Vla72, Vla76, Vla79, Voo74a, Voo74b, WDB95, WZS12, Wan15, YLX14, ZF15]. Lotkaian [Egg05d, Egg05a, Egg05c, Egg09, Egg10a]. Lottery [Sim98]. Loyalty [KM11]. LRU [JK08]. Lucas [Was81, Wlo71]. Luck [Wea63, Wea82]. lügen [Vog00]. Luhn [Los01]. lying [Vog00].

M. [Ano93]. M32 [KR81]. MA [Mel06]. Machine [CS11, GCB^{+}05, HM65, Kle17, GCB^{+}05]. Machines [Meb06a]. MacRoberts [Sti87]. Macroeconomic [GGP09, NM07b, Bah15]. Made [Ban13]. Madoff [Hen11]. magnetic [dLKHS04a]. magnitude [BH74]. Maintenance [Lef82]. major [Ora17]. make [GT99]. makes [Hil99b]. Making [Die12, HRJB02a, HRJB02b, ANRV07, CS09a, HCS09, Wal02]. makroökonomischen [Bah15]. malicious [ME22]. Malo [NKZ12]. Management [Hay08b, LP15, Van02, Vuk08, AG08, AA11, Car15, DZ03, LW14, Nig05]. Managers [Rob97, SGW04, AVV08]. managing [WYTD93, WDY93]. Mandelbrot [And11, Aus14, Lef82, MVW^{+}03, ARMG03, Baw11, BP84, Cha11, Dev57, Egg99a, Egg99b, Fai69, Fai05, GMV09, Izs06, dLKHS04b, dLKHS06, Lef82, MPM02, Mas05a, Mas05d, Mas06c, Mon01, Orl70, Orl76, TW11, Val11, You13]. Manifestation [MG07b]. Manipulation [AHH08, RGE13, AMKMS12, LWFW14]. Manipulationen [Moc02]. manipulations [Moc02]. Manner [DCH^{+}17]. Mantegna [MBG^{+}96]. Mantissa [Kon65, Sch88b, SM10a, EMS11, EMS13, Sch73, Sch81, Sch87a]. mantissae [BS07]. Mantissas [BHM08, Sch88c, Sch86]. Mantisse [Sch73]. many [MBS99, Tay05]. Map [Sch75a, OR07, Sch75c]. maps [Bro80c]. Marey [WHN13]. Market [HH13, RGE13, BCC05, GGPS03, Mas05a, Mas05d]. Marketing [Lu07].

Markets [LK12, BS00, BCM02, CEZ10, DDS98, DK09, LW14, RC18]. Markov [BHKR11, KK95, KN90, KN91, KBHR10, NNN89]. Markovian [Rou76b, Rou78]. markoviennes [Rou76b, Rou78]. Marsaglia [Bac98]. Maslov [Mas06e, Mas06d]. Mass [Ano50]. massive [SS12]. Math [Ber95, Dub01, Wel05]. Mathematical [Egg07, Iud12, Nig99b, Whi72, Yul25, AI99, BH11b, BH15a, BBE10, Che12a, Cou10, Hum08, Jin09b, Mac22, MM99, MM04, MM14, STA11, Szé90, Wap12]. Mathematician [Key36]. Mathematics [Hün07, LS11, Olv14, Ano11, BH21, CFH^{+}20]. Mathematik [Hün07]. mathematisch [BBE10]. mathematisch-statistischer [BBE10]. mathematischer [Szé90]. maths [Aro13]. matrices [GN07]. Matrix [Pet66]. Matthew [VB08]. Maximal [Liu08, TN87]. Maximum [HT01, Izs06, Ano02]. May [NSGP05]. McAllister [TGD05]. me [Olv14, Poi12]. Mean [MK10, Moc02]. meaning [iCV18]. meaning-frequency [iCV18]. measure [BMM84]. Measurement [Bor20, MM87, Sti87]. Measurements [Moh79]. Measures [MM86, BE69, CI96, CZ04, Dia73, Jec92, Per10, Sch89]. Measuring [KK20, Möl09, LSS00]. Mechanism [MSSvK08]. media [GTC^{+}07, ME22]. Medical [Ism04, GBA04, SHC^{+}12]. Medicine [Sch75b]. Medieval [CNS15]. medizinischen [SHC^{+}12]. MEDLINE [Fed82b]. members [BOT06]. membrane [MWS99]. Memo [Meb10d]. Mental [Bor20, Hsü48]. merge [WDY93]. merkwürdiges [Sch06]. metabolomics [D’A20]. metals [DB14]. Metcalfe [BOT06]. Method [BH07, KNT89, SHC^{+}12, Est16, SF11, Tic87]. Methode [SHC^{+}12, Est16]. Methoden [BBE10]. Methodical [Möl09]. methodische [Die02]. methodological [Die02]. methodology [dGGd21]. Methods [Cig64, Fle81, FLP03, WG68, BFS03, BCP21, BPS11, BBE10, HSR^{+}08a, HSR^{+}08b, Ham62, Her10, MS10, MBG^{+}95, Nig11e]. metric [Yua84]. metrical [Fur46, Gei48]. metrische [Fur46, Gei48]. Microarray [LY02, HRJB02a, HRJB02b]. microarrays [LCC^{+}05]. Microcystis [CLRTFM08]. microelectronics [NKZ12]. microparticle [dLKHS06]. microparticles [dLKHS04a]. Microscopy [CNS15]. MIDI [MPM02]. MIDI-Encoded [MPM02]. Might [MS13]. Military [RGL14]. millennium [ACM00]. Miller [Krä15]. Millions [MSA^{+}11]. Mimicking [KN91, KN90]. minification [CGJ10, Ris08]. Minimizing [MZLH15a, MZLH15b]. Mining [Bar11, Ole10, Ole14, DG10, Gra10, GD14, Kri10, RRRB11]. MINO [NKZ12]. Misconception [Sta10]. misleading [SM11]. misreporting [GBF21]. miss [JK08]. Mixtures [Rod04, Bal12, Bal15]. Mobile [RBXM15]. Mobilizations [MK13]. Mod [Ant91, Kui69, Sch91a, Wey15]. mode [LZL^{+}12]. mode-based [LZL^{+}12]. Model [BMLRV10, Fed82b, LB05, MZLH15a, MZLH15b, RVAN15, Sol00, ZM97, Aus14, BBL07, CW97, CZ04, Che12c, FWW85, FP04, GGO15, Has07, HW15, Hub02, IMS15, Liu15, MZ98a, MG15, MG07b, OKO06, Per82, SM96, SS09, VM04, Yeh90]. Model-based [MZLH15a, MZLH15b]. Modeling [BFS03, NP02, Che89, HS08, KR12, Nic89, ZDCW12]. modelli [Bra76].

Modelling [MHS95, MRSCCS11, NKZ12]. Models [JKK^{+}08, LW92, Mit03, Moh79, BCD23, BCC05, Bra76, Cha11, CM11, CZ98, DM19, HV01, Izs06, JKK^{+}09, Los01, OV97, RMB^{+}14, Yeh07]. Modern [MTB06, Sta99]. Modified [She68, cTnHwH17, WSY11]. Modular [ARS10, ARS11]. Module [AH09]. Modulo [BD70, Kem75, Loy73, MN08a, Sze10, Too92, MN06a]. monetären [Bah15]. monetary [Bah15]. Money [TS94]. monitoring [CSC14]. monkey [BBM11b]. Monographs [Wor75]. Mont [NKZ12]. mortality [GBF21]. Most [Ber15a, Ioa05, Liv02, Wel05, AS68, Adh69, Boy94, MPS11]. motif [BP06, BGP07]. motifs [GÁVRRL14]. motion [LF16]. movies [KSH11]. MPEG [CS09b]. Mr. [Cra10]. much [BOT06]. Muhasebe [AA10]. Multi [Ber05c, CZ04, JHTS10, Boo94, CGPT15, Egg99a, Egg99b, FC14]. Multi-dimensional [Ber05c]. multi-disciplinary [Boo94]. Multi-fractal [CZ04]. multi-objective [CGPT15, FC14]. Multi-size [JHTS10]. multi-word [Egg99a, Egg99b]. multiauthorship [Rou92]. multiband [HG14]. Multicore [LHK^{+}15]. multifragmentation [CK05, PBP07]. multinomial [Izs06]. Multiobjective [BR15, RZS^{+}15]. Multiple [BRS38, SID13, SID15, Fed84, Hop16, LACL14]. Multiple-Choice [SID13, SID15]. Multiplication [Mor10]. Multiplicative [Uze10, CLPR23, MZ99]. multipliers [Egg90a]. multispecies [HJ15]. Multivariate [HHG14, MF21, Yeh02, Yeh07]. Museum [RBXM15]. Music [MPM02, MVW^{+}03, MRM^{+}05, MF21, NWNG22]. mutation [Ber15b]. mutualism [HJ15]. mystérieuse [GD08]. Mysterious [Wel05, GD08]. Mystery [Kru96a].

Name [YSC13, Pet08, Rou76a]. names [ZM01]. nanoelectronics [NKZ12]. Naranan [Egg10b, Egg10c, Hub76]. National [Sim98]. Natural [AS05, DS05, GJARPGAC10, BK07, Cha11, Che91, EMS11, EMS13, LCJ10, Man62, New81, STJ10]. Naturally [Gil05, Gil07, Liu08]. Nature [Kaf09a, Pér07, PTTV98, Bak96, Car10, Cou10, GN02, PTTV01, RH02]. Negative [Lin08, Mas06b]. n’est [GD08]. Net [ŽD11]. Netherlands [AW89]. network [ACCG13, HKL^{+}06, SS09]. Networks [BW98, CT22, IHWZ17, LEC^{+}15, BA99, BGP07, BOT06, MPPH06, TV21, VK90, ZS11]. neue [Szé90]. Neuere [CH61]. Neural [BW98, IHWZ17, MPPH06]. neuromuscular [dSFSL20]. neutrosophic [BKS05]. Newby [Bur04]. Newcomb [Doo10, PT04, Pos10, BCCP18, CBC^{+}19, FG10, For10, Haf13, HV01, Kre03, Lac18, MdSPZ06, MdSSZ09, PT04, PT11, Pos10, Rav08, RR09, Smi13, Tor06, dSFSL20]. Newforms [JTY14, JTY16]. Newton [BH07]. next [CGPT15]. nichts [Vog00]. NMR [BSM14]. No [BHB^{+}96, BH11b]. No. [Bro98]. NOEC [dVM13]. Noise [IKC96, LFY15, BTW87, Che12b, LC15, MS09, MS83, NS96]. nom [Rou76a]. nombres [FL94, FL96, San02c, Var99]. nominal [eSHK05]. Non [RM11, SD78, CI96, CW97, GJS07, HJ15, Kim12, LF16, Liu08, Liu15, Lol08, Sim07, SR01, LS11]. non-autonomous [HJ15, Liu15]. non-equilibrium [Kim12]. non-existence [Lol08, Sim07]. non-homogeneous [CW97]. Non-Linear [RM11]. non-parametric [GJS07]. NON-Shor [LS11]. non-stationarity [LF16]. non-stationary [SR01]. non-symmetric [Liu08]. non-uniformly [CI96]. Non-Unique [SD78]. nonautonomous [Ber03, BS07]. Noncoding [BHB^{+}96, IKC96, MBG^{+}94, Vos96, MBG^{+}95]. Nonextensive [SM10c, SM10d]. nonlinear [Pop02]. Normal [Bor20, Bra76, Per10]. normalize [LCC^{+}05]. normalized [Abe02]. not-Benford [Pos04a]. Note [Dun69, HM65, Meb09, New81, Sli98, Egg90a, Hil96, Man59, NWNG22, SN91]. Notes [Kon65, Whi72, Val11, She94]. Notion [Ant91]. Novel [JHTS10, SF01, CSC14, LSS00]. November [And11]. NPO [Van15a, Van15b]. Nuclear [HJJYM11, FG10, JSZ11, MNW^{+}05, PBP07]. nuclei [NR08]. Null [SBB01, FP04]. Number [Ano13, Ber95, Can03a, CKT^{+}09, FMS10, Kry09, Liv02, LB07, LS11, MTB06, Nig99b, Nig11c, Bac98, BFLB10, BH74, BS00, Can03b, CS22, Car10, CLRTFM08, DM92, Dia73, Hil88, Lem86, Lin86, MS12, Nig00b, Sav04, Tic86, WHP99]. number-theoretic [Tic86]. number-theoretical [CS22]. Number-Theory [LS11]. Numberphile [Ano13]. Numbers [BS10a, Ben38, Car88, CKT^{+}09, Dun67, FM95, FH45, GF44, Ham70, HW75a, Hsü48, JP01, Kui69, LK12, Mur23, NM07b, PTTV98, San02a, Vog00, Was81, Web75, Wel05, Wlo71, Aro14, BHS11, BK07, BS92, CI96, CK84, D’A20, Dia74, DMO06, EMS11, EMS13, FL94, FL96, GSR12, Gin57, Her57, Hum06, Hum18, IMS15, KR04, LOR06, Lin86, LS14, LL09, MS15, Mir11, New81, PTTV01, Sav04, SF08, Sta05, Ste93, SB93, Tic88, TFGS07, Tsu52, Var99, Wey15, San02b, San02c, Har09]. numeri [Gin57, Her57]. numeric [Bah15]. Numerical [ALT16, Ham62, JP01, Kub77, FMS10, KR12]. nuovi [Bra76].

Obey [BHKR11, Wlo71, BS92, FC20, KBHR10, SB93]. Obeys [BH07]. Obituary [And11]. Objective [Bro80c, CGPT15, FC14]. objectives [LACL14]. Observation [Mil15c, Mil15d, Her10, dLKHS04b, Sar73]. Observations [AG70, AG72, Vla72, Bur04]. observe [Ano93, Tra92]. observed [Gin57]. Observes [NP12]. Occupational [Lan03]. Occur [Hay08a]. occurrence [DMO06]. Occurring [Gil05, Kos13, Gil07]. Occurs [Hay08b]. October [And11, GCB^{+}05]. Odds [CKT^{+}09]. officers [Lac18]. oil [BCK12]. Omnidirectional [NKRS15]. One [AP10, BBH05, BD70, KP04, LLD15, Mat99, Yeh90, Car10, CEZ17, CEZ18, Hum18, MS08, SCD01, Too92, WHP99, Wey15]. One-dimensional [BBH05, MS08, SCD01]. One-sided [KP04]. Online [BH09]. ONLY [LS11]. Open [MSSvK08, NGJ03]. Openings [BT09]. OpenStreetMap [Moc21]. operational [Ano93, Tra92]. Optimal [LACL14, SJP^{+}15, Zom14, BCIS09, Kre73, Kre75]. Optimale [Kre73, Kre75]. Optimality [MYP14]. Optimization [BR15, LP15]. optimizing [GÁVRRL14]. Optimum [McK80]. optoelectronics [NKZ12]. Order [Haf13, MN06b, MN08b, NM09, Nig11d, CSC14, SN91, TA12, TH13, Yua84].

Oregon [CLT20]. organization [BP84, LF16]. organized [BTW87, Bak96, SM96]. Organizes [PMdM^{+}99]. Organizing [Kru96b, SAG^{+}00]. Oriented [Car88, Asa81, Tic88]. origin [BCC05, CSV96, WSB15, WSB16, ZS11]. origins [iCV18, FS03]. oscillators [Mas06a]. osservati [Gin57]. Other [Nig99b, NM07a, MPPH06, MS83, NM06, RRZR09, Ree01, TN87, TMS02]. Outils [Lef82]. Outlier [Meb08a]. outliers [DN11, Sco13]. output [CSC14, Per10, VM04]. outputs [vZ13]. overcome [SHC^{+}12]. Overflow [FT86]. Oversight [AA10]. oversimplifying [Da 22]. Overview [Piv13]. overwhelming [Mar12]. ovvero [CEZ17]. Oxley [AG08].

P [Mas06d]. Pacific [BCM02]. packages [XLW14]. Packet [CL05]. Pages [NKL04, Ale05, MA18, NKL03, Sch06]. Paintings [CNS15]. Panel [CNS15, Sch10]. Paper [Kui69, KsS73, Bur04, Man59]. papers [GBA04, HZKS12, HSM17, MG07b, Sco13, Kon65]. paradigm [BSM15, Bro81, Poc06]. paradox [Fel17, RRZR09, Nov10]. Paradoxa [Szé90]. Paradoxes [Szé90]. paradoxical [BBM11b]. Parallel [WDY93, Nic88, NT89, WYTD93]. Parallelizing [GÁVRRL14, WAT15]. Parameter [AMP06, CT07, HW87, Ism04, CZ04, Gar11, TW11]. Parameters [DS01, ADAI13, AS15, Izs06, KS10, Nic87a, Nic87b, Sco13, TN87, TKP14]. parametric [GJS07]. Pareto [Bal66, Gre86, AMP06, AAMAE03, ADAI13, AS15, ANRV07, Afi10, AG70, AG72, Ami07, AFS04, AP89, Aus14, BVT96, BB76, Blu70, BT37, CH97, CT07, CGPT15, CGJ10, DW15, DN10, DN11, FWW85, FC14, Fer13, FDA^{+}04, FAD^{+}04, GGO15, GÁVRRL14, Gun15, GN07, HZZ14, HW87, HDV13, HHG14, Ism04, Jee97, Key36, KS10, KA84, KPF15, KP09, Lip09, LACL14, Mac22, MPS11, Man60, Man62, MYP14, Mas05a, McK80, Mei09, MN14, MTTT08, NK07, New05, OJM15, Orl76, Per82, RAV10, RRR04, Ree01, Ree03, RZS^{+}15, Ris08, SZVHA10, San87a, San87b, San88, San92, SS06, Sco13, SJP^{+}15, SM08, Sol00, SR01, SS12, Sta14, Tao09, TA12, TH13, TKP14, VAGE09, Wil06, WC03, Wu08, XP89, ZSB^{+}12, Zom14, dCS15, vZ13, Sha89]. Pareto-Efficient [RZS^{+}15]. Pareto-Optimal [SJP^{+}15, Zom14]. Pareto-type [HDV13]. Part [Bro80a, Bro80b, Bro80c, Bro81, Sha05]. Partial [QBP^{+}04, RME98, WCP^{+}07, HS08]. Partial-Volume [QBP^{+}04]. partially [AS15]. particular [FGZZ11a]. Partisan [Lee18]. Partition [ARS10, ARS11]. partitioning [Lip09]. Partitions [KN91, KN90, NNN89]. Parts [Pav82, Giu11, Sur93]. Passive [LYY09, BM13, Her10]. Patent [YLX14, HHC^{+}14]. path [Art94, FC14]. Pathology [Hay08a, Hay08b]. paths [RC18]. Pathway [GM08]. Patient [Hay08b]. Pattern [Mur23]. Patterns [DKM07, Mil15c, Mil15d, Rod04, SGW04, IRU83, MHS95, MS08, NGJ03, Nig99c, Nig00b, RDFM10, Tho89]. Payment [KŽ08a, ŽK09]. Peculiar [Gor12a, Ley96, Nig99c, Rai69a]. Peer [HKL^{+}06, ANFF16, HS08]. peer-to-peer [HS08]. pendant [Poi12]. People [BM08, Bur10a, Bur09, MWE95]. Percentages [DF79, MYZ67]. percolation [Ber15b, PBP07, Wat96a, Wat96b]. Performance [Egg09, Fed87, LP15, MZLH15a, MZLH15b, Vla79, NHD16, WR03]. Periodic [HJ15]. periodical [Pop75, SN91]. Periodicity [She68]. Periods [WDB95]. permanence [LAL15]. persistence [Sjo14]. person [MTTT08]. Personal [Pet08]. Personalized [KA15, LCY^{+}15]. Perspective [Bat08, JJ21, CS22, Eli13]. persuasive [KA15]. Pervades [SG00]. PET [QBP^{+}04, RME98, WCP^{+}07]. PET/SPECT [QBP^{+}04]. Petch [DB14]. Petiet [WHN13]. Phantomware [AH09]. Phase [Mas07, RMB^{+}14, SS11]. Phenomena [Kos06, MZS99, Vla79, Fur46, Gei48, Sta99]. Phenomenon [Hil98, LG78, Nig99b, Rai85, Bas94, Coh76, CK84, Hil88, Hil95b, HHC^{+}14, MS11, MS12, MS15, OV93, Val12, WZS12]. Phi [Liv02]. philology [Zip35, Zip65]. Philosophical [Bro80a]. phonetic [Zip29]. phonograms [NG04]. Photorealistic [XWLD11]. phrases [Egg99a, Egg99b]. Physical [ALT16, Ben43, GSW92, Kos13, BK91]. Physics [HJJYM11, LS11, PSM10, Sud10, JSZ11, SM10e]. pillars [Sta99]. pioneers [RC23]. pitfall [Urz11]. Pitfalls [Gou77, Goo16]. PL [Els78]. PL/I [Els78]. placement [FMS10]. planar [Wan15]. planetary [Aro13]. planning [FC14]. plant [Shi89]. Play [LCY^{+}15]. Players [Sim98]. Pleasant [MVW^{+}03]. Plots [BVT96, Hav95, SBH^{+}95]. Plotting [WG68]. PODS’09 [PS09]. Poincaré [TW71]. Point [BB85, HM65, Kre73, Kre75, Sch88c, Ste74, FT86, Gar11, Sch73, Sch86]. Points [Van02, AVV08, HKWE14, HDV13]. Poisson [BG15, Eli13]. polar [PGLL10]. Polignac [CL04b]. Political [NM07b, Par97]. politique [Par97]. pollutant [Bro05]. polluted [LC15]. Polyhedral [GML15]. polymers [BSWB16]. Polytope [GML15]. Popularity [BT09, Aus14]. population [Aue13, BS00, GBF21, Hil70]. Populations [San02a, San02b, San02c]. portfolio [PK06]. Portland [CLT20]. Porto [GCB^{+}05]. Portugal [GCB^{+}05]. Posamentier [Har09]. positivity [BG15]. possessing [Nic88, NT89]. possibility [TR91]. Possible [AR10, AR11, CSV96]. possibly [SCE03]. poster [PQ00]. potential [Bar11, DdA^{+}09]. potentials [NH07]. pour [Est16, FL94, FL96, GD10]. Pourquoi [GD08]. Power [BS00, CSN07, CSN09a, CSN09b, CBP12, Egg05d, EC12, HS01, LHK^{+}15, MBS99, Mat99, Mit03, New05, NBH^{+}98, Ree06, Sch91b, Ber03, Ber05b, BCC05, CSV96, EKJ^{+}20, Egg05a, Egg05b, GGPS03, Hür06b, Hür09, LB06a, Per05, Ree01, RH02, RDFM10, Rou05, RC23]. Power-Aware [LHK^{+}15]. Power-Law [CSN09b, CSN07, CSN09a, EC12, MBS99, Ree06, CSV96, GGPS03, RC23]. power-like [Ber03, Ber05b]. powerful [Hür15b, MPS11]. powers [EMS11, EMS13, Hür04, Hür15b, MM50, Wój13a, Wój13b]. pp [Ano50]. Practical [NH07]. Practice [Vuk08, ER12, Mül03, Pos04c, Pos10]. Practices [GP13]. Pragmatic [JP01]. Praxis [Pos05, Pos04c, Pos10]. pre [AA11, NHD16]. pre-processing [NHD16]. pre-SEC [AA11]. precinct [Meb10d]. Precise [AI99, SBB01, Sti87]. precision [OR07]. predict [BFLB10]. Predicted [HSM11, MAC14b]. Prediction [Fai69, Fai05, RAV10, Sol00, AP89, Bac98]. Predictions [LK12]. Preference [Hay08a, Hay08b, Hay09, LACL14, XLW14]. Preferences [Sal97]. preferential [Ree06]. preferred [Hum18]. prefrontal [BFLB10]. Preliminary [Ans09, KS10]. premier [FL84, FL94, FL96, Hil99b]. Premiers [Var99, FL94, FL96]. presence [DN11, SZVHA10, Sco13]. presentation [KA15]. Preservation [NKRS15]. preserved [CCC09]. Presidential [Meb10c, Her10, Meb08c, Rou09, Ans09, GM22, Meb06b, Meb08b, Meb09, Rou14]. Press [Ale05, Mel06]. preventative [Bec14]. prevented [Mül11]. Price [HKWE14, eSHK05, BCM02, BM08, KH95]. Prices [Gil05, BCK12, Gil07, KŽ08b]. pricing [BCK12, Shi89, eSHK05]. primary [Meb10d]. Prime [CK84, Wel05, EMS11, EMS13, FL94, FL96, LL09]. primer [AP99, BH21, LOR06]. Primes [Whi72, LL08]. primitivité [DJ10a, DJ13]. primitivity [DJ10a, DJ13]. Princeton [Ale05]. Principle [Kaf09a, Kaf09b, Nov10, RME98, Ano50, HPV^{+}09, Zip32, Zip49, San87a, San87b, San88, San92]. Principles [PS09, SAG^{+}00]. Printer [JHTS10, FNP^{+}15]. prior [Hil70]. private [YD15]. Proactively [Lan03]. Probabilistic [FM95, MZLH15a, MZLH15b, BFS03, GJ10, Loy73]. Probabilités [Poi12, Jan12]. Probabilities [LB09, NKL04, Poi12, NKL03]. Probability [Dur91, Dur05, Fel50, Fel57, Fel66, Fel68, GSR12, Gor12b, Gou77, Hor09b, Wea63, WG68, BKS05, Dia74, FC20, Fle66, GLSW96, Hai69, Jan12, SM06b, Szé90, Tij04, Tij07, Tij12, Wea82]. Probable [Ben43]. Probably [Bur10b, LX05]. Problem [AdSPG^{+}76, Dun69, KBRV04, Lag85, LW92, Lin08, Boy94, BE69, CGPT15, FL84, FL94, FL96, KNRS88, KM05, Lag03, Lag10, Mac22, MK05, Rai76, Sim07, Sin03]. problème [FL84, FL94, FL96]. Problems [Gor12b, JS09, RGL14, Str12, JS06]. procaryotic [RV98]. Procedure [BW98, Pao85]. Procedures [NM97, DN00, SH10]. proceedings [AW89, CLT20, NKZ12, ACM00, PS09, Rai03, GCB^{+}05]. process [Egg05d, Ris08, Yeh90]. Processes [Bal66, KK95, Tor06, Ber15b, CGJ10, FM89, Kim12, KPH10, LSS00, MZ99, MN14, Nic89, Nic88, NT89, PT11, Sch08, Sch15, Zel58, ZB70]. Processing [LP15, WAT15, NHD16, Nig00b]. Product [Dur06]. production [BH74, Egg05d]. Productivity [Coi77, Lot26, Egg09, Row05, MM82]. Products [MN06a, GTM12, HSM11, MN08a]. professé [Par97]. professées [Poi12]. Profession [Sta12, KMN09]. Profiles [EC11]. program [BMM84]. Programming [Ben84, Gut16, Che91, NGJ03]. programs [Els78, dGGd21]. Progress [Fai69, Fai05, MPM02]. progressive [AS15, WC03]. progressively [RAV10, TH13]. proliferation [Dur06]. promises [Goo16]. prone [SCE03]. proof [Egg10b, Egg10c, Pet81b, Sim07, Spe09]. propagation [CL04a]. proper [TRB^{+}12]. Properties [Iud12, AAMAE03, Ber11, KZ15, RRZR09, Rou76a, Sta05, Str10, TN87, Yeh02]. Property [KC86, Kun87, Ana03, BBM11b, Rou92, Lim06]. proportional [HS08].

Proportionally [HG14]. Proportions [Fle81, FLP03, MPPH06]. propos [Bel88, Fra17]. Propriétés [Rou76a]. protein [LX05, RV98]. Proteins [MdSSZ09]. Providence [PS09]. Prüfung [BBJ11]. Prüfungspraxis [Mül03]. Psychiatric [PMdM^{+}99]. Psycho [Mel06, Zip35, Zip65]. Psycho-biology [Mel06, Zip35, Zip65]. Psychological [DK09, Kry09, DDS98, LV94, LG10, eSHK05]. public [YD15]. publication [Egg94]. Publications [Jin09b]. Published [Ioa05]. publishing [KH95, TV21]. Puerto [PT11]. Pulsars [SM10a]. Punchworks [CNS15]. Pure [LS11]. puzzle [Sch06, ZIWU14]. Python [Gut16].

Qualität [Die02]. Qualitative [PL86]. Qualitätstest [GTM12]. Quality [DCH^{+}17, GGP09, Moh79, Möl09, Van15a, Van15b, ANFF16, Die02, DdA^{+}09, GTM12, GSB24, Hol14, HNY20, Sto16, Wal02]. Quantifying [ANFF16, LF16, WCP^{+}07]. Quantile [BVT96, HW87, SS12]. Quantitative [BPS11, MM86, MSA^{+}11, Mon01, Bro80b, Bro80c]. Quantities [Kos13, Kos15]. quantization [Mas05d, Mas07]. quantized [Mas06b]. quantum [RMB^{+}14, SS11]. quark [GM97b]. Quarterly [GR07]. quasiperiodic [ZF15]. Quebec [AH09]. quella [Bra76]. Quelques [BMM10, Jan12]. Questioned [MDDK02]. questionnaire [TMS02]. questions [Zen11]. queue [Bel88]. Quick [PT11, Mil15e].

R [Hor09b]. R. [Kui69]. radiation [CSC14]. radiology [CSC14]. Radix [Lin20]. Raising [GD10]. Random [BGTM^{+}11, iCE10, Hil88, JD11, JD12, Li92, LB07, MS12, MWE95, Sol00, AS68, Bac98, BH74, BA99, BBC^{+}18, BT18, CGJ10, Dia73, EL03, FMSV24, FS02, Fle66, GSR12, Giu11, HS05, Hol69, Hür15c, Kem75, KR13, Lip09, MN14, Per96, Rou76a, Sch73, Sch81, Sta05, Sze10, UP85, Val12, Wój13a, Wój13b, WC03, Yeh90, Bal66]. Randomized [Die12]. randomly [Zör15]. Randomness [Nil10, Zen11, McA03]. ranges [Est16]. Rank [BGVV99, iCE10, KUY05, BG78, Hai69, Hil74, MM06]. rank-frequency [Hil74]. rank-size [Hai69]. Ranked [CT07, ADAI13]. Ranking [Bro84, GLSW96, Jin09b]. Rankings [Jin09a]. rapidly [FMS10]. rare [Ber15b]. Rate [GM08, AMVBJ11, Car15]. Rates [Fle81, FLP03, ZHS10, Ano93, Tra92]. Ratio [Liv02]. rational [BK07]. Rationales [KLC84]. ratios [Poc98]. Rätsel [Sch06]. Re [MM82, YLX14]. Re-applicability [YLX14]. Re-evaluation [MM82]. reactions [LZNR11]. Reading [SF08, Yul27]. Real [iCE10, Kem75, Nil10]. really [GTC^{+}07]. Reals [Lin20]. reason [DMZ05]. recall [PT04, Tay05]. received [NH07]. recently [SM06b]. Rechnungslegungsdaten [QW03]. Rechnungswesen [Moc02]. reciprocal [FGPM16]. Recognizing [MPM02]. Recommender [RZS^{+}15]. Reconcilable [McK80]. Reconstruction [SM06a, BBK^{+}15]. Recording [AH09]. Records [SD78, ARMG03]. recovering [LCJ10]. Recurrence [FLM^{+}21, Sli98, DJ10a, DJ13, Jol05, Jol09, NS87, Nil10, DJ10a, DJ13, Jol05, Jol09]. recurrences [SN91]. Recurrent [Sin74]. Recursive [BD70, KsS73, RM11]. red [RR03]. red-flag [RR03]. redistributions [Ree06]. reduced [Sch83b, Sch84]. Reduction [Nig96a]. Reference [Van02, AVV08, FRB05, VM04]. References [Sud10]. Referendum [PT04, Her10, Tay05]. Refinement [Mas05b, Mas06a, Mas06c]. Refining [Mas05c]. Regarding [Lon92]. regimes [HS01]. region [LZL^{+}12]. regions [Fer13, MAC14b]. Regression [BG09, BG10, BG11a, BVT96, Fed84]. regressions [Pen10]. Regular [AG05, GGA03, Pet66, RC18]. Regularities [ACC16, Bur10a, Bur09, Kry09, Pai13, DM92]. Regularity [GD11, HS05, GD09a]. Reihen [GTM12]. reinforcement [Jas10, LB06b]. Related [Vla79, UP85, Yeh02]. Relation [For10, Hay09, Egg09, Egg13, dLKHS04a, dLKHS04b, dLKHS06, Mas06b]. relational [WR03]. Relations [DJ10a, DJ13, Egg05b, FLM^{+}21, Jol05, Jol09, LWFW14, RV09]. relationship [CL86a, CL86b, Che12a, Egg94, Hub78, Irm97, LAL15, TN87]. Relative [Kos13, Zip29, Kos15, MPPH06, RRZR09, Sch81, Zip32]. relativistic [ZF15]. Release [dMH06, CGPT15]. Relevance [Kry09, NM07a, NM06, Sea02b]. Reliability [Ano93, JJ21, Wan11b, CP23, CW97, Jon02, dVM13]. religions [Mir12b]. religious [MG15, Mir14]. rely [AVV08]. Remark [Kui69, KsS73]. Remarks [Ale09, Dav76]. removals [WC03]. removes [Par75]. renormalization [Sta99]. Rényi [Abe02]. Reply [MM87, MBG^{+}96, McA05, TGD05]. report [Her98]. Reported [SGW04, dMH06, HSR^{+}08a, DZ03, HSR^{+}08b, Tho89, ZIWU14]. Reporting [Hay08b, KO20]. Reports [Lan03, HSD12]. Representation [Kry09, Sch73, ZA95]. representations [Tic88]. Republic [BM08]. request [SM06b]. Research [Bas92, CL86b, CCT95, Egg86, Egg90b, Egg90a, Egg94, Egg99b, GT09, Hür15a, Ioa05, KR01, Nic89, Pao86b, Rou92, Die02, GD14, MG07b, Nor13, Per10, Row05, vZ13]. researches [NKZ12]. reset [MZ99]. Resistance [Ylv77]. resizing [SLS^{+}16]. resource [HG14]. Resources [Bar11, Man62]. Response [Die12, Kub77, Che95]. restricted [TKP14]. Result [LB09]. Resultate [Tic88]. Results [Lin08, NM07a, Adh69, Loy73, NM06, SM11, TA12, Tic88, Die04a]. Retrieval [BMLRV10, Ben75, Hea78, QZHC11]. Return [BGVV99, BS92]. returns [Art94, GMV09, Pos04c]. Revelations [WHN13, Tur84]. Review [Ale05, Bal66, BH02, BW98, Gei48, Har09, Lan00, Lim06, Mel06, Sha89, Ano50, SDK14]. reviewers [ANFF16]. reviewing [And09]. Reviews [Gre86]. revised [Wan11a]. Revising [Wil06]. Revisited [Sud10, ER02, Egg13, FAD^{+}04, Pot81, SEOV04]. Revocatorio [PT04]. Rhode [PS09]. Rich [YSC13]. riches [Aro13]. Rico [PT11]. Riemann [LL08, LL09]. rife [Cou10]. Risk [Nig96a, ACCG13]. risks [BMM10]. risques [BMM10]. Robert [Rou05]. Robust [MWX^{+}20]. robustness [Rou92]. Role [Kos12, ZM97]. Rolling [Pen10]. root [DB14]. rough [RV09]. roulette [Haf69]. Round [JP01, LK12, SGW04]. Rounded [MYZ67]. Rounding [DZ03, DF79]. Rounding-up [DZ03]. Roundoff [BB85, Tsa74]. roundup [YD15]. rule [McG11]. rules [Cha11, RDFM10, Tij04, Tij07, Tij12]. Russell [Nov10]. Russia [Meb08a, MK09, MK10]. Russian [KM11, McG11].

S [Krä15, AB03, BG11a]. S&P [CEZ10]. sa [Hil99b]. Saint [NKZ12]. Saint-Michel [NKZ12]. Sal [NNF21]. Sales [AH09]. Sample [Sol00, Afi10, Pen10, Rou76a, Wu08]. Samples [GC13, KS10, RAV10]. Sampling [CT07, Sun28, TN87, ADAI13, BGP07, CCC09]. Sarbanes [AG08]. SARS [GBF21]. SARS-CoV-2 [GBF21]. satisfies [CLRTFM08]. Satisfying [LSE00, Kaz23, Pos04b, TRB^{+}12]. Savvy [Sta10]. Say [BS10a, LLD15]. Scale [Bat08, BHM08, GM08, Poc98, SAG^{+}00, Whi83, Yul27, ZM97, ADAI13, DB14, Poc06, SM96, SS09, TKP14, Vol20, Wój14]. Scale-Distortion [BHM08]. scale-free [SS09]. scale-invariance [Vol20, Wój14]. Scale-Invariant [GM08, Whi83]. scaled [BH74]. scales [Egg10a]. Scaling [MZS99, NBH^{+}98, RV09, SL96, Sta99, VB08, ARMG03, BA99, BS00, BMT96, CI96, Che12a, FS03, NS96, RMB^{+}14]. scam [Ano12]. Scarrott [Ben75]. Scatter [GD09a, GD11]. Scattering [BRS38, Sud10, DG78, HN05, HHC^{+}14]. Schäden [HB18]. Schedule [WHN13]. schedules [EKJ^{+}20]. Scheme [JHTS10]. schemes [Afi10]. Scholar [Ber95]. Scholes [Sch08]. Schools [RVAN15]. Schwerpunkt [Mül03]. Science [Han07, Jud04, Kle17, MM86, OM11, RK79, RVAN15, Sch74, Voo74b, ACC16, Bak96, Bro80a, Bro80b, Bro80c, Bro81, Die02, Hor09b, Pop75, Voo74a]. sciences [Bro79, STJ10]. Scientific [Coi77, Die04b, Die07, HZKS12, Hür15a, Lot26, MM82, Moo36, Vla72, Vla79, WDB95, AYS14, CC11, DG78, HSM17, Per10, SHC^{+}12]. Scientists [MAdV05, Ham62]. Scopus [AYS14]. scores [Egg10a]. Scoring [BMLRV10]. screening [Bro05, Bro07]. Seam [SG16]. Seam-Carving [SG16]. Search [Ban13, LB09, SD78, BGP07, DMJH11, LACL14]. Searching [BG11b, LDJ89]. seascape [MS08, SM08]. SEC [AA11, HSD12]. Second [LLD15, Meb06b, Meb08b, MK10, Meb10b, Meb12, MK13, NM09, Nig11d, CSC14, Meb08c, Meb10d, Poi12, SN91, Kaf08]. Second-Digit [MK10, Meb12, Meb06b, Meb08b, Meb10b, Meb08c, Meb10d]. Second-Order [NM09, Nig11d]. Secondary [CG93]. secondo [Bra76]. sections [LZNR11, MG22]. Sector [OB13]. seine [BBJ11]. Seismic [SPGM12]. Seiten [Sch06]. Selected [Zip32]. Selecting [GC13, WSU08]. Selection [CKT^{+}09, WDB95, IG08, ME22, RRRB11]. Selections [MK09]. selective [ZDCW12]. Self [BTW87, Kru96b, dMH06, Bak96, Baw11, Egg05a, SM96, ZIWU14]. Self-organized [BTW87, Bak96, SM96]. Self-Organizing [Kru96b]. Self-Reported [dMH06, ZIWU14]. self-similar [Egg05a]. self-similarity [Baw11]. semester [Poi12]. semestre [Poi12]. Semi [ABBM10, ZHS10, CGJ10, PGLL10, Ris08]. Semi-Automated [ABBM10]. semi-Burr [CGJ10]. Semi-Fragile [ZHS10, PGLL10]. semi-Pareto [CGJ10, Ris08]. Seminumerical [Knu69, Knu98]. semiotic [Mas06e, Mas06d]. semiotics [MM06]. Senate [Meb10d]. sense [HRJB02a, HRJB02b]. Sensitivity [Bur10a, Bur09, Kry09]. Sentance [SB93]. sentence [SEOV04]. Separable [LP15]. Sequence [BD70, Kui69, KsS73, Whi72, JK08]. Sequences [BHS11, BD70, KN91, KC86, KsS73, MBG^{+}94, ML84, RM11, She68, Sli98, Vos96, BNS10a, BNS10b, Ber11, CHL19, CPS16, CMHS95, Den97, Giu11, GÁVRRL14, Got92, Hol69, Hür06b, Hür09, Jol05, Jol09, KN90, KN74, KN06, LX05, MBG^{+}95, MS12, MS15, NS87, Pet03, Pos04b, Sch83a, Sch88d, Sch89, Sin74, Sta05, Too92]. Sequential [LDJ89]. Series [SPGM12, AI99, Boy94, GTM12, Har49, Har91, VA99, Yeh90, Yeh07]. Serjev [WHN13]. server [HKL^{+}06, PQ00]. server-less [HKL^{+}06]. servers [KSH11]. service [SPGC10]. Services [Möl09]. session [PQ00]. Set [CT07, YSC13, ADAI13, Mas05a]. Sets [BHM08, Can03a, DW15, HHG14, Kos13, Mil15c, Mil15d, AG05, Can03b, Gai06, GW04, GGA03, LCJ10, TGD05, WR03]. Several [FC20, KG72a, KG72b, Sch81]. Severe [CL23, CYG07, FT86]. Shape [Bat08, Ism04, ADAI13]. shareholders [Bar11]. sharing [HKL^{+}06]. Shen [Bar11]. shift [TW11]. shifted [ER12]. Shifter [WAT15]. shock [eSHK05]. Shor [LS11]. Short [Kon65, Pet81b, Ben19, BH15a, Egg10b, Egg10c]. shuffling [Che12c]. SHZ [Bar11]. sic [SW10]. Sicily [CP23]. SID [LHK^{+}15]. SID-Based [LHK^{+}15]. sided [KP04]. Sierpiński [PR22]. SIGACT [PS09]. SIGART [PS09]. sight [BH11b]. SIGMOD [PS09]. SIGMOD-SIGACT-SIGART [PS09]. signal [SM11]. Signaling [KM11]. Significance [LK12, ZM08, FG10, FT86, OC83]. significands [BT18]. Significant [ABBM10, FH45, Gor12a, GF44, HH13, Pin61, Rai69b, Rod04, Tsa74, Var99, Won69, AS68, Adh69, AP99, Boy94, Hil95b, Hil95c, Hil99b, HS05, HW86, Hür15c, KR04, LCJ10, Pha13, Poc06, Rai69a, Sar73, SM10e, Tur82]. significant-digit [Hil95b, Hil95c]. significatif [Hil99b]. significatifs [Var99]. significativo [AP99]. Signs [BHB^{+}96, SG00]. Silence [Mil57]. similar [Egg05a]. Similarity [BMLRV10, Baw11]. similarly [SS11]. Simon [Che91, Doo10, Gar11, Man59]. Simple [Few09, GT99, Ism04, Mil15c, Mil15d, Sta10, BH11b, Egg11, GM97b, Sim07, TR91, Urz00]. Simplification [HHG14]. Simpson [RRZR09]. simulation [MG07a, NKZ12]. Simulations [FM89]. Simulator [LHK^{+}15]. Single [ABBM10]. Single-Significant-Digit [ABBM10]. sinners [Vog00]. sites [BS00]. Situation [RVAN15]. Six [Yeh02]. Size [Bat08, BGVV99, DKM07, KNT89, Sol00, BBL07, Bra76, CZ04, CM11, DB14, Egg06, GLS06, Hai69, JHTS10, KLS23, LL08, RC23, Sem08, SBH^{+}95, TN87]. size-dependent [LL08]. Sizes [Axt01, FGPM12, LZZ13, RH02]. Skew [Sim55, YZ16, Man59, WYTD93, WDY93]. sleeping [MA18]. sleuth [RR03]. slide [Haf69]. Slovenia’s [Per10]. Small [DCH^{+}17, Rou15]. Smallest [Fer13]. Smart [LEC^{+}15, GT99]. SMEs [Moc17]. Smirnov [Lon92, Lou92].

Smoothed [BR15, ANRV07]. Snap [LCY^{+}15]. SNARC [FMS10, SF08]. Social [McK80, RGBK14, Bro79, Die02, Fur46, Gei48, ME22, ZS11, dGGd21]. Socio [Sch10]. Socio-Economic [Sch10]. Sociological [Kaf08]. sociologist [Key36]. SOEP [Sch10, SSMW05]. Software [ACM00, Ban13, Lan03, Lan04, Moh79, QBP^{+}04, Wal03, Ano93, CL04a, CW97, LSS00, NGJ03, Tra92]. soll [Sch06]. Solution [Kle17, HJ15, LWY10, MAE06]. Solutions [BE14a, KZ15, Whi82, ZF15]. Solvable [FP04]. Some [Adh69, Ber11, BMM10, Dav76, Dur06, FM95, For10, Got92, Jan12, LW14, Loy73, Mil57, Mil04, Pha13, Sch87b, TA12, Yeh07, GN07, Str10, Too92, Zen11, Bur04, MS13]. Sons [Lim06]. Sorry [Sav04]. sort [WDY93]. sorted [Lag03]. Source [IHWZ17, MSSvK08, NGJ03, Rou76b]. Sources [Bra34, Bra85, MdSPZ06, Rou76b, Rou78]. sous [Rou76a]. South [Meb10d]. sozialen [Fur46, Gei48]. sozialwissenschaftlichen [Die02]. Space [BCIS09, LB07, BFLB10, SF08, VB08]. Space-optimal [BCIS09]. Spain [NSGP05]. Spanish [AP99, PT04]. Spatial [RDFM10, FMS10, NHD16]. spatial-numerical [FMS10]. spatio [BBK^{+}15]. spatio-temporal [BBK^{+}15]. Special [Egg94]. Specific [Bra34, Bra85, KSG^{+}16]. SPECT [QBP^{+}04]. spectra [BSWB16, Pai08, Pai13]. spectroscopic [BSM15]. spectrum [KA84]. speculations [Lin86]. speed [Est16]. spillovers [Dur06]. Spontaneity [Kub77]. Spontaneous [SL96]. Spot [Sta10]. Spotlight [BK08]. Spread [Ber10, Wag10]. Spur [Vog00]. spurious [GLS06]. square [DB14]. Squares [Ros12]. SRM [AH09]. Stability [Abe02, SR01, Wan15]. Stable [KP05, KP04, MBS99]. Standard [ZM08]. startles [Her98]. state [BHKR11, KBHR10, MG15, Whi82]. Statement [Van15a, Van15b, GD14, RRRB11]. Statements [GR07]. States [HKKS06, KLS23, LW14, MK13]. Static [LEC^{+}15, KSH11]. stationarity [LF16]. stationary [SR01]. statistic [Cam24]. Statistical [BCD23, Bat09, BG54, BH02, Bur10a, Bur09, Fle81, FLP03, GF44, GHP09, Kry09, LS11, Mac14a, Man53, MWS99, OR07, PSM10, Sin03, Sto16, TMS02, Tic88, ZM08, BCP21, BBE10, CM21, CG07, DMF09, GTM12, Her57, Hil95c, HV01, IMS15, Lan00, MPPH06, MBG^{+}95, Mar12, Mül11, OC83, Poc98, Poc06, Reg82, Sch06, SM10e, Sti45, SH10]. Statistical-Physics [LS11]. statistici [Her57]. Statistician [Kos12]. Statistics [CS09b, Con28, HR10, Man62, Meb07, MN06b, NCT10, NCT13, Nig97, Nig01, NM07b, RGBK14, SM10c, Fur46, GBA04, Gei48, GD10, Has07, Hol14, MM99, MM04, MN08b, MM14, Nig00b, Nig00c, SM10d, Szé90, TA12, TH13]. Statistik [Fur46, Gei48, Szé90]. statistique [GD10]. statistische [GTM12, Tic88]. statistischer [BBE10]. statistisches [Sch06]. steady [Whi82]. steady-state [Whi82]. steganography [AOT13]. Steindl [Bal66]. Stenographic [Est16]. Sténographiques [Est16]. step [Bro07]. Steuer [Vog00]. Steuer- [Vog00]. Steuerbetrüger [Sch06]. Steuererklärungen [Pos04c]. Steuersündern [Hün07]. stiess [Sch06]. Stigler [LCJ10, MM87, WHN13]. Stochastic [LW92, MZ99, SCD01, SL96, HJ15, HW15, Liu15, LC15, MBS99, OV97].

Stock [Ley96, Sav06, SW10, Tor06, BCM02, BM08, DD06, DDS98, DK09, KŽ08b, Mas05a, Mas05d]. stopping [FMSV24, MG22]. Storage [Ben75]. stores [MTTT08]. Story [Liv02]. strange [Sch06]. Strategic [Meb10a, Meb10b, Sjo14]. Strategically [MS13]. Strategies [Meb12, MK13]. Strategy [AB03, Ped94, XS09]. streaming [CCC09, MF21]. streams [LWY10, MAE06]. strength [DB14]. strikes [BH11b]. String [IRU83]. Strong [Dia74, Lee18, Per05, BCIS09]. Stronger [HW75b]. Structural [MdSSZ09, Els78]. Structure [CG77, HJJYM11, Man53, NBH^{+}98, AR10, AR11, Che12c, JSZ11, Pai13, SS09, TST97, Tul85]. Structured [Ben84, LDJ89]. structures [CL04a]. struggle [Mir11]. Studies [QBP^{+}04, MDJH05, Zip32, Wan11b]. Study [Bal66, CG77, Hsü48, MM86, Ora17, DG10, Els78, Gra10, Kri10, LB14, OR07, Sud10]. studying [Nor13]. Subgraph [BP06, BGP07]. Subgraphs [CT22]. subject [HN05]. Subjective [NKL04, Bro80c, NKL03]. Subjects [Bra34, Bra85]. submarines [McG11]. Subsequences [Smi13]. subtraction [FT86]. Sufficient [BSO10]. suites [Jol05, Jol09]. suivent [San02c]. suivent-elles [San02c]. Sulla [Gin57, Her57]. Sum [All97]. summability [Cig64, Dur74, Sch83a]. Summation [Nig11d, Tic87]. Summierungsverfahren [Tic87]. Sums [Blu70, MYZ67, Sch83b, Sch84, Sch86, Sch87a, Sch88a, Sch91a, Sze10]. Supervised [CS11]. Supply [KV14, CS09a, HCS09]. support [BXK11, ME22]. Supreme [Lee18]. Sure [Sch88c]. Surfaces [BB76, RV09]. surfers [LBL01]. Surprises [ALT16, Szé90]. surprising [CFH^{+}20, MAC14b]. Survey [JS09, BM13, JS06, MS11, Pos04b, SCE03]. Surveyor [Dun80, Moh80]. surveys [Sch10]. Survival [LSE00]. sus [PT04]. Susceptibility [ZSB^{+}12]. sustainable [Ano02]. swaps [ACC16]. symbolic [FRB05, KUY05]. symmetric [Liu08]. symmetries [Pai13, Poc98, Poc06]. Symposium [CLT20, PS09]. Synaesthesia [BEJ07]. Synonymy [Man08]. syntactic [WZS12]. syntax [FRB05, VB08]. Synthetic [AS05, DS05, GJARPGAC10]. System [Fed82b, KŽ08a, ŽK09, NBH^{+}98, Pér07, BXK11, BMT96, Egg06, KZ15, MG07a, Wan15, Wir98]. Systematic [Da 20, MBG^{+}95, LB14]. systemic [ACCG13]. Systems [ALY19, BE14b, LHK^{+}15, RZS^{+}15, SJP^{+}15, SL96, Ada11, Ber03, Ber05b, Ber05c, BBH05, BE16, BS00, DS01, Egg05a, Egg09, HKL^{+}06, HS08, HG14, HS01, KUY05, LC15, MNW^{+}05, MBS99, Mas06e, Mas06d, Nic88, NT89, Nil10, PS09, SCD01, SAG^{+}00, TBL00].

Tables [BG09, BG10, BG11a, GF44, NKL04, SD78, Fra17, MA18, NKL03, Sti45, TW71, WR03]. Taches [Lef82]. Tail [BVT96, SS06, Bel88, GGO15]. tailed [KLS23, Yeh90]. tails [MS83]. Taiwan [LW14]. Taking [Bec14]. tale [D’A20]. tampered [LZL^{+}12]. target [BCK12]. Targets [WSU08]. Tasks [Lef82]. Tax [AH09, BS92, Her98, WSU08, Hün07, Nig92, Pos04c, Sch06, Vog00, She94]. taxes [MAC14b]. taxonomy [GD14]. Taxpayer [Nig96b]. taxpayers [Her98]. teaches [Hün07]. teams [MG07b]. Technical [Kon65]. Technique [Die12, Yor00]. techniques [BM13, BL95, Bro84, Fed84, Nig11e, RRRB11, TMS02]. Technology [Sta12, AH09, KMN09, TB00]. Tell [Meb10b]. Temelli [AA10]. Temporal [DD06, BBK^{+}15]. Tend [Wlo71, BS92, LL08, SB93]. Term [Los01]. Terminal [FM95, Hay08a, Hay08b, Hay09, MDDK02]. Terrorist [CYG07]. Test [AP10, BS10a, GBD^{+}15, GBTB^{+}16, Lon92, Meb06b, Meb08b, Mor06, SM11, Ylv77, Bra76, FG10, GBA04, GTM12, HB18, KS10, Lou92, MPS11, Meb08c, MF21, Pos04a, SZVHA10, Urz00, VAGE09, ZS11]. Testbanks [SID13, SID15]. Testing [BCCP18, HSR^{+}08a, CS09a, HSR^{+}08b, HCS09, HNY20, Kos12, MPS11, Pet08, SBB01, Urz11, CL23, Mei09, Pao85]. Tests [Ami07, BCCP22, Con10, MSSvK08, Meb08a, MK10, Meb10b, Meb12, Nig97, Nig01, NM09, Nig11c, Nig11d, AS15, BG11b, Meb10d, Nig00b, Nig00c, OJM15, Zör15, HB18]. Testverfahren [Pos05]. text [Bak07, BP84, Che89, FMS10, IG08, LAL15, Tul85]. Textbook [SID13, SID15]. Texts [iCE10, FS02, Li92]. textual [LAL15]. Theft [Meb06a, Ana03, Lim06]. Theil [CA12]. Their [DKM07, Haf13, Ber95, Ber15b, KR12, KG72a, KG72b, LS14, Mas06b, MS11, OJM15, Yeh02]. Theorem [Ban13, CL04b, Egg10b, Egg10c, Hür15b, Mas05a, MN08a, Sze10, MN06a, Per96, TbG98]. theoretic [Tic86]. Theoretical [Fed82a, LLD15, Bha12, CS22, Hea78]. Theorie [CH61, Pos04c, Pos05, Pos10, Jan12, Zel58, ZB70]. Theories [Kry09, Pra88]. Theory [BG54, Bro77, Can03a, Can03b, ER12, Fel50, Fel57, Fel66, Fel68, GM97a, HW75a, KA84, Krä15, LS11, Man53, MTB06, SMS10, Wea63, Yul25, BH11a, BH15a, CH61, Dia74, DMJH11, Dur91, Dur05, Egg90b, Egg94, Egg05c, FM89, GGPS03, Jan12, Kos15, dLKHS04a, dLKHS06, Loy73, Mas05a, McG11, Mil15a, PBP07, Pos04c, Pos10, Wea82, Zel58, ZB70]. There [CBP12, LV94]. thermodynamic [AR10, AR11]. Thermodynamics [AR09, Den97]. theses [Sud10]. things [Bro80b, Lem86]. Thirty [LLD15]. Thirty-One [LLD15]. Thirty-Second [LLD15]. Thought [McK80, Hol18]. thoughts [Lin86, Mil04]. thousands [Her98]. Three [Sta99, CZ04, MG15]. three-parameter [CZ04]. three-state [MG15]. thresholds [DZ03]. ties [MG07b, Tay05]. Time [BE14b, SPGM12, Vla76, WDB95, AI99, BE16, Has07, Yeh90, Yeh07]. Times [SD78]. timing [SPGC10]. timing-channel [SPGC10]. tissue [KSG^{+}16]. tolerance [You13]. Too [Mac14a, Tay05]. tool [DdA^{+}09, HSM17, Lan00, RR03]. Tools [Lan04, Lef82, BL95]. Top [ŽD11, LWY10, MAE06, XLW14, ZCW09]. top- [LWY10, MAE06, XLW14]. Topics [ZDCW12]. topology [MS08]. tourism [CP23, NNF21]. tournament [CGPT15]. tournaments [GÁVRRL14]. toxicant [LC15]. Toxics [dMH06]. Trachtenberg [Ano93]. tracing [Pos04c]. track [ARMG03, Vog00]. Tracking [AMVBJ11, Car15, DM19, PK06]. trade [CBC^{+}19, GTM12, Lac18, dGL13]. traditional [HSR^{+}08a, HSR^{+}08b, KCCT23, SH10, Szé90]. Traffic [HS08, Pér07, AJ14, GTC^{+}07]. Trailing [FG76]. Train [WHN13]. Training [YSC13]. Transactions [Ano93]. transcriptomes [OKO06]. transform [CA12, TRB^{+}12]. Transformation [KR12, LB06a, PGLL10]. Transformations [Pet66, CI96]. Transient [MZS99]. transitions [Mas07, RMB^{+}14, SS11]. transmission [ZM01, dSFSL20]. transparency [Sjo14]. Tree [LDJ89, CL05, SS09]. Tree-Structured [LDJ89]. trees [Ber15b]. Trends [Bha02]. Triage [OW12]. trials [Ora17]. Trip [Sta10]. triumphant [McG11]. True [DMZ05, Mac14a, FDA^{+}04, Hil96]. Truncated [AMP06, KPF15, RS19]. trust [Hen11]. Tsallis [Abe02, Den97]. Tuning [KNT89]. tuple [CSV96]. turbulent [Bia15, LF16]. Turkish [AA10, OB13]. Turn [RR03]. turnover [KŽ08b]. Twardy [McA05]. twenty [PS09]. twenty-eighth [PS09]. Two [Hai69, LW92, Nig11c, Ora17, ER02, Has07, Hub78, McG11, Sco13, cTnHwH17]. two-dimensional [Has07]. two-exponent [cTnHwH17]. Type [CT05, SZVHA10, CM11, GGO15, HJ15, HDV13, KG72a, KG72b, Per82, SM06b, Uze10, Wu08, RAV10, WC03]. Type-II [RAV10]. typology [SM08].

U.S. [Axt01, Ley96]. Überraschungen [Szé90]. Ubiquity [FGPM16, Ber04, Ber05a]. ucenje [Jas10]. UK [Rai03, Sim98]. ultimate [Lag10]. Unambiguous [Rav08]. unbiased [MPS11]. Uncertainty [GBD^{+}15, GBTB^{+}16, MM87]. uncitedness [Egg13]. Uncover [Nig99b]. Uncovering [Lu07]. underflow [FT86]. Underlying [AR09, Ami07, KC86, AR10, AR11]. Understanding [BGVV99, Gut16, KM11, Kos06, Tij04, Tij07, Tij12, WCCK09]. Uneven [PTTV98, PTTV01]. Unexpected [Mur23, Wap12]. Ungleichungen [Tic86]. Unified [LDJ89, ZA95, BCD23, Mei09, TA12, ZF15, BG11b]. Uniform [Ant91, BD70, Dun67, JD11, JD12, Kui69, KN74, KN06, Rod04, Tic85, Vol20, BT18, CH61, Cig64, Dia77, Hol69, Hür15c, JdlR04, Loy73, Sch83b, Sch83a, Sch84, Sch87b, Sch88d, Sch91a, Shi89, Too92, Tsu52, Wey15, Tic86, Tic87]. uniformity [LL08]. uniformly [CI96, MPS11, Sch89]. unifying [GJ10]. Unique [SD78]. unit [Per96]. United [KLS23, LW14, MK13]. Units [NBH^{+}98, BP84]. Unity [MM84]. Universal [MM87]. Universality [CMS10, KH02, Sta99, SAG^{+}00, VB08]. Universe [NKL04, NKL03, AL14a]. université [Par97]. universities [Jin09b]. University [Ale05, Jin09a, Par97]. unknown [ZHS09]. Unsolved [Str12]. unstable [NR08]. Unsupervised [CT22]. Untangling [Egg07]. Untersuchung [Fur46, Gei48]. Unusual [SGW04, Tho89]. unwanted [Par75]. unzipped [BBM11a]. Unzipping [Ada11]. Update [Mar12, Ped94]. upper [Hür15c]. Urban [Kru96a, MZS99, ZM97, Che12c, GB13, MHS95, MML01, MZ98a, RDFM10]. urn [KH95]. USA [CLT20, PT11]. Use [BG09, BG10, BG11a, Da 20, Die12, DHP04, Lef82, Lin08, NM97, OMNO10, San87a, San87b, San88, San92, TV21, Ale09, Ano05, Ban13, BBE10, Bro07, KCCT23, MDJH05, New81, RR03, Sjo14]. used [dLKHS04a, dLKHS06, Nig93, SM06b, Wal02]. useful [BSM15]. Uses [Ber95]. Using [BW98, CS11, CT05, Cod99, FGZZ11b, GC13, GP13, Gut16, IHWZ17, KV14, Lan00, Lan03, LY02, LP15, MS09, MK10, MSA^{+}11, Nig94, Nig01, NM09, RGE13, RGL14, SD78, Sav06, TB00, Vuk08, WCP^{+}07, WSY11, ŽD11, ZHS10, ADAI13, BP06, BXK11, BM13, BMP93, DN00, FG10, Fed84, GW04, GMV09, GSB24, HB18, HHG14, Hür06b, Hür09, JS06, JS09, KSG^{+}16, LZL^{+}12, LFCR04, LBC06, LB06b, LACL14, MPPH06, MBG^{+}95, MRSCCS11, ME22, Nig00b, Nig00c, Nig15, OW12, OM11, QZH10, QZHC11, RVAN15, RRRB11, SM06a, Sch10, SCE03, ZHS09, Die04b, Die07]. Utilizing [Sta12].

V [Gei48]. V. [Mas06d]. Validation [RME98, LD04, dSFSL20]. Validity [Ami07, CM21, Nic89]. valuable [HSM17]. value [BOT06, DMF09, Nig99a, Per82, TN87]. Values [SBB01, GBA04, GMV09, KM05, MK05]. Variable [Lin20, Vla72, BS00, Kem75, Rou76a, Sch73, Sch98, TN87, Wój13a, Wój13b, Yua84]. Variable-Radix [Lin20]. Variables [BGTM^{+}11, AS68, BBC^{+}18, BT18, EL03, Giu11, HS05, Hol69, KR12, KR13, Sch81, UP85, Val12, Yeh90]. Variates [Blu70]. variational [HPV^{+}09]. variational-principle [HPV^{+}09]. vectors [MN14]. vehicle [HB18]. Venezuela [PT04, PT11]. Venezuelan [Tay05]. Verde [NNF21]. Verifica [Bra76]. verified [HSM11]. Versicherungsbetrug [HB18]. versus [Hil96, RGBK14, SM08, SH10]. Verteilung [Hum06, Mör01, Sch73, Die04a]. Vertical [ZM01]. Very [AW89]. via [CEZ10, GB13, LYY09, Lip09, LS11, MWX^{+}20, XLW14]. via-à-vis [GB13]. Video [LP15, KSH11]. view [NH07]. Vilfredo [Key36]. vindication [RR09]. vis [GB13]. Visit [RBXM15]. Visitor [RBXM15]. Visual [WHN13]. Visualization [NSGP05]. vitesse [Est16]. Vocabulary [Sor12, Con28]. voids [Gai06]. volcanology [GM12]. Volterra [BG15, HJ15, HW15, KZ15, Liu15, LC15, Wan15, ZF15]. Volume [Haf13, QBP^{+}04, RME98, WCP^{+}07]. Vom [SHC^{+}12]. Vor [Sch06]. Vote [Meb06a, Meb06c, MK10, Meb15]. Voter [Meb12, MK13]. Votes [GM22]. Voting [Ans09, Lee18, Meb06a, Meb10a, Meb10b]. vs [Haf13]. vulnerability [WCCK09].

Wahrscheinlichkeitsrechnung [Szé90]. walked [Hol18]. want [Spe09]. Ward [PMdM^{+}99]. Warehouse [KV14]. warum [Hum18]. Was [Moc02]. watermarked [ZHS09]. Watermarking [ZHS10, PGLL10]. Watkins [LS11]. waves [Wan15]. way [Val12]. Weak [CI96, Dia74, Giu11, Che80, MG07b, Per05]. weakly [Per82]. wealth [FWW85]. web [IG08, BFS03, BMLRV10, DMO06, HA99, LBL01, OM11, PQ00]. Weibull [CLM15, GGO15, NG04, dSFSL20]. Weibull-type [GGO15]. weighted [MS11]. welfare [dGGd21, dGL13]. Well [Da 20]. Well-Founded [Da 20].

Wesley [Ano50]. Where [DK09, VM04]. which [Par75]. Who [Haf13]. whose [AS68]. Wide [DMO06, Lee18, HA99]. Width [SM10b, SM09]. Wie [Hün07]. Wikileaks [Con10]. Wiley [Lim06]. will [Sch06]. Willis [Yul25]. Wilson [CL04b]. wireless [HG14]. Wissenschaftsbetrug [SHC^{+}12]. within [Kos13]. wizard [Hen11]. Wlodarski [SB93]. Word [LB09, SEOV04, Egg99a, Egg99b, FP04, Li92, WZS12]. words [DM92, KUY05, TR91]. Working [Dun80, Moh80]. workload [XS09]. works [Bak96]. Workshop [NSGP05]. World [Liv02, SPP17, ARMG03, Mir12b, dGL13, DMO06, HA99]. World-Wide [HA99]. worldwide [Mir12a, Mir16a]. worn [Sch06]. Worth [Wal03]. WoS [OM11]. would [McG11]. Write [YZ16]. wrong [BOT06, Sav04].

XBRL [HSD12]. xxiii [Ale05].

year [Kri77]. years [Per10, Sch06]. yield [Has76]. Yule [Ber15b, Che91, Gar11, MRSCCS11, RRZR09, XP89]. Yule-Simon [Gar11].

Zahlen [Hum18, Vog00, Wey15, Tic88]. zahlentheoretische [Tic86]. Zahn [Hum06]. zakon [Jas10]. Zaman [Bac98]. Zapper [AH09]. Zeckendorf [BDE^{+}14, BDE^{+}17]. zeros [LL08, LL09]. zeroth [Mas07]. zeta [LL08, LL09]. Zhou [Bar11]. Ziffer [Hum06, Hum18]. Ziffern [Hün07, Mör01]. Ziffernanalyse [Bah15, Mül03, Pos04c, Pos05, Pos10]. Zipf [Fai69, Fai05, Mel06, Ada11, AR09, AR10, AR11, ARMG03, Ano50, Ano93, Ano02, Ano05, Ano07, AFS04, AI99, AB03, Aus14, ANFF16, Axt01, BBM11a, Bak07, Bel88, BBL07, Ben75, Ben84, BT09, BP84, BGVV99, Bra76, BMLRV10, Bro07, BMT96, CK05, Cha11, Che78, Che80, CL86a, CL86b, CL87b, Che89, Che91, CW97, CZ04, Che12a, Che12b, Che12c, CM11, CMS10, CMFS11, CBP12, CMHS95, CSV96, DMZ05, Den97, Dep08, Dev57, DS01, Dur06, EKJ^{+}20, EC11, Egg99a, Egg99b, Egg06, Egg10b, Egg10c, Egg11, Eli11, EC12, Fed82a, iCE10, iCV18, FS02, FRB05, FDA^{+}04, FAD^{+}04, Fuj04, FK03, Gab99, Gai06, GLS06, GB13, GS01, GM97a, GMV09, GSW92, GLSW96, GTC^{+}07, Hai69, HT05a, HT05b, HT06, Hav95, HPV^{+}09, Hil70, Hil74]. Zipf [HW75b, HSM11, HYYH08, IO03, Irm97, IKC96, Izs06, Jin09a, Jin09b, KK95, KN90, KN91, KH02, KUY05, KPH10, Knu01, KA84, KLS23, KG72a, KG72b, dLKHS04a, dLKHS04b, dLKHS06, LAL15, Lee18, Lef82, LBL01, LY02, LX05, Li11, Liu08, Los01, LCC^{+}05, LZZ13, MSSvK08, MPM02, MVW^{+}03, MRM^{+}05, MPPH06, Man08, MG07a, MZ98b, Mas05a, Mas05c, Mas05b, Mas05d, Mas06a, Mas06b, Mas06c, MM06, Mas07, Mas06e, Mas06d, MDJH05, MS08, Mon01, MR06, NG04, Nal03, NP12, New05, Nic87a, Nic87b, Nic88, NNN89, NT89, Nit05, OW12, OKO06, OTT99, Orl70, Orl76, OC83, PBP07, Pen10, Per10, Per82, Per96, Pet73, PTTV01, PMdM^{+}99, Pop02, Pra88, Pus04, RV98, Ree01, Rob11, Rou76a, Rou76b, Rou78, RC23, SMS10, SS06, Sem08, SM08, Sha05, Shi89, SEOV04, SR01]. Zipf [Soo05, Soo13, Sor12, SBH^{+}95, SM06b, Tag70, TN87, TW11, Tao09, Tra92, TbG98, TR91, TST97, Tul85, Urz00, Urz11, VW06, VA99, VW22, VB08, Wat96a, Wat96b, WH75, YZ16, Yav74, Yeh90, Yeh02, Yeh07, You13, ZDCW12, ZCW09, ZS11, ZA95, Zör15, dGL13]. Zipf-Like [BMLRV10, DS01, GTC^{+}07, RV98]. Zipf-Mandelbrot [MPM02]. Zipf-scaling [BMT96]. Zipf-type [CM11, KG72a, KG72b, SM06b]. ZipfAllocation [KSH11]. Zipfian [Egg05c, Fed82b]. Zipf’s [Li92]. Zipf’slaw-like [Li92]. zipped [Nit05]. zone [BCK12]. Zufallsgröße [Sch73]. zum [SHC^{+}12, Tic87]. zur [Pos05, Sch73].

References

Alagoz:2010:MDB

[AA10] Ali Alagöz and Mustafa Ay. Muhasebe denetiminde Benford kanunu temelli dijital analiz. (Turkish) [Benford Law-based digital analysis of accounting oversight],

January 8, 2010. URL www.alialagoz.com.tr/doc-dr-
alialagoz-makaleleri/muhasebe_denetiminde_benfrod_
kanunu_temelli_dijital_analiz.pdf.

Archambault:2011:EMA

[AA11] J. Archambault and M. Archambault. Earnings management among firms during the pre-SEC era: a Benford’s law analysis. Accounting Historians Journal, 38(2):145–170, ???? 2011. ISSN 0148-4184.

Abdel-All:2003:GPP

[AAMAE03] N. H. Abdel-All, M. A. W. Mahmoud, and H. N. Abd-Ellah. Geometrical properties of Pareto distribution. Applied Mathematics and Computation, 145(2–3):321–339, December 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ausloos:2003:SIZ

[AB03] M. Ausloos and Ph. Bronlet. Strategy for investments from Zipf law(s). Physica A: Statistical Mechanics and its Applications, 324(1–2):30–37, June 1, 2003. CODEN PHYADX. ISSN 0378-4371 (print), 1873-2119 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0378437102018459. International Econophysics Conference IEC2002 (Bali).

Abourbih:2010:SSD

[ABBM10] Jonathan A. Abourbih, Luke Blaney, Alan Bundy, and Fiona McNeill. A single-significant-digit calculus for semi-automated guesstimation. Lecture Notes in Computer Science, 6173: 354–368, 2010. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/
content/pdf/10.1007/978-3-642-14203-1_31.

Abe:2002:STE

[Abe02] Sumiyoshi Abe. Stability of Tsallis entropy and instabilities of Rényi and normalized Tsallis entropies: a basis for q-exponential distributions. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 66(??):046134, October 24, 2002. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-2376 (electronic). URL http://link.

aps.org/doi/10.1103/PhysRevE.66.046134.

Ausloos:2016:RDC

[ACC16] Marcel Ausloos, Rosella Castellano, and Roy Cerqueti. Regularities and discrepancies of credit default swaps: a data science approach through Benford’s law. Chaos, Solitons & Fractals, 90:8–17, September 2016. CODEN CSFOEH. ISSN 0960-0779 (print), 1873-2887 (electronic).

Abergel:2013:ESR

[ACCG13] Frédéric Abergel, Bikas K. Chakrabarti, Anirban Chakraborti, and Asim Ghosh, editors. Econophysics of systemic risk and network dynamics. New Econ. Windows. Springer Italia, Milan, Italy, 2013. ISBN 88-470-2553-2, 88-470-2552-4. 119–129 pp.

ACM:2000:PIC

[ACM00] ACM, editor. Proceedings of the 2000 International Conference on Software Engineering: ICSE 2000, the new millennium: June 4–11, 2000, Limerick, Ireland. ACM Press, New York, NY 10036, USA, 2000. ISBN 1-58113-206-9, 0-7803-6325-6 (casebound), 0-7803-6326-4 (microfiche). LCCN QA76 N271 2000.

Adamic:2011:CSU

[Ada11] Lada Adamic. Complex systems: Unzipping Zipf’s law. Nature, 474(7350):164–165, June 8, 2011. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://

www.nature.com/nature/journal/v474/n7350/full/474164a.
html.

Abu-Dayyeh:2013:ESS

[ADAI13] Walid Abu-Dayyeh, Aissa Assrhani, and Kamarulzaman Ibrahim. Estimation of the shape and scale parameters of Pareto distribution using ranked set sampling. Statistical Papers, 54(1):207–225, February 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http://link.

springer.com/article/10.1007/s00362-011-0420-3.

Adhikari:1969:SRD

[Adh69] A. K. Adhikari. Some results on the distribution of the most significant digit. Sankhyā (Indian Journal of Statistics), Series B. Methodological, 31(??):413–420, ???? 1969. CODEN SANBBV. ISSN 0581-5738.

Allison:1976:LLP

[AdSPG^{+}76] Paul D. Allison, Derek de Solla Price, Belver C. Griffith, Michael J. Moravcsik, and John A. Stewart. Lotka’s Law: a problem in its interpretation and application. Social Studies of Science, 6(2):269–276, May 1, 1976. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631277600600205.

Afify:2010:EEP

[Afi10] W. M. Afify. On estimation of the exponentiated Pareto distribution under different sample schemes. Statistical Methodology, 7(2):77–83, March 2010. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1572312709000653.

Aoyama:2004:KDP

[AFS04] Hideaki Aoyama, Yoshi Fujiwara, and Wataru Souma. Kinematics and dynamics of Pareto–Zipf’s law and Gibrat’s law. Physica A: Statistical Mechanics and its Applications, 344(1–2):117–121, December 1, 2004. CODEN PHYADX. ISSN 0378-4371 (print), 1873-2119 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378437104009185.

Aigner:1970:EPL

[AG70] Dennis J. Aigner and Arthur S. Goldberger. Estimation of Pareto’s Law from grouped observations. Journal of the American Statistical Association, 65(330):712–723, June 1970. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2284582.

Aigner:1972:CEP

[AG72] Dennis J. Aigner and Arthur S. Goldberger. Corrigenda: Estimation of Pareto’s Law from grouped observations. Journal of the American Statistical Association, 67(337):252, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2284755.

Antonini:2005:RSC

[AG05] Rita Giuliano Antonini and Georges Grekos. Regular sets and conditional density: an extension of Benford’s law. Colloquium Mathematicum, 103(2):173–192, 2005. CODEN CQMAAQ. ISSN 0010-1354 (print), 1730-6302 (electronic).

Aono:2008:ISO

[AG08] June Y. Aono and Liming Guan. The impact of Sarbanes– Oxley act on cosmetic earnings management. Research in Accounting Regulation, 20(0):205–215, October 2008. CO-DEN ???? ISSN 1052-0457 (print), 2467-9895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1052045707002123.

Ainsworth:2009:QSR

[AH09] Richard Thompson Ainsworth and Urs Hengartner. Quebec’s Sales Recording Module (SRM): Fighting the Zapper, Phantomware, and tax fraud with technology. Canadian Tax Journal, 57(4):715–761, ???? 2009. CODEN ???? ISSN 0068-

9823. URL https://www.ctf.ca/ctfweb/Documents/PDF/
2009ctj/09ctj4-ainsworth.pdf.

Alvarez:2008:EFD

[AHH08] R. Michael Alvarez, Thad E. (Thad Edward) Hall, and Susan D. Hyde, editors. Election Fraud: Detecting and Deterring Electoral Manipulation. Brookings Institution Press, Washington, DC, USA, 2008. ISBN 0-8157-0138-1 (hardcover), 0-8157-0139-X (paperback). xiii + 255 pp. LCCN

JK1994 .E43 2008. URL http://www.loc.gov/catdir/toc/
ecip0810/2008005719.html.

Ausloos:2014:BBL

[AHI14] M. Ausloos, C. Herteliu, and B. Ileanu. Breakdown of Benford’s Law for birth data. arXiv.org, ??(??):17, October 6,

2014. URL http://arxiv.org/abs/1410.1755v1.

Ausloos:2015:BBL

[AHI15] M. Ausloos, C. Herteliu, and B. Ileanu. Breakdown of Benford’s Law for birth data. Physica A: Statistical Mechanics and its Applications, 419(0):736–745, 2015. CO-DEN PHYADX. ISSN 0378-4371 (print), 1873-2119 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378437114008796.

Ausloos:1999:PMK

[AI99] M. Ausloos and K. Ivanova. Precise (m, k)-Zipf diagram analysis of mathematical and financial time series when m = 6, k = 2. Physica A: Statistical Mechanics and its Applications, 270(3–4):526–542, August 15, 1999. CO-DEN PHYADX. ISSN 0378-4371 (print), 1873-2119 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378437199001788.

Arshadi:2014:BLB

[AJ14] Laleh Arshadi and Amir Hossein Jahangir. Benford’s Law behavior of Internet traffic. Journal of Network and Computer Applications, 40(0):194–205, 2014. CO-DEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804513001951.

Alexopoulos:2014:BLU

[AL14a] Theodoros Alexopoulos and Stefanos Leontsinis. Benford’s Law and the Universe. arXiv.org, ??(??):6, January 23, 2014.

URL http://arxiv.org/abs/1401.5794v2.

Alexopoulos:2014:BLA

[AL14b] Theodoros Alexopoulos and Stefanos Leontsinis. Benford’s law in astronomy. Journal of Astrophysics and Astronomy, 35(4):639–648, 2014. CODEN JASRD7. ISSN 0250-6335 (print), 0973-7758 (electronic). URL http://link.

springer.com/article/10.1007/s12036-014-9303-z. See

follow-up [HF16].

Alexanderson:2005:BRG

[Ale05] Gerald L. Alexanderson. Book review: Gamma: Exploring Euler’s Constant, by Julian Havil. Princeton, Princeton University Press 2003. xxiii + 266 pages. US $29.95. ISBN 0-691-09983-9. The Mathematical Intelligencer, 27(1):86–88, December 2005. CODEN MAINDC. ISSN 0343-6993 (print),

1866-7414 (electronic). URL http://link.springer.com/
article/10.1007/BF02984820.

Alexander:2009:RUB

[Ale09] James C. Alexander. Remarks on the use of Benford’s Law. Working paper, Department of Mathematics and Cognitive Science, Case Western Reserve University, Cleveland, OH,

USA, November 23, 2009. URL http://papers.ssrn.com/
sol3/papers.cfm?abstract_id=1505147.

Allaart:1997:ISC

[All97] Pieter C. Allaart. An invariant-sum characterization of Benford’s Law. Journal of Applied Probability, 34(1):288– 291, March 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3215195.

Amir:2016:SNE

[ALT16] Ariel Amir, Mikhail Lemeshko, and Tadashi Tokieda. Surprises in numerical expressions of physical constants. American Mathematical Monthly, 123(6):609–612, June 2016. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/10.4169/
amer.math.monthly.123.6.609.

Aleksandrova:2019:BLD

[ALY19] Elena B. Aleksandrova, Daria S. Lavrova, and A. V. Yarmak. Benford’s Law in the detection of DoS attacks on industrial systems. Automatic Control and Computer Sciences, 53(8): 954–962, ???? 2019. CODEN ACCSCE. ISSN 0146-4116 (print), 1558-108X (electronic). URL https://dblp.org/db/

journals/accs/accs53.html#AleksandrovaLY19.

Amin:2007:TVA

[Ami07] Zeinab H. Amin. Tests for the validity of the assumption that the underlying distribution of life is Pareto. Journal of Applied Statistics, 34(2):195–201, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Abrantes-Metz:2012:LM

[AMKMS12] Rosa M. Abrantes-Metz, Michael Kraten, Albert D. Metz, and Gim S. Seow. Libor manipulation? Journal of Banking & Finance, 36(1):136–150, 2012. ISSN 0378-4266 (print),

1872-6372 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378426611002032.

Aban:2006:PET

[AMP06] Inmaculada B. Aban, Mark M. Meerschaert, and Anna K. Panorska. Parameter estimation for the truncated Pareto distribution. Journal of the American Statistical Association, 101 (473):270–277, March 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Abrantes-Metz:2011:TLR

[AMVBJ11] R. M. Abrantes-Metz, S. B. Villas-Boas, and G. Judge. Tracking the Libor rate. Applied Economics Letters, 18(10):893– 899, ???? 2011. ISSN 1350-4851 (print), 1466-4291 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/13504851.2010.515197.

Anastasi:2003:NFI

[Ana03] Joe Anastasi. The new forensics: investigating corporate fraud and the theft of intellectual property. Wiley, New York, NY, USA, 2003. ISBN 0-471-26994-8 (hardcover). xiv + 270 pp. LCCN HV8079.W47 A5 2003. URL http://www.loc.

gov/catdir/bios/wiley044/2003001697.html;   http:/
/www.loc.gov/catdir/description/wiley037/2003001697.
html;   http://www.loc.gov/catdir/toc/wiley031/2003001697.
html.

Anderson:2009:CRC

[And09] Thomas Anderson. Conference reviewing considered harmful. Operating Systems Review, 43(2):108–116, April 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Andrew:2011:OBM

[And11] Alex M. Andrew. Obituary: Benoît Mandelbrot: 20 November 1924–14 October 2010. Kybernetes, 40(1– 2):??, ???? 2011. CODEN KBNTA3. ISSN 0368-492X (print), 1758-7883 (electronic). URL http://en.

wikipedia.org/wiki/ernst_von_glasersfeld;   http:/
/www.boston.com/news/world/europe/articles/2010/11/
12/intellectual_ernst_von_glasersf;   http://www.emeraldinsight.
com/journals.htm?issn=0368-492X&volume=40&issue=1/
2&articleid=1921825;   http://www.guardian.co.uk/science/
2010/oct/17/benoit-mandelbrot-obituary;   http://www.
math.yale.edu/mandelbrot/;   http://www.telegraph.
co.uk/news/obituaries/science-obituaries/8069558/benoit-
mandelbrot.html.

Ausloos:2016:QQP

[ANFF16] Marcel Ausloos, Olgica Nedic, Agata Fronczak, and Piotr Fronczak. Quantifying the quality of peer reviewers through Zipf’s law. Scientometrics, 106(1):347–368, January 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/s11192-015-1704-5.

Anonymous:1932:FL

[Ano32] Anonymous. Fishpole lamp. Web site, April 30, 1932. URL

http://scienceservice.si.edu/pages/012020.htm. This

Web site reports that General Electric researcher Frank Benford invented the electric light pointer, the forerunner of the modern laser pointer.

Anonymous:1950:BRZ

[Ano50] Anonymous. Book review: Zipf, George K., Human behavior and the principle of least effort. Cambridge, (Mass.): Addison-Wesley, 1949, pp. 573. Journal of Clinical Psychology, 6(3): 306, 1950. CODEN JCPYAO. ISSN 0021-9762 (print), 1097-4679 (electronic).

Anonymous:1993:WFR

[Ano93] Anonymous. Why failure rates observe Zipf’s law in operational software: M. Trachtenberg. IEEE Transactions on Reliability, 41(3), 386 (1992). Microelectronics and Reliability, 33(13):2062, October 1993. CODEN MCR-LAS. ISSN 0026-2714 (print), 1872-941X (electronic).

URL http://www.sciencedirect.com/science/article/
pii/002627149390406O.

Anonymous:2002:ZLM

[Ano02] Anonymous. Zipf’s law and maximum sustainable growth. Statistics & Probability Letters, 56(2):??, January 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165188913000341.

Anonymous:2005:UZL

[Ano05] Anonymous. The use of Zipf’s law in animal communication analysis. Animal Behaviour, 69(1):??, January 20, 2005. CODEN ANBEA8. ISSN 0003-3472 (print), 1095-

8282 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0003347204003471.

Anonymous:2007:ZL

[Ano07] Anonymous. Zipf’s law, 2007. URL http://en.wikipedia.

org/wiki/Zipf's_law.

Anonymous:2011:BLH

[Ano11] Anonymous. Benford’s Law — how mathematics can detect fraud! Web video lecture (10m47s)., April 9, 2011. URL

https://www.youtube.com/watch?v=vIsDjbhbADY.

Anonymous:2012:SB

[Ano12] Anonymous. The scam busters. The Economist, ??(??):??, December 15, 2012. CODEN EONOEH. ISSN 0013-0613 (print),

1476-8860 (electronic). URL http://www.economist.com/
news/finance-and-economics/21568364-how-antitrust-
economists-are-getting-better-spotting-cartels-scam-
busters.

Anonymous:2013:NBL

[Ano13] Anonymous. Number 1 and Benford’s Law — numberphile. Web video lecture., January 20, 2013. URL https://www.

youtube.com/watch?v=XXjlR2OK1kM.

Ackermann:2007:DMB

[ANRV07] Heiner Ackermann, Alantha Newman, Heiko Röglin, and Berthold Vöcking. Decision-making based on approximate and smoothed Pareto curves. Theoretical Computer Science, 378(3):253–270, June 9, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ansari:2009:PAV

[Ans09] Ali Ansari. Preliminary analysis of the voting figures in Iran’s 2009 Presidential election. Report MENAP PP 2009/01, Chatham House and the Institute of Iranian Studies, University of St Andrews, St Andrews, UK, June 21,

2009. URL https://www.chathamhouse.org/sites/files/
chathamhouse/public/Research/Middle%20East/iranelection0609.
pdf.

Antonini:1991:NUD

[Ant91] Rita Giuliano Antonini. On the notion of uniform distribution mod 1. Fibonacci Quarterly, 29(3):230–234, August 1991. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.

math.ca/Scanned/29-3/antonini.pdf.

Andriotis:2013:JSD

[AOT13] Panagiotis Andriotis, George Oikonomou, and Theo Tryfonas. JPEG steganography detection with Benford’s Law. Digital Investigation, 9(3–4):246–257, 2013. ISSN 1742-2876 (print),

1873-202X (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1742287613000066.

Arnold:1989:BEP

[AP89] Barry C. Arnold and S. James Press. Bayesian estimation and prediction for Pareto data. Journal of the American Statistical Association, 84(408):1079–1084, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

AyllonBurguillo:1999:PDS

[AP99] Juan David Ayllón Burguillo and Manuel Perera Domínguez. El primer dígito significativo. (Spanish) [the first significant digit]. Epsilon: Revista de la Sociedad Andaluza de Educación Matemática “Thales”, 15(3(45)):339–352, ???? 1999. ISSN 1131-9321.

Aldous:2010:WCO

[AP10] David Aldous and Tung Phan. When can one test an explanation? Compare and contrast Benford’s Law and the fuzzy CLT. The American Statistician, 64(3):221–227, August 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/abs/
10.1198/tast.2010.09098.

Altamirano:2009:GTU

[AR09] Carlo Altamirano and Alberto Robledo. Generalized thermodynamics underlying the laws of Zipf and Benford. In J. Zhou, editor, Complex Sciences, volume 5, pages 2232–?? Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 3-642-02468-8. URL http://adsabs.

harvard.edu/abs/2009cosc.conf.2232A.

Altamirano:2010:PTS

[AR10] Carlo Altamirano and Alberto Robledo. Possible thermodynamic structure underlying the laws of Zipf and Benford. ArXiv e-prints, August 10, 2010. URL http://arxiv.org/

abs/1008.1614.

Altamirano:2011:PTS

[AR11] C. Altamirano and A. Robledo. Possible thermodynamic structure underlying the laws of Zipf and Benford. European Physical Journal B: Condensed Matter and Complex Systems, 81(3):345–351, June 2011. CODEN EPJBFY. ISSN 1434-6028 (print), 1434-6036 (electronic). URL http://adsabs.

harvard.edu/abs/2011EPJB...81..345A.

Alali:2013:BLA

[AR13] Fatima A. Alali and Silvia Romero. Benford’s Law: Analyzing a decade of financial data. Journal of Emerging Technologies in Accounting, 10(1):1–39, December 2013. ISSN 1554-1908 (print), 1558-7940 (electronic). URL http://aaajournals.

org/doi/pdf/10.2308/jeta-50749.

Alvarez-Ramirez:2003:ZMS

[ARMG03] Jose Alvarez-Ramirez, Monica Meraz, and Gustavo Gallegos. Zipf–Mandelbrot scaling law for world track records. Physica A: Statistical Mechanics and its Applications, 328 (3–4):545–560, October 15, 2003. CODEN PHYADX. ISSN 0378-4371 (print), 1873-2119 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S037843710300579X.

Aron:2013:CFM

[Aro13] Jacob Aron. Crime-fighting maths law confirms planetary riches. New Scientist, 220(2945):12–, 2013. CO-DEN NWSCAL. ISSN 0262-4079 (print), 1364-8500 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S026240791362775X.

Aron:2014:LLN

[Aro14] Jacob Aron. Looking for life in numbers. New Scientist, 222(2968):46–47, 2014. CODEN NWSCAL. ISSN 0262-4079 (print), 1364-8500 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0262407914609258.

Anderson:2010:BLC

[ARS10] Theresa Anderson, Larry Rolen, and Ruth Stoehr. Benford’s law for coefficients of modular forms and partition functions. ArXiv e-prints, September 3, 2010. URL http://arxiv.org/

abs/1009.0780.

Anderson:2011:BLC

[ARS11] Theresa C. Anderson, Larry Rolen, and Ruth Stoehr. Benford’s law for coefficients of modular forms and partition functions. Proceedings of the American Mathematical Society, 139(5):1533–1541, 2011. CODEN PAMYAR. ISSN 0002-9939 (print), 1088-6826 (electronic). URL http://adsabs.

harvard.edu/abs/2010arXiv1009.0780A.

Arthur:1994:IRP

[Art94] W. Brian Arthur. Increasing returns and path dependence in the economy. Economics, cognition, and society. University of Michigan Press, Ann Arbor, MI, USA, 1994. ISBN 0-472-09496-3, 0-472-06496-7 (paperback). xx + 201 pp. LCCN HD69.S5 A77 1994. With a foreword by Kenneth J. Arrow.

Adhikari:1968:DMS

[AS68] A. K. Adhikari and B. P. Sarkar. Distribution of most significant digit in certain functions whose arguments are random variables. Sankhyā (Indian Journal of Statistics), Series B. Methodological, 30(??):47–58, ???? 1968. CODEN SANBBV. ISSN 0581-5738.

Acebo:2005:BLN

[AS05] E. Acebo and M. Sbert. Benford’s Law for natural and synthetic images. In L. Neumann, M. Sbert, B. Gooch, and W. Purgathofer, editors, Conference: Computational Aesthetics 2005: Eurographics Workshop on Computational Aesthetics in Graphics, Visualization and Imaging 2005, Girona, Spain, May 18–20, 2005, pages 169–176. ????, ????, 2005.

Abushal:2015:EPP

[AS15] Tahani A. Abushal and Ahmed A. Soliman. Estimating the Pareto parameters under progressive censoring data for constant-partially accelerated life tests. Journal of Statistical Computation and Simulation, 85(5):917–934, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Asai:1981:GFB

[Asa81] Isao Asai. A general formulation of Bradford’s distribution: the graph-oriented approach. Journal of the American Society for Information Science, 32(2):113–119, March 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Auerbach:1913:GBG

[Aue13] Felix Auerbach. Das Gesetz der Bevölkerungskonzentration. (German) [The law of population concentration]. Petermann’s Geographische Mitteilungen, 59(1):74–76, ???? 1913. CODEN PGGMA3. ISSN 0031-6229. URL http://hdl.handle.net/

2027/mdp.39015035583528?urlappend=%3Bseq=122. This

paper is reported in [New05] to be the first discovery of Zipf’s Law in connection with city populations (see [Zip32]). See also [Est16].

Ausloos:2014:ZMP

[Aus14] Marcel Ausloos. Zipf–Mandelbrot–Pareto model for coauthorship popularity. Scientometrics, 101(3):1565–1586, December 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1302-y.

Aerts:2008:CDD

[AVV08] Walter Aerts, Geert Van Campenhout, and Tom Van Caneghem. Clustering in dividends: Do managers rely on cognitive reference points? Journal of Economic Psychology, 29(3):276–284, 2008. ISSN 0167-4870 (print), 1872-7719 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167487007000864.

Apers:1989:VLD

[AW89] P. M. G. (Petrus Maria Gerardus) Apers and Gio Wiederhold, editors. Very large data bases: proceedings: proceedings of the Fifteenth International Conference on Very Large Data Bases, August 22–25, 1989, Amsterdam, The Netherlands. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1989. ISBN 1-55860-101-5. LCCN QA 76.9 D3 I61 1989.

Axtell:2001:ZDU

[Axt01] R. L. Axtell. Zipf distribution of U.S. firm sizes. Science, 293(5536):1818–1820, September 7, 2001. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http:/

/www.sciencemag.org/content/293/5536/1818.

Alves:2014:BLA

[AYS14] Alexandre Donizeti Alves, Horacio Hideki Yanasse, and Nei Yoshihiro Soma. Benford’s Law and articles of scientific journals: comparison of JCR(R) and Scopus data. Scientometrics, 98(1):173–184, January 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/s11192-013-
1030-8.pdf.

Alves:2016:ABI

[AYS16] Alexandre Donizeti Alves, Horacio Hideki Yanasse, and Nei Yoshihiro Soma. An analysis of bibliometric indicators to JCR according to Benford’s Law. Scientometrics, 107(3):1489–1499, June 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/s11192-016-
1908-3.pdf.

Barabasi:1999:ESR

[BA99] Albert-László Barabási and Réka Albert. Emergence of scaling in random networks. Science, 286(5439):509–512, October 15, 1999. CODEN SCIEAS. ISSN 0036-8075 (print),

1095-9203 (electronic). URL http://www.sciencemag.org/
content/286/5439/509.

Bach:1998:EPM

[Bac98] Eric Bach. Efficient prediction of Marsaglia–Zaman random number generators. IEEE Transactions on Information Theory, 44(3):1253–1257, 1998. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Bahmann:2015:AMM

[Bah15] Paul Bahmann. Eine Analyse von monetären, makroökonomischen Daten der BRD und DDR: Ziffernanalyse auf Grundlage des Benford’s Law. (German) [An analysis of monetary, macroeconomic data of the FRG and the GDR: numeric analysis on the basis of Benford’s Law]. AV Akademikerverlag, Saarbrücken, Germany, 2015. ISBN 3-639-80882-7. 76 pp. LCCN ????

Bak:1996:HNW

[Bak96] Per Bak. How nature works: the science of self-organized criticality. Copernicus, New York, NY, USA, 1996. ISBN 0-387-94791-4 (hardcover). xiii + 212 + 8 pp. LCCN QC173.4.C74 B34 1996.

Bakulina:2007:AZL

[Bak07] M. P. Bakulina. Application of Zipf’s law for text compression. Diskretn. Anal. Issled. Oper. Ser. 2, 14(2):3–13, 2007. CODEN ???? ISSN 1560-9901.

Ball:1966:BRB

[Bal66] R. J. Ball. Book review: Random Processes and the Growth of Firms: A Study of the Pareto Law, by Josef Steindl. Journal of the Royal Statistical Society. Series A (General), 129(4): 600–601, ???? 1966. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/2982275.

Balanzario:2012:BLM

[Bal12] E. P. Balanzario. Benford’s law for mixtures. Preprint, ????, ????, ???? 2012.

Balanzario:2015:BLM

[Bal15] Eugenio P. Balanzario. Benford’s Law for mixtures. Communications in Statistics: Theory and Methods, 44(4):698–709, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Banyard:2000:NAF

[Ban00] Peter Banyard. A new aid for fraud detection. Credit Management: journal of the Institute of Credit Management, ?? (??):32–33, March 2000. ISSN 0265-2099.

Banks:2013:BLM

[Ban13] David G. Banks. Benford’s Law made easy: Common software eases use of fraud search theorem. Fraud Magazine, ??(??):??, January/February 2013. ISSN 1553-6645. URL http://www.

fraud-magazine.com/article.aspx?id=4294968485.

Barnes:2011:LCS

[Bar11] K. Barnes. Letter to China Shen Zhou Mining and Resources (AMEX: SHZ) current shareholders and potential investors from Absaroka Capital. Absaroka Capital Management, LLP, Wyoming, USA., 2011.

Basu:1992:RHD

[Bas92] Aparna Basu. Research: Hierarchical distributions and Bradford’s Law. Journal of the American Society for Information Science, 43(7):494–500, August 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Basu:1994:LEE

[Bas94] Aparna Basu. Letters to the Editor: The elusive phenomenon of Bradford’s Law. Journal of the American Society for Information Science, 45(1):59–60, January 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Batty:2008:PSS

[Bat08] Michael Batty. Perspective: The size, scale, and shape of cities. Science, 318(5864):769–771, February 8, 2008. CO-DEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic). URL http://www.sciencemag.org/content/319/
5864/769.abstract.

Battersby:2009:SAH

[Bat09] Stephen Battersby. Statistical analyses hint at fraud in Iranian election. New Scientist, 202(2714):10–??, June 24, 2009. CODEN NWSCAL. ISSN 0262-4079 (print), 1364-

8500 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0262407909616821.

Bawden:2011:BMS

[Baw11] David Bawden. Benoît Mandelbrot and the self-similarity of information. Journal of Documentation, 67(2):??, ????

2011. CODEN JDOCAS. ISSN 0022-0418 (print), 1758-

7379 (electronic).  URL http://www.emeraldinsight.
com/journals.htm?issn=0022-0418&volume=67&issue=2&
articleid=1911717.

Billera:1976:PSC

[BB76] Louis J. Billera and Robert E. Bixby. Pareto surfaces of complexity 1. SIAM Journal on Applied Mathematics, 30(1):81– 89, January 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Barlow:1985:RED

[BB85] J. L. Barlow and E. H. Bareiss. On roundoff error distributions in floating point and logarithmic arithmetic. Computing: Archiv für Informatik und Numerik, 34(4):325–347, December 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Becker:2018:BLC

[BBC^{+}18] Thealexa Becker, David Burt, Taylor C. Corcoran, Alec Greaves-Tunnell, Joseph R. Iafrate, Joy Jing, Steven J. Miller, Jaclyn D. Porfilio, Ryan Ronan, Jirapat Samranvedhya, Frederick W. Strauch, and Blaine Talbut. Benford’s law and continuous dependent random variables. Annals of Physics, 388: 350–381, 2018. CODEN APNYA6. ISSN 0003-4916 (print), 1096-035x (electronic).

Brahler:2010:EMS

[BBE10] Gernot Brähler, Markus Bensmann, and Anna-Lena Emke. Der Einsatz mathematisch-statistischer Methoden in der digitalen Betriebsprüfung. (German) [The use of mathematical and statistical methods in the digital audit]. In Ilmenauer Schriften zur Betriebswirtschaftslehre. (German) [Illmenauer writings on economics], volume 4/2010, page 48. Verlag proWiWi e.V., Ilmenau, Germany, April 2010. ISBN 3-940882-23-2. ISSN 1866-2145 (print), 2192-4643 (elec-

tronic). LCCN ???? URL https://www.econstor.eu/
dspace/bitstream/10419/55680/1/665348576.pdf.

Berger:2005:ODD

[BBH05] Arno Berger, Leonid A. Bunimovich, and Theodore P. Hill. One-dimensional dynamical systems and Benford’s law. Transactions of the American Mathematical Society, 357(1):

197–219, 2005. CODEN TAMTAM. ISSN 0002-9947 (print), 1088-6850 (electronic).

Brahler:2011:BGS

[BBJ11] Gernot Brähler, Markus Bensmann, and Hans-Ralph Jakobi. Das Benfordsche Gesetz und seine Anwendbarkeit bei der digitalen Prüfung von Fahrtenbüchern. (German) [The Benford law and its applicability in the digital examination of logbooks]. In ????, editor, Ilmenauer Schriften zur Betriebswirtschaftslehre, volume 3/2011, page ?? Verlag proWiWi e.V., Ilmenau, Germany, March 2011. ISBN 3-940882-28-3. ISSN 1866-2145 (print), 2192-4643 (electronic). LCCN

???? URL https://www.econstor.eu/dspace/bitstream/
10419/55703/1/662518055.pdf.

Bermano:2015:DST

[BBK^{+}15] Amit Bermano, Thabo Beeler, Yeara Kozlov, Derek Bradley, Bernd Bickel, and Markus Gross. Detailed spatio-temporal reconstruction of eyelids. ACM Transactions on Graphics, 34 (4):44:1–44:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Benguigui:2007:DMC

[BBL07] Lucien Benguigui and Efrat Blumenfeld-Lieberthal. A dynamic model for city size distribution beyond Zipf’s law. Physica A: Statistical Mechanics and its Applications, 384(2):613–627, October 15, 2007. CODEN PHYADX. ISSN 0378-4371 (print), 1873-2119 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378437107006061.

Baek:2011:ZLU

[BBM11a] Seung Ki Baek, Sebastian Bernhardsson, and Petter Minnhagen. Zipf’s law unzipped. New Journal of Physics, 13(4):043004, ???? 2011. CODEN NJOPFM. ISSN 1367-2630. URL http:

//iopscience.iop.org/1367-2630/13/4/043004.

Bernhardsson:2011:PPM

[BBM11b] Sebastian Bernhardsson, Seung Ki Baek, and Petter Minnhagen. A paradoxical property of the monkey book. Journal of Statistical Mechanics: Theory and Experiment, 2011(??):P07013, July 2011. CODEN JSMTC6. ISSN 1742-5468. URL http:/

/arxiv.org/abs/1103.2681;   http://iopscience.iop.
org/1742-5468/2011/07/P07013/.

Bhattacharyya:2005:COP

[BCC05] P. Bhattacharyya, A. Chatterjee, and B. K. Chakrabarti. A common origin of the power law distributions in models of market and earthquake. ArXiv e-prints, 2005. URL http://

arxiv.org/abs/physics/0510038.

Barabesi:2018:GFT

[BCCP18] Lucio Barabesi, Andrea Cerasa, Andrea Cerioli, and Domenico Perrotta. Goodness-of-fit testing for the Newcomb– Benford Law with application to the detection of customs fraud. Journal of Business & Economic Statistics, 36(2): 346–358, 2018. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2016.1172014.

Barabesi:2022:CTB

[BCCP22] Lucio Barabesi, Andrea Cerasa, Andrea Cerioli, and Domenico Perrotta. On characterizations and tests of Benford’s Law. Journal of the American Statistical Association, 117(540):1887–1903, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Barabesi:2023:SMB

[BCD23] Lucio Barabesi, Andrea Cerioli, and Marco Di Marzio. Statistical models and the Benford hypothesis: a unified framework. TEST, 32(4):1479–1507, December 2023. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-023-00881-y.

Berinde:2009:SOH

[BCIS09] Radu Berinde, Graham Cormode, Piotr Indyk, and Martin J. Strauss. Space-optimal heavy hitters with strong error bounds. In Paredaens and Su [PS09], pages 157–166. ISBN 1-60558-553-X. LCCN ????

Bharati:2012:CCO

[BCK12] Rakesh Bharati, Susan J. Crain, and Vincent Kaminski. Clustering in crude oil prices and the target pricing zone hypothesis. Energy Economics, 34(4):1115–1123, 2012. ISSN 0140-9883 (print), 1873-6181 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0140988311002258.

Brown:2002:ICF

[BCM02] Philip Brown, Angeline Chua, and Jason Mitchell. The influence of cultural factors on price clustering: Evidence from Asia Pacific stock markets. Pacific-Basin Finance Journal, 10(3):307–332, 2002. ISSN 0927-538X (print), 1879-

0585 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0927538X02000495. 13th Annual

PACAP/FMA Finance Conference.

Barabesi:2021:FBL

[BCP21] Lucio Barabesi, Andrea Cerioli, and Domenico Perrotta. Forum on Benford’s law and statistical methods for the detection of frauds [Editorial]. Stat. Methods Appl., 30(3):767–778, 2021. ISSN 1618-2510.

Brown:1970:MOU

[BD70] J. L. Brown, Jr. and R. L. Duncan. Modulo one uniform distribution of the sequence of logarithms of certain recursive sequences. Fibonacci Quarterly, 8(5):482–486, December 1970. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.

math.ca/Scanned/8-5/brown.pdf. See remark [KsS73].

Best:2014:BBZ

[BDE^{+}14] Andrew Best, Patrick Dynes, Xixi Edelsbrunner, Brian Mc-Donald, Steven J. Miller, Kimsy Tor, Caroline Turnage-Butterbaugh, and Madeleine Weinstein. Benford behavior of Zeckendorf decompositions. Fibonacci Quarterly, 52(5): 35–??, December 2014. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Papers1/52-5/Best-
Benford.pdf.

Best:2017:BBG

[BDE^{+}17] Andrew Best, Patrick Dynes, Xixi Edelsbrunner, Brian Mc-Donald, Steven J. Miller, Kimsy Tor, Caroline Turnage-Butterbaugh, and Madeleine Weinstein. Benford behavior of generalized Zeckendorf decompositions. In M. Nathanson, editor, Springer Proceedings in Mathematics & Statistics, pages 25–37. Springer International Publishing, Cham, Switzerland,

Bumby:1969:FAM

[BE69] R. Bumby and E. Ellentuck. Finitely additive measures and the first digit problem. Fundamenta mathematicae, 65(??): 33–42, ???? 1969. CODEN ???? ISSN 0016-2736 (print), 1730-6329 (electronic).

Berger:2014:BSL

[BE14a] Arno Berger and Gideon Eshun. Benford solutions of linear difference equations. In Ziyad AlSharawi, Jim M. Cushing, and Saber Elaydi, editors, Theory and Applications of Difference Equations and Discrete Dynamical Systems: ICDEA, Muscat, Oman, May 26–30, 2013, volume 102 of Springer Proceedings in Mathematics & Statistics, pages 23–60. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. ISBN 3-662-44139-X, 3-662-44140-3 (e-book). LCCN QA218 .T446 2014.

Berger:2014:CBL

[BE14b] Arno Berger and Gideon Eshun. A characterization of Benford’s Law in discrete-time linear systems. Journal of Dynamics and Differential Equations, ??(??):1–39, ???? 2014. CODEN JDDEEH. ISSN 1040-7294 (print), 1572-9222 (electronic).

Berger:2016:CBL

[BE16] Arno Berger and Gideon Eshun. A characterization of Benford’s Law in discrete-time linear systems. Journal of Dynamics and Differential Equations, 28(2):431–469, 2016. CODEN JDDEEH. ISSN 1040-7294 (print), 1572-9222 (electronic).

Beckett:2014:TPA

[Bec14] Phil Beckett. Taking a preventative approach towards the Bribery Act. Computer Fraud & Security, 2014(6): 16–18, 2014. ISSN 1361-3723 (print), 1873-7056 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1361372314705020.

Beeli:2007:FCG

[BEJ07] Gian Beeli, Michaela Esslen, and Lutz Jäncke. Frequency correlates in grapheme-color synaesthesia. Psychological Science, 18(9):788–792, 2007. URL http://pss.sagepub.com/

content/18/9/788.abstract.

Belevitch:1988:PQD

[Bel88] Vitold Belevitch. À propos de la queue de la distribution de Zipf. (French) [On the tail of the Zipf distribution]. Ann. Soc. Sci. Bruxelles Sér. I, 102(3):73–85 (1989), 1988. CODEN ASSBAH. ISSN 0037-959X.

Benford:1938:LAN

[Ben38] Frank Benford. The law of anomalous numbers. Proceedings of the American Philosophical Society, 78(4):551–572, March 1938. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL http://links.jstor.org/sici?sici=
0003-049X%2819380331%2978%3A4%3C551%3ATLOAN%3E2.0.

CO%3B2-G. See comments about Benford’s biased rounding practices [DF79].

Benford:1943:PAG

[Ben43] Frank Benford. The probable accuracy of the general physical constants. Physical Review, 63(5–6):212, March 1943. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://link.aps.org/doi/10.
1103/PhysRev.63.212.

Bennett:1975:SDI

[Ben75] John M. Bennett. Storage design for information retrieval: Scarrott’s conjecture and Zipf’s law. In Int. Comput. Symp. 1975, Proc., Antibes, (1975), pages 233–237. ????, ????, 1975.

Bennett:1984:ZLS

[Ben84] J. M. Bennett. Zipf’s Law, structured programming and creativity. Australian Computer Journal, 16(4):122–129, 1984. CODEN ACMJB2. ISSN 0004-8917.

Benjamin:2019:LSB

[Ben19] Arthur T. Benjamin. The long and the short of Benford’s law. Math Horizons, 26(4):8–9, 2019. ISSN 1072-4117.

Berton:1995:HGT

[Ber95] L. Berton. He’s got their number. Scholar uses math to foil financial fraud. Wall Street Journal, ??(??):??, July 10, 1995. CODEN WSJOAF. ISSN 0099-9660.

Berger:2003:BLP

[Ber03] Arno Berger. Benford’s law in power-like nonautonomous dynamical systems. Preprint, Vienna University of Technology, Vienna, Austria, 2003.

Berger:2004:DDU

[Ber04] A. Berger. Dynamics and digits: on the ubiquity of Benford’s Law. Preprint, Institute of Mechanics, Vienna University of Technology, Vienna, Austria, November 19, 2004.

3 pp. URL http://www.math.leidenuniv.nl/~verduyn/
EQUADIFF03/MS14/berger_eqdiff03.pdf.

Berger:2005:DDU

[Ber05a] A. Berger. Dynamics and digits: on the ubiquity of Benford’s law. In EQUADIFF 2003, pages 693–695. World Sci. Publ., Hackensack, NJ, 2005.

Berger:2005:BLP

[Ber05b] Arno Berger. Benford’s law in power-like dynamical systems. Stochastics and Dynamics, 5(4):587–607, 2005. CODEN ???? ISSN 0219-4937.

Berger:2005:MDD

[Ber05c] Arno Berger. Multi-dimensional dynamical systems and Benford’s law. Discrete and Continuous Dynamical Systems. Series A, 13(1):219–237, 2005. CODEN ???? ISSN 1078-0947 (print), 1553-5231 (electronic).

Berger:2010:LSD

[Ber10] A. Berger. Large spread does not imply Benford’s Law. Preprint, Department of Mathematical and Statistical Sciences, University of Alberta, Edmonton, AB, Canada,

2010.  URL http://www.math.ualberta.ca/~aberger/
Publications.html.

Berger:2011:SDP

[Ber11] Arno Berger. Some dynamical properties of Benford sequences. Journal of Difference Equations and Applications, 17(2):137–159, 2011. CODEN JDEAEA. ISSN 1023-6198 (print), 1563-5120 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1080/10236198.2010.
549012.

Berger:2015:MLF

[Ber15a] Arno Berger. Most linear flows on R^{d} are Benford. Journal of Differential Equations, 259(5):1933–1957, 2015. CO-DEN JDEQAK. ISSN 0022-0396 (print), 1090-2732 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022039615001527.

Berzunza:2015:YPR

[Ber15b] Gabriel Berzunza. Yule processes with rare mutation and their applications to percolation on b-ary trees. Electronic Journal of Probability, 20(??):43:1–43:23, ???? 2015. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/
view/3789.

Bradley:2009:WBL

[BF09] Jonathan R. Bradley and David L. Farnsworth. What is Benford’s Law? Teaching Statistics, 31(1):2–6, January 2009. CO-DEN ???? ISSN 0141-982X (print), 1467-9639 (electronic).

Bachmann:2010:APC

[BFLB10] Valérie Bachmann, Martin H. Fischer, Hans-Peter Landolt, and Peter Brugger. Asymmetric prefrontal cortex functions predict asymmetries in number space. Brain and Cognition, 74(3):306–311, December 2010. CODEN ???? ISSN 0278-2626 (print), 1090-2147 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0278262610001272.

Baldi:2003:MIW

[BFS03] Pierre Baldi, Paolo Frasconi, and Padhraic Smyth. Modeling the Internet and the Web: probabilistic methods and algorithms. Wiley, New York, NY, USA, 2003. ISBN 0-470-86492-3 (e-book), 0-470-84906-1. xix + 285 pp. LCCN TK5105.875.I57 B35 2003eb.

Blackwell:1954:TGS

[BG54] David Blackwell and Meyer A. Girshick. Theory of Games and Statistical Decisions. Wiley, New York, NY, USA, 1954. xi + 355 pp. LCCN QA269 .B5. URL http://catalog.

hathitrust.org/Record/000577721.

Brookes:1978:FRD

[BG78] Bertram C. Brookes and José M. Griffiths. Frequencyrank distributions. Journal of the American Society for Information Science, 29(1):5–13, January 1978. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (elec-

tronic). URL http://onlinelibrary.wiley.com/doi/10.
1002/asi.4630290104/abstract.

Bauer:2009:DDF

[BG09] J. Bauer and J. Gross. Difficulties detecting fraud? The use of Benford’s Law on regression tables. Report ??, Institute of Sociology, LMU Munich, Munich, Germany, ???? 2009.

Bauer:2010:DDF

[BG10] J. Bauer and J. Gross. Difficulties detecting fraud? The use of Benford’s Law on regression tables. Report, Institute of Sociology, Ludwig-Maximilians-Universität, München, Germany, 2010. ???? pp.

Bauer:2011:DDF

[BG11a] Johannes Bauer and Jochen Gross. Difficulties detecting fraud? The use of Benford?s Law on regression tables. Journal of Economics and Statistics (Jahrbücher für Nationalökonomie und Statistik), 231(5–6):733–748, Novem-

ber 2011. URL http://ideas.repec.org/a/jns/jbstat/
v231y2011i5-6p733-748.html.

Breunig:2011:SEI

[BG11b] Christian Breunig and Achim Goerres. Searching for electoral irregularities in an established democracy: Applying Benford’s Law tests to Bundestag elections in Unified Germany. Electoral Studies, 30(3):534–545, September 2011. CODEN ???? ISSN 0261-3794 (print), 1873-6890 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0261379411000370. Special Symposium on

the Politics of Economic Crisis.

Beck:2015:PPI

[BG15] Mélanie Beck and Martin J. Gander. On the positivity of Poisson integrators for the Lotka–Volterra equations. BIT Numerical Mathematics, 55(2):319–340, June 2015. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125

(electronic). URL http://link.springer.com/article/10.
1007/s10543-014-0505-1.

Baskerville:2007:GAS

[BGP07] Kim Baskerville, Peter Grassberger, and Maya Paczuski. Graph animals, subgraph sampling, and motif search in large networks. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 76(??):036107, September 11, 2007. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-2376 (electronic). URL http://link.aps.org/

doi/10.1103/PhysRevE.76.036107.

Becker:2011:BLC

[BGTM^{+}11] T. Becker, A. Greaves-Tunnell, Steven J. Miller, R. Ronan, and F. W. Strauch. Benford’s Law and continuous dependent random variables. ArXiv e-prints,

November 2011. CODEN ????  ISSN ????  URL

http://adsabs.harvard.edu/abs/2011arXiv1111.0568B;
 http://arxiv.org/abs/1111.0568.

Brakman:1999:RZT

[BGVV99] Steven Brakman, Harry Garretsen, Charles Van Marrewijk, and Marianne Van Den Berg. The return of Zipf: Towards a further understanding of the rank–size distribution. Journal of Regional Science, 39(1):183–213, 1999. ISSN 0022-4146 (print), 1467-9787 (electronic).

Banks:1974:AMN

[BH74] William P. Banks and David K. Hill. The apparent magnitude of number scaled by random production. Journal of Experimental Psychology: General, 102(2):353–376, February 1974. CODEN JPGEDD. ISSN 0096-3445 (print), 1939-2222 (electronic). URL http://psycnet.apa.org/

journals/xge/102/2/353/;   http://psycnet.apa.org/
psycinfo/1974-24305-001.

Bolton:2002:SFD

[BH02] Richard J. Bolton and David J. Hand. Statistical fraud detection: A review. Statistical Science, 17(3):235–255, August 2002. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1042727940.

Berger:2007:NMO

[BH07] Arno Berger and Theodore P. Hill. Newton’s method obeys Benford’s law. American Mathematical Monthly, 114(7):588– 601, August/September 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/27642274.

Berger:2009:BOB

[BH09] A. Berger and T. P. Hill. Benford online bibliography. Technical report, Department of Mathematical and Statistical Sciences, University of Alberta, Edmonton, AB, Canada, 2009.

URL http://www.benfordonline.net.

Berger:2010:FFFa

[BH10a] A. Berger and T. P. Hill. Fundamental flaws in Feller’s classical derivation of Benford’s Law. Preprint, Department of Mathematical and Statistical Sciences, University of Alberta, Edmonton, AB, Canada, 2010. 8 pp. URL http://www.

benfordonline.net/ARTICLES/BergerHill2010.pdf.

Berger:2010:FFFb

[BH10b] Arno Berger and Theodore P. Hill. Fundamental flaws in Feller’s classical derivation of Benford’s Law. ArXiv e-

prints, May 2010. CODEN ????  ISSN ????  URL

http://adsabs.harvard.edu/abs/2010arXiv1005.2598B;
 http://arxiv.org/abs/1005.2598.

Berger:2011:BTB

[BH11a] Arno Berger and Theodore P. Hill. A basic theory of Benford’s Law. Probability Surveys, 8(??):1–126, ???? 2011. CODEN

???? ISSN 1549-5787. URL http://projecteuclid.org/
euclid.ps/1311860830.

Berger:2011:BLS

[BH11b] Arno Berger and Theodore P. Hill. Benford’s law strikes back: no simple explanation in sight for mathematical gem. The Mathematical Intelligencer, 33(1):85–91, March 2011. CODEN MAINDC. ISSN 0343-6993 (print), 1866-7414

(electronic). URL http://link.springer.com/article/10.
1007/s00283-010-9182-3. See letter [Hil11].

Berger:2015:SIM

[BH15a] Arno Berger and T. P. Hill. A short introduction to the mathematical theory of Benford’s Law. In Miller [Mil15a], pages 23– 67. ISBN 0-691-14761-2 (hardcover), 1-4008-6659-6 (e-book).

LCCN QA273.6 .B46 2015. URL http://press.princeton.
edu/titles/10527.html.

Berger:2015:IBL

[BH15b] Arno Berger and Theodore P. Hill. An introduction to Benford’s Law. Princeton University Press, Princeton, NJ, USA, 2015. ISBN 0-691-16306-5, 1-4008-6658-8 (ebook). viii + 248 pp. LCCN QA273.6 .B474 2015.

URL http://site.ebrary.com/id/11040167;   http://
www.degruyter.com/viewbooktoc/product/465875.

Berger:2017:WLB

[BH17] Arno Berger and Theodore P. Hill. What is … Benford’s Law? Notices of the American Mathematical Society, 64(2):132–134, February 2017. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL http:/

/www.ams.org/publications/journals/notices/201702/
rnoti-p132.pdf.

Berger:2021:MBL

[BH21] Arno Berger and Theodore P. Hill. The mathematics of Benford’s law: a primer. Stat. Methods Appl., 30(3):779–795, 2021. ISSN 1618-2510.

Bhattacharya:2002:KBT

[Bha02] S. Bhattacharya. From Kautilya to Benford — trends in forensic and investigative accounting. In ????, editor, 9th World Congress of Accounting Historians, Deakin University, Melbourne, Australia July 30–August 3, 2002, page ?? Accounting Association of Australia and New Zealand, ????, 2002. ISBN ???? LCCN ????

Bharath:2012:BLT

[Bha12] H. M. Bharath. Benford’s Law: A theoretical explanation for base 2. arXiv.org, November 24, 2012. URL http://adsabs.

harvard.edu/abs/2012arXiv1211.7008B.

Bonhoeffer:1996:NSH

[BHB^{+}96] Sebastian Bonhoeffer, Andreas V. M. Herz, Maarten C. Boerlijst, Sean Nee, Martin A. Nowak, and Robert M. May. No signs of hidden language in noncoding DNA. Physical Review Letters, 76(11):1977, March 1996. CO-DEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (elec-

tronic), 1092-0145. URL http://link.aps.org/doi/10.
1103/PhysRevLett.76.1977.

Berger:2011:FSM

[BHKR11] Arno Berger, Theodore P. Hill, Bahar Kaynar, and Ad Ridder. Finite-state Markov chains obey Benford’s Law. SIAM Journal on Matrix Analysis and Applications, 32(3):665–684, ???? 2011. CODEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic). URL http://epubs.siam.org/

simax/resource/1/sjmael/v32/i3/p665_s1;   https://
dblp.org/db/journals/siammax/siammax32.html#BergerHKR11.

Berger:2008:SDI

[BHM08] Arno Berger, Theodore P. Hill, and Kent E. Morrison. Scaledistortion inequalities for mantissas of finite data sets. Journal of Theoretical Probability, 21(1):97–117, March 2008. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/s10959-007-0112-z.

Barrale:2011:SID

[BHS11] Tom Barrale, R. Hendel, and Michael Sluys. Sequences of the initial digits of Fibonacci numbers. In Proceedings of the 14th international conference on Fibonacci numbers and their applications, Morelia, Mexico, July 5–9, 2010, pages 25–42. Sociedad Matemática Mexicana, México, DF, México, 2011. ISBN 607-02-2543-0 (paperback).

Biau:2015:FDF

[Bia15] Damien Biau. The first-digit frequencies in data of turbulent flows. Physica A: Statistical Mechanics and its Applications, 440(??):147–154, December 15, 2015. CODEN PHYADX. ISSN 0378-4371 (print), 1873-2119 (electronic).

Bohachevsky:2006:DBL

[BJS06] Ihor O. Bohachevsky, Mark E. Johnson, and Myron L. Stein. A derivation of Benford’s law. American Journal of Mathematical and Management Sciences, 26(3–4):355–370, 2006. CODEN AMMSDX. ISSN 0196-6324.

Burke:1991:BLP

[BK91] John Burke and Eric Kincanon. Benford’s law and physical constants: The distribution of initial digits. American Journal of Physics, 59(10):952, October 1991. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

URL http://adsabs.harvard.edu/abs/1991AmJPh..59..
952B;   http://ajp.aapt.org/resource/1/ajpias/v59/
i10/p952_s1.

Bernik:2007:BLA

[BK07] V. I. Bernik and N. I. Kalosha. Benford’s law and the approximation of logarithms of natural numbers by rational numbers. Vestsī Nats. Akad. Navuk Belarusī Ser. Fīz.-Mat. Navuk, ?? (3):68–73, 127, 2007. CODEN ???? ISSN 0002-3574.

Bhattacharya:2008:FAB

[BK08] S. Bhattacharya and K. Kumar. Forensic accounting and Benford’s Law [in the spotlight]. IEEE Signal Processing Magazine, 25(2):152–150, March 2008. CODEN ISPRE6. ISSN 1053-5888 (print), 1558-0792 (electronic). URL http:/

/adsabs.harvard.edu/abs/2008ISPM...25..152B.

Bhattacharya:2005:CPA

[BKS05] Sukanto Bhattacharya, Kuldeep Kumar, and Florentin Smarandache. Conditional probability of actually detecting a financial fraud — a neutrosophic extension to Benford’s law. International Journal of Applied Mathematics, 17(1): 7–14, 2005. CODEN ???? ISSN 1311-1728. URL http://

arxiv.org/abs/math/0504520.

Bologna:1995:FAF

[BL95] Jack Bologna and Robert J. Lindquist. Fraud auditing and forensic accounting: new tools and techniques. Wiley, New York, NY, USA, second edition, 1995. ISBN 0-471-10646-1 (hardcover). xv + 249 pp. LCCN HV8079.W47 B65

1995. URL http://www.loc.gov/catdir/bios/wiley042/
95005510.html;   http://www.loc.gov/catdir/description/
wiley033/95005510.html;   http://www.loc.gov/catdir/
toc/onix04/95005510.html.

Blum:1970:SID

[Blu70] M. Blum. On the sums of independently distributed Pareto variates. SIAM Journal on Applied Mathematics, 19(1):191– 198, July 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Brown:2008:CSP

[BM08] Philip Brown and Jason Mitchell. Culture and stock price clustering: Evidence from The Peoples’ Republic of China. Pacific-Basin Finance Journal, 16(1 2):95–120, 2008. ISSN 0927-538X (print), 1879-0585 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0927538X0700025X. Behavioral Finance in Asia.

Birajdar:2013:DIF

[BM13] Gajanan K. Birajdar and Vijay H. Mankar. Digital image forgery detection using passive techniques: A survey. Digital Investigation, 10(3):226–245, 2013. ISSN 1742-2876 (print),

1873-202X (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1742287613000364.

Bravo-Marquez:2010:HLM

[BMLRV10] Felipe Bravo-Marquez, Gaston L’Huillier, Sebastián A. Ríos, and Juan D. Velásquez. Hypergeometric language model and Zipf-like scoring function for Web document similarity retrieval. Lecture Notes in Computer Science, 6393:303– 308, 2010. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/
content/pdf/10.1007/978-3-642-16321-0_32.

Bunt:1984:MPL

[BMM84] Richard B. Bunt, Jennifer M. Murphy, and Shikharesh Majumdar. A measure of program locality and its application. ACM SIGMETRICS Performance Evaluation Review, 12(3): 28–40, August 1984. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bonache:2010:DFL

[BMM10] A. B. Bonache, K. Moris, and J. Maurice. Détection de fraudes et loi de Benford: Quelques risques associés. (French) [Fraud detection and Benford’s law: Some risks]. Revue Française de Comptabilité, ??(431):24–27, April 2010. CODEN ???? ISSN 0484-8764. URL ????.

Buck:1993:IBF

[BMP93] B. Buck, A. C. Merchant, and S. M. Perez. An illustration of Benford’s first digit law using alpha decay half lives. European Journal of Physics, 14(2):59–63, ???? 1993. CODEN EJPHD4. ISSN 0143-0807 (print), 1361-6404 (electronic).

URL http://adsabs.harvard.edu/abs/1993EJPh...14...
59B.

Burgos:1996:ZSB

[BMT96] Javier D. Burgos and Pedro Moreno-Tovar. Zipf-scaling behavior in the immune system. Biosystems (A6E), 39(3): 227–232, November 15, 1996. CODEN BSYMBO. ISSN 0303-2647 (print), 1872-8324 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0303264796016188.

Balaz:2010:BLD

[BNS10a] V. Baláž, K. Nagasaka, and O. Strauch. Benford’s law and distribution functions of sequences in (0, 1). Mathematical Notes of the Academy of Sciences of the USSR = Matematicheskie Zametki, 88(3–4):449–463, October 2010. CODEN MTHNB2. ISSN 0001-4346 (print), 1573-8876

(electronic). URL http://link.springer.com/article/10.
1134/S0001434610090178.

Balazh:2010:BLD

[BNS10b] V. Balazh, K. Nagasaka, and O. Shtraukh. Benford’s law and distribution functions of sequences in (0, 1). Matematicheskie Zametki, 88(4):485–501, 2010. ISSN 0025-567X.

Bookstein:1994:TMD

[Boo94] A. Bookstein. Towards a multi-disciplinary Bradford Law. Scientometrics, 30(1):353–361, May 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02017233.

Boring:1920:LNL

[Bor20] Edwin G. Boring. The logic of the normal law of error in mental measurement. The American Journal of Psychology, 31(1):1–33, January 1920. CODEN AJPCAA. ISSN 0002-9556 (print), 1939-8298 (electronic). URL http://www.

jstor.org/stable/1413989.

Briscoe:2006:MLW

[BOT06] Bob Briscoe, Andrew Odlyzko, and Benjamin Tilly. Metcalfe’s Law is wrong — communications networks increase in value as they add members — but by how much? IEEE Spectrum, 43(7):34–39, July 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Boyle:1994:AFS

[Boy94] Jeff Boyle. An application of Fourier series to the most significant digit problem. American Mathematical Monthly, 101(9): 879–886, November 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Boroda:1984:ZML

[BP84] M. G. Boroda and A. A. Polikarpov. The Zipf–Mandelbrot law and units of different levels of text organization. Tartu Riikl. Ül. Toimetised, ??(689):35–60, 1984. CODEN ???? ISSN ????

Baskerville:2006:SEM

[BP06] Kim Baskerville and Maya Paczuski. Subgraph ensembles and motif discovery using an alternative heuristic for graph isomorphism. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 74(??):051903, November 3, 2006. CODEN PLEEE8. ISSN 1539-3755 (print),

1550-2376 (electronic). URL http://link.aps.org/doi/10.
1103/PhysRevE.74.051903.

Barabesi:2020:GBL

[BP20] Lucio Barabesi and Luca Pratelli. On the Generalized Benford law. Statistics & Probability Letters, 160(??):Article 108702, May 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715220300055.

Bose:2011:QMD

[BPS11] Indranil Bose, Selwyn Piramuthu, and Michael J. Shaw. Quantitative methods for detection of financial fraud. Decision Support Systems, 50(3):557–558, February 2011. CO-DEN DSSYDK. ISSN 0167-9236 (print), 1873-5797 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167923610001296. On quantitative methods

for detection of financial fraud.

Brunsch:2015:ISA

[BR15] Tobias Brunsch and Heiko Röglin. Improved smoothed analysis of multiobjective optimization. Journal of the ACM, 62(1): 4:1–4:??, February 2015. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bradford:1934:SIS

[Bra34] Samuel C. Bradford. Sources of information on specific subjects. Engineering: An Illustrated Weekly Journal (London), 137(??):85–86, January 26, 1934. CODEN ???? ISSN

???? URL http://en.wikipedia.org/wiki/Bradford%27s_

law. This paper (reprinted in [Bra85]) introduces an observation later known as Bradford’s Law: there are exponentiallydiminishing returns of extending a search for references in journals.

Brambilla:1976:DDA

[Bra76] Francisco Brambilla. La distribuzione delle aziende secondo le dimensioni. Verifica empirica della legge logonormale e di quella di Zipf e nuovi modelli interpretativi. (Italian) [The distribution of firms by size. Empirical test of the log-normal law and Zipf’s law and new models of interpretation]. Giornale degli Economisti e Annali de Economia (Nuova Serie), 35(3– 4):131–153, 235, March/April 1976. CODEN ???? ISSN 0017-

0097. URL http://www.jstor.org/stable/23242834.

Brady:1978:MBL

[Bra78] W. G. Brady. More on Benford’s law. Fibonacci Quarterly, 16(1):51–52, February 1978. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/16-1/brady.
pdf.

Bradford:1985:SIS

[Bra85] Samuel C. Bradford. Sources of information on specific subjects. Journal of Information Science, Principles and Practice, 10(4):173–180, October 1985. CODEN JISCDI. ISSN 0165-5515 (print), 1741-6485 (electronic). Reprint of [Bra34], the source of Bradford’s Law.

Brookes:1977:TBL

[Bro77] Bertram C. Brookes. Theory of the Bradford Law. Journal of Documentation, 33(3):180–209, ???? 1977. CODEN JDO-CAS. ISSN 0022-0418 (print), 1758-7379 (electronic).

Brookes:1979:LEB

[Bro79] B. C. Brookes. Letters to the Editor: The Bradford Law: a new calculus for the social sciences? Journal of the American Society for Information Science, 30(4):233–234, July 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Brookes:1980:FISa

[Bro80a] Bertram C. Brookes. The foundations of information science. Part I. Philosophical aspects. Journal of Information Science, Principles and Practice, 2(3–4):125–133, June 1980. CODEN JISCDI. ISSN 0165-5515 (print), 1741-6485 (electronic). URL

http://jis.sagepub.com/content/2/3-4/125.abstract.

Brookes:1980:FISb

[Bro80b] Bertram C. Brookes. The foundations of information science. Part II. Quantitative aspects: classes of things and the challenge of human individuality. Journal of Information Science, Principles and Practice, 2(5):209–221, October 1980. CODEN JISCDI. ISSN 0165-5515 (print), 1741-6485 (electronic). URL

http://jis.sagepub.com/content/2/5/209.abstract.

Brookes:1980:FISc

[Bro80c] Bertram C. Brookes. The foundations of information science. Part III. Quantitative aspects: Objective maps and subjective landscapes. Journal of Information Science, Principles and Practice, 2(6):269–275, December 1980. CODEN JIS-CDI. ISSN 0165-5515 (print), 1741-6485 (electronic). URL

http://jis.sagepub.com/content/2/6/269.abstract.

Brookes:1981:FIS

[Bro81] Bertram C. Brookes. The foundations of information science. Part IV. Information science: the changing paradigm. Journal of Information Science, Principles and Practice, 3 (1):3–12, February 1981. CODEN JISCDI. ISSN 0165-5515 (print), 1741-6485 (electronic). URL http://jis.sagepub.

com/content/3/1/3.

Brookes:1984:RTE

[Bro84] Bertram C. Brookes. Ranking techniques and the empirical log law. Information Processing and Management, 20(1–2):37– 46, ???? 1984. CODEN IPMADK. ISSN 0306-4573 (print),

1873-5371 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0306457384900384.

Browne:1998:FBL

[Bro98] Malcolm W. Browne. Following Benford’s Law, or looking out for No. 1. New York Times, ??(??):F4, August 4, 1998. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL http://search.proquest.com/
hnpnewyorktimes/docview/109868161/13C4AACC02923FE97A/

1. Science section.

Brown:2005:BLS

[Bro05] Richard J. C. Brown. Benford’s law and the screening of analytical data: the case of pollutant concentrations in ambient air. The Analyst, 130(9):1280–1285, ???? 2005. CODEN ANALAO. ISSN 0003-2654 (print), 1364-5528 (electronic).

URL http://www.ncbi.nlm.nih.gov/pubmed/16096674.

Brown:2007:UZL

[Bro07] R. J. C. Brown. The use of Zipf’s law in the screening of analytical data: a step beyond Benford. The Analyst, 132 (4):344, 2007. CODEN ANALAO. ISSN 0003-2654 (print),

1364-5528 (electronic). URL http://adsabs.harvard.edu/
abs/2007Ana...132..344B.

Bethe:1938:MSE

[BRS38] H. A. Bethe, M. E. Rose, and L. P. Smith. The multiple scattering of electrons. Proceedings of the American Philosophical Society, 78(4):573–585, March 1938. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http://www. jstor.org/stable/984803. This paper, which immediately follows Benford’s [Ben38] in this journal issue, is reported in [LG78, p. 197] to have been of considerable interest to scientists involved in secret nuclear physics work in World War II. That is how Benford’s paper “in a journal of rather limited circulation and not usually read by mathematicians” came to be noticed by physicists.

Busta:1992:TRN

[BS92] B. Busta and R. Sundheim. Tax return numbers tend to obey Benford’s law. Working Paper W93-106-94, Center for Business Research, St. Cloud State University, St. Cloud, MN, USA, 1992.

Blank:2000:PLC

[BS00] Aharon Blank and Sorin Solomon. Power laws in cities population, financial markets and Internet sites (scaling in systems with a variable number of components). Physica A: Statistical Mechanics and its Applications, 287(1–2):279–288, November 15, 2000. CODEN PHYADX. ISSN 0378-4371 (print),

1873-2119 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378437100004647.

Berger:2007:DMN

[BS07] A. Berger and S. Siegmund. On the distribution of mantissae in nonautonomous difference equations. Journal of Difference Equations and Applications, 13(8–9):829–845, ???? 2007. CODEN JDEAEA. ISSN 1023-6198 (print), 1563-5120 (elec-

tronic). URL http://www.math.ualberta.ca/~aberger/
benford_bibliography/berger_siegmund_07.pdf. Second

Memorial Issue in Honour of Bernd Aulbach.

Beber:2009:DD

[BS09] B. Beber and A. Scacco. The devil is in the digits. Washington Post, ??(??):??, June 20, 2009. ISSN 0190-8286.

Beber:2010:WNS

[BS10a] Bernd Beber and Alexandria Scacco. What the numbers say: a digit-based test for election fraud. Political Analysis, 20(2):211–234, Spring 2010. ISSN 1047-1987 (print), 1476-

4989 (electronic). URL http://pan.oxfordjournals.org/
content/20/2/211.abstract.

Block:2010:GEB

[BS10b] Henry W. Block and Thomas H. Savits. A general example for Benford data. The American Statistician, 64(4):335–339, November 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bhole:2014:BDN

[BSM14] Gaurav Bhole, Abhishek Shukla, and T. S. Mahesh. Benford distributions in NMR. arXiv.org, ??(??):6, June 27, 2014.

URL http://arxiv.org/abs/1406.7077v1.

Bhole:2015:BAU

[BSM15] Gaurav Bhole, Abhishek Shukla, and T. S. Mahesh. Benford analysis: A useful paradigm for spectroscopic analysis. Chemical Physics Letters, 639(??):36–40, 2015. CO-DEN CHPLBC. ISSN 0009-2614 (print), 1873-4448 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0009261415006582.

Balanzario:2010:SCB

[BSO10] Eugenio P. Balanzario and Jorge Sánchez-Ortiz. Sufficient conditions for Benford’s Law. Statistics & Probability Letters, 80(23–24):1713–1719, December 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210002087.

Bormashenko:2016:BLA

[BSWB16] Ed. Bormashenko, E. Shulzinger, G. Whyman, and Ye. Bormashenko. Benford’s Law, its applicability and breakdown in the IR spectra of polymers. Physica A: Statistical Mechanics and its Applications, 444(??):524–529, 2016. CO-DEN PHYADX. ISSN 0378-4371 (print), 1873-2119 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378437115009498.

Bresciani-Turroni:1937:PL

[BT37] C. Bresciani-Turroni. On Pareto’s law. Journal of the Royal Statistical Society, 100(3):421–432, ???? 1937. ISSN 0952-

8385. URL http://www.jstor.org/stable/2980528.

Blasius:2009:ZLP

[BT09] Bernd Blasius and Ralf Tönjes. Zipf’s Law in the popularity distribution of chess openings. Physical Review Letters, 103(21):218701, November 16, 2009. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

URL http://link.aps.org/doi/10.1103/PhysRevLett.
103.218701.

Berger:2018:SUR

[BT18] Arno Berger and Isaac Twelves. On the significands of uniform random variables. Journal of Applied Probability, 55(2):353– 367, June 2018. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL https://www.cambridge.org/
core/journals/journal-of-applied-probability/article/
on-the-significands-of-uniform-random-variables/0D08D0EEC38F52E431CF41290651603A.

Bak:1987:SOC

[BTW87] Per Bak, Chao Tang, and Kurt Wiesenfeld. Self-organized criticality: An explanation of the 1/f noise. Physical Review Letters, 59(4):381–384, July 1987. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145. URL

http://link.aps.org/doi/10.1103/PhysRevLett.59.381.

Bul:2011:BGG

[Bul11] K. Bul. Das Benford’sche Gesetz. (German) [Benford’s Law]. Working paper, Max-von-Laue-gymnasium, Koblenz, Germany, 2011.

Burrell:1991:BDG

[Bur91] Q. L. Burrell. The Bradford distribution and the Gini index. Scientometrics, 21(2):181–194, June 1991. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02017568.

Burrell:2004:LEF

[Bur04] Quentin L. Burrell. Letters to the Editor: Fitting Lotka’s Law: Some cautionary observations on a recent paper by Newby et al. (2003). Journal of the American Society for Information Science and Technology, 55(13):1209–1210, November 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic). See [NGJ03].

Burns:2009:SSR

[Bur09] B. D. Burns. Sensitivity to statistical regularities: People (largely) follow Benford’s Law. In N. Taatgen and H. van Rijn, editors, Proceedings of the Thirty-First Annual Conference of the Cognitive Science Society, Cognitive Science Society, Austin, TX, pages 2872–2877. ????, ????, 2009. ISBN ???? LCCN ????

Burns:2000:SSR

[Bur10a] B. D. Burns. Sensitivity to statistical regularities: People (largely) follow Benford’s Law. In Niels A. Taatgen and Hedderik van Rijn, editors, Proceedings of the 31st Annual Conference of the Cognitive Science Society 2009 (CogSci 2009): Amsterdam, Netherlands, 29 July–1 August 2009, pages 2872– 2877. Cognitive Science Society, Austin, TX, USA, 2010. ISBN 1-61567-407-1. LCCN ????

Burns:2010:PL

[Bur10b] John Burns. Probably likely. New Scientist, 208(2786):32– , 2010. CODEN NWSCAL. ISSN 0262-4079 (print), 1364-

8500 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0262407910628102.

Beirlant:1996:TIE

[BVT96] Jan Beirlant, Petra Vynckier, and Jozef L. Teugels. Tail index estimation, Pareto quantile plots, and regression diagnostics. Journal of the American Statistical Association, 91 (436):1659–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Busta:1998:UBL

[BW98] Bruce Busta and Randy Weinberg. Using Benford’s Law and neural networks as a review procedure. Managerial Auditing Journal, 13(6):356–366, ???? 1998. CODEN ???? ISSN 0268-6902 (print), 1758-7735 (electronic).

Berger:2018:BFA

[BX18] Arno Berger and Chuang Xu. Best finite approximations of Benford’s Law. Journal of Theoretical Probability, ??(??):??, April 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Berger:2019:BFA

[BX19] Arno Berger and Chuang Xu. Best finite approximations of Benford’s Law. Journal of Theoretical Probability, 32(3):1525– 1553, September 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Bhattacharya:2011:ABA

[BXK11] Sukanto Bhattacharya, Dongming Xu, and Kuldeep Kumar. An ANN-based auditor decision support system using Benford’s law. Decision Support Systems, 50(3):576–584, February 2011. CODEN DSSYDK. ISSN 0167-9236 (print), 1873-

5797 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167923610001351. On quantita-

tive methods for detection of financial fraud.

Clippe:2012:BLT

[CA12] Paulette Clippe and Marcel Ausloos. Benford’s Law and Theil transform of financial data. Physica A: Statistical Mechanics and its Applications, 391(24):6556–6567, December 15, 2012. CODEN PHYADX. ISSN 0378-4371 (print), 1873-

2119 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378437112007546.

Cali:2013:FA

[Cal13] J. Cali. Forensic accounting. In Jay A. Siegel, Pekka J. Saukko, and Max M. Houck, editors, Encyclopedia of Forensic Sciences, pages 423–431. Academic Press, Waltham, 2013.

ISBN 0-12-382166-5. URL http://www.sciencedirect.com/
science/article/pii/B978012382165200218X.

Campanelli:2024:EDS

[Cam24] Leonardo Campanelli. On the Euclidean distance statistic of Benford’s law. Communications in Statistics: Theory and Methods, 53(2):451–474, 2024. CODEN CST-MDC. ISSN 0361-0926,1532-415X. URL https://doi.org/

10.1080/03610926.2022.2082480.

Canessa:2003:TAFa

[Can03a] Enrique Canessa. Theory of analogous force on number sets.

ArXiv e-prints, 2003. URL http://arxiv.org/abs/cond-
mat/0307703.

Canessa:2003:TAFb

[Can03b] Enrique Canessa. Theory of analogous force on number sets. Physica A: Statistical Mechanics and its Applications, 328(1–2):44–52, October 2003. CODEN PHYADX. ISSN 0378-4371 (print), 1873-2119 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378437103005260.

Carslaw:1988:AIN

[Car88] C. A. Carslaw. Anomalies in income numbers: Evidence of goal-oriented behavior. The Accounting Review, 63(2): 321–327, April 1988. CODEN ACRVAS. ISSN 0001-4826 (print), 1558-7967 (electronic). URL http://www.jstor.

org/stable/248109.

Cartlidge:2010:NNO

[Car10] Edwin Cartlidge. In nature, number one dominates. Institute of Physics, Environmental Research Letters, ??(??):??, Oc-

tober 25, 2010. URL http://environmentalresearchweb.
org/cws/article/news/44124.

Carrera:2015:TER

[Car15] César Carrera. Tracking exchange rate management in Latin America. Review of Financial Economics, 25(??): 35–41, 2015. ISSN 1058-3300 (print), 1873-5924 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S105833001500018X. Special Issue on Chang-

ing Dynamics in Financial Economics.

Cerioli:2019:NBL

[CBC^{+}19] Andrea Cerioli, Lucio Barabesi, Andrea Cerasa, Mario Menegatti, and Domenico Perrotta. Newcomb–Benford law and the detection of frauds in international trade. Proceedings of the National Academy of Sciences of the United States of America, 116(1):106–115, 2019. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic).

Cristelli:2012:TMT

[CBP12] Matthieu Cristelli, Michael Batty, and Luciano Pietronero. There is more than a power law in Zipf. Scientific Reports, 2(??):1–7, November 2012. CODEN SRCEC3. ISSN 2045-

2322. URL http://www.nature.com/srep/2012/121108/
srep00812/full/srep00812.html.

Campanario:2011:BLC

[CC11] Juan Campanario and María Coslado. Benford’s law and citations, articles and impact factors of scientific journals. Scientometrics, 88(2):421–432, August 2011. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0387-9.

Chuang:2009:FPS

[CCC09] Kun-Ta Chuang, Hung-Leng Chen, and Ming-Syan Chen. Feature-preserved sampling over streaming data. ACM Transactions on Knowledge Discovery from Data (TKDD), 2(4): 15:1–15:??, January 2009. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Chen:1995:RDB

[CCT95] Ye-Sho Chen, P. Pete Chong, and Morgan Y. Tong. Research: Dynamic behavior of Bradford’s Law. Journal of the American Society for Information Science, 46(5):370–383, June 1995.

CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Corazza:2010:CFM

[CEZ10] Marco Corazza, Andrea Ellero, and Alberto Zorzi. Checking financial markets via Benford’s law: the S&P 500 case. In Mathematical and statistical methods for actuarial sciences and finance, pages 93–102. Springer Italia, Milan, Italy,

2010. URL http://link.springer.com/chapter/10.1007/
978-88-470-1481-7_10.

Corazza:2017:LET

[CEZ17] Marco Corazza, Andrea Ellero, and Alberto Zorzi. L’importanza di essere “UNO” (ovvero la legge di Benford). (Italian) [the importance of being “one” (i.e., the Benford law)]. Lettera Matematica Pristem, 103(1):31–38, December 2017. ISSN 1593-5884 (print), 1970-6820 (electronic).

Corazza:2018:IBO

[CEZ18] Marco Corazza, Andrea Ellero, and Alberto Zorzi. The importance of being “one” (or Benford’s law). Lettera Matematica, 6 (1):33–39, March 01, 2018. ISSN 2281-5937 (print), 2281-6917 (electronic). Translated from the Italian.

Cai:2020:SAB

[CFH^{+}20] Zhaodong Cai, Matthew Faust, A. J. Hildebrand, Junxian Li, and Yuan Zhang. The surprising accuracy of Benford’s law in mathematics. American Mathematical Monthly, 127(3):217– 237, 2020. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Clark:1977:ESL

[CG77] Douglas W. Clark and C. Cordell Green. An empirical study of list structure in Lisp. Communications of the ACM, 20(2): 78–87, February 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Christian:1993:NES

[CG93] C. W. Christian and S. Gupta. New evidence on “secondary evasion”. Journal of the American Taxation Association, 15 (??):72–93, ???? 1993. ISSN 0198-9073 (print), 1558-8017 (electronic).

Cho:2007:BBL

[CG07] Wendy K. Tam Cho and Brian J. Gaines. Breaking the (Benford) law: statistical fraud detection in campaign finance. The American Statistician, 61(3):218– 223, August 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://pubs.amstat.

org/doi/pdfplus/10.1198/000313007X223496;   http:/
/www.tandfonline.com/doi/abs/10.1198/000313007X223496.

Cifarelli:2010:GSP

[CGJ10] D. Michele Cifarelli, R. P. Gupta, and K. Jayakumar. On generalized semi-Pareto and semi-Burr distributions and random coefficient minification processes. Statistical Papers, 51 (1):193–208, January 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http://link.

springer.com/article/10.1007/s00362-008-0132-5.

Chaves-Gonzalez:2015:DEP

[CGPT15] José M. Chaves-González and Miguel A. Pérez-Toledano. Differential evolution with Pareto tournament for the multiobjective next release problem. Applied Mathematics and Computation, 252(??):1–13, February 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0096300314016257.

Cigler:1961:NET

[CH61] J. Cigler and G. Helmberg. Neuere Entwicklungen der Theorie der Gleichverteilung. (German) [Recent developments in the theory of uniform distribution]. Jber. Deutsch. Math.–Verein, 64(??):1–50, ???? 1961. CODEN ???? ISSN ????

Castillo:1997:FGP

[CH97] Enrique Castillo and Ali S. Hadi. Fitting the generalized Pareto distribution to data. Journal of the American Statistical Association, 92(440):1609–1620, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.amstat.org/publications/
jasa/abstracts_97/december/CASTILLO.HTM; http://www.
jstor.org/stable/2965432.

Chang:2011:CGR

[Cha11] Harry M. Chang. Constructing n-gram rules for natural language models through exploring the limitation of the Zipf–Mandelbrot law. Computing: Archiv für Informatik und Numerik, 91(3):241–264, March 2011. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0010-485X&volume=91&issue=3&spage=241.

Chen:1978:ZL

[Che78] Wen-Chen Chen. On Zipf’s law. Thesis (Ph.D.), University of Michigan, Ann Arbor, MI, USA, 1978. 114 pp. URL http:/

/gateway.proquest.com/openurl?url_ver=Z39.88-2004&
rft_val_fmt=info:ofi/fmt:kev:mtx:dissertation&res_
dat=xri:pqdiss&rft_dat=xri:pqdiss:7822871.

Chen:1980:WFZ

[Che80] Wen Chen Chen. On the weak form of Zipf’s law. Journal of Applied Probability, 17(3):611–622, 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3212955.

Chen:1989:ZLT

[Che89] Ye-Sho Chen. Zipf’s laws in text modeling. International Journal of General Systems. Methodology, Applications, Education, 15(3):233–252, 1989. CODEN IJGSAX. ISSN 0308-1079, 1026-7492, 1563-5104.

Chen:1991:ZLN

[Che91] Ye-Sho Chen. Zipf’s law in natural languages, programming languages, and command languages: the Simon–Yule approach. International Journal of Systems Science, 22(11): 2299–2312, 1991. CODEN IJSYA9. ISSN 0020-7721 (print), 1464-5319 (electronic).

Chen:1995:LED

[Che95] Ye-Sho Chen. Letters to the Editor: On the dynamic behavior of Bradford’s Law — response. Journal of the American Society for Information Science, 46(10):799–800, December 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Chen:2012:MRB

[Che12a] Yanguang Chen. The mathematical relationship between Zipf’s law and the hierarchical scaling law. Physica A: Statistical Mechanics and its Applications, 391(11):3285–3299, June 1, 2012. CODEN PHYADX. ISSN 0378-4371 (print),

1873-2119 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378437111009678.

Chen:2012:ZLF

[Che12b] Yanguang Chen. Zipf’s law, 1/f noise, and fractal hierarchy. Chaos, Solitons & Fractals, 45(1):63–73, January 15, 2012. CODEN CSFOEH. ISSN 0960-0779 (print), 1873-

2887 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0960077911001901.

Chen:2012:ZLH

[Che12c] Yanguang Chen. Zipf’s law, hierarchical structure, and cardsshuffling model for urban development. Discrete Dynamics in Nature and Society, page 480196, 2012. CODEN DDNSFA. ISSN 1026-0226 (print), 1607-887X (electronic).

Cai:2019:LBL

[CHL19] Zhaodong Cai, A. J. Hildebrand, and Junxian Li. A local Benford Law for a class of arithmetic sequences. International Journal of Number Theory, 15(3):613–638, April 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https://www.

worldscientific.com/doi/10.1142/S1793042119500325.

Campanino:1996:IIM

[CI96] Massimo Campanino and Stefano Isola. Infinite invariant measures for non-uniformly expanding transformations of [0, 1]: Weak law of large numbers with anomalous scaling. Forum Mathematicum, 8(1):71–92, 1996. CODEN FOMAEF. ISSN 0933-7741 (print), 1435-5337 (electronic).

Cigler:1964:MSU

[Cig64] J. Cigler. Methods of summability and uniform distribution mod1. Compositio Mathematica, 16(??):44–51, 1964. CO-DEN CMPMAF. ISSN 0010-437X (print), 1570-5846 (electronic).

Ciofalo:2005:EBF

[Cio05] Michele Ciofalo. Entropy, Benford’s first digit law, and the distribution of everything. Report, Dipartimento di Ingegneria Nucleare, Università degli Studi di Palermo, Palermo, Italy,

2005. 34 pp. URL http://www.din.unipa.it/Struttura/
Personale/Profili/ciofalo_allegati/Paper_Benford.pdf.

Cohen:1984:PNF

[CK84] D. I. A. Cohen and T. M. Katz. Prime numbers and the first digit phenomenon. Journal of Number Theory, 18(3):261–268, June 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Campi:2005:ZLM

[CK05] X. Campi and H. Krivine. Zipf’s law in multifragmentation. Physical Review C (Nuclear Physics), 72(??):057602, November 30, 2005. CODEN PRVCAN. ISSN 0556-2813 (print),

1089-490X, 1538-4497. URL http://link.aps.org/doi/10.
1103/PhysRevC.72.057602.

Chou:2009:BLN

[CKT^{+}09] Mabel C. Chou, Qingxia Kong, Chung-Piaw Teo, Zuozheng Wang, and Huan Zheng. Benford’s Law and number selection in fixed-odds numbers game. Journal of Gambling Studies, 25(4):503–521, December 2009. CODEN JGSTEM. ISSN 1050-5350 (print), 1573-3602 (electronic). URL http://link.

springer.com/article/10.1007/s10899-009-9145-9.

Chen:1986:RBL

[CL86a] Ye-Sho Chen and Ferdinand F. Leimkuhler. A relationship between Lotka’s Law, Bradford’s Law, and Zipf’s Law. Journal of the American Society for Information Science, 37(5): 307–314, 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Chen:1986:RRB

[CL86b] Ye-Sho Chen and Ferdinand F. Leimkuhler. Research: A relationship between Lotka’s Law, Bradford’s Law, and Zipf’s Law. Journal of the American Society for Information Science, 37(5):307–314, September 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Chen:1987:BLI

[CL87a] Ye-Sho Chen and F. F. Leimkuhler. Bradford’s Law: An index approach. Scientometrics, 11(3–4):183–198, March 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02016591.

Chen:1987:AZL

[CL87b] Ye-Sho Chen and Ferdinand F. Leimkuhler. Analysis of Zipf’s law: An index approach. Information Processing and Management, 23(3):171–182, April 1, 1987. CO-DEN IPMADK. ISSN 0306-4573 (print), 1873-5371 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0306457387900021.

Challet:2004:BPD

[CL04a] Damien Challet and Andrea Lombardoni. Bug propagation and debugging in asymmetric software structures. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 70(4):046109, October 2004. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-2376

(electronic). URL http://link.aps.org/doi/10.1103/
PhysRevE.70.046109;   http://pre.aps.org/abstract/
PRE/v70/i4/e046109.

Cong:2004:WTP

[CL04b] Lin Cong and Zhipeng Li. On Wilson’s theorem and Polignac conjecture. ArXiv e-prints, August 2, 2004. URL http://

arxiv.org/abs/math.NT/0408018.

Cohen:2005:PCL

[CL05] Edith Cohen and Carsten Lund. Packet classification in large ISPs: design and evaluation of decision tree classifiers. ACM SIGMETRICS Performance Evaluation Review, 33(1):73–84, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Cerqueti:2023:STB

[CL23] Roy Cerqueti and Claudio Lupi. Severe testing of Benford’s law. TEST, 32(2):677–694, June 2023. CO-DEN ???? ISSN 1133-0686 (print), 1863-8260 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11749-023-00848-z.

Cuff:2015:WDB

[CLM15] Victoria Cuff, Allison Lewis, and Steven J. Miller. The Weibull distribution and Benford’s Law. Involve, 8(5):859–874, 2015. ISSN 1944-4176 (print), 1944-4184 (electronic).

Chandee:2023:BLM

[CLPR23] Vorrapan Chandee, Xiannan Li, Paul Pollack, and Akash Singha Roy. On Benford’s law for multiplicative functions. Proceedings of the American Mathematical Society, 151(11):4607– 4619, 2023. CODEN PAMYAR. ISSN 0002-9939,1088-6826.

URL https://doi.org/10.1090/proc/16480.

Costas:2008:NCC

[CLRTFM08] Eduardo Costas, Victoria López-Rodas, F. Javier Toro, and Antonio Flores-Moya. The number of cells in colonies of the cyanobacterium Microcystis aeruginosa satisfies Benford’s law. Aquatic Botany, 89(3):341–343, October 2008. CO-DEN AQBODS. ISSN 0304-3770 (print), 1879-1522 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304377008000533.

Cornea:2020:ISC

[CLT20] Marius Cornea, Weiqiang Liu, and Arnaud Tisserand, editors. 2020 27th IEEE Symposium on Computer Arithmetic: ARITH 2020: proceedings: Portland, Oregon, USA, 7–10 June 2020. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2020. ISBN 1-72817-120-2, 1-72817-121-0. LCCN ???? URL https://

ieeexplore.ieee.org/servlet/opac?punumber=9146973.

Chida:2011:LSI

[CM11] Satoshi Chida and Naoto Miyoshi. Limiting size index distributions for ball-bin models with Zipf-type frequencies. Annals of the Institute of Statistical Mathematics (Tokyo), 63(4): 745–768, 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Cerqueti:2021:DVS

[CM21] Roy Cerqueti and Mario Maggi. Data validity and statistical conformity with Benford’s Law. Chaos, Solitons & Fractals, 144:Paper No. 110740, 7, 2021. CODEN CSFOEH. ISSN 0960-0779 (print), 1873-2887 (electronic).

Corominas-Murtra:2011:EZL

[CMFS11] Bernat Corominas-Murtra, Jordi Fortuny, and Ricard V. Solé. Emergence of Zipf’s law in the evolution of communication. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 83(1):036115, March 28, 2011. CODEN PLEEE8. ISSN 1539-3755 (print),

1550-2376 (electronic). URL http://link.aps.org/doi/10.
1103/PhysRevE.83.036115.

Czirok:1995:CBS

[CMHS95] Andras Czirók, Rosario N. Mantegna, Shlomo Havlin, and H. Eugene Stanley. Correlations in binary sequences and a generalized Zipf analysis. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 52(??):446–??, July 1, 1995. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-2376 (electronic). URL http://link.aps.

org/doi/10.1103/PhysRevE.52.446.

Corominas-Murtra:2010:UZL

[CMS10] Bernat Corominas-Murtra and Ricard V. Solé. Universality of Zipf’s law. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 82(1): 011102, July 1, 2010. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-2376 (electronic). URL http://link.aps.org/

doi/10.1103/PhysRevE.82.011102.

Cacciari:2015:DMC

[CNS15] I. Cacciari, P. Nieri, and S. Siano. 3D digital microscopy for characterizing punchworks on medieval panel paintings. Journal on Computing and Cultural Heritage (JOCCH), 7(4): 19:1–19:??, February 2015. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Coderre:1999:FDU

[Cod99] David Coderre. Fraud detection using digital analysis. ED-PACS, 27(3):1–8, ???? 1999. CODEN EDPCDF. ISSN 0736-6981 (print), 1936-1009 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1201/1079/43249.27.
3.19990901/30268.1.

Cohen:1976:EFD

[Coh76] Daniel I. A. Cohen. An explanation of the first digit phenomenon. Journal of Combinatorial Theory (Series A), 20 (3):367–370, May 1976. CODEN JCBTA7. ISSN 0097-3165 (print), 1096-0899 (electronic).

Coile:1977:LFD

[Coi77] Russell C. Coile. Lotka’s frequency distribution of scientific productivity. Journal of the American Society for Information Science, 28(6):366–370, November 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Condon:1928:SV

[Con28] E. U. Condon. Statistics of vocabulary. Science, 67(1733): 300–??, March 16, 1928. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Conway:2010:CFM

[Con10] D. Conway. Benford’s Law tests for Wikileaks data. Zero Intelligence Agents website., August 1, 2010. URL http://

www.drewconway.com/zia/?p=2234.

Courtland:2010:CML

[Cou10] Rachel Courtland. Curious mathematical law is rife in nature. New Scientist, 208(2782):10, October 16, 2010. CO-DEN NWSCAL. ISSN 0262-4079 (print), 1364-8500 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0262407910624979.

Cerqueti:2023:BLE

[CP23] Roy Cerqueti and Davide Provenzano. Benford’s law for economic data reliability: the case of tourism flows in Sicily. Chaos, Solitons & Fractals, 173:1–10, 2023. CODEN CS-FOEH. ISSN 0960-0779,1873-2887. URL https://doi.org/

10.1016/j.chaos.2023.113635.

Chen:2016:LBS

[CPS16] Evan Chen, Peter S. Park, and Ashvin A. Swaminathan. On logarithmically Benford sequences. Proceedings of the American Mathematical Society, 144(11):4599–4608, 2016. CODEN PAMYAR. ISSN 0002-9939 (print), 1088-6826 (electronic).

Crato:2010:MB

[Cra10] Nuno Crato. Mr. Benford. In Figuring It Out: Entertaining Encounters with Everyday Math, pages 173–178. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. CODEN LNMAA2. ISBN 3-642-04832-3 (hard-

cover). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-04833-3_41.pdf.

Chakravorty:2009:TBL

[CS09a] S. Chakravorty and V. Sridharan. Testing Benford’s Law for improving supply chain decision-making: a field experiment. Drug Discovery Today, 122(2):606–618, December 2009. CODEN DDTOFS. ISSN 1359-6446 (print), 1878-

5832 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0925527309002163.

Chen:2009:DDM

[CS09b] Wen Chen and Yun Q. Shi. Detection of double MPEG compression based on first digit statistics. In Digital Watermarking, pages 16–30. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. URL http://

link.springer.com/chapter/10.1007/978-3-642-04438-
0_2.

Cantu:2011:FDA

[CS11] F. Cantu and S. M. Saiegh. Fraudulent democracy? An analysis of Argentina’s infamous decade using supervised machine learning. Political Analysis, 19(4):409–433, Autumn 2011. ISSN 1047-1987 (print), 1476-4989 (electronic). URL http:/

/pan.oxfordjournals.org/content/19/4/409.abstract.

Cangiotti:2022:BLN

[CS22] Nicolò Cangiotti and Mattia Sensi. Benford’s law: a numbertheoretical perspective. Palestine Journal of Mathematics, 11 (3):379–385, 2022. ISSN 2219-5688.

Cournane:2014:NAB

[CSC14] S. Cournane, N. Sheehy, and J. Cooke. The novel application of Benford’s second order analysis for monitoring radiation output in interventional radiology. Physica Medica, 30 (4):413–418, 2014. ISSN 1120-1797 (print), 1724-191X (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1120179713004328.

Clauset:2007:PLDa

[CSN07] Aaron Clauset, Cosma Rohilla Shalizi, and M. E. J. Newman. Power-law distributions in empirical data. ArXiv Physics e-

prints, June 7, 2007. URL http://arxiv.org/abs/0706.

1062v1. arXiv:0706.1062.

Clauset:2007:PLDb

[CSN09a] Aaron Clauset, Cosma Rohilla Shalizi, and M. E. J. Newman. Power-law distributions in empirical data. ArXiv Physics e-prints, February 2, 2009. URL http://arxiv.org/abs/ 0706.1062v1. arXiv:0706.1062.

Clauset:2009:PLDc

[CSN09b] Aaron Clauset, Cosma Rohilla Shalizi, and M. E. J. Newman. Power-law distributions in empirical data. SIAM Review, 51 (4):661–703, ???? 2009. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Czirok:1996:POP

[CSV96] András Czirók, H. Eugene Stanley, and Tamás Vicsek. Possible origin of power-law behavior in n-tuple Zipf analysis. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 53(??):6371–??, June 1, 1996. CODEN PLEEE8. ISSN 1539-3755 (print),

1550-2376 (electronic). URL http://link.aps.org/doi/10.
1103/PhysRevE.53.6371.

Cleary:2005:ADA

[CT05] Richard Cleary and Jay C. Thibodeau. Applying digital analysis using Benford’s Law to detect fraud: The dangers of type I errors. Auditing: A Journal of Practice & Theory, 24(1):77– 81, May 2005. CODEN ???? ISSN 0278-0380 (print), 1558-

7991 (electronic). URL http://aaapubs.org/loi/ajpt;
http://link.aip.org/link/AJPTXX/v24/i1/p77/s1.

Chacko:2007:EPB

[CT07] Manoj Chacko and P. Yageen Thomas. Estimation of a parameter of bivariate Pareto distribution by ranked set sampling. Journal of Applied Statistics, 34(6):703–714, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chen:2022:ASU

[CT22] Tianyi Chen and Charalampos E. Tsourakakis. AntiBenford subgraphs: Unsupervised anomaly detection in financial networks. CoRR, ??(??):??, ???? 2022. CODEN ????

ISSN ???? URL https://dblp.org/db/journals/corr/
corr2205.html#abs-2205-13426.

Tseng:2017:MBL

[cTnHwH17] Hsiang chi Tseng, Wei neng Huang, and Ding wei Huang. Modified Benford’s Law for two-exponent distributions. Scientometrics, 110(3):1403–1413, March 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2217-6.

Chen:1997:NHS

[CW97] Weishing Chen and Tai-Hsi Wu. A non-homogeneous software reliability model based on Zipf’s law. International Journal of Quality & Reliability Management, 14(4):409–431, ???? 1997. ISSN 0265-671X (print), 1758-6682 (electronic).

Clauset:2007:FST

[CYG07] Aaron Clauset, Maxwell Young, and Kristian Skrede Gleditsch. On the frequency of severe terrorist events. Journal of Conflict Resolution, 51(1):58–87, February 2007. CODEN JCFRAL. ISSN 0022-0027 (print), 1552-8766

(electronic). URL http://jcr.sagepub.com;   http://
jcr.sagepub.com/content/51/1/58.abstract;   http:/
/www.jstor.org/journals/00220027.html.

Chongde:1998:EMB

[CZ98] Wang Chongde and Wang Zhe. Evaluation of the models for Bradford’s Law. Scientometrics, 42(1):89–95, May 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02465014.

Chen:2004:MFM

[CZ04] Yanguang Chen and Yixing Zhou. Multi-fractal measures of city-size distributions based on the three-parameter Zipf model. Chaos, Solitons & Fractals, 22(4):793–805, November 1, 2004. CODEN CSFOEH. ISSN 0960-0779 (print),

1873-2887 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0960077904001031.

DaSilva:2020:BBS

[Da 20] Stéphane Blondeau Da Silva. Benford or not Benford: A systematic but not always well-founded use of an elegant law in experimental fields. Communications in Mathematics and Statistics, 8(2):167–201, June 2020. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://link.

springer.com/article/10.1007/s40304-018-00172-1.

DaSilva:2022:AOB

[Da 22] Stéphane Blondeau Da Silva. An alternative to the oversimplifying Benford’s law in experimental fields. Sankhyā (Indian Journal of Statistics), Series B. Methodological, 84(2):778– 808, 2022. CODEN SANBBV. ISSN 0976-8386,0976-8394.

URL https://doi.org/10.1007/s13571-022-00287-0.

DAlessandro:2020:BLM

[D’A20] Angelo D’Alessandro. Benford’s law and metabolomics: A tale of numbers and blood. Transfusion and Apheresis Science, 59 (6):103019, December 2020. ISSN 1473-0502 (print), 1878-1683 (electronic).

Davis:1976:SRI

[Dav76] Basil Davis. Some remarks on initial digits. Fibonacci Quarterly, 14(1):13–14, February 1976. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Scanned/14-1/
davis.pdf.

Dunstan:2014:GSD

[DB14] D. J. Dunstan and A. J. Bushby. Grain size dependence of the strength of metals: The Hall–Petch effect does not scale as the inverse square root of grain size. International Journal of Plasticity, 53(0):56–65, 2014. CO-DEN IJPLER. ISSN 0749-6419 (print), 1879-2154 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074964191300140X.

Daniels:2017:BLQ

[DCH^{+}17] Jeremy Daniels, Samantha-Jo Caetano, Dirk Huyer, Andrew Stephen, John Fernandes, Alice Lytwyn, and Fred M. Hoppe. Benford’s Law for quality assurance of manner of death counts in small and large databases. Journal of Forensic Sciences, 62 (5):1326–1331, June 2017. ISSN 0022-1198 (print), 1556-4029 (electronic).

delCastillo:2015:LIG

[dCS15] Joan del Castillo and Isabel Serra. Likelihood inference for generalized Pareto distribution. Computational Statistics & Data Analysis, 83(??):116–128, March 2015. CO-DEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167947314003065.

Dallacasa:2006:TBS

[DD06] F. Dallacasa and Valerio Dallacasa. Temporal behaviour of stock data from Benford’s law. In ????, editor, 5th International Conference: APLIMAT 2006, Slovak University of Technology, Bratislava, pages 425–432. ????, ????, 2006. ISBN

???? LCCN ???? US$. URL http://archiv.aplimat.com/
2006/Aplimat/Proceedings/Communications/Financial_
Actuary_math/Dallasca-Dallasca.pdf.

Docampo:2009:BLA

[DdA^{+}09] Silvia Docampo, María del Mar Trigo, María Aira, Baltasar Cabezudo, and Antonio Flores-Moya. Benford’s law applied to aerobiological data and its potential as a quality control tool. Aerobiologia, 25(4):275–283, December 2009. CO-DEN AROBFT. ISSN 0393-5965 (print), 1573-3025 (elec-

tronic). URL http://www.ingentaconnect.com/content/
klu/aero/2009/00000025/00000004/00009132.

DeCeuster:1998:HPB

[DDS98] Marc J. K. De Ceuster, Geert Dhaene, and Tom Schatteman. On the hypothesis of psychological barriers in stock markets and Benford’s Law. Journal of Empirical Finance, 5(3):263–279, September 1998. CODEN JE-FIEC. ISSN 0927-5398 (print), 1879-1727 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0927539897000248.

Deakin:1993:ADB

[Dea93] Michael A. B. Deakin. Another derivation of Benford’s law. Australian Mathematical Society Gazette, 20(5):162–163, 1993. CODEN ???? ISSN 0311-0729 (print), 1326-2297 (electronic).

Denisov:1997:FBS

[Den97] S. Denisov. Fractal binary sequences: Tsallis thermodynamics and the Zipf law. Physics Letters A, 235(5):447–451, November 17, 1997. CODEN PYLAAG. ISSN 0375-9601 (print),

1873-2429 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0375960197006889.

Depken:2008:BZB

[Dep08] Craig A. Depken II. Benford, Zipf and the blogosphere. Applied Economics Letters, 15(9):689–692, ???? 2008. ISSN 1350-4851 (print), 1466-4291 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/13504850600735270.

Devooght:1957:LZM

[Dev57] J. Devooght. Sur la loi de Zipf–Mandelbrot. (French) [on the Zipf–Mandelbrot law]. Acad. Roy. Belg. Bull. Cl. Sci. (5), 43: 244–251, 1957. CODEN ???? ISSN 0001-4141.

Diaconis:1979:RP

[DF79] Persi Diaconis and David Freedman. On rounding percentages. Journal of the American Statistical Association, 74(366): 359–364, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor. org/stable/2284288. The authors extend prior work on correctness of sums of rounded percentages [MYZ67], and criticize biased rounding practices in [Ben38].

Drott:1978:EEB

[DG78] M. Carl Drott and Belver C. Griffith. An empirical examination of Bradford’s Law and the scattering of scientific literature. Journal of the American Society for Information Science, 29(5):238–246, September 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Debreceny:2010:DMJ

[DG10] Roger S. Debreceny and Glen L. Gray. Data mining journal entries for fraud detection: An exploratory study. International Journal of Accounting Information Systems, 11(3):157–181, September 2010. CODEN IJAIA7. ISSN 1467-0895 (print), 1873-4723 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1467089510000540;   http://www.sciencedirect.
com/science/journal/14670895. See discussion [Gra10,

Kri10].

Dorrell:2012:FFB

[DG12] Darrell D. Dorrell and Gregory A. Gadawski. Financial forensics body of knowledge. Wiley finance. Wiley, New York, NY, USA, 2012. ISBN 0-470-88085-6 (hardcover), 1-118-21897-3 (e-book), 1-118-21896-5 (e-book), 1-118-21898-1 (e-book), 1-119-20073-3 (e-book). xvi + 541 pp. LCCN

HV6769 .D675 2012. URL http://catalogimages.wiley.
com/images/db/jimages/9780470880852.jpg;   http://
site.ebrary.com/id/10534000.

daSilvaAzevedo:2021:BLB

[dGGd21] Caio da Silva Azevedo, Rodrigo Franco Gonçalves, Vagner Luiz Gava, and Mauro de Mesquita Spinola. A Benford’s Law based methodology for fraud detection in social welfare programs: Bolsa Familia analysis. Physica A: Statistical Mechanics and its Applications, 567:125626, April 2021. CODEN PHYADX. ISSN 0378-4371 (print), 1873-2119 (electronic).

diGiovanni:2013:FET

[dGL13] Julian di Giovanni and Andrei A. Levchenko. Firm entry, trade, and welfare in Zipf’s world. Journal of International Economics, 89(2):283–296, March 15, 2013. CO-DEN JIECBE. ISSN 0022-1996 (print), 1873-0353 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022199612001365.

Durst:2020:IGD

[DHL^{+}20] R. F. Durst, C. Huynh, A. Lott, S. J. Miller, E. A. Palsson, W. Touw, and G. Vriend. The inverse gamma distribution and Benford’s law. PUMP J. Undergrad. Res., 3:95–109, 2020. ISSN 2765-8724.

Durtschi:2004:EUB

[DHP04] C. Durtschi, W. Hillison, and C. Pacini. The effective use of Benford’s Law to assist in detecting fraud in accounting data. Journal of Forensic Accounting, 5(1):17–34, ???? 2004. CODEN ???? ISSN 1524-5586.

Diaconis:1973:LMI

[Dia73] Persi Diaconis. Limits of measures of the integers with applications to random number generators and the distribution of leading digits. Memorandum NS-211, Department of Statistics, Harvard University, Cambridge, MA, USA, March 22, 1973.

Diaconis:1974:WSA

[Dia74] Persi Diaconis. Weak and Strong Averages in probability and the theory numbers. Ph.D. thesis, Department of Statistics, Harvard University, Cambridge, MA, USA, 1974. xvii + 117 pp.

Diaconis:1977:DLD

[Dia77] Persi Diaconis. The distribution of leading digits and uniform distribution mod1. Annals of Probability, 5(1):72–81, February 1977. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/
euclid.aop/1176995891.

Diekmann:2002:DFM

[Die02] Andreas Diekmann. Diagnose von Fehlerquellen und methodische Qualität in der sozialwissenschaftlichen Forschung. (German) [Diagnosis of errors and methodological quality in social science research]. Report ITA-02-04, Institut für Technikfolgen-Abschätzung, Vienna, Austria, 2002. ???? pp.

Diekmann:2004:DEE

[Die04a] Andreas Diekmann. Datenfälschung. Ergebnisse aus Experimenten mit der Benford Verteilung. (German) [Data falsification. Results from experiments with the Benford distribution]. Report, ETH Zürich, Zürich, Switzerland, 2004.

Diekmann:2004:FDU

[Die04b] Andreas Diekmann. Not the first digit! Using Benford’s Law to detect fraudulent scientific data. Report, ETH Zürich, Zürich, Switzerland, October 2004. 26 pp. URL http://

rfe.cs.oswego.edu/eps/othr/papers/0507/0507001.pdf.

Diekmann:2007:FDU

[Die07] Andreas Diekmann. Not the first digit! Using Benford’s Law to detect fraudulent scientific data. Journal of Applied Statistics, 34(3–4):321–329, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/02664760601004940.

Diekmann:2012:MUB

[Die12] Andreas Diekmann. Making use of “Benford’s Law” for the randomized response technique. Sociological Methods and Research, 41(2):325–334, May 2012. ISSN 0049-1241 (print), 1552-8294 (electronic). URL http://smr.sagepub.com/

content/41/2/325;   http://www.socio.ethz.ch/news/
Diekmann_Benford_RRT_SMR_2012.pdf.

Deligny:2010:RRL

[DJ10a] H. Deligny and P. Jolissaint. Relations de récurrence linéaires, primitivité et loi de Benford. (French) [Linear recurrence relations, primitivity, and Benford’s Law]. ArXiv e-prints, July 23, 2010. CODEN ???? ISSN ???? URL

http://adsabs.harvard.edu/abs/2010arXiv1007.5349D;
 http://arxiv.org/abs/1007.5349.

Diekmann:2010:BLF

[DJ10b] Andreas Diekmann and Ben Jann. Benford’s Law and fraud detection: Facts and legends. German Economic Review, 11 (3):397–401, August 2010. CODEN ???? ISSN 1465-6485 (print), 1468-0475 (electronic).

Deligny:2013:RRL

[DJ13] Hugues Deligny and Paul Jolissaint. Relations de récurrence linéaires, primitivité et loi de Benford. (French) [Linear recurrence relations, primitivity, and Benford’s law]. Elemente der Mathematik, 68(1):9–21, 2013. ISSN 0013-6018 (print), 1420-8962 (electronic).

Dorfleitner:2009:PBE

[DK09] Gregor Dorfleitner and Christian Klein. Psychological barriers in European stock markets: Where are they? Global Finance Journal, 19(3):268–285, October 2009. CO-DEN ???? ISSN 1044-0283 (print), 1873-5665 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1044028308000549.

Decker:2007:GPC

[DKM07] Ethan H. Decker, Andrew J. Kerkhoff, and Melanie E. Moses. Global patterns of city size distributions and their fundamental drivers. PLoS One, 2(9):e934, September 2007. CODEN POLNCL. ISSN 1932-6203. URL http://www.plosone.org/

article/info:doi/10.1371/journal.pone.0000934.

Dumbgen:2007:EBA

[DL07] L. Dümbgen and C. Leuenberger. Explicit bounds for the approximation error in Benford’s law. ArXiv e-prints, May 2007.

CODEN ???? ISSN ???? URL http://adsabs.harvard.
edu/abs/2007arXiv0705.4488D.

Dumbgen:2008:EBA

[DL08] Lutz Dümbgen and Christoph Leuenberger. Explicit bounds for the approximation error in Benford’s law. Electronic Communications in Probability, 13:99–112, 2008. CO-DEN ???? ISSN 1083-589X. URL http://arxiv.org/

abs/0705.4488;   http://weber.math.washington.edu/
~ejpecp/ECP/index.php.

Kristiansen:2004:BTZ

[dLKHS04a] Kai de Lange Kristiansen, Geir Helgesen, and Arne T. Skjeltorp. Braid theory and Zipf relation used in dynamics of magnetic microparticles. Economics Letters, 30(1):2387–2388, May 1, 2004. CODEN ECLEDS. ISSN 0165-1765 (print),

1873-7374 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304885303019127.

Kristiansen:2004:EOZ

[dLKHS04b] Kai de Lange Kristiansen, Geir Helgesen, and Arne T. Skjeltorp. Experimental observation of Zipf–Mandelbrot relation. Physica A: Statistical Mechanics and its Applications, 335(3–4):413–420, April 15, 2004. CODEN PHYADX. ISSN 0378-4371 (print), 1873-2119 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378437103011804.

Kristiansen:2006:BTZ

[dLKHS06] K. de Lange Kristiansen, G. Helgesen, and A. T. Skjeltorp. Braid theory and Zipf–Mandelbrot relation used in microparticle dynamics. European Physical Journal B: Condensed Matter and Complex Systems, 51(3):363–371, June 2006. CODEN EPJBFY. ISSN 1434-6028 (print), 1434-6036 (electronic).

Dehaene:1992:CLR

[DM92] Stanislas Dehaene and Jacques Mehler. Cross-linguistic regularities in the frequency of number words. Cognition, 43 (1):1–29, 1992. CODEN CGTNAU. ISSN 0010-0277 (print),

1873-7838 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001002779290030L.

Durst:2019:BLB

[DM19] Rebecca F. Durst and Steven J. Miller. Benford’s law beyond independence: tracking Benford behavior in copula models. Involve, 12(7):1193–1218, 2019. ISSN 1944-4176.

Dlugosz:2009:VLD

[DMF09] Stephan Dlugosz and Ulrich Müller-Funk. The value of the last digit: statistical fraud detection with digit analysis. Advances in Data Analysis and Classification, 3(3):281–290, December

2009. URL http://link.springer.com/article/10.1007/
s11634-009-0048-5.

deMarchi:2006:AAS

[dMH06] Scott de Marchi and James Hamilton. Assessing the accuracy of self-reported data: an evaluation of the toxics release inventory. Journal of Risk and Uncertainty, 32(1):57–76, January 2006. CODEN JRUNEN. ISSN 0895-5646 (print), 1573-0476 (electronic).

Doyle:2011:ITA

[DMJH11] Laurance R. Doyle, Brenda McCowan, Simon Johnston, and Sean F. Hanser. Information theory, animal communication, and the search for extraterrestrial intelligence. Acta Astronautica, 68(3–4):406–417, February/March 2011. CO-DEN AASTCF. ISSN 0094-5765 (print), 1879-2030 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0094576509005682. SETI Special Edition.

Dorogovtsev:2006:FON

[DMO06] S. N. Dorogovtsev, J. F. F. Mendes, and J. G. Oliveira. Frequency of occurrence of numbers in the World Wide Web. Physica A: Statistical Mechanics and its Applications, 360(2): 548–556, February 2006. CODEN PHYADX. ISSN 0378-4371 (print), 1873-2119 (electronic). URL http://arxiv.org/

abs/physics/0504185;   http://www.sciencedirect.com/
science/article/pii/S0378437105006874.

Deckert:2010:IBL

[DMO10] Joseph Deckert, Mikhail Myagkov, and Peter C. Ordeshook. The irrelevance of Benford’s Law for detecting fraud in elections. CALTECH working paper 9, California Institute of Technology, Pasadena, CA, USA, March 9, 2010. 27 pp. URL

http://vote.caltech.edu/drupal/node/327.

Deckert:2011:BLD

[DMO11] Joseph Deckert, Mikhail Myagkov, and Peter C. Ordeshook. Benford’s Law and the detection of election fraud. Political Analysis, 19(3):245–268, Summer 2011. ISSN 1047-1987 (print), 1476-4989 (electronic). URL http://pan.

oxfordjournals.org/content/19/3/245.abstract. See

[Meb11] for a rebuttal.

Dahui:2005:TRZ

[DMZ05] Wang Dahui, Li Menghui, and Di Zengru. True reason for Zipf’s law in language. Physica A: Statistical Mechanics and its Applications, 358(2–4):545–550, December 15, 2005. CO-DEN PHYADX. ISSN 0378-4371 (print), 1873-2119 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378437105004085.

Drake:2000:CAA

[DN00] Philip D. Drake and Mark J. Nigrini. Computer assisted analytical procedures using Benford’s Law. Journal of Accounting Education, 18(2):127–146, Spring 2000. CO-DEN ???? ISSN 0748-5751 (print), 1873-1996 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0748575100000087;   http://www.sciencedirect.
com/science/journal/07485751.

Dixit:2010:EEP

[DN10] U. J. Dixit and M. Jabbari Nooghabi. Efficient estimation in the Pareto distribution. Statistical Methodology, 7(6):687– 691, November 2010. CODEN ???? ISSN 1572-3127 (print),

1878-0954 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1572312710000419.

Dixit:2011:EEP

[DN11] U. J. Dixit and M. Jabbari Nooghabi. Efficient estimation in the Pareto distribution with the presence of outliers. Statistical Methodology, 8(4):340–355, July 2011. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1572312711000141.

Doolittle:1910:SNF

[Doo10] C. L. Doolittle. Simon Newcomb, F.R.S., LL.D., D.C.L. Proceedings of the American Philosophical Society, 49(197):iii– xviii, October/December 1910. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/984092.

Dolgikh:2001:PCS

[DS01] Dmitry G. Dolgikh and Andrei M. Sukhov. Parameters of cache systems based on a Zipf-like distribution. Computer Networks (Amsterdam, Netherlands: 1999), 37(6):711– 716, December 15, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.elsevier.

nl/gej-ng/10/15/22/67/36/31/abstract.html;   http:
//www.sciencedirect.com/science/article/pii/S1389128601002432.

DelAcebo:2005:BLN

[DS05] Esteve Del Acebo and Mateu Sbert. Benford’s Law for natural and synthetic images. In Neumann et al. [NSGP05], pages 169–176. ISBN 3-905673-27-4. ISSN 1816-0859. URL http://

www.eg.org/EG/DL/WS/COMPAESTH/COMPAESTH05/169-176.
pdf.

daSilva:2020:VNB

[dSFSL20] A. J. da Silva, S. Floquet, D. O. C. Santos, and R. F. Lima. On the validation of the Newcomb–Benford law and the Weibull distribution in neuromuscular transmission. Physica A: Statistical Mechanics and its Applications, 553:124606, 15, 2020. CODEN PHYADX. ISSN 0378-4371 (print), 1873-2119 (electronic).

Dubinsky:2001:MFF

[Dub01] Bruce G. Dubinsky. Math formula fights fraud. Legal Times of Washington, XXIV(9):??, February 26, 2001. ISSN 0732-7536.

Duncan:1967:AUD

[Dun67] R. L. Duncan. An application of uniform distributions to the Fibonacci numbers. Fibonacci Quarterly, 5(2):137–140, April 1967. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/5-2/duncan.pdf.

Duncan:1969:NID

[Dun69] R. L. Duncan. Note on the initial digit problem. Fibonacci Quarterly, 7(5):474–475, December 1969. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/7-
5/duncan.pdf.

Dunn:1980:SFW

[Dun80] Stanley Dunn. Surveyor’s forum: Working on interpretations. ACM Computing Surveys, 12(2):255, June 1980. CODEN CMSVAN. ISSN 0010-4892. See [Moh79, Moh80].

Duran:1974:ACS

[Dur74] J. Peter Duran. Almost convergence, summability and ergodicity. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):372–387, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Durrett:1991:PTE

[Dur91] Richard Durrett. Probability: theory and examples. The Wadsworth and Brooks/Cole statistics/probability series. Wadsworth and Brooks/Cole Advanced Books and Software, Pacific Grove, CA, USA, 1991. ISBN 0-534-13206-5. ix + 453 pp. LCCN QA273 .D865 1991. US$49.95.

Durrett:2005:PTE

[Dur05] Richard Durrett. Probability: theory and examples. Duxbury advanced series. Thomson Brooks/Cole, Belmont, CA, USA, third edition, 2005. ISBN 0-534-42441-4. xi + 497 pp. LCCN QA273 .D865 2005.

Duranton:2006:SFZ

[Dur06] Gilles Duranton. Some foundations for Zipf’s law: Product proliferation and local spillovers. Regional Science and Urban Economics, 36(4):542–563, July 1, 2006. CO-DEN RSUEDM. ISSN 0166-0462 (print), 1879-2308 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166046206000354.

deVries:2013:CLN

[dVM13] Pepijn de Vries and Albertinka J. Murk. Compliance of LC50 and NOEC data with Benford’s Law: An indication of reliability? Ecotoxicology and Environmental Safety, 98 (0):171–178, 2013. ISSN 0147-6513 (print), 1090-2414 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0147651313003692.

Dandurand:2015:DCP

[DW15] Brian Dandurand and Margaret M. Wiecek. Distributed computation of Pareto sets. SIAM Journal on Optimization, 25 (2):1083–1109, ???? 2015. CODEN SJOPE8. ISSN 1052-6234 (print), 1095-7189 (electronic).

Das:2003:RRE

[DZ03] Somnath Das and Huai Zhang. Rounding-up in reported EPS, behavioral thresholds, and earnings management. Journal of Accounting and Economics, 35(1):31– 50, 2003. ISSN 0165-4101 (print), 1879-1980 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165410102000964.

Edwards:2011:LFP

[EC11] Roderick Edwards and Laura Collins. Lexical frequency profiles and Zipf’s Law. Language Learning, 61(1):1–30, 2011. ISSN 0023-8333 (print), 1467-9922 (electronic).

Eliazar:2012:PLC

[EC12] Iddo I. Eliazar and Morrel H. Cohen. Power-law connections: From Zipf to Heaps and beyond. Information Processing and Management, 23(3):56–74, May 1, 2012. CO-DEN IPMADK. ISSN 0306-4573 (print), 1873-5371 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0003491613000171.

Egghe:2012:AGL

[EG12a] Leo Egghe and Raf Guns. Applications of the generalized law of Benford to informetric data. Journal of the American Society for Information Science and Technology, 63(8):1662– 1665, August 2012. CODEN JASIEF. ISSN 1532-2882 (print),

1532-2890 (electronic). URL http://onlinelibrary.wiley.
com/doi/10.1002/asi.22690/abstract.

Egghe:2012:BCA

[EG12b] Leo Egghe and Raf Guns. Brief communication: Applications of the generalized law of Benford to informetric data. Journal of the American Society for Information Science and Technology, 63(8):1662–1665, August 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Egghe:1985:CLL

[Egg85] L. Egghe. Consequences of Lotka’s Law for the Law of Bradford. Journal of Documentation, 41(3):173–189, ???? 1985. CODEN JDOCAS. ISSN 0022-0418 (print), 1758-7379 (electronic).

Egghe:1986:RDB

[Egg86] L. Egghe. Research: The dual of Bradford’s Law. Journal of the American Society for Information Science, 37(4):246– 255, July 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Egghe:1990:RND

[Egg90a] L. Egghe. Research: A note on different Bradford multipliers. Journal of the American Society for Information Science, 41 (3):204–209, April 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Egghe:1990:RAT

[Egg90b] L. Egghe. Research: Applications of the theory of Bradford’s Law to the calculation of Leimkuhler’s Law and to the completion of bibliographies. Journal of the American Society for Information Science, 41(7):469–492, October 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Egghe:1994:RSF

[Egg94] L. Egghe. Research: Special features of the author– publication relationship and a new explanation of Lotka’s Law based on convolution theory. Journal of the American Society for Information Science, 45(6):422–427, July 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Egghe:1999:LZM

[Egg99a] L. Egghe. On the law of Zipf–Mandelbrot for multi-word phrases. Journal of the American Society for Information Science, 50(3):233–241, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Egghe:1999:RLZ

[Egg99b] L. Egghe. Research: On the law of Zipf–Mandelbrot for multiword phrases. Journal of the American Society for Information Science, 50(3):233–241, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Egghe:2005:PPL

[Egg05a] L. Egghe. The power of power laws and an interpretation of Lotkaian informetric systems as self-similar fractals. Journal of the American Society for Information Science and Technology, 56(7):669–675, May 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Egghe:2005:RBC

[Egg05b] L Egghe. Relations between the continuous and the discrete Lotka power function. Journal of the American Society for Information Science and Technology, 56(7):664–668, May 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Egghe:2005:ZLC

[Egg05c] L. Egghe. Zipfian and Lotkaian continuous concentration theory. Journal of the American Society for Information Science and Technology, 56(9):935–945, July 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Egghe:2005:PLI

[Egg05d] Leo Egghe. Power laws in the information production process: Lotkaian informetrics. Elsevier, Amsterdam, The Netherlands, 2005. ISBN 0-12-088753-3, 0-08-048011-X (e-book). xvii + 427 pp. LCCN Z669.8 .E44 2005. See pages 204–205 for Benford’s Law discussion.

Egghe:2006:DHL

[Egg06] L. Egghe. The dependence of the height of a Lorenz curve of a Zipf function on the size of the system. Mathematical and Computer Modelling, 43(7–8):870–879, April 15, 2006. CO-DEN MCMOEG. ISSN 0895-7177 (print), 1872-9479 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S089571770500508X.

Egghe:2007:UHL

[Egg07] Leo Egghe. Untangling Herdan’s Law and Heaps’ Law: Mathematical and informetric arguments. Journal of the American Society for Information Science and Technology, 58(5):702– 709, March 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Egghe:2009:PRP

[Egg09] Leo Egghe. Performance and its relation with productivity in Lotkaian systems. Scientometrics, 81(2):567–585, November 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/s11192-008-2226-1.

Egghe:2010:CSS

[Egg10a] L. Egghe. Characteristic scores and scales in a Lotkaian framework. Scientometrics, 83(2):455–462, May 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/s11192-009-0009-y.

Egghe:2010:BCN

[Egg10b] Leo Egghe. Brief communication: A new short proof of Naranan’s theorem, explaining Lotka’s Law and Zipf’s Law. Journal of the American Society for Information Science and Technology, 61(12):2581–2583, December 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Egghe:2010:NSP

[Egg10c] Leo Egghe. A new short proof of Naranan’s theorem, explaining Lotka’s Law and Zipf’s Law. Journal of the American Society for Information Science and Technology, 61(12): 2581–2583, 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Egghe:2011:BLS

[Egg11] Leo Egghe. Benford’s law is a simple consequence of Zipf’s law. ISSI Newsletter, 7(3):55–56, September 2011. URL http://

www.issi-society.info/archives/newsletter27.pdf.

Egghe:2013:FRB

[Egg13] L. Egghe. The functional relation between the impact factor and the uncitedness factor revisited. Journal of Informetrics, 7(1):183–189, 2013. ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157712000909.

Edelsbrunner:2014:BLC

[EHM^{+}14] X. Edelsbrunner, K. Huan, B. Mackall, S. J. Miller, J. Powell, C. Turnage-Butterbaugh, and M. Weinstein. Benford’s Law, the Cauchy distribution and financial data. Report, ????, ????, 2014.

Ectors:2020:ZPL

[EKJ^{+}20] Wim Ectors, Bruno Kochan, Davy Janssens, Tom Bellemans, and Geert Wets. Zipf’s power law in activity schedules and the effect of aggregation. Future Generation Computer Systems, 107(??):1014–1025, June 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167739X17321891.

Engel:2003:BLE

[EL03] Hans-Andreas Engel and Christoph Leuenberger. Benford’s law for exponential random variables. Statistics & Probability Letters, 63(4):361–365, July 15, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203001019.

Eliazar:2011:LLZ

[Eli11] Iddo Eliazar. Limit laws for Zipf’s law. Journal of Physics A (Mathematical and General), 44(2):022001, 2011. CODEN JPHAC5. ISSN 1751-8113 (print), 1751-8121 (electronic).

Eliazar:2013:BLP

[Eli13] Iddo I. Eliazar. Benford’s Law: A Poisson perspective. Physica A: Statistical Mechanics and its Applications, 392(16):3360– 3373, 2013. CODEN PHYADX. ISSN 0378-4371 (print),

1873-2119 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378437113002902.

Elshoff:1978:SSC

[Els78] James L. Elshoff. A study of the structural composition of PL/I programs. ACM SIGPLAN Notices, 13(6):29–37, June 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Eliahou:2011:MDP

[EMS11] Shalom Eliahou, Bruno Massé, and Dominique Schneider. On the mantissa distribution of powers of natural and prime numbers. Preprint, Université du Littoral Côte d’Opale, Nord-Pas-de-Calais, France, November 15,

2011. 14 pp. URL http://fr.wikipedia.org/wiki/
Universit%C3%A9_du_Littoral_C%C3%B4te_d%27Opale;
http://www-lmpa.univ-littoral.fr/publications/articles/
lmpa447.pdf.

Eliahou:2012:MDP

[EMS13] Shalom Eliahou, Bruno Massé, and Dominique Schneider. On the mantissa distribution of powers of natural and prime numbers. Acta Mathematica Hungarica, 139(1–2):1–5, April 2013. ISSN 0001-5954, 0236-5294, 1588-2632. URL http://link.

springer.com/article/10.1007/s10474-012-0244-1.

Egghe:2002:DRC

[ER02] L. Egghe and I. K. Ravichandra Rao. Duality revisited: Construction of fractional frequency distributions based on two dual Lotka laws. Journal of the American Society for Information Science and Technology, 53(10):789–801, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Egghe:2012:TPS

[ER12] Leo Egghe and Ronald Rousseau. Theory and practice of the shifted Lotka function. Scientometrics, 91(1):295–301, April 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/s11192-011-0539-y.

elSehity:2005:PDA

[eSHK05] Tarek el Sehity, Erik Hoelzl, and Erich Kirchler. Price developments after a nominal shock: Benford’s Law and psychological pricing after the euro introduction. International Journal of Research in Marketing, 22(4):471–480, December 2005. CODEN IJRME6. ISSN 0167-8116 (print), 1873-

8001 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167811605000522.

Estoup:1916:GSM

[Est16] J. B. Estoup. Gammes Sténographiques: méthode & exercices pour l’acquisition de la vitesse (French) [Stenographic ranges: method and exercises for achieving speed]. Institut Sténographique de France, Paris, France, fourth edition, 1916. 151 pp. LCCN ???? This on shorthand writing is the earliest known publication of the power-law distribution of word frequencies, generally credited to [Zip32].

Fujiwara:2004:GPZ

[FAD^{+}04] Yoshi Fujiwara, Hideaki Aoyama, Corrado Di Guilmi, Wataru Souma, and Mauro Gallegati. Gibrat and Pareto–Zipf revisited with European firms. Physica A: Statistical Mechanics and its Applications, 344(1–2):112–116, December 1, 2004. CODEN PHYADX. ISSN 0378-4371 (print), 1873-

2119 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378437104009173.

Fairthorne:1969:PDE

[Fai69] Robert A. Fairthorne. Progress in documentation: Empirical hyperbolic distributions (Bradford–Zipf–Mandelbrot) for bibliometric description and prediction. Journal of Documentation, 25(4):313–343, December 1969. CODEN JDOCAS. ISSN 0022-0418 (print), 1758-7379 (electronic). Reprinted in [Fai05].

Fairthorne:2005:PDE

[Fai05] Robert A. Fairthorne. Progress in documentation: Empirical hyperbolic distributions (Bradford–Zipf–Mandelbrot) for bibliometric description and prediction. Journal of Documentation, 61(2):171–193, ???? 2005. CODEN JDOCAS. ISSN 0022-0418 (print), 1758-7379 (electronic). Reprint of [Fai69].

Ferariu:2014:PGP

[FC14] L. Ferariu and C. Cimpanu. Pareto genetic path planning hybridized with multi-objective Dijkstra’s algorithm. In 2014 18th International Conference System Theory, Control and Computing (ICSTCC), pages 341–346. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring,

MD 20910, USA, 2014. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=6982439.

Fang:2020:SCP

[FC20] Guojun Fang and Qihong Chen. Several common probability distributions obey Benford’s law. Physica A: Statistical Mechanics and its Applications, 540:123129, 8, 2020. CODEN PHYADX. ISSN 0378-4371 (print), 1873-2119 (electronic).

Fujiwara:2004:DPZ

[FDA^{+}04] Yoshi Fujiwara, Corrado Di Guilmi, Hideaki Aoyama, Mauro Gallegati, and Wataru Souma. Do Pareto–Zipf and Gibrat laws hold true? An analysis with European firms. Physica A: Statistical Mechanics and its Applications, 335(1–2):197–216, April 1, 2004. CODEN PHYADX. ISSN 0378-4371 (print),

1873-2119 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378437103011294.

Fedorowicz:1982:TFZ

[Fed82a] Jane Fedorowicz. The theoretical foundation of Zipf’s Law and its application to the bibliographic database environment. Journal of the American Society for Information Science, 33 (5):285–293, 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Fedorowicz:1982:ZMA

[Fed82b] Jane Fedorowicz. A Zipfian model of an automatic bibliographic system: An application to MEDLINE. Journal of the American Society for Information Science, 33(4):223–232, July 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Fedorowicz:1984:DEU

[Fed84] Jane Fedorowicz. Database evaluation using multiple regression techniques. SIGMOD Record (ACM Special Interest Group on Management of Data), 14(2):70–76, 1984. CODEN SRECD8. ISBN 0-89791-128-8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Fedorowicz:1987:DPE

[Fed87] Jane Fedorowicz. Database performance evaluation in an indexed file environment. ACM Transactions on Database Systems, 12(1):85–110, March 1987. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http:

//www.acm.org/pubs/articles/journals/tods/1987-12-

1/p85-fedorowicz/p85-fedorowicz.pdf;   http://www.
acm.org/pubs/citations/journals/tods/1987-12-1/p85-
fedorowicz/;   http://www.acm.org/pubs/toc/Abstracts/
tods/13675.html.

Feller:1950:IPT

[Fel50] William Feller. Introduction to Probability Theory, volume 2 of Wiley series in probability and mathematical statistics. Probability and mathematical statistics. Wiley, New York, NY, USA, 1950. ???? pp. LCCN QA273 .F37.

Feller:1957:IPT

[Fel57] William Feller. Introduction to Probability Theory, volume 2. Wiley, New York, NY, USA, 1957. 461 (vol. 1), 626 (vol. 2)

pp. LCCN QA273 .F3712. URL http://catdir.loc.gov/
catdir/toc/onix01/57010805.html.

Feller:1966:IPT

[Fel66] William Feller. Introduction to Probability Theory, volume 2 of Wiley series in probability and mathematical statistics. Probability and mathematical statistics. Wiley, New York, NY, USA, second edition, 1966. ISSN 0271-6232. xviii + 626 pp. LCCN QA273 .F3727 1966.

Feller:1968:IPT

[Fel68] William Feller. Introduction to Probability Theory and its Applications, volume 2. Wiley, New York, NY, USA, third edition, 1968. ISBN 0-471-25708-7. xviii + 509 pp. LCCN QA273 .F3713. See [BH10a, BH10b] for a discussion of flaws in the derivation of Benford’s Law in this book.

Fellman:2017:BP

[Fel17] Johan Fellman. Benford paradox. Arkhimedes, ??(4):26–33, 2017. CODEN AKMDA5. ISSN 0004-1920.

Fernandez:2013:SPC

[Fer13] Arturo J. Fernández. Smallest Pareto confidence regions and applications. Computational Statistics & Data Analysis, 62(??):11–25, June 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167947312004501.

Fewster:2009:SEB

[Few09] Rachel M. Fewster. A simple explanation of Benford’s Law. The American Statistician, 63(1):26–32, February 2009. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://pubs.amstat.org/doi/pdfplus/10.
1198/tast.2009.0005.

Feldstein:1976:CED

[FG76] Alan Feldstein and Richard Goodman. Convergence estimates for the distribution of trailing digits. Journal of the ACM, 23 (2):287–297, April 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Farkas:2010:SUN

[FG10] Janos Farkas and György Gyurky. The significance of using the Newcomb–Benford law as a test of nuclear halflife calculations. Acta Physica Polonica B, 41(6):1213–1221, 2010. CODEN APOBBB. ISSN 0587-4254 (print), 1509-

5770 (electronic). URL http://adsabs.harvard.edu/abs/
2010arXiv1006.3615F;   http://arxiv.org/abs/1006.3615.

Friar:2012:GSB

[FGPM12] James L. Friar, Terrance Goldman, and Juan Pérez-Mercader. Genome sizes and the Benford distribution. PLoS One, 7(5): e36624:1–e36624:9, May 2012. CODEN POLNCL. ISSN 1932-

6203. URL http://journals.plos.org/plosone/article?
id=10.1371/journal.pone.0036624.

Friar:2016:UBL

[FGPM16] J. L. Friar, T. Goldman, and J. Pérez-Mercader. Ubiquity of Benford’s Law and emergence of the reciprocal distribution. Physics Letters A, 380(22–23):1895–1899, 2016. CODEN PY-LAAG. ISSN 0375-9601 (print), 1873-2429 (electronic).

Farbaniec:2011:AFD

[FGZZ11a] Marzena Farbaniec, Tadeusz Grabiński, Bartłomiej Zabłocki, and Wacław Zaja̧c. Application of the first digit law in credibility evaluation of the financial accounting data based on particular cases. Lecture slides for 10th International Congress on Internal Control, Internal Audit, Fraud and Anti-Corruption Issues, Kraków, September 14–16, 2011., 2011. URL http:/

/benford.pl/documents/benford_pikw.pdf.

Farbaniec:2011:UBL

[FGZZ11b] Marzena Farbaniec, Tadeusz Grabiński, Bartłomiej Zabłocki, and Wacław Zaja̧c. Using Benford’s law to detect irregularities in the collection of data. In ????, editor, 10th International Congress on Internal Control, Internal Audit, Fraud and Anti-Corruption Issues, Kraków, September 14–16, 2011, page ?? ????, ????, 2011. ISBN ???? LCCN ????

Furry:1945:DND

[FH45] W. H. Furry and Henry Hurwitz. Distribution of numbers and distribution of significant figures. Nature, 155(3924):52– 53, January 13, 1945. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v155/n3924/pdf/155052a0.pdf.

Furusawa:2003:ZLG

[FK03] Chikara Furusawa and Kunihiko Kaneko. Zipf’s Law in gene expression. Physical Review Letters, 90(??):088102, February 26, 2003. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145. URL http://link.aps.org/

doi/10.1103/PhysRevLett.90.088102.

Fuchs:1984:PDP

[FL84] A. Fuchs and G. Letta. Sur le problème du premier chiffre décimal. (French) [On the problem of the first decimal digit]. Bollettino della Unione Matemàtica Italiana. Serie VI. B, 3: 451–461, 1984. ISSN 0392-4041.

Fuchs:1994:PDP

[FL94] Aimé Fuchs and Giorgio Letta. Le problème du premier chiffre décimal pour les nombres premiers. (French) [the problem for the first decimal digit of prime numbers]. Rendiconti della Accademia Nazionale delle Scienze detta dei XL. Serie V. Memorie di Matemàtica, 18(1):81–87, ????

1994. URL http://www.combinatorics.org/ojs/index.
php/eljc/article/download/v3i2r25/pdf.

Fuchs:1996:PDP

[FL96] Aimé Fuchs and Giorgio Letta. Le problème du premier chiffre décimal pour les nombres premiers. (French) [the problem for the first decimal digit of prime numbers]. Electronic Journal of Combinatorics, 3(2):R25:1–R25:7, ???? 1996.

ISSN 1077-8926 (print), 1097-1440 (electronic). URL http:/

/www.combinatorics.org/ojs/index.php/eljc/article/

download/v3i2r25/pdf. The Foata Festschrift volume.

Flehinger:1966:PRI

[Fle66] B. J. (Betty J.) Flehinger. On the probability that a random integer has initial digit A. American Mathematical Monthly, 73(10):1056–1061, December 1966. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2314636.

Fleiss:1981:SMR

[Fle81] Joseph L. Fleiss. Statistical Methods for Rates and Proportions. Wiley series in probability and mathematical statistics. Wiley, New York, NY, USA, second edition, 1981. ISBN 0-471-06428-9. xviii + 321 pp. LCCN QA279 .F58 1981. URL http://www.loc.gov/

catdir/description/wiley032/80026382.html;   http:
//www.loc.gov/catdir/enhancements/fy0607/80026382-
b.html;   http://www.loc.gov/catdir/toc/onix02/80026382.
html.

Farris:2021:RRB

[FLM^{+}21] Madeleine Farris, Noah Luntzlara, Steven J. Miller, Lily Shao, and Mengxi Wang. Recurrence relations and Benford’s law. Stat. Methods Appl., 30(3):797–817, 2021. ISSN 1618-2510.

Fleiss:2003:SMR

[FLP03] Joseph L. Fleiss, Bruce A. Levin, and Myunghee Cho Paik. Statistical Methods for Rates and Proportions. Wiley series in probability and statistics. Wiley, New York, NY, USA, third edition, 2003. ISBN 0-471-52629-0 (hardcover). xxvii + 760 pp. LCCN QA279 .F58 2003. URL http://www.loc.

gov/catdir/bios/wiley042/2002191005.html;   http:/
/www.loc.gov/catdir/description/wiley034/2002191005.
html;   http://www.loc.gov/catdir/toc/wiley031/2002191005.
html.

Family:1989:KDG

[FM89] Fereydoon Family and Paul Meakin. Kinetics of droplet growth processes: Simulations, theory, and experiments. Physical Review A (Atomic, Molecular, and Optical Physics), 40(7):3836–3854, October 1989. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-4519. URL

http://link.aps.org/doi/10.1103/PhysRevA.40.3836.

Filipponi:1995:SPA

[FM95] Piero Filipponi and Renato Menicocci. Some probabilistic aspects of the terminal digits of Fibonacci numbers. Fibonacci Quarterly, 33(4):325–331, August 1995. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/
33-4/filipponi.pdf.

Fischer:2010:HCS

[FMS10] Martin H. Fischer, Richard A. Mills, and Samuel Shaki. How to cook a SNARC: Number placement in text rapidly changes spatial-numerical associations. Brain and Cognition, 72(3): 333–336, April 2010. CODEN ???? ISSN 0278-2626 (print),

1090-2147 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0278262609002024.

Fang:2024:BLR

[FMSV24] Xinyu Fang, Steven J. Miller, Maxwell Sun, and Amanda Verga. Benford’s law and random integer decomposition with congruence stopping condition. Journal of Number Theory, 264(??):307–356, November 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022314X24001367.

Ferreira:2015:LPA

[FNP^{+}15] Anselmo Ferreira, Luiz C. Navarro, Giuliano Pinheiro, Jefersson A. dos Santos, and Anderson Rocha. Laser printer attribution: Exploring new features and beyond. Forensic Science International, 247(0):105–125, 2015. ISSN 0379-0738 (print),

1872-6283 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0379073814005064. See also

[JHTS10].

Formann:2010:NBL

[For10] Anton K. Formann. The Newcomb–Benford Law in its relation to some common distributions. PLoS One, 5(5):e10541:1– e10541:13, May 2010. CODEN POLNCL. ISSN 1932-

6203. URL http://adsabs.harvard.edu/abs/2010PLoSO.
..510541F.

Fontanari:2004:SNM

[FP04] J. F. Fontanari and L. I. Perlovsky. Solvable null model for the distribution of word frequencies. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 70(??):042901, October 25, 2004. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-2376 (electronic). URL http:/

/link.aps.org/doi/10.1103/PhysRevE.70.042901.

Franel:1917:PTL

[Fra17] J. Franel. À propos des tables de logarithmes. (French) [On tables of logarithms]. Festschrift Naturforschenden der Gesellschaft in Zürich Vierteljahrsschrift, 62(??):286–295, ???? 1917. CODEN ???? ISSN ????

FerreriCancho:2005:CZL

[FRB05] Ramon Ferrer i Cancho, Oliver Riordan, and Béla Bollobás. The consequences of Zipf’s law for syntax and symbolic reference. Proceedings of the Royal Society of London. Series B. Biological sciences, 272(1562):561–565, March 7, 2005. CO-DEN PRLBA4. ISSN 0962-8452 (print), 1471-2954 (elec-

tronic). URL http://rspb.royalsocietypublishing.org/
content/272/1562/561.full.

Friedberg:1984:DFD

[Fri84] Stephen H. Friedberg. The distribution of first digits. College Mathematics Journal, 15(2):120–125, March 1984. CO-DEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic).

URL http://www.jstor.org/stable/2686516;   http://
www.tandfonline.com/doi/abs/10.1080/00494925.1984.
11972760.

Frunza:2016:CBL

[Fru16] Marius-Christian Frunza. Benford’s Law. In Marius-Christian Frunza, editor, Solving Modern Crime in Financial Markets, chapter 2K, pages 233–245. Academic Press, New York, USA, 2016. ISBN 0-12-804494-2 (hardcover), 0-12-804532-9

(e-book). URL http://www.sciencedirect.com/science/
article/pii/B9780128044940000176.

FerreriCancho:2002:ZLR

[FS02] Ramon Ferrer i Cancho and Ricard V. Solé. Zipf’s law and random texts. Advances in Complex Systems, 5(1):1–6, 2002. CODEN ???? ISSN 0219-5259.

FerreriCancho:2003:LEO

[FS03] Ramon Ferrer i Cancho and Ricard V. Solé. Least effort and the origins of scaling in human language. Proceedings of the National Academy of Sciences of the United States of America, 100(3):788–791, February 4, 2003. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic).

Fu:2007:GBL

[FSS07] D. Fu, Y. Q. Shi, and W. Su. A generalized Benford’s law for JPEG coefficients and its applications in image forensics. In Edward J. Delp III and Ping Wah Wong, editors, Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, volume 6505 of Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series. SPIE Optical Engineering Press, Bellingham, WA, USA, February 2007. ISSN 0277-786X (print), 1996-756X (electronic). URL http:

//adsabs.harvard.edu/abs/2007SPIE.6505E..55F.

Feldstein:1986:OUS

[FT86] Alan Feldstein and Peter Turner. Overflow, underflow, and severe loss of significance in floating-point addition and subtraction. IMA Journal of Numerical Analysis, 6(2):241–251, April 1986. CODEN IJNADH. ISSN 0272-4979 (print), 1464-3642 (electronic).

Fujiwara:2004:ZLF

[Fuj04] Yoshi Fujiwara. Zipf law in firms bankruptcy. Physica A: Statistical Mechanics and its Applications, 337(1–2):219–230, June 1, 2004. CODEN PHYADX. ISSN 0378-4371 (print),

1873-2119 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378437104001165.

Furlan:1946:HSU

[Fur46] Luigi Vladimir Furlan. Das Harmoniegesetz der Statistik: Eine Untersuchung über die metrische Interdependenz der sozialen Erscheinungen. (German) [The Law of Harmony in statistics: an investigation of the metrical interdependence of social phenomena], volume 2 of Beihefte zum Assekuranzjahrbuch. Verlag für Recht und Gesellschaft A.-G., Basel, Switzerland, 1946. xii + 504 pp. LCCN HA29 .F85.

Faber:1985:AML

[FWW85] Vance Faber, Andrew B. White, Jr., and G. Milton Wing. Analysis of a model that leads to the Pareto law of wealth distribution. Journal of Mathematical Analysis and Applications, 112(2):579–594, 1985. CODEN JMANAK. ISSN 0022-247X (print), 1096-0813 (electronic).

Gabaix:1999:ZLC

[Gab99] Xavier Gabaix. Zipf’s law for cities: An explanation. The Quarterly Journal of Economics, 114(3):739–767, 1999. CO-DEN QJECAT. ISSN 0033-5533 (print), 1531-4650 (electronic).

Gaite:2006:CSZ

[Gai06] José Gaite. Cut-out sets and the Zipf law for fractal voids. Physica. D, Nonlinear phenomena, 223(2):248–255, November 15, 2006. CODEN PDNPDT. ISSN 0167-2789 (print),

1872-8022 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167278906003848.

Garcia:2011:FPA

[Gar11] Juan Manuel Garcia Garcia. A fixed-point algorithm to estimate the Yule-Simon distribution parameter. Applied Mathematics and Computation, 217(21):8560–8566, July 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-

5649 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0096300311004607.

Gonzalez-Alvarez:2014:POH

[GÁVRRL14] David L. González-Álvarez, Miguel A. Vega-Rodríguez, and Álvaro Rubio-Largo. Parallelizing and optimizing a hybrid differential evolution with Pareto tournaments for discovering motifs in DNA sequences. The Journal of Supercomputing, 70(2):880–905, November 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http://link.

springer.com/article/10.1007/s11227-014-1266-y.

Gangopadhyay:2013:EZL

[GB13] Kausik Gangopadhyay and Banasri Basu. Evolution of Zipf’s law for Indian urban agglomerations via-à-vis Chinese urban agglomerations. In Abergel et al. [ACCG13], pages 119–129. ISBN 88-470-2553-2, 88-470-2552-4.

Garcia-Berthou:2004:IBT

[GBA04] Emili García-Berthou and Carles Alcaraz. Incongruence between test statistics and P values in medical papers. BMC Medical Research Methodology, 4(1):1–5, 2004. ISSN 1471-2288.

Garrabrant:2015:ALU

[GBD^{+}15] Scott Garrabrant, Siddharth Bhaskar, Abram Demski, Joanna Garrabrant, George Koleszarik, and Evan Lloyd. Asymptotic logical uncertainty and the Benford test. CoRR, ??(??): ??, ???? 2015. CODEN ???? ISSN ???? URL http:

//arxiv.org/abs/1510.03370;   https://dblp.org/db/
journals/corr/corr1510.html#GarrabrantBDGKL15.

Galveas:2021:FAS

[GBF21] Daniel Galvêas, Fernando Barros, and Carlos Alessandro Fuzo. A forensic analysis of SARS-CoV-2 cases and COVID-19 mortality misreporting in the Brazilian population. Public Health, May 2021. CODEN PUHEAE. ISSN 0033-3506 (print), 1476-5616 (electronic).

Garrabrant:2016:ALU

[GBTB^{+}16] Scott Garrabrant, Tsvi Benson-Tilsen, Siddharth Bhaskar, Abram Demski, Joanna Garrabrant, George Koleszarik, and Evan Lloyd. Asymptotic logical uncertainty and the Benford test. In Artificial General Intelligence (AGI 2016), volume 9782 of Lecture Notes in Computer Science, pages 202–211. Springer International Publishing, Cham, Switzerland, 2016. ISBN 3-319-41648-0 (print), 3-319-41649-9 (e-book).

GomesdaSilva:2013:SAS

[GC13] Carlos Gomes da Silva and Pedro M. R. Carreira. Selecting audit samples using Benford’s Law. Auditing: A Journal of Practice & Theory, 32(2):53–65, May 2013. ISSN 0278-0380 (print), 1558-7991 (electronic). URL http://aaajournals.

org/doi/pdf/10.2308/ajpt-50340.

Gama:2005:MLE

[GCB^{+}05] Joãao Gama, Rui Camacho, Pavel Brazdil, Alípio Mário Jorge, and Luís Torgo, editors. Machine learning — ECML 2005: 16th European Conference on Machine Learning, Porto, Portugal, October 3–7, 2005: proceedings, volume 3720 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 3-540-29243-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q325.5 .E26 2005. URL http://

www.loc.gov/catdir/enhancements/fy0663/2005933045-
d.html;   http://www.springerlink.com/openurl.asp?
genre=issue\%26issn=0302-9743\%26volume=3720.

Gauvrit:2008:PLB

[GD08] Nicolas Gauvrit and Jean-Paul Delahaye. Pourquoi la loi de Benford n’est pas mystérieuse. (French) [Why Benford’s law is not mysterious]. Mathématiques et Sciences Humaines = Mathematics and Social Sciences, 46(182):7–15, 2008. CODEN ???? ISSN 0987-6936 (print), 1950-6821

(electronic). URL http://www.ehess.fr/revue-msh/pdf/
N182R1280.pdf.

Gauvrit:2009:SRI

[GD09a] N. Gauvrit and J.-P. Delahaye. Scatter and regularity imply Benford’s law … and more. ArXiv e-prints,

October 2009. CODEN ????  ISSN ????  URL

http://adsabs.harvard.edu/abs/2009arXiv0910.1359G;
 http://arxiv.org/abs/0910.1359.

Gauvrit:2009:LBG

[GD09b] Nicolas Gauvrit and Jean-Paul Delahaye. Loi de Benford générale. (French) [General Benford’s law]. Math. Sci. Hum. Math. Soc. Sci., ??(186):5–15, 2009. CODEN ???? ISSN 0987-6936 (print), 1950-6821 (electronic). URL http://msh.

revues.org/document11034.html.

Genolini:2010:EIP

[GD10] Christophe Genolini and Tarak Driss. Eveiller l’intérêt pour la statistique par l’exemple. (French) [Raising interest in statistics by example]. Statistique et Enseignement, 1(2):49–57, ???? 2010. CODEN ???? ISSN

2108-6745. URL http://publications-sfds.math.cnrs.
fr/index.php/StatEns/article/view/16.

Gauvrit:2011:SRI

[GD11] N. Gauvrit and J-P. Delahaye. Scatter and regularity implies Benford’s Law… and more. In Zenil [Zen11], pages 53–69. ISBN 981-4327-74-3. LCCN QA274 .R363 2011.

Gray:2014:TGR

[GD14] Glen L. Gray and Roger S. Debreceny. A taxonomy to guide research on the application of data mining to fraud detection in financial statement audits. International Journal of Accounting Information Systems, 15(4):357–380, 2014. CODEN IJAIA7. ISSN 1467-0895 (print), 1873-4723 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1467089514000323. 2013 Research Sympo-

sium on Information Integrity & Information Systems Assurance.

Geiringer:1948:RVF

[Gei48] Hilda Geiringer. Review of L. V. Furlan’s book: Das Harmoniegesetz der Statistik: Eine Untersuchung über die metrische Interdependenz der sozialen Erscheinungen. (German) [The Law of Harmony in statistics: an investigation of the metrical interdependence of social phenomena]. Journal of the American Statistical Association, 43(242):325–328, June 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2280379.

See [Fur46].

Goudsmit:1944:SFN

[GF44] S. A. Goudsmit and W. H. Furry. Significant figures of numbers in statistical tables. Nature, 154(3921):800–801, December 23, 1944. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v154/n3921/pdf/154800a0.pdf. See

later work [LG78], and severe criticism [Rai76, Rai85].

Grekos:2003:RSC

[GGA03] G. Grekos and R. Giuliano-Antonini. Regular sets and conditional density: an extension of Benford’s law. In ????, editor, Journées Arithmétiques XXIII, Graz, July 6–12, 2003, page ?? ????, ????, 2003. ISBN ???? LCCN ????

Goegebeur:2015:ETI

[GGO15] Yuri Goegebeur, Armelle Guillou, and Michael Osmann. An estimator for the tail index of an integrated conditional Pareto–Weibull-type model. Statistics & Probability Letters, 103(??):8–16, August 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715215001182.

Gonzales-Garcia:2009:BLM

[GGP09] J. Gonzales-Garcia and G. Pastor. Benford’s Law and macroeconomic data quality. Working paper, International Monetary Fund, Washington, DC, USA, 2009. URL http://papers.

ssrn.com/sol3/papers.cfm?abstract_id=1356437.

Gabaix:2003:TPL

[GGPS03] Xavier Gabaix, Parameswaran Gopikrishnan, Vasiliki Plerou, and H. Eugene Stanley. A theory of power-law distributions in financial market fluctuations. Nature, 423(6937):267–270, May 15, 2003. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v423/n6937/full/nature01624.html.

Graham:2009:SFD

[GHP09] Scott D. J. Graham, John Hasseldine, and David Paton. Statistical fraud detection in a commercial lobster fishery. New Zealand Journal of Marine and Freshwater Research, 43(1): 457–463, 2009.

Gibrat:1931:IE

[Gib31] Robert Gibrat. Les Inégalités économiques. Librairie du Recueil Sirey, Paris, France, 1931. 286 pp. LCCN HB251.GIB.

URL http://en.wikipedia.org/wiki/Gibrat%27s_law. Ap-

plications: aux inégalités des richesses, à la concentration des entreprises, aux populations des villes, aux statistiques des familles, etc. d’une loi nouvelle. La loi de l’effet proportionnel.

Giles:2005:BLN

[Gil05] David E. Giles. Benford’s Law and naturally occurring prices in certain ebaY auctions. Econometrics Working Paper EWP0505, Department of Economics, University of Victoria, Victoria, BC, Canada, May 2005. URL http://web.uvic.

ca/econ/ewp0505.pdf.

Giles:2007:BLN

[Gil07] D. E. Giles. Benford’s Law and naturally occurring prices in certain eBay auctions. Applied Economics Letters, 14(3): 157–161, 2007. CODEN ???? ISSN 1350-4851 (print), 1466-

4291 (electronic). URL http://www.tandfonline.com/toc/
rael20/current.

Gini:1957:SFD

[Gin57] Corrado Gini. Sulla frequenza delle cifre iniziali dei numeri osservati. (Italian) [On the frequency of initial digits of observed numbers]. Bulletin de l’Institut International de Statistique, 35(2):57–76, 1957. CODEN BISQA3. ISSN ???? 29th session, 2nd delivery, Rio de Janerio.

Giuliano:2011:WCS

[Giu11] Rita Giuliano. Weak convergence of sequences from fractional parts of random variables and applications. Theory of Probability and Mathematical Statistics, 83 (??):49–58, ???? 2011. CODEN TPMSCO. ISSN 0094-9000 (print), 1547-7363 (electronic). URL http:

//www.ams.org/journals/tpms/2011-83-00/S0094-9000-
2012-00841-7/S0094-9000-2012-00841-7.pdf.

Giuliano:2010:UPI

[GJ10] Rita Giuliano and Élise Janvresse. A unifying probabilistic interpretation of Benford’s law. Uniform Distribution Theory, 5(2):169–182, 2010. CODEN ???? ISSN 1336-

913X. URL http://www.boku.ac.at/MATH/udt/vol05/no2/
91GiulJan10-2.pdf.

Gonzalez-Jimenez:2010:BLN

[GJARPGAC10] D. Gonzalez-Jimenez, E. Argones-Rua, F. Perez-Gonzalez, and J. L. Alba-Castro. Benford’s Law for natural and synthetic images. In IEEE, editor, IEEE International Conference on Image Processing, pages 1245–1248. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2010.

Grendar:2007:ENP

[GJS07] Marian Grendar, George Judge, and Laura Schechter. An empirical non-parametric likelihood family of data-based Benford-like distributions. Physica A: Statistical Mechanics and its Applications, 380(0):429–438, July 1, 2007. CODEN PHYADX. ISSN 0378-4371 (print), 1873-2119 (electronic).

URL http://adsabs.harvard.edu/abs/2007PhyA..380..
429G;   http://www.sciencedirect.com/science/article/
pii/S0378437107001963.

Gan:2006:ZLS

[GLS06] Li Gan, Dong Li, and Shunfeng Song. Is the Zipf law spurious in explaining city-size distributions? Economics Letters, 92(2):256–262, August 1, 2006. CODEN ECLEDS. ISSN 0165-1765 (print), 1873-7374 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165176506000772.

Gunther:1996:ZLE

[GLSW96] R. Günther, L. Levitin, B. Schapiro, and P. Wagner. Zipf’s law and the effect of ranking on probability distributions. International Journal of Theoretical Physics, 35(2):395–417, 1996. CODEN IJTPBM. ISSN 0020-7748 (print), 1572-9575 (electronic).

Gell-Mann:1997:ETZ

[GM97a] Murray Gell-Mann. Empirical theory — Zipf’s law. In The quark and the jaguar: adventures in the simple and the complex [GM97b], pages 92–97. ISBN 0-7167-2725-0 (paperback). LCCN QC774.G45 A3 1994. URL http://www.

gbv.de/dms/bowker/toc/9780716725817.pdf;   http://
www.zentralblattmath.org/zmath/en/search/?an=0833.
00011.

Gell-Mann:1997:QJA

[GM97b] Murray Gell-Mann. The quark and the jaguar: adventures in the simple and the complex. W. H. Freeman and Company, New York, NY, USA, 1997. ISBN 0-7167-2725-0 (paperback). xviii + 392 pp. LCCN QC774.G45

A3 1994.  URL http://www.gbv.de/dms/bowker/toc/
9780716725817.pdf;   http://www.zentralblattmath.org/
zmath/en/search/?an=0833.00011.

Grandison:2008:BPK

[GM08] S. Grandison and R. J. Morris. Biological pathway kinetic rate constants are scale-invariant. Bioinformatics, 24(6):741–743, ???? 2008. ISSN 1367-4803 (print), 1367-4811 (electronic).

URL http://www.ncbi.nlm.nih.gov/pubmed/18238786.

Geyer:2012:ABL

[GM12] A. Geyer and J. Martí. Applying Benford’s law to volcanology. Geology, 40(4):327–330, April 2012. URL http://geology.

gsapubs.org/content/40/4/327.

Groharing:2022:BLC

[GM22] Brooks Groharing and David McCune. Benford’s Law and county-level votes in US Presidential elections. Chance: a magazine for people interested in the analysis of data, 35(2): 4–10, 2022. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-

2280 (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/09332480.2022.2066408.

Goel:2015:PCA

[GML15] Gagan Goel, Vahab Mirrokni, and Renato Paes Leme. Polyhedral clinching auctions and the AdWords polytope. Journal of the ACM, 62(3):18:1–18:??, June 2015. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gisin:2009:EEV

[GMV09] Vladimir Gisin, Andrey Markov, and Igor Vinukov. Estimation of extreme values of returns using the Zipf–Mandelbrot flow. International Journal of Pure and Applied Mathematics, 50(2):245–250, 2009. CODEN ???? ISSN 1311-8080 (print), 1314-3395 (electronic).

Gottwald:2002:NBL

[GN02] Georg A. Gottwald and Matthew Nicol. On the nature of Benford’s Law. Physica A: Statistical Mechanics and its Applications, 303(3–4):387–396, January 15, 2002. CODEN PHYADX. ISSN 0378-4371 (print), 1873-2119 (electronic).

URL http://adsabs.harvard.edu/abs/2002PhyA..303..
387G;   http://www.sciencedirect.com/science/article/
pii/S0378437101004976.

Gupta:2007:IMS

[GN07] Arjun K. Gupta and Saralees Nadarajah. Information matrices for some bivariate Pareto distributions. Applied Mathematics and Computation, 184(2):1069–1079, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Good:1965:LE

[Goo65] I. J. Good. Letter to the editor. The American Statistician, 19(3):43, June 1965. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681423.

Goodman:2016:PPB

[Goo16] William Goodman. The promises and pitfalls of Benford’s law. Significance (Oxford, England), 13(3):38–41, June 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Gorroochurn:2012:BPB

[Gor12a] Prakash Gorroochurn. Benford and the peculiar behavior of the first significant digit (1938). In Classic Problems of Probability [Gor12b], chapter 27, pages 233–239. ISBN 1-118-06325-2 (paperback), 1-118-31432-8, 1-118-31433-6 (e-book), 1-118-31434-4 (e-book), 1-118-31431-X (e-book). LCCN QA273.A4.

Gorroochurn:2012:CPP

[Gor12b] Prakash Gorroochurn. Classic Problems of Probability. Wiley, New York, NY, USA, 2012. ISBN 1-118-06325-2 (paperback), 1-118-31432-8, 1-118-31433-6 (e-book), 1-118-31434-4 (e-book), 1-118-31431-X (e-book). xi + 314 pp. LCCN QA273.A4.

Goto:1992:SEB

[Got92] Kazuo Gotô. Some examples of Benford sequences. Math. J. Okayama Univ., 34:225–232 (1994), 1992. CODEN MJOKAP. ISSN 0030-1566.

Goudsmit:1977:PEP

[Gou77] Samuel Goudsmit. Pitfalls in elementary probability. Proceedings of the American Philosophical Society, 121(2):188– 189, April 29, 1977. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http://www.jstor. org/stable/986527. Based on a lecture presented November 11, 1976.

Grammatikos:2013:UBL

[GP13] T. Grammatikos and N. I. Papanikolaou. Using Benford’s Law to detect fraudulent practices in banking. SSRN Scholarly Paper 2352775, Social Science Research Network, Rochester, NY, USA, 2013.

Gava:2007:IQF

[GR07] Alexandre Majola Gava and Luiz Roque de Souza Vitiello, Jr. Inflation, quarterly financial statements and fraud: Benford’s Law and the Brazilian case. In EnAN-PAD 2007: XXXI Encontro da ANPAD. Rio de Janeiro / RJ-22 a 26 de setembro de 2007, pages 1–13. ????,

????, 2007. URL http://www.anpad.org.br/diversos/
trabalhos/EnANPAD/enanpad_2007/CONT/2007_CONA819.pdf.

Grabski:2010:DDM

[Gra10] Severin Grabski. Discussion of “Data mining journal entries for fraud detection: An exploratory study”. International Journal of Accounting Information Systems, 11(3):182–185, September 2010. CODEN IJAIA7. ISSN 1467-0895 (print),

1873-4723 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1467089510000394. See [DG10,

Kri10].

Green:1986:BRP

[Gre86] M. W. Green. Book reviews: Pareto Distributions, by B. C. Arnold. Applied Statistics, 35(2):215, 1986. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Groos:1967:BCB

[Gro67] Ole V. Groos. Brief communications: Bradford’s Law and the Keenan–Atherton data. American Documentation, 18(1):46, January 1967. CODEN AMDOA7. ISSN 0096-946X (print), 1936-6108 (electronic).

Gelbukh:2001:ZHL

[GS01] Alexander Gelbukh and Grigori Sidorov. Zipf and Heaps Laws’ coefficients depend on language. Lecture Notes in Computer Science, 2004:332–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/0558/bibs/2004/
20040332.htm;   http://link.springer-ny.com/link/service/
series/0558/papers/2004/20040332.pdf.

Gupta:2024:ADQ

[GSB24] Solanki Gupta, Vivek Kumar Singh, and Sumit Kumar Banshal. Altmetric data quality analysis using Benford’s law. Scientometrics, 129(7):4597–4621, July 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05061-9.

Gambini:2012:PDD

[GSR12] Alessandro Gambini, Giovanni Mingari Scarpello, and Daniele Ritelli. Probability of digits by dividing random numbers: A ψ and ζ functions approach. Expositiones Mathematicae, 30(3): 223–238, ???? 2012. CODEN ???? ISSN 0723-0869 (print),

1878-0792 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0723086912000230.

Gunther:1992:PCZ

[GSW92] R. Günther, B. Schapiro, and P. Wagner. Physical complexity and Zipf’s law. International Journal of Theoretical Physics, 31(3):525–543, 1992. CODEN IJTPBM. ISSN 0020-7748 (print), 1572-9575 (electronic).

Gigerenzer:1999:SHM

[GT99] Gerd Gigerenzer and Peter M. Todd. Simple heuristics that make us smart. Evolution and cognition. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1999. ISBN 0-19-512156-2. xv + 416 pp. LCCN

BD260 .G54 1999. URL http://www.loc.gov/catdir/
enhancements/fy0640/98051084-d.html;   http://www.
loc.gov/catdir/enhancements/fy0640/98051084-t.html;
 http://www.loc.gov/catdir/enhancements/fy0726/98051084-
b.html.

Gunnel:2009:DBL

[GT09] S. Günnel and K.-H. Tödter. Does Benford’s Law hold in economic research and forecasting? Empirica, 36(3):273– 292, August 2009. CODEN ???? ISSN 0340-8744 (print), 1573-6911 (electronic). URL http://www.springerlink.

com/content/100261/.

Guo:2007:DIM

[GTC^{+}07] Lei Guo, Enhua Tan, Songqing Chen, Zhen Xiao, and Xiaodong Zhang. Does Internet media traffic really follow Zipflike distribution? ACM SIGMETRICS Performance Evaluation Review, 35(1):359–360, June 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Guettler:2012:BGQ

[GTM12] S. Guettler, F. Thiemann, and R. A. E. Mueller. Benfords Gesetz: Ein Qualitätstest für statistische Reihen angewendet auf Handelsdaten für Agrarprodukte. (German) [Benford’s Law: a quality test applied for statistical series on trade data for agricultural products]. In M. Clasen, G. Fröhlich, H. Bernhardt, K. Hildebrand, and B. Theuvsen, editors, Informationstechnologie für eine nachhaltige Landbewirtschaftung. Fokus Forstwirtschaft. Referate der 32. GIL-Jahrestagung, 29. Februar–1. März 2012, Freising. (German) [Information technology for sustainable land management. Focus on forestry. Papers of the 32nd GIL Annual Conference 29, February–1 March 2012, Freising], volume 194 of Lecture Notes in Informatics, pages 111–114. Gesellschaft für Informatik., Bonn, Germany, 2012. ISBN 3-88579-288-5. ISSN 1617-5468. LCCN

???? URL http://www.gil-net.de/Publikationen/24_
111.pdf.

Gunasekera:2015:GIR

[Gun15] Sumith Gunasekera. Generalized inferences of R = Pr(X > Y) for Pareto distribution. Statistical Papers, 56(2):333– 351, May 2015. CODEN STPAE4. ISSN 0932-5026 (print),

1613-9798 (electronic). URL http://link.springer.com/
article/10.1007/s00362-014-0584-8.

Guttag:2016:ICP

[Gut16] John Guttag. Introduction to Computation and Programming Using Python: with Application to Understanding Data. MIT Press, Cambridge, MA, USA, second edition, 2016. ISBN 0-262-52962-9 (paperback), 0-262-33738-X (e-book). xvii + 447 pp. LCCN QA76.73.P98 G88 2016.

Gent:2001:BL

[GW01] Ian Gent and Toby Walsh. Benford’s Law. Report, School of Computer Science, University of St. Andrews, and Department of Computer Science, University of York, St. Andrews, Scotland and York, England, February 2, 2001. 7

pp. URL http://www.dcs.st-and.ac.uk/~apes/reports/
apes-25-2001.pdf.

Geyer:2004:DFD

[GW04] Christina Lynn Geyer and Patricia Pepple Williamson. Detecting fraud in data sets using Benford’s Law. Communications in Statistics: Simulation and Computation, B33(1): 229–246, ???? 2004. CODEN CSSCDB. ISSN 0361-0918.

Huberman:1999:IGD

[HA99] Bernardo A. Huberman and Lada A. Adamic. Internet: Growth dynamics of the World-Wide Web. Nature, 401(6749): 131, September 9, 1999. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Hafner:1969:CSR

[Haf69] Everett M. Hafner. Circular slide roulette. Unpublished but dated 1969. Presented in 1975 in an address at Hampshire College, Amherst, MA, USA, 1969.

Haferkorn:2013:HVA

[Haf13] Martin Haferkorn. Humans vs. algorithms — who follows Newcomb–Benford’s Law better with their order volume? In Fethi Rabhi and Peter Gomber, editors, Enterprise applications and services in the finance industry: 6th International Workshop, FinanceCom 2012, Barcelona, Spain, June 10, 2012. Revised papers, pages 61–70. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 3-642-36219-2. LCCN HG173 .I58

2012. URL http://link.springer.com/content/pdf/10.
1007/978-3-642-36219-4_4.

Haight:1969:TPD

[Hai69] F. A. Haight. Two probability distributions connected with Zipf’s rank-size conjecture. Zastosowania Matematyki, 10: 225–228, 1969. CODEN ZAMTAK. ISSN 0044-1899.

Hamming:1962:NMS

[Ham62] R. W. (Richard Wesley) Hamming. Numerical methods for scientists and engineers. International Series in Pure and Applied Mathematics. McGraw-Hill, New York, NY, USA, 1962. xvii + 411 pp. LCCN QA297 .H28.

Hamming:1970:DN

[Ham70] R. W. Hamming. On the distribution of numbers. The Bell System Technical Journal, 49(8):1609–1625, October 1970. CODEN BSTJAN. ISSN 0005-8580. URL http://bstj.

bell-labs.com/BSTJ/images/Vol49/bstj49-8-1609.pdf;

 http://www.alcatel-lucent.com/bstj/vol49-1970/articles/
bstj49-8-1609.pdf.

Hand:2007:DDD

[Han07] David Hand. Deception and dishonesty with data: Fraud in science. Significance (Oxford, England), 4(1):22–25, March 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Hardy:1949:DS

[Har49] G. H. (Godfrey Harold) Hardy. Divergent series. Clarendon Press, Oxford, UK, 1949. xvi + 396 pp. LCCN QA295 .H29.

Hardy:1991:DS

[Har91] G. H. (Godfrey Harold) Hardy. Divergent series. Chelsea Pub. Co., New York, NY, USA, second edition, 1991. ISBN 0-8284-0334-1. xvi + 396 pp. LCCN QA295 .H29 1991.

Hartshorn:2009:BRB

[Har09] Kevin Hartshorn. Book review: The fabulous Fibonacci numbers by Alfred Posamentier and Ingmar Lehmann. Journal of Mathematics and the Arts, 3(2):113–116, ???? 2009. CODEN ???? ISSN 1751-3472 (print), 1751-3480 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/17513470902897569.

Haspers:1976:YFB

[Has76] Jan H. Haspers. The yield formula and Bradford’s Law. Journal of the American Society for Information Science, 27(5): 281–287, September 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hasan:2002:ADA

[Has02] Bassam Hasan. Assessing data authenticity with Benford’s law. Information Systems Control Journal, 6:??, 2002. CODEN ISYJFS. ISSN 1526-7407. URL http://www.

isaca.org/Journal/Past-Issues/2002/Volume-6/Pages/
Assessing-Data-Authenticity-With-Benfords-Law.aspx.

Hassan:2003:EDA

[Has03] Bassam Hassan. Examining data accuracy and authenticity with leading digit frequency analysis. Industrial Management + Data Systems, 103(2):121–125, ???? 2003. CODEN MDSD8F. ISSN 0263-5577 (print), 1758-5783 (electronic).

Hasumi:2007:ITS

[Has07] Tomohiro Hasumi. Interoccurrence time statistics in the two-dimensional Burridge–Knopoff earthquake model. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 76(??):026117, August 29, 2007. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-

2376 (electronic). URL http://link.aps.org/doi/10.
1103/PhysRevE.76.026117.

Havlin:1995:DBZ

[Hav95] Shlomo Havlin. The distance between Zipf plots. Physica A: Statistical Mechanics and its Applications, 216(1–2):148–150, June 1, 1995. CODEN PHYADX. ISSN 0378-4371 (print),

1873-2119 (electronic). URL http://www.sciencedirect.
com/science/article/pii/037843719500069J.

Havil:2003:GEE

[Hav03] Julian Havil. Gamma: Exploring Euler’s Constant. Princeton University Press, Princeton, NJ, USA, 2003. ISBN 0-691-09983-9. xxiii + 266 pp. LCCN QA41 .H23 2003. US$29.95.

URL http://www.loc.gov/catdir/description/prin031/
2002192453.html.

Hayes:2008:LED

[Hay08a] S. J. Hayes. Letter to the Editor: Does terminal digit preference occur in pathology? Journal of Clinical Pathology, 61 (8):975–976, ???? 2008. CODEN JCPAAK. ISSN 0021-9746 (print), 1472-4146 (electronic). URL http://jcp.bmj.com/

content/61/8/975.2.

Hayes:2008:TDP

[Hay08b] S. J. Hayes. Terminal digit preference occurs in pathology reporting irrespective of patient management implication. Journal of Clinical Pathology, 61(9):1071–1072, ???? 2008. CODEN JCPAAK. ISSN 0021-9746 (print), 1472-

4146 (electronic). URL http://jcp.bmj.com/content/61/
9/1071.extract.

Hayes:2009:CBL

[Hay09] S. J. Hayes. Correspondence: Benford’s Law in relation to terminal digit preference. Journal of Clinical Pathology, 62 (6):574–575, ???? 2009. CODEN JCPAAK. ISSN 0021-9746 (print), 1472-4146 (electronic). URL http://jcp.bmj.com/

content/62/6/575.2.extract.

Hartmann:2018:AVK

[HB18] Sandro Hartmann and Daniel Brinkert. Aufdeckung von Versicherungsbetrug bei Kfz-Schäden mit Hilfe des Benford-Tests. (German) [Detection of insurance fraud for vehicle damage using the Benford test]. Zeitschrift für die gesamte Versicherungswissenschaft, 107(1):41–59, January 2018. ISSN 0044-2585 (print), 1865-9748 (electronic).

Hales:2009:TBL

[HCS09] Douglas N. Hales, Satya S. Chakravorty, and V. Sridharan. Testing Benford’s Law for improving supply chain decisionmaking: A field experiment. International Journal of Production Economics, 122(2):606–618, December 2009. CO-DEN JPCEYE. ISSN 0925-5273 (print), 1873-7579 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0925527309002163.

Hubert:2013:DID

[HDV13] Mia Hubert, Goedele Dierckx, and Dina Vanpaemel. Detecting influential data points for the Hill estimator in Pareto-type distributions. Computational Statistics & Data Analysis, 65(??):13–28, September 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016794731200285X.

Heaps:1978:IRC

[Hea78] Harold Stanley Heaps. Information retrieval, computational and theoretical aspects. Library and information science. Academic Press, New York, USA, 1978. ISBN 0-12-335750-0. xii + 344 pp. LCCN Z699 .H38.

Henriques:2011:WLB

[Hen11] Diana B. Henriques. The wizard of lies: Bernie Madoff and the death of trust. Times Books/Henry Holt, New York, NY, USA, 2011. ISBN 0-8050-9134-3. xxiv + 419 pp. LCCN HV6692.M33 H46 2011.

Herzel:1957:SDD

[Her57] Amato Herzel. Sulla distribuzione della cifre iniziali dei numeri statistici. (Italian) [On the frequency of initial digits of statistical numbers]. Atti dell XV e XVII Riunione, Societa Italiana di Statistica, ??(??):205–228, 1957. CODEN ???? ISSN ????

Herman:1998:TRI

[Her98] Tom Herman. Tax report: An IRS blooper startles thousands of taxpayers. Wall Street Journal, ??(??):1, February 18, 1998. CODEN WSJOAF. ISSN 0099-9660. URL http://online.

wsj.com/article/SB887753996209804500.html.

Herron:2010:EPO

[Her10] Erik S. Herron. The effect of passive observation methods on Azerbaijan’s 2008 presidential election and 2009 referendum. Electoral Studies, 29(3):417–424, 2010. ISSN 0261-3794 (print), 1873-6890 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0261379410000259.

Special Symposium: Voters and Coalition Governments.

Hill:2016:HLI

[HF16] Theodore P. Hill and Ronald F. Fox. Hubble’s Law implies Benford’s Law for distances to galaxies. Journal of Astrophysics and Astronomy, 37(1):??, February 2016. CODEN JASRD7. ISSN 0250-6335 (print), 0973-7758 (electronic). See [AL14b].

Hou:2014:PFD

[HG14] I-Hong Hou and Piyush Gupta. Proportionally fair distributed resource allocation in multiband wireless systems. IEEE/ ACM Transactions on Networking, 22(6):1819–1830, December 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Hofmarcher:2013:FSD

[HH13] P. Hofmarcher and K. Hornik. First significant digits and the credit derivative market during the financial crisis. Contemporary Economics, 7(2):21–29, ????

2013. URL http://papers.ssrn.com/sol3/papers.cfm?
abstract_id=2286063.

Huang:2014:GSP

[HHC^{+}14] Mu-Hsuan Huang, Wei-Tzu Huang, Cheng-Ching Chang, Dar-Zen Chen, and Chang-Pin Lin. The greater scattering phenomenon beyond Bradford’s law in patent citation. Journal of the Association for Information Science and Technology, 65(9):1917–1928, September 2014. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Huettenberger:2014:DSM

[HHG14] Lars Huettenberger, Christian Heine, and Christoph Garth. Decomposition and simplification of multivariate data using Pareto sets. IEEE Transactions on Visualization and Computer Graphics, 20(12):2684–2693, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic),

2160-9306. URL http://csdl.computer.org/csdl/trans/
tg/2014/12/06875963-abs.html.

Hill:1970:ZLP

[Hil70] Bruce M. Hill. Zipf’s law and prior distributions for the composition of a population. Journal of the American Statistical Association, 65(331):1220–1232, September 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284288.

Hill:1974:RFF

[Hil74] Bruce M. Hill. The rank-frequency form of Zipf’s law. Journal of the American Statistical Association, 69(348):1017–1026, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hill:1988:RNG

[Hil88] Theodore Hill. Random number guessing and the first digit phenomenon. Psychological Reports, 62(3):967–971, June 1988. CODEN PYRTAZ. ISSN 0033-2941 (print), 1558-691X

(electronic). URL http://www.amsciepub.com/doi/abs/10.
2466/pr0.1988.62.3.967.

Hill:1995:BII

[Hil95a] Theodore P. Hill. Base-invariance implies Benford’s Law. Proceedings of the American Mathematical Society, 123(3): 887–895, March 1995. CODEN PAMYAR. ISSN 0002-9939 (print), 1088-6826 (electronic). URL http://www.jstor.

org/stable/2160815.

Hill:1995:SDP

[Hil95b] Theodore P. Hill. The significant-digit phenomenon. American Mathematical Monthly, 102(4):322–327, April 1995. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Hill:1995:SDS

[Hil95c] Theodore P. Hill. A statistical derivation of the significantdigit law. Statistical Science, 10(4):354–363, 1995. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL 2-P&origin=MSN;   http://links.jstor.org/sici?
sici=0883-4237(199511)10:4<354:ASDOTS>2.0.CO.

Hill:1996:NDT

[Hil96] Theodore P. Hill. A note on distributions of true versus fabricated data. Preprint, Department of Mathematics, Georgia Institute of Technology, Atlanta, GA, USA,

1996. 4 pp. URL http://digitalcommons.calpoly.edu/
cgi/viewcontent.cgi?article=1046&context=rgp_rsr.

Hill:1997:BL

[Hil97] Theodore Hill. Benford’s Law. In Michiel Hazewinkel, editor, Encyclopedia of Mathematics, Supplement I, volume 1, pages 102–103. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 90-481-4896-0. LCCN ????

Hill:1998:FDP

[Hil98] Theodore P. Hill. The first digit phenomenon. American Scientist, 86(4):358–363, July/August 1998. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (electronic). URL http:

//people.math.gatech.edu/~hill/publications/PAPER%
20PDFS/TheFirstDigitPhenomenonAmericanScientist1996.
pdf;   http://www.math.gatech.edu/~hill/publications/
cv.dir/1st-dig.pdf.

Hill:1999:DFD

[Hil99a] Theodore P. Hill. The difficulty of faking data. Chance: a magazine for people interested in the analysis of data, 12(3): 27–31, ???? 1999. CODEN CNDCE4. ISSN 0933-2480 (print),

1867-2280 (electronic). URL http://amstat.tandfonline.
com/doi/abs/10.1080/09332480.1999.10542154.

Hill:1999:PCS

[Hil99b] Theodore P. Hill. Le premier chiffre significatif fait sa loi. (French) [The first significant digit makes its law]. La Recherche: L’actualité des sciences, 2(316):72–76, January 1999. CODEN RCCHBV. ISSN 0029-5671 (print), 1625-9955

(electronic). URL http://www.larecherche.fr/savoirs/
autre/premier-chiffre-significatif-fait-sa-loi-01-
01-1999-70970.

Hill:2002:RAB

[Hil02] T. P. Hill. Recent applications of Benford’s law. In ????, editor, Conference, Leiden University, page ?? ????, ????, 2002. ISBN ???? LCCN ????

Hill:2011:LEB

[Hil11] Theodore P. Hill. Letter to the editor: Benford’s Law blunders. The American Statistician, 65(2):141, May 2011. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/pdf/10.
1198/tast.2011.11011.

Han:2015:PSS

[HJ15] Qixing Han and Daqing Jiang. Periodic solution for stochastic non-autonomous multispecies Lotka–Volterra mutualism type ecosystem. Applied Mathematics and Computation, 262(??):204–217, July 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0096300315005007.

Hui:2011:BLN

[HJJYM11] Jiang Hui, Shen Jia-Jie, and Zhao Yu-Min. Benford’s Law in nuclear structure physics. Chinese Physics Letters, 28(3): 32101–32104, March 2011. CODEN CPLEEU. ISSN 0256-307X (print), 1741-3540 (electronic).

Horn:2006:DSA

[HKKS06] B. Horn, M. Kreuzer, E. Kochs, and G. Schneider. Different states of anesthesia can be detected by Benford’s Law. Journal of Neurosurgical Anesthesiology, 18(4):328–329, October 2006. ISSN 0898-4921 (print), 1537-1921 (electronic).

Handurukande:2006:PSB

[HKL^{+}06] S. B. Handurukande, A.-M. Kermarrec, F. Le Fessant, L. Massoulié, and S. Patarin. Peer sharing behaviour in the eDonkey network, and implications for the design of server-less file sharing systems. Operating Systems Review, 40(4):359–371, October 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hackl:2014:CPP

[HKWE14] Franz Hackl, Michael E. Kummer, and Rudolf Winter-Ebmer. 99 cent: Price points in e-commerce. Information Economics and Policy, 26(0):12–27, 2014. ISSN 0167-6245 (print), 1873-

5975 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167624513000450.

Hamming:1965:NLB

[HM65] R. W. Hamming and W. L. Mammel. A note on the location of the binary point in a computing machine. IEEE Transactions on Electronic Computers, EC-14(2):260–261, April 1965. CO-DEN IEECA8. ISSN 0367-7508. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=4038414.

Hjorland:2005:BLS

[HN05] Birger Hjørland and Jeppe Nicolaisen. Bradford’s law of scattering: ambiguities in the concept of “subject”. In Fabio Crestani and Ian Ruthven, editors, Context: nature, impact, and role: 5th International Conference on Conceptions of Library and Information Sciences, CoLIS 2005, Glasgow, UK, June 4–8, 2005; proceedings, volume 3507 of Lecture Notes in Computer Science, pages 96–106. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 3-540-26178-8. LCCN Z672.5 .I5616 2005.

Huang:2020:TFL

[HNY20] Yasheng Huang, Zhiyong Niu, and Clair Yang. Testing firmlevel data quality in China against Benford’s law. Economics Letters, 192:109182, 4, 2020. CODEN ECLEDS. ISSN 0165-1765 (print), 1873-7374 (electronic).

Holewijn:1969:UDS

[Hol69] Dr. P. J. Holewijn. On the uniform distribution of sequences of random variables. Zeitschrift für Wahrscheinlichkeitstheorie und verwandte Gebiete, 14(2):89–92, ???? 1969. CODEN ZWVGAA. ISSN 0044-3719. URL http://link.springer.

com/article/10.1007/BF00537514.

Holz:2014:QCG

[Hol14] Carsten A. Holz. The quality of China’s GDP statistics. China Economic Review, 30(0):309–338, 2014. ISSN 1043-951X (print), 1873-7781 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1043951X14000753.

Holt:2018:WEW

[Hol18] Jim Holt. When Einstein walked with Gödel: excursions to the edge of thought. Farrar, Strauss, and Giroux, New York, NY, USA, 2018. ISBN 0-374-14670-5 (hardcover), 0-374-71784-2 (e-book). xi + 368 pp. LCCN PS3608.O4943595 A6 2018.

Hoppe:2016:BLD

[Hop16] Fred M. Hoppe. Benford’s Law and distractors in multiple choice exams. International Journal of Mathematical Education in Science and Technology, 47(4):606–612, 2016. CODEN IJMEBM. ISSN 0020-739x (print), 1464-5211 (electronic).

Horgan:2009:BL

[Hor09a] Jane M. Horgan. Benford’s Law. In Probability with R: an introduction with computer science applications [Hor09b], chapter 9.4, pages 142–144. ISBN 0-470-28073-5 (cloth). LCCN QA76.9.M35 H863 2009. URL http://www.loc.gov/

catdir/toc/ecip0819/2008022817.html;   http://www.
wiley.com/WileyCDA/WileyTitle/productCd-0470280735.
html.

Horgan:2009:PRI

[Hor09b] Jane M. Horgan. Probability with R: an introduction with computer science applications. Wiley, New York, NY, USA, 2009. ISBN 0-470-28073-5 (cloth). xviii + 393 pp. LCCN QA76.9.M35 H863 2009. URL http://www.loc.gov/

catdir/toc/ecip0819/2008022817.html;   http://www.
wiley.com/WileyCDA/WileyTitle/productCd-0470280735.
html.

Hernando:2009:ZLF

[HPV^{+}09] A. Hernando, D. Puigdomènech, D. Villuendas, C. Vesperinas, and A. Plastino. Zipf’s law from a Fisher variationalprinciple. Physics Letters A, 374(1):18–21, December 14, 2009. CODEN PYLAAG. ISSN 0375-9601 (print), 1873-

2429 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0375960109012894.

Hickman:2010:DAC

[HR10] Matthew Hickman and Stephen Rice. Digital analysis of crime statistics: Does crime conform to Benford’s Law? Journal of Quantitative Criminology, 26(3):333–349, September 2010. CODEN JQCRE6. ISSN 0748-4518 (print), 1573-7799 (electronic).

Hoyle01042002

[HRJB02a] David C. Hoyle, Magnus Rattray, Ray Jupp, and Andrew Brass. Making sense of microarray data distributions. Bioinformatics, 18(4):576–584, 2002. ISSN 1367-4803 (print), 1367-4811 (electronic). URL http://bioinformatics.

oxfordjournals.org/content/18/4/576.abstract.

Hoyle:2002:MSM

[HRJB02b] David C. Hoyle, Magnus Rattray, Ray Jupp, and Andrew Brass. Making sense of microarray data distributions. Bioinformatics, 18(4):576–584, April 2002. ISSN 1367-4803 (print), 1367-4811 (electronic).

Huang:2001:PLE

[HS01] Z. F. Huang and S. Solomon. Power, Lévy, exponential and Gaussian-like regimes in autocatalytic financial systems. European Physical Journal B: Condensed Matter and Complex Systems, 20(4):601–607, April 2001. CODEN EPJBFY. ISSN 1434-6028 (print), 1434-6036 (electronic). URL http://link.

springer.com/article/10.1007/PL00011114.

Hill:2005:RDS

[HS05] Theodore P. Hill and Klaus Schürger. Regularity of digits and significant digits of random variables. Stochastic Processes and Their Applications, 115(10):1723–1743, October 2005. CODEN STOPB7. ISSN 0304-4149 (print), 1879-

209X (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304414905000657.

Hefeeda:2008:TMP

[HS08] Mohamed Hefeeda and Osama Saleh. Traffic modeling and proportional partial caching for peer-to-peer systems. IEEE/ ACM Transactions on Networking, 16(6):1447–1460, December 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Henselmann:2012:ABL

[HSD12] Klaus Henselmann, Elisabeth Scherr, and Dominik Ditter. Applying Benford’s Law to individual financial reports: An empirical investigation on the basis of SEC XBRL filings. Working Papers in Accounting Valuation Auditing, 2012-1:

1–23, 2012. URL http://hdl.handle.net/10419/88418;
http://www.econstor.eu/bitstream/10419/88418/1/773918388.
pdf.

Hisano:2011:PVD

[HSM11] Ryohei Hisano, Didier Sornette, and Takayuki Mizuno. Predicted and verified deviations from Zipf’s law in ecology of competing products. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 84(2): 026117, August 24, 2011. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-2376 (electronic). URL http://link.aps.org/

doi/10.1103/PhysRevE.84.026117.

Hullemann:2017:ABL

[HSM17] S. Hüllemann, G. Schüpfer, and J. Mauch. Application of Benford’s law: a valuable tool for detecting scientific papers with fabricated data? Der Anaesthesist, 66(10):795–802, June 2017. ISSN 0003-2417 (print), 1432-055X (electronic).

Chakravorty:2008:TAE

[HSR^{+}08a] Douglas N. Hales, V. Sridharan, Abirami Radhakrishnan, Satya S. Chakravorty, and Samia M. Siha. Testing the accuracy of employee-reported data: An inexpensive alternative approach to traditional methods. European Journal of Operational Research, 189(3):583–593, September 16, 2008. CODEN EJORDT. ISSN 0377-2217 (print), 1872-6860 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0377221706011702.

Hales:2008:TAE

[HSR^{+}08b] Douglas N. Hales, V. Sridharan, Abirami Radhakrishnan, Satya S. Chakravorty, and Samia M. Siha. Testing the accuracy of employee-reported data: An inexpensive alternative approach to traditional methods. European Journal of Operational Research, 189(3):583–593, 2008. CO-DEN EJORDT. ISSN 0377-2217 (print), 1872-6860 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0377221706011702.

Hsu:1948:ESM

[Hsü48] E. H. Hsü. An experimental study on “mental numbers” and a new application. The Journal of General Psychology, 38(1): 57–67, ???? 1948. CODEN JGPSAY. ISSN 0022-1309 (print),

1940-0888 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00221309.1948.9711768.

Harremoes:2001:MEF

[HT01] Peter Harremoës and Flemming Topsøe. Maximum entropy fundamentals. Entropy, 3(3):191–226, September 2001. CO-DEN ENTRFG. ISSN 1099-4300. URL http://www.mdpi.

com/1099-4300/3/3/191.

Harremoes:2005:ZLHa

[HT05a] Peter Harremoës and Flemming Topsoe. Zipf’s law, hyperbolic distributions and entropy loss. Physica A: Statistical Mechanics and its Applications, 328(3–4):315–318, August 1, 2005. CODEN PHYADX. ISSN 0378-4371 (print), 1873-

2119 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1571065305051450.

Harremoes:2005:ZLHb

[HT05b] Peter Harremoës and Flemming Topsoe. Zipf’s law, hyperbolic distributions and entropy loss. In Rudolf Ahlswede et al., editors, General theory of information transfer and combinatorics., volume 21 of Electronic Notes in Discrete Mathematics, pages 315–318. Elsevier, Amsterdam, The Netherlands, 2005.

Harremoes:2006:ZLH

[HT06] P. Harremoës and F. Topsoe. Zipf’s law, hyperbolic distributions and entropy loss. In Rudolf Ahlswede et al., editors, General theory of information transfer and combinatorics, volume 4123 of Lecture Notes in Computer Science, pages 788–792. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006.

Hubert:1976:NIB

[Hub76] John J. Hubert. On the Naranan interpretation of Bradford’s Law. Journal of the American Society for Information Science, 27(5):339–341, September 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hubert:1977:LEL

[Hub77] John J. Hubert. Letters to the Editor: Lotka’s Law in the humanities. Journal of the American Society for Information Science, 28(1):66, January 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hubert:1978:BCR

[Hub78] John J. Hubert. Brief communications: A relationship between two forms of Bradford’s Law. Journal of the American Society for Information Science, 29(3):159–161, May 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Huber:2002:NMG

[Hub02] John C. Huber. A new model that generates Lotka’s Law. Journal of the American Society for Information Science and Technology, 53(3):209–219, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Humenberger:2006:BGV

[Hum06]  Hans Humenberger.  Das ,,Benford-Gesetz” über die

Verteilung der ersten Ziffer von Zahn. (German) [Benford’s law on the distribution of the first digit of numbers]. Report, Institut für Mathematik und Angewandte Statistik, Universität für Bodenkultur, Gregor Mendel-Straße 33, A-1180, Wien, Austria, August 2006. URL http://www.mathematik.

uni-kassel.de/didaktik/HomePersonal/fischer/StochScan/
1996-3_Hum.pdf.

Humenberger:2008:EBB

[Hum08] H. Humenberger. Eine elementarmathematische Begründung des Benford-Gesetzes. (German) [An elementary mathematical derivation of the Benford Law]. Der Mathematikunterricht, 54(??):24–34, ???? 2008. CODEN ???? ISSN 0025-

5807. URL http://www.oemg.ac.at/DK/Didaktikhefte/
2008%2520Band%252041/VortragHumenberger.pdf.

Humenberger:2018:BGW

[Hum18] Hans Humenberger. Das Benford-Gesetz“ — warum ist die Eins als führende Ziffer von Zahlen bevorzugt?. (German) [The “Benford Law” — why is one preferred as the leading digit of numbers?]. In Hans-Stefan Siller, Gilbert Greefrath, and Werner Blum, editors, Neue Materialien für einen realitätsbezogenen Mathematikunterricht 4: 25 Jahre ISTRON-Gruppe - eine Best-of-Auswahl aus der ISTRON-Schriftenreihe, chapter 12, pages 161–176. Springer Fachmedien Wiesbaden, Wiesbaden, 2018. ISBN 978-36581-7-5-9-9-3.

Hungerbuhler:2007:BGF

[Hün07] Norbert Hüngerbühler. Benfords Gesetz über führende Ziffern. Wie die Mathematik Steuersündern das Fürchten lehrt [Benford’s Law on leading digits. mathematics teaches tax evaders to fear]. Report, EducETH, Zürich, Switzerland,

March 2, 2007. 24 pp. URL http://www.educ.ethz.ch/unt/
um/mathe/ana/benford/Benford_Fuehrende_Ziffern.pdf.

Hurlimann:2003:GBL

[Hür03] Werner Hürlimann. A generalized Benford law and its application. Advances and Applications in Statistics, 3(3):217–228, 2003. CODEN ???? ISSN 0972-3617.

Hurlimann:2004:IPB

[Hür04] Werner Hürlimann. Integer powers and Benford’s law. International Journal of Pure and Applied Mathematics, 11(1): 39–46, 2004. CODEN ???? ISSN 1311-8080 (print), 1314-3395 (electronic).

Hurlimann:2006:BLB

[Hür06a] Werner Hürlimann. Benford’s Law from 1881 to 2006: a bibliography. Report, Feldstrasse 145, CH-8004 Zürich, Switzerland, July 5, 2006. 15 pp. URL http://arxiv.org/pdf/

math.ST/0607168.

Hurlimann:2006:GBL

[Hür06b] Werner Hürlimann. Generalizing Benford’s law using power laws: application to integer sequences. ArXiv Mathematics e-prints, July 2006. CODEN ???? ISSN ???? URL http://

adsabs.harvard.edu/abs/2006math......7166H.

Hurlimann:2009:GBL

[Hür09] Werner Hürlimann. Generalizing Benford’s law using power laws: application to integer sequences. International Journal of Mathematics and Mathematical Sciences, page 10, 2009. CODEN ???? ISSN 0161-1712 (print), 1687-0425 (elec-

tronic). URL http://arxiv.org/abs/math/0607166. Article

ID 970284.

Hurlimann:2015:BLS

[Hür15a] Werner Hürlimann. Benford’s Law in scientific research. International Journal of Scientific and Engineering Research, 6 (7):143–148, July 2015. ISSN 2229-5518. URL http://www.

ijser.org/onlineResearchPaperViewer.aspx?Benfords-
Law-in-Scientific-Research.pdf.

Hurlimann:2015:FDT

[Hür15b] Werner Hürlimann. A first digit theorem for powerful integer powers. SpringerPlus, 4(1):576:1–576:??, October 6, 2015. CODEN SPRIFB. ISSN 2193-1801. URL http://link.

springer.com/article/10.1186/s40064-015-1370-3.

Hurlimann:2015:URU

[Hür15c] Werner Hürlimann. On the uniform random upper bound family of first significant digit distributions. Journal of Informetrics, 9(2):349–358, 2015. ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157715000255.

Hobza:2001:NBL

[HV01] Tomáš Hobza and Igor Vajda. On the Newcomb–Benford law in models of statistical data. Revista Matemática Complutense, 14(2):407–420, 2001. CODEN ???? ISSN 1139-1138, 1696-8220, 1988-2807.

Hardy:1975:ITN

[HW75a] Godfrey H. Hardy and Edward M. Wright. An Introduction to the Theory of Numbers. Clarendon Press, Oxford, UK, fourth edition, 1975. ISBN 0-19-853310-1. 421 pp. LCCN ????

Hill:1975:SFZ

[HW75b] Bruce M. Hill and Michael Woodroofe. Stronger forms of Zipf’s law. Journal of the American Statistical Association, 70(349):

212–219, March 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hoare:1986:DFS

[HW86] G. T. Q. Hoare and E. Wright. The distribution of first significant digits. The Mathematical Gazette, 70(451):34–37, March 1986. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL http://www.jstor.org/stable/3615826.

Hosking:1987:PQE

[HW87] J. R. M. Hosking and J. R. Wallis. Parameter and quantile estimation for the generalized Pareto distribution. Technometrics, 29(3):339–349, ???? 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Hu:2015:EES

[HW15] Yangzi Hu and Fuke Wu. Exponential extinction of a stochastic Lotka–Volterra model with expectations in coefficients. IMA Journal of Applied Mathematics, 80(4):1219–1234, August 2015. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-

3634 (electronic). URL http://imamat.oxfordjournals.
org/content/80/4/1219.

Huang:2008:IZL

[HYYH08] Shi-Ming Huang, David C. Yen, Luen-Wei Yang, and Jing-Shiuan Hua. An investigation of Zipf’s Law for fraud detection (DSS# 06-10-1826R(2)). Decision Support Systems, 46(1):70–83, December 2008. CODEN DSSYDK. ISSN 0167-9236 (print), 1873-5797 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167923608001073.

Hein:2012:SFF

[HZKS12] J. Hein, R. Zobrist, C. Konrad, and G. Schuepfer. Scientific fraud in 20 falsified anesthesia papers. Der Anaesthesist, 61(6):543–549, June 2012. CODEN ANATAE. ISSN 0003-2417 (print), 1432-055X (electronic). URL http://link.

springer.com/article/10.1007/s00101-012-2029-x.

He:2014:EPD

[HZZ14] Hui He, Na Zhou, and Ruiming Zhang. On estimation for the Pareto distribution. Statistical Methodology, 21(??):49– 58, November 2014. CODEN ???? ISSN 1572-3127 (print),

1878-0954 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1572312714000288.

Ferrer-i-Cancho:2010:RTD

[iCE10] Ramon Ferrer i Cancho and Brita Elvevåg. Random texts do not exhibit the real Zipf’s Law-like rank distribution. PLoS One, 5(3):e9411:1–e9411:10, March 2010. CODEN POL-

NCL. ISSN 1932-6203. URL http://journals.plos.org/
plosone/article?id=10.1371/journal.pone.0009411.

Ferrer-i-Cancho:2018:OZM

[iCV18] Ramon Ferrer i Cancho and Michael S. Vitevitch. The origins of Zipf’s meaning-frequency law. Journal of the Association for Information Science and Technology, 69(11):1369– 1379, November 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Ipeirotis:2008:CAH

[IG08] Panagiotis G. Ipeirotis and Luis Gravano. Classification-aware hidden-web text database selection. ACM Transactions on Information Systems, 26(2):6:1–6:??, March 2008. CODEN ATISET. ISSN 1046-8188.

Iorliam:2014:DBI

[IHPS14] A. Iorliam, A. T. S. Ho, N. Poh, and Y. Q. Shi. Do biometric images follow Benford’s Law? In 2014 International Workshop on Biometrics and Forensics (IWBF), pages 1–6. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, March 2014.

Iorliam:2017:UBL

[IHWZ17] Aamo Iorliam, Anthony T. S. Ho, Adrian Waller, and Xi Zhao. Using Benford’s Law divergence and neural networks for classification and source identification of biometric images. In Y. Shi, H. Kim, F. Perez-Gonzalez, and F. Liu, editors, Digital Forensics and Watermarking (IWDW 2016), volume 10082 of Lecture Notes in Computer Science, pages 88–105. Springer International Publishing, Cham, Switzerland, 2017.

Israeloff:1996:CZD

[IKC96] N. E. Israeloff, M. Kagalenko, and K. Chan. Can Zipf distinguish language from noise in noncoding DNA? Physical Review Letters, 76(11):1976–??, March 11, 1996. CO-DEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (elec-

tronic), 1092-0145. URL http://link.aps.org/doi/10.
1103/PhysRevLett.76.1976.

Iafrate:2015:EDN

[IMS15] Joseph R. Iafrate, Steven J. Miller, and Frederick W. Strauch. Equipartitions and a distribution for numbers: A statistical model for Benford’s Law. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 91(6):062138:1–062138:6, June 2015. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-2376 (electronic). URL http://link.

aps.org/doi/10.1103/PhysRevE.91.062138.

Ioannides:2003:ZLC

[IO03] Yannis M. Ioannides and Henry G. Overman. Zipf’s law for cities: an empirical examination. Regional Science and Urban Economics, 33(2):127–137, March 1, 2003. CO-DEN RSUEDM. ISSN 0166-0462 (print), 1879-2308 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166046202000066.

Ioannidis:2005:WMP

[Ioa05] John P. A. Ioannidis. Why most published research findings are false. PLoS Medicine, 2(8):696–701, August 2005. CODEN PMLEAC. ISSN 1549-1277 (print), 1549-1676

(electronic). URL http://www.plosmedicine.org/article/
info:doi/10.1371/journal.pmed.0020124.

Irmay:1997:RBZ

[Irm97] Shragga Irmay. The relationship between Zipf’s law and the distribution of first digits. Journal of Applied Statistics, 24(4):383–393, August 1, 1997. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/02664769723594.

Iyengar:1983:SPL

[IRU83] S. Sitharama Iyengar, A. K. Rajagopal, and V. R. R. Uppuluri. String patterns of leading digits. Applied Mathematics and Computation, 12(4):321–337, July 1983. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0096300383900450.

Ismail:2004:SES

[Ism04] Sanaa Ismaïl. A simple estimator for the shape parameter of the Pareto distribution with economics and medical applications. Journal of Applied Statistics, 31(1):3–13, January 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Iudica:2012:BLM

[Iud12] Federico Iudica. Benford’s Law: Mathematical properties and forensic accounting applications. Master’s thesis, Facoltà di Impresa e Management, Cattedra di Matematica, Luiss Guido Carli University, Rome, Italy, 2012. iv + 56 pp. URL http:/

/tesi.eprints.luiss.it/8189/1/iudica-tesi-2012.pdf.

Izsak:2006:MLE

[Izs06] F. Izsák. Maximum likelihood estimation for constrained parameters of multinomial distributions — application to Zipf–Mandelbrot models. Computational Statistics & Data Analysis, 51(3):1575–1583, December 1, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016794730600154X.

Jamain:2001:BL

[Jam01] Adrien Jamain. Benford’s Law. Master’s thesis, Department of Mathematics, Imperial College of London and ENSIMAG, London, UK, 2001. ???? pp. URL http://www.math.

ualberta.ca/~aberger/benford_bibliography/jamain_thesis01.

pdf. Not found in Imperial College Library or COPAC catalogs on 16 February 2013. URL link is broken too.

Janvresse:2012:QCP

[Jan12] Élise Janvresse. Quelques contributions aux probabilités et à la théorie ergodique. (French) [Some contributions to probability theory and ergodic theory]. Document de synthèse présenté pour l’Habilitation à Diriger des Recherches, l’Université de Rouen, Rouen, France, February 17, 2012.

67 pp. URL http://www.univ-rouen.fr/LMRS/Persopage/
Janvresse/hdr_Janvresse.pdf.

Jasak:2010:BZR

[Jas10] Zoran Jasak. Benfordov zakon i reinforcement učenje. (Bosnian) [Benford’s Law and reinforcement learning]. Magistarski Rad (M.Sc. thesis), Prirodno-Matematicki Fakultet, Odsjek za matematiku, Ju Univerzitet u Tuzla, Tuzla, Bosnia and Herzegovina, December 2010. 124 pp.

URL http://www.math.ualberta.ca/~aberger/benford_
bibliography/jasak_10.pdf.

Janvresse:2011:AAU

[JD11] Élise Janvresse and Thierry De La Rue. Averaging along uniform random integers. ArXiv e-prints, 2011. URL http://

arxiv.org/abs/1103.1719.

Janvresse:2012:AAU

[JD12] Élise Janvresse and Thierry De La Rue. Averaging along uniform random integers. Uniform Distribution Theory, 7(2): 35–60, ???? 2012. ISSN 1336-913X. URL http://www.boku.

ac.at/MATH/udt/vol07/no2/03JanRue16-11.pdf.

Janvresse:2004:UDB

[JdlR04] Élise Janvresse and Thierry de la Rue. From uniform distributions to Benford’s law. Journal of Applied Probability, 41(4): 1203–1210, 2004. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://projecteuclid.org/
getRecord?id=euclid.jap/1101840566;   http://www.jstor.
org/stable/4141393;   https://www.univ-rouen.fr/LMRS/
Persopage/Delarue/Publis/PDF/uniform_distribution_
to_Benford_law.pdf.

Janvresse:2009:BL

[JdlR09] Elise Janvresse and Thierry de la Rue. Benford’s law. Butl. Soc. Catalana Mat., 24(1):5–12, 2009. CODEN ???? ISSN 0214-316X. Translated by Frederic Utzet.

Jech:1992:LDL

[Jec92] Thomas Jech. The logarithmic distribution of leading digits and finitely additive measures. Discrete Mathematics, 108(1– 3):53–57, ???? 1992. CODEN DSMHA4. ISSN 0012-365X (print), 1872-681X (electronic).

Jeevanand:1997:BEB

[Jee97] E. S. Jeevanand. Bayes estimation of P (X_{2} < X_{1}) for a bivariate Pareto distribution. Journal of the Royal Statistical Society. Series D (The Statistician), 46(1):93–99, ???? 1997.

CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/2988496.

Jiang:2010:NMS

[JHTS10] Weina Jiang, Anthony T. S. Ho, Helen Treharne, and Yun Q. Shi. A novel multi-size block Benford’s Law scheme for printer identification. Lecture Notes in Computer Science, 6297: 643–652, 2010. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/
content/pdf/10.1007/978-3-642-15702-8_59.

Jin:2009:AUR

[Jin09a] Jang C. Jin. Asian university rankings in international and development economics: An application of Zipf’s Law. Review of International Economics, 17(1):137–143, 2009. ISSN 0965-7576 (print), 1467-9396 (electronic).

Jin:2009:PME

[Jin09b] Jang C. Jin. Publications in mathematical economics and econometrics: ranking of Asian universities and an application of Zipf’s Law. Asian-Pacific Economic Literature, 23(2):116– 122, 2009. ISSN 0818-9935 (print), 1467-8411 (electronic).

Jianu:2021:RFI

[JJ21] Ionel Jianu and Iulia Jianu. Reliability of financial information from the perspective of Benford’s Law. Entropy, 23(5):557–??, May 2021. CODEN ENTRFG. ISSN

????  URL https://dblp.org/db/journals/entropy/
entropy23.html#JianuJ21.

Jelenkovic:2008:CMS

[JK08] Predrag R. Jelenković and Xiaozhu Kang. Characterizing the miss sequence of the LRU cache. ACM SIGMETRICS Performance Evaluation Review, 36(2):119–121, September 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jang:2008:CDH

[JKK^{+}08] Dennis Jang, Jung Uk Kang, Alex Kruckman, Jun Kudo, and Steven J. Miller. Chains of distributions, hierarchical Bayesian models and Benford’s Law. arxiv.org, pages 1–15, May 27, 2008. CODEN ???? ISSN ???? URL

http://adsabs.harvard.edu/abs/2008arXiv0805.4226J;
 http://arxiv.org/abs/0805.4226.

Jang:2009:CDH

[JKK^{+}09] D. Jang, J. U. Kang, A. Kruckman, J. Kudo, and Steven J. Miller. Chains of distributions, hierarchical Bayesian models and Benford’s Law. Journal of Algebra, Number Theory: Advances and Applications, 1(1):37–60, March 2009. CODEN

???? ISSN 0972-5555. URL http://adsabs.harvard.edu/
abs/2008arXiv0805.4226J.

Jager:2008:DAD

[JL08] Hendrik Jager and Pierre Liardet. Distributions arithmétiques des dénominateurs de convergents de fractions continues. (French) [arithmetic distributions of the denominators of continued fractions]. Proceedings of the Koninklijke Nederlandse Akademie van Wetenschappen, Series A, Indagationes mathematicae, 91(2):181–197, June 20, 2008. CO-DEN ???? ISSN 1385-7258 (print), 1878-5972 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S138572588880026X.

Jolion:2001:IBL

[Jol01] Jean-Michel Jolion. Images and Benford’s law. Journal of Mathematical Imaging and Vision: JMIV, 14(1):73–81, 2001. CODEN JMIVEK. ISSN 0924-9907 (print), 1573-7683

(electronic). URL http://link.springer.com/article/10.

1023/A%3A1008363415314. See comments and extensions in [Mil15a, Chapter 19].

Jolissaint:2005:LBR

[Jol05] Paul Jolissaint. Loi de Benford, relations de récurrence et suites équidistribuées. (French) [Benford’s law, recurrence relations, and equidistributed sequences]. Elemente der Mathematik, 60(1):10–18, 2005. CODEN ???? ISSN 0013-6018 (print), 1420-8962 (electronic).

Jolissaint:2009:LBR

[Jol09] Paul Jolissaint. Loi de Benford, relations de récurrence et suites équidistribuées. II. (French) [Benford’s law, recurrence relations, and equidistributed sequences. II]. Elemente der Mathematik, 64(1):21–36, 2009. CODEN ???? ISSN 0013-6018 (print), 1420-8962 (electronic).

Jones:2002:LDR

[Jon02] B. K. Jones. Logarithmic distributions in reliability analysis. Microelectronics and Reliability, 42(4–5):779–786, April/ May 2002. CODEN MCRLAS. ISSN 0026-2714 (print), 1872-

941X (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0026271402000318.

Jansen:2001:RNP

[JP01] C. J. M. Jansen and M. M. W. Pollmann. On round numbers: Pragmatic aspects of numerical expressions. Journal of Quantitative Linguistics, 8(3):187–201, ???? 2001. ISSN 0929-6174 (print), 1744-5035 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1076/jqul.8.3.187.4095.

Judge:2006:DPS

[JS06] George Judge and Laura Schechter. Detecting problems in survey data using Benford’s law. Preprint, University of California at Berkeley and University of Wisconsin at Madison, Berkeley, CA, USA and Madison, WI, USA, November 1,

2006. 30 pp. URL http://www.aae.wisc.edu/lschechter/
benford.pdf.

Judge:2009:DPS

[JS09] G. Judge and L. Schechter. Detecting problems in survey data using Benford’s Law. Journal of Human Resources, 44(1):1–24, Winter 2009. CODEN JHREA9. ISSN 0022-166X (print), 1548-8004 (electronic). URL http://jhr.

uwpress.org/content/44/1/1.abstract;   http://www.
ssc.wisc.edu/jhr/2009ab/judge1.htm.

Jiang:2011:BLN

[JSZ11] Hui Jiang, Jia-Jie Shen, and Yu-Min Zhao. Benford’s law in nuclear structure physics. Chinese Physics Letters, 28(3): 032101, ???? 2011. CODEN CPLEEU. ISSN 0256-307X (print), 1741-3540 (electronic). See also [FNP^{+}15].

Jameson:2014:BLC

[JTY14] Marie Jameson, Jesse Thomer, and Lynelle Ye. Benford’s Law for coefficients of newforms. arxiv.org, ??(??):??, July 7, 2014.

URL http://arxiv.org/abs/1407.1577.

Jameson:2016:BLC

[JTY16] Marie Jameson, Jesse Thorner, and Lynnelle Ye. Benford’s Law for coefficients of newforms. International Journal of Number Theory, 12(2):483–494, March 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https://www.

worldscientific.com/doi/10.1142/S1793042116500299.

Judson:2004:GBF

[Jud04] Horace Freeland Judson. The Great Betrayal: Fraud in Science. Harcourt, Orlando, FL, USA, 2004. ISBN 0-15-100877-

9. xiv + 463 pp. LCCN Q175.37 .J84 2004. URL http:/

/www.loc.gov/catdir/bios/har051/2004005906.html;
http://www.loc.gov/catdir/description/har051/2004005906.
html;   http://www.loc.gov/catdir/samples/har051/2004005906.
html;   http://www.loc.gov/catdir/toc/ecip0416/2004005906.
html.

Kohyama:1984:TIC

[KA84] Tamotsu Kohyama and Yōji Aizawa. Theory of the intermittent chaos. 1/f spectrum and the Pareto–Zipf law. Progress of Theoretical Physics, 71(5):917–929, 1984. CODEN PTPKAV. ISSN 0033-068X (print), 1347-4081 (electronic).

Khataei:2015:PPB

[KA15] Amirsam Khataei and Ali Arya. Personalized presentation builder for persuasive communication. Communication Design Quarterly Review, 3(3):25–32, May 2015. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Kafri:2008:SIS

[Kaf08] Oded Kafri. Sociological inequality and the Second Law.

ArXiv e-prints, 2008. URL http://arxiv.org/abs/0805.
3206.

Kafri:2009:DNE

[Kaf09a] Oded Kafri. The distributions in nature and entropy principle. ArXiv e-prints, July 28, 2009. URL http://arxiv.org/abs/

0907.4852.

Kafri:2009:EPD

[Kaf09b] Oded Kafri. Entropy principle in direct derivation of Benford’s Law. ArXiv e-prints, January 2009. CODEN ???? ISSN ????

URL http://adsabs.harvard.edu/abs/2009arXiv0901.
3047K;   http://arxiv.org/abs/0901.3047.

Kazemitabar:2023:GFC

[Kaz23] Javad Kazemitabar. A general framework for constructing distributions satisfying Benford’s law. Communications in Statistics: Simulation and Computation, 52(12):6160–6167, 2023. CODEN CSSCDB. ISSN 0361-0918.

Kumar:2003:BLA

[KB03] Kuldeep Kumar and Sukanto Bhattacharya. Benford’s law and its application in financial fraud detection. In Cheng-Few Lee, editor, Advances in Financial Planning and Forecasting, volume 11, pages 57–70. Elsevier, Amsterdam, The Netherlands, 2003. ISBN 0-7623-1016-2. ISSN 1046-5847. LCCN ????

Kumar:2007:DDD

[KB07] Kuldeep Kumar and Sukanto Bhattacharya. Detecting the dubious digits: Benford’s law in forensic accounting. Significance (Oxford, England), 4(2):81–83, June 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Kaynar:2010:FSM

[KBHR10] B. Kaynar, A. Berger, T. P. Hill, and A. Ridder. Finitestate Markov chains obey Benford’s Law. ArXiv e-prints,

March 2, 2010. CODEN ????  ISSN ????  URL

http://adsabs.harvard.edu/abs/2010arXiv1003.0562K;
 http://arxiv.org/abs/1003.0562.

Kotz:2004:FDP

[KBRV04] Samuel Kotz, N. Balakrishnan, Campbell B. Read, and Brani Vidakovic. First-digit problem. In Samuel Kotz, N. Balakrishnan, Campbell B. Read, and Brani Vidakovic, editors, Encyclopedia of Statistical Sciences, page ?? Wiley, New York, NY, USA, second edition, 2004. ISBN 0-471-66719-6 (e-book).

Katz:1986:FDP

[KC86] Talbot M. Katz and Daniel I. A. Cohen. The first digit property for exponential sequences is independent of the underlying distribution. Fibonacci Quarterly, 24(1):2–7, February 1986. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/24-1/katz.pdf.

Kim:2023:EFA

[KCCT23] Phuoc Nguyen Kim, Jonatan Contreras, Martine Ceberio, and Nguyen Ngoc Thach. Economic and financial applications of Benford’s Law: from traditional use in audits to help in deep learning. Internat. J. Uncertain. Fuzziness Knowledge-Based Systems, 31:197–207, 2023. ISSN 0218-4885,1793-6411. URL

https://doi.org/10.1142/S0218488523400111.

Kemperman:1975:BDM

[Kem75] J. H. B. Kemperman. Bounds on the discrepancy modulo 1 of a real random variable. Bulletin — Institute of Mathematical Statistics, 4(1):138, ???? 1975. CODEN SMBCVA. ISSN 0146-3942. Abstract no. 75t-47.

Keyser:1936:VFD

[Key36] C. J. Keyser. Vilfredo Federico Damaso Pareto. Mathematician, economist, sociologist. Scripta Mathematica, 4(??):5–23, ???? 1936. ISSN 0036-9713.

Krevitt:1972:BCC

[KG72a] Beth Krevitt and Belver C. Griffith. Brief communications: A comparison of several Zipf-type distributions in their goodness of fit to language data. Journal of the American Society for Information Science, 23(3):220–221, May 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Krevitt:1972:CSZ

[KG72b] Beth Krevitt and Belver C. Griffith. A comparison of several Zipf-type distributions in their goodness of fit to language data. Journal of the American Society for Information Science, 23(3):220–221, 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Koenig:1995:BCL

[KH95] Michael Koenig and Toni Harrell. Brief communications: Lotka’s Law, Price’s urn, and electronic publishing. Journal of the American Society for Information Science, 46(5):386– 388, June 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kawamura:2002:UZL

[KH02] Kenji Kawamura and Naomichi Hatano. Universality of Zipf’s law. Journal of the Physical Society of Japan, 71(5):1211– 1213, 2002. CODEN JUPSAU. ISSN 0031-9015 (print), 1347-4073 (electronic). URL http://jpsj.ipap.jp/.

Kim:2012:BLN

[Kim12] Sangrak Kim. Benford’s Law in non-equilibrium processes: Droplet collisions case. Physica A: Statistical Mechanics and its Applications, 391(20):4970–4975, October 15, 2012. CO-DEN PHYADX. ISSN 0378-4371 (print), 1873-2119 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378437112004141.

Kanter:1995:MPL

[KK95] I. Kanter and D. A. Kessler. Markov processes: Linguistics and Zipf’s Law. Physical Review Letters, 74(??):4559–??, May 29, 1995. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145. URL http://link.aps.org/

doi/10.1103/PhysRevLett.74.4559.

Kazemitabar:2020:MCD

[KK20] Jalil Kazemitabar and Javad Kazemitabar. Measuring the conformity of distributions to Benford’s Law. Communications in Statistics: Theory and Methods, 49(14):3530–3536, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1590599.

Karmeshu:1984:RBL

[KLC84] Karmeshu, N. C. Lind, and V. Cano. Rationales for Bradford’s Law. Scientometrics, 6(4):233–241, July 1984. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02279358.

Klepac:2017:CDS

[Kle17] Goran Klepac. Cognitive data science automatic fraud detection solution, based on Benford’s Law, fuzzy logic with elements of machine learning. In A. Sangaiah, A. Thangavelu, and Sundaram V. Meenakshi, editors, Cognitive Computing for Big Data Systems Over IoT, Lecture Notes on Data Engineering and Communications Technologies, pages 79–95. Springer International Publishing, Cham, Switzerland, December 2017. ISBN 3-319-70687-X (print), 3-319-70688-8 (electronic).

Kondo:2023:HTZ

[KLS23] Illenin O. Kondo, Logan T. Lewis, and Andrea Stella. Heavy tailed but not Zipf: Firm and establishment size in the United States. Journal of Applied Econometrics, 38(5):767–785, August 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kontorovich:2005:BLV

[KM05] Alex V. Kontorovich and Steven J. Miller. Benford’s law, values of L-functions and the 3x + 1 problem. Acta Arithmetica, 120(3):269–297, 2005. CODEN AARIA9. ISSN 0065-1036 (print), 1730-6264 (electronic). URL

http://adsabs.harvard.edu/abs/2005AcAri.120..269K;
 http://arxiv.org/abs/math/0412003.

Kalinin:2011:UEF

[KM11] K. Kalinin and Walter R. Mebane, Jr. Understanding electoral frauds through evolution of Russian federalism: From “bargaining loyalty” to “signaling loyalty”. In ????, editor, Proceedings of the 2011 Annual Meeting of the Midwest Political Science Association, Chicago, IL, March 31–April 2, 2011, page ?? ????, ????, 2011.

Kim:2009:ITA

[KMN09] Hyo-Jeong Kim, Michael Mannino, and Robert J. Nieschwietz. Information technology acceptance in the internal audit profession: Impact of technology features and complexity. International Journal of Accounting Information Systems, 10 (4):214–228, 2009. CODEN IJAIA7. ISSN 1467-0895 (print),

1873-4723 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1467089509000360. Ninth In-

ternational Research Symposium on Accounting Information Systems (IRSAIS).

Kuipers:1974:UDS

[KN74] Lauwerens Kuipers and Harald Niederreiter. Uniform distribution of sequences. Pure and applied mathematics. Wiley, New York, NY, USA, 1974. ISBN 0-471-51045-9. xiv + 390 pp. LCCN QA292 .K84.

Katsikas:1990:CDG

[KN90] A. A. Katsikas and J. S. Nicolis. Chaotic dynamics of generating Markov partitions and linguistic sequences mimicking Zipf’s law. Nuovo Cimento D (1), 12(2):177–195, 1990. CODEN NCSDDN. ISSN 0392-6737 (print), 1826-9893 (electronic).

Katsikas:1991:CDG

[KN91] A. A. Katsikas and J. S. Nicolis. Chaotic dynamics of generating Markov partitions, and linguistic sequences mimicking Zipf’s law. Lecture Notes in Computer Science, 565:335–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kuipers:2006:UDS

[KN06] Lauwerens Kuipers and Harald Niederreiter. Uniform distribution of sequences. Dover Publications, Inc., New York, NY, USA, 2006. ISBN 0-486-45019-8 (paperback). xviii + 390 pp. LCCN QA292 .K84 2006. URL http://

www.loc.gov/catdir/enhancements/fy0625/2005056064-
d.html;   http://www.loc.gov/catdir/toc/fy0612/2005056064.
html.

Kanemitsu:1988:BLF

[KNRS88] Shigeru Kanemitsu, Kenji Nagasaka, Gérard Rauzy, and Jau-Shyong Shiue. On Benford’s law: the first digit problem. In Probability theory and mathematical statistics (Kyoto, 1986), volume 1299 of Lecture Notes in Math., pages 158–169. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1988.

Kitsuregawa:1989:EBS

[KNT89] Masaru Kitsuregawa, Masaya Nakayama, and Mikio Takagi. The effect of bucket size tuning in the dynamic hybrid GRACE hash join method. In Apers and Wiederhold [AW89], pages 257–266. CODEN VLDBDP. ISBN 1-55860-101-5. LCCN QA 76.9 D3 I61 1989.

Knuth:1969:SA

[Knu69] Donald E. Knuth. Seminumerical Algorithms, volume 2 of The Art of Computer Programming. Addison-Wesley, Reading, MA, USA, 1969. ISBN 0-201-03802-1. xi + 624 pp. LCCN QA76.5 .K57. US$19.75.

Knuth:1998:SA

[Knu98] Donald E. Knuth. Seminumerical Algorithms, volume 2 of The Art of Computer Programming. Addison-Wesley, Reading, MA, USA, third edition, 1998. ISBN 0-201-89684-2. xiii + 762 pp. LCCN QA76.6 .K64 1997. US$52.75. See section 4.2.4V, The Fraction Parts, pages 254–262, for a discussion of Benford’s Law.

Knudsen:2001:ZLC

[Knu01] Thorbjørn Knudsen. Zipf’s law for cities and beyond: the case of Denmark. The American Journal of Economics and Sociology, 60(1):123–146, January 2001. CODEN AJESA3. ISSN 0002-9246 (print), 1536-7150 (electronic). URL http:/

/www.jstor.org/stable/3487947.

Koch:2020:BLC

[KO20] Christoffer Koch and Ken Okamura. Benford’s Law and COVID-19 reporting. Economics Letters, 196:109573, November 2020. CODEN ECLEDS. ISSN 0165-1765 (print), 1873-7374 (electronic).

Konheim:1965:MDT

[Kon65] Alan G. Konheim. Mantissa distributions (in Technical Notes and Short Papers). Mathematics of Computation, 19(89):143– 144, April 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kossovsky:2006:TBU

[Kos06] Alex Ely Kossovsky. Towards a better understanding of the leading digits phenomena. ArXiv Mathematics e-prints, 2006.

URL http://arxiv.org/abs/math/0612627.

Kossovsky:2012:SNR

[Kos12] Alex Ely Kossovsky. Statistician’s new role as a detective — testing data for fraud. Ciencias Económica, 30(2): 179–200, ???? 2012. ISSN 0252-9521 (print), 2215-3489

(electronic). URL http://revistas.ucr.ac.cr/index.php/
economicas/article/view/8015/7634.

Kossovsky:2013:RQO

[Kos13] A. Ely Kossovsky. On the relative quantities occurring within physical data sets. arxiv.org, May 2013. URL http://

adsabs.harvard.edu/abs/2013arXiv1305.1893E.

Kossovsky:2015:BLT

[Kos15] Alex Ely Kossovsky. Benford’s Law: theory, the general law of relative quantities, and forensic fraud detection applications. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2015. ISBN 981-4583-68-5 (hardcover), 981-4583-69-3 (e-book). xxi + 649 pp. LCCN HV8079.F7 K67 2015.

Kulikova:2004:OSS

[KP04] A. A. Kulikova and Yu. V. Prokhorov. One-sided stable distributions and Benford’s law. Teor. Veroyatn. Primen., 49(1): 178–184, 2004. CODEN ???? ISSN 0040-361X.

Kulikova:2005:CAS

[KP05] A. A. Kulikova and Yu. V. Prokhorov. Completely asymmetric stable laws and Benford’s Law. Theory of Probability and its Applications, 49(1):163–169, March 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:/

/epubs.siam.org/sam-bin/dbq/article/98094.

Krishna:2009:DBD

[KP09] Hare Krishna and Pramendra Singh Pundir. Discrete Burr and discrete Pareto distributions. Statistical Methodology, 6(2):177–188, March 2009. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S157231270800052X.

Kozubowski:2015:DTP

[KPF15] Tomasz J. Kozubowski, Anna K. Panorska, and Matthew L. Forister. A discrete truncated Pareto distribution. Statistical Methodology, 26(??):135–150, September 2015. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1572312715000313.

Klar:2010:ZLL

[KPH10] B. Klar, P. R. Parthasarathy, and N. Henze. Zipf and Lerch limit of birth and death processes. Probability in the Engineering and Informational Sciences, 24(1):129–144, 2010. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic).

Kulikova:2005:HFD

[KPK05] A. A. Kulikova, Yu. V. Prokhorov, and V. I. Khokhlov. H. F. D. (H-function distribution) and the Benford law. I. Teor. Veroyatn. Primen., 50(2):366–371, 2005. CODEN ???? ISSN 0040-361X.

Kulikova:2006:HFD

[KPK06] A. A. Kulikova, Yu. V. Prokhorov, and V. I. Khokhlov. H.F.D. (H-function distribution) and Benford’s Law. I. Theory of Probability and its Applications, 50(2):311–315, January 2006. CODEN TPRBAU. ISSN 0040-585X (print),

1095-7219 (electronic). URL http://epubs.siam.org/tvp/
resource/1/tprbau/v50/i2/p311_s1.

Kennard:1981:MDF

[KR81] Robert W. Kennard and John E. Reith. M32. On the distribution of first digits. Communications in Statistics: Simulation and Computation, 10(1):97–98, 1981. CODEN CSSCDB.

ISSN 0361-0918. URL http://www.tandfonline.com/doi/
abs/10.1080/03610918108812195.

Kretschmer:2001:RAI

[KR01] Hildrun Kretschmer and Ronald Rousseau. Research: Author inflation leads to a breakdown of Lotka’s Law. Journal of the American Society for Information Science and Technology, 52 (8):610–614, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kollath-Romano:2004:DSD

[KR04] P. Kollath-Romano. On the distribution of significant digits in numbers. Preprint, Rensselaer Polytechnic Institute, Troy, NY, USA, 2004.

Khosravani:2012:TIB

[KR12] A. Khosravani and C. Rasinariu. Transformation invariance of Benford variables and their numerical modeling. In Niola et al. [NKZ12], pages 57–61. ISBN 1-61804-080-4 (paper-

back). LCCN ???? URL http://www.wseas.us/e-library/
conferences/2012/SaintMalo/ACMIN/ACMIN-06.pdf.

Khosravani:2013:DBD

[KR13] Azar Khosravani and Constantin Rasinariu. n-digit Benford distributed random variables. Advances and Applications in Statistics, 36(2):119–130, October 2013. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/7887.
htm.

Khosravani:2015:DBC

[KR15] Azar Khosravani and Constantin Rasinariu. n-digit Benford converges to Benford. International Journal of Mathematics and Mathematical Sciences, pages Art. ID 123816, 4, 2015. ISSN 0161-1712 (print), 1687-0425 (electronic).

Kramer:2015:MJE

[Krä15] Walter Krämer. Miller, S. J. (ed.): Benford’s law. Theory and applications. Statistical Papers, 56(4):1251–1252, November 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s00362-015-0717-8.

Kreifelts:1973:OBG

[Kre73] Thomas Kreifelts. Optimale Basiswahl für eine Gleitkomma-Arithmetik. (German) [Optimal choice of basis for a floatingpoint arithmetic]. Computing: Archiv für Informatik und Numerik, 11(4):353–363, December 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See correction [Kre75].

Kreifelts:1975:OBF

[Kre75] T. Kreifelts. Optimale Basiswahl für eine Gleitkomma-Arithmetik (Berichtigung) (German) Optimal Basis Choice for a Floating-Point Arithmetic (Correction). Computing: Archiv für Informatik und Numerik, 14(3):313–314, September 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See [Kre73].

Kreiner:2002:FDL

[Kre02] W. A. Kreiner. First digit law. Preprint, Abteilung Chemische Physik, Arbeitsgruppe Laseranwendungen, Universität Ulm, Ulm, Germany, May 3, 2002. 7 pp. URL http://vts.uni-

ulm.de/docs/2002/1441/vts_1441.pdf.

Kreiner:2003:NBL

[Kre03] W. A. Kreiner. On the Newcomb–Benford law. Zeitschrift für Naturforschung, 58a(11):618–622, ???? 2003. CODEN

ZNTFA2. ISSN 0372-9516. URL http://www.znaturforsch.
com/aa/v58a/s58a0618.pdf.

Krisciunas:1977:LEL

[Kri77] Kevin Krisciunas. Letters to the Editor: Lotka’s Law — year by year. Journal of the American Society for Information Science, 28(1):65–66, January 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kriel:2010:DDM

[Kri10] Eckhardt Kriel. Discussion of “Data mining journal entries for fraud detection: An exploratory study”. International Journal of Accounting Information Systems, 11(3):186–188, September 2010. CODEN IJAIA7. ISSN 1467-0895 (print), 1873-

4723 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1467089510000400. See [DG10,

Gra10].

Krugman:1996:CMU

[Kru96a] Paul Krugman. Confronting the mystery of urban hierarchy. Journal of the Japanese and International Economies, 10(4):399–418, December 1996. CODEN JJIEBD. ISSN 0889-1583 (print), 1095-8681 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0889158396900234.

Krugman:1996:SOE

[Kru96b] Paul R. Krugman. The Self-Organizing Economy. Blackwell Publishers, Cambridge, MA, USA, 1996. ISBN 1-55786-699-6 (paperback). vi + 122 pp. LCCN HB199 .K75 1996.

Krygier:2009:PRB

[Kry09] J. Krygier. Psychological relevance of Benford’s Law: Sensitivity to statistical regularities and implications for theories of number representation. Honours thesis, University of Sydney, Sydney, NSW, Australia, 2009.

Kibria:2010:PTE

[KS10] B. M. Golam Kibria and A. K. Md. E. Saleh. Preliminary test estimation of the parameters of exponential and Pareto distributions for censored samples. Statistical Papers, 51(4): 757–773, December 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http://link.

springer.com/article/10.1007/s00362-008-0163-y.

Karthik:2016:ETS

[KSG^{+}16] Deepak Karthik, Gil Stelzer, Sivan Gershanov, Danny Baranes, and Mali Salmon-Divon. Elucidating tissue specific genes using the Benford distribution. BMC Genomics, 17(1), August 2016. ISSN 1471-2164.

Kim:2011:ZAS

[KSH11] SeongKi Kim, HaYoon Song, and SangYong Han. ZipfAllocation: an algorithm for static allocation of movies in a cluster of video servers. Software — Practice and Experience, 41 (6):695–716, May 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kuipers:1973:RPD

[KsS73] L. Kuipers and Jau shyong Shiue. Remark on a paper by Duncan and Brown on the sequence of logarithms of certain recursive sequences. Fibonacci Quarterly, 11(3):292–293, October 1973. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/11-3/kuipers.pdf. See [BD70].

Kubovy:1977:RAA

[Kub77] M. Kubovy. Response availability and the apparent spontaneity of numerical choices. Journal of Experimental Psychology: Human Performance and Performance, 3(2): 359–364, May 1977. ISSN 0096-1523 (print), 1939-1277

(electronic). URL http://psycnet.apa.org/index.cfm?fa=
search.displayrecord&uid=1977-30946-001.

Kuipers:1969:RPR

[Kui69] L. Kuipers. Remark on a paper by R. L. Duncan concerning the uniform distribution mod 1 of the sequence of the logarithms of the Fibonacci numbers. Fibonacci Quarterly, 7(5):465–466, December 1969. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/7-
5/kuipers-a.pdf.

Kunoff:1987:FDP

[Kun87] Sharon Kunoff. N! has the first digit property. Fibonacci Quarterly, 25(4):365–367, November 1987. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/
25-4/kunoff.pdf.

Kechedzhi:2005:RDW

[KUY05] K. E. Kechedzhi, O. V. Usatenko, and V. A. Yampol′skii. Rank distributions of words in correlated symbolic systems and the Zipf law. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 72(4): 046138, October 28, 2005. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-2376 (electronic). URL http://link.aps.org/

doi/10.1103/PhysRevE.72.046138.

Kraus:2014:DWD

[KV14] Cornelia Kraus and Raul Valverde. A data warehouse design for the detection of fraud in the supply chain by using the Benford’s Law. American Journal of Applied Sciences, 11(9): 1507–1518, September 2014. CODEN AJASCJ. ISSN 1546-9239 (print), 1554-3641 (electronic).

Krakar:2008:ABL

[KŽ08a] Zdravko Krakar and Mario Žgela. Application of Benford’s Law in payment system auditing. In Boris Aurer and Miroslav Baća, editors, Proceedings of the 19th Central European Conference on Information and Intelligent Systems (CE-CIS 2008), pages 187–193. Fakultet organizacije i informatike, Varaždin, Croatia, 2008.

Krakar:2008:EBL

[KŽ08b] Zdravko Krakar and Mario Žgela. Evaluation of Benford’s Law: application in stock prices and stock turnover. In Mario Plenković, editor, Society and Technology 2008, pages 57– 71. Hrvatsko komunikološko društvo, Zagreb, Croatia, 2008. ISBN ???? LCCN ????

Kudryashov:2015:APE

[KZ15] Nikolay A. Kudryashov and Anastasia S. Zakharchenko. Analytical properties and exact solutions of the Lotka– Volterra competition system. Applied Mathematics and Computation, 254(??):219–228, March 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0096300314017743.

Lacasa:2018:NBL

[Lac18] Lucas Lacasa. Newcomb–Benford law helps customs officers to detect fraud in international trade. Proceedings of the National Academy of Sciences of the United States of America, 116(1):11–13, December 2018. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic). URL https://www.

pnas.org/cgi/doi/10.1073/pnas.1819470116.

Lu:2014:ODE

[LACL14] Lu Lu, Christine M. Anderson-Cook, and Dennis K. J. Lin. Optimal designed experiments using a Pareto front search for focused preference of multiple objectives. Computational Statistics & Data Analysis, 71(??):1178–1192, March 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167947313001382.

Lagarias:1985:PG

[Lag85] Jeffrey C. Lagarias. The 3x + 1 problem and its generalizations. American Mathematical Monthly, 92(1):3–23, January 1985. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lagarias:2003:PAB

[Lag03] Jeffrey C. Lagarias. The 3x + 1 problem: An annotated bibliography (1963–1999) (sorted by author). arxiv.org, ??

(??), September 2003. URL http://arxiv.org/abs/math/
0309224.

Lagarias:2010:UCP

[Lag10] Jeffrey C. Lagarias, editor. The ultimate challenge: the 3x + 1 problem. American Mathematical Society, Providence, RI, USA, 2010. ISBN 0-8218-4940-9. xiv + 344 pp. LCCN QA292 .U48 2010.

Lafouge:2015:DTP

[LAL15] Thierry Lafouge, Abdellatif Agouzal, and Genevieve Lallich. The deconstruction of a text: the permanence of the generalized Zipf law — the inter-textual relationship between entropy and effort amount. Scientometrics, 104(1):193–217, July 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/s11192-015-1600-z.

Lanza:2000:UDA

[Lan00] R. B. Lanza. Using digital analysis to detect fraud: Review of the DATAS(R) statistical analysis tool. Journal of Forensic Accounting, 1(??):291–296, ???? 2000. CODEN

???? ISSN 1524-5586. URL http://www.rtedwards.com/
journals/JFA/.

Lanza:2003:FSP

[Lan03] Richard B. Lanza. Fear not the software: Proactively detecting occupational fraud using computer audit reports. Fraud Magazine, ??(??):??, September/February 2003. ISSN 1553-

6645. URL http://www.fraud-magazine.com/article.
aspx?id=4294967868.

Lanza:2004:CBS

[Lan04] Richard B. Lanza. Comparing best software for fraud examinations: Fraud data interrogation tools. Fraud Magazine, ??(??):??, November/December 2004. ISSN 1553-6645.

URL http://www.fraud-magazine.com/article.aspx?id=
4294967837.

Lu:2005:DFH

[LB05] Fletcher Lu and J. Efrim Boritz. Detecting fraud in health insurance data: Learning to model incomplete Benford’s Law distributions. In Gama et al. [GCB^{+}05], pages 633– 640. ISBN 3-540-29243-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q325.5 .E26 2005. URL http://link.

springer.com/chapter/10.1007/11564096_63.

Leydesdorff:2006:CPL

[LB06a] Loet Leydesdorff and Stephen Bensman. Classification and power laws: the logarithmic transformation. Journal of the American Society for Information Science and Technology, 57 (11):1470–1486, September 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lu:2006:BCF

[LB06b] F. Lu and J. E. Boritz. Building a case for fraud using Benford’s law with reinforcement learning. In ????, editor, Uncertainty in Artificial Intelligence: Proceedings of the 22nd Conference, page ?? ????, ????, 2006.

Loetscher:2007:ENS

[LB07] T. Loetscher and P. Brugger. Exploring number space by random digit generation. Experimental Brain Research, 180 (4):655–665, July 2007. CODEN EXBRAP. ISSN 0014-4819 (print), 1432-1106 (electronic). URL http://link.

springer.com/article/10.1007/s00221-007-0889-0.

Lansey:2009:ISR

[LB09] Jonathan C. Lansey and Bruce Bukiet. Internet search result probabilities: Heaps’ Law and word associativity. Journal of Quantitative Linguistics, 16(1):40–66, ???? 2009. CODEN ???? ISSN 0929-6174 (print), 1744-5035 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09296170802514153.

Leemann:2014:SAS

[LB14] Lucas Leemann and Daniel Bochsler. A systematic approach to study electoral fraud. Electoral Studies, 35(0): 33–47, 2014. ISSN 0261-3794 (print), 1873-6890 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0261379414000390.

Lu:2006:AFD

[LBC06] Fletcher Lu, J. Efrim Boritz, and Dominic Covvey. Adaptive fraud detection using Benford’s law. In Advances in artificial intelligence, volume 4013 of Lecture Notes in Computer Science, pages 347–358. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006.

Levene:2001:ZLW

[LBL01] Mark Levene, José Borges, and George Loizou. Zipf’s law for Web surfers. Knowledge and Information Systems, 3(1): 120–129, 2001. CODEN KISNCR. ISSN 0219-1377.

Liu:2015:DSD

[LC15] Qun Liu and Qingmei Chen. Dynamics of stochastic delay Lotka–Volterra systems with impulsive toxicant input and Lévy noise in polluted environments. Applied Mathematics and Computation, 256(??):52–67, April 1, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0096300315000235.

Lu:2005:CZL

[LCC^{+}05] Tim Lu, Christine M. Costello, Peter J. P. Croucher, Robert Häsler, Günther Deuschl, and Stefan Schreiber. Can Zipf’s law be adapted to normalize microarrays? BMC Bioinformatics, 6(??):37–49, ???? 2005. CODEN BBMIC4. ISSN 1471-2105.

URL http://link.springer.com/article/10.1186/1471-
2105-6-37;   http://www.biomedcentral.com/1471-2105/
6/37.

Lee:2010:SAR

[LCJ10] Joanne Lee, Wendy K. Tam Cho, and George G. Judge. Stigler’s approach to recovering the distribution of first significant digits in natural data sets. Statistics & Probability Letters, 80(2):82–88, January 15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715209003563.

Lee:2015:GBL

[LCJ15] Joanne Lee, Wendy K. Tam Cho, and George Judge. Generalizing Benford’s Law. In Miller [Mil15a], pages 304–316. ISBN 0-691-14761-2 (hardcover), 1-4008-6659-6 (e-book). LCCN

QA273.6 .B46 2015. URL http://press.princeton.edu/
titles/10527.html.

Liu:2015:SPA

[LCY^{+}15] Si Liu, Qiang Chen, Shuicheng Yan, Changsheng Xu, and Hanqing Lu. Snap & Play: Auto-generated personalized findthe-difference game. ACM Transactions on Intelligent Systems and Technology (TIST), 5(4):65:1–65:??, January 2015. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Lehner:2004:DVG

[LD04] B. Lehner and P. Doll. Development and validation of a global database of lakes. Journal of Hydrology, 296(1–4):1–22, August 20, 2004. CODEN JHYDA7. ISSN 0022-1694 (print),

1879-2707 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022169404001404.

Lang:1989:UAB

[LDJ89] Sheau-Dong Lang, James R. Driscoll, and Jiann H. Jou. A unified analysis of batched searching of sequential and treestructured files. ACM Transactions on Database Systems, 14(4):604–618, December 1989. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

www.acm.org/pubs/articles/journals/tods/1989-14-4/

p604-lang/p604-lang.pdf;   http://www.acm.org/pubs/
citations/journals/tods/1989-14-4/p604-lang/;   http:
//www.acm.org/pubs/toc/Abstracts/tods/76908.html.

Lewis:2015:SDA

[LEC^{+}15] Peter R. Lewis, Lukas Esterle, Arjun Chandra, Bernhard Rinner, Jim Torresen, and Xin Yao. Static, dynamic, and adaptive heterogeneity in distributed smart camera networks. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 10(2):8:1–8:??, June 2015. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Leemis:2015:BLG

[Lee15] Lawrence Leemis. Benford’s Law geometry. In Miller [Mil15a], pages 109–118. ISBN 0-691-14761-2 (hardcover), 1-4008-6659-6 (e-book). LCCN QA273.6 .B46 2015. URL http://press.

princeton.edu/titles/10527.html.

Lee:2018:PIB

[Lee18] Edward D. Lee. Partisan intuition belies strong, institutional consensus and wide Zipf’s Law for voting blocs in US Supreme Court. Journal of Statistical Physics, 173(6):1722–1733, December 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Lefort:1982:LOA

[Lef82] B. Lefort. L’emploi des outils au cours de taches d’entretien et la loi de Zipf–Mandelbrot. (French) [the use of tools in maintenance tasks and the Zipf–Mandelbrot Law]. Le Travail Humain, 45(2):307–316, ???? 1982. CODEN TRHUAH. ISSN 0041-1868 (print), 2104-3663 (electronic). URL http://www.

jstor.org/stable/40657805.

Leimkuhler:1980:EFB

[Lei80] Ferdinand F. Leimkuhler. An exact formulation of Bradford’s Law. Journal of Documentation, 36(4):285–292, ???? 1980. CODEN JDOCAS. ISSN 0022-0418 (print), 1758-7379 (electronic).

Lemons:1986:NTD

[Lem86] Don S. Lemons. On the number of things and the distribution of first digits. American Journal of Physics, 54(9):816–817, September 1986. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://ajp.aapt.org/.

Ley:1996:PDU

[Ley96] Eduardo Ley. On the peculiar distribution of the U.S. stock indexes’ digits. The American Statistician, 50(4):311– 313, November 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684926.

Li:2016:QNS

[LF16] Qinglei Li and Zuntao Fu. Quantifying non-stationarity effects on organization of atmospheric turbulent eddy motion by Benford’s law. Communications in Nonlinear Science and Numerical Simulation, 33(??):91–98, April 2016. ISSN 1007-5704 (print), 1878-7274 (electronic).

Lindsay:2004:DFD

[LFCR04] D. H. Lindsay, P. S. Foote, A. Campbell, and D. R. Reilly. Detecting fraud in the data using automatic intervention detection. Fraud Magazine, ??(??):??, January/February 2004. ISSN 1553-6645.

Li:2015:BBL

[LFY15] Qinglei Li, Zuntao Fu, and Naiming Yuan. Beyond Benford’s Law: Distinguishing noise from chaos. PLoS One, 10 (6):e0129161:1–e0129161:11, June 1, 2015. CODEN POL-

NCL. ISSN 1932-6203. URL http://journals.plos.org/
plosone/article?id=10.1371/journal.pone.0129161.

Logan:1978:FDP

[LG78] Jonothan L. Logan and Samuel A. Goudsmit. The first digit phenomenon. Proceedings of the American Philosophical Society, 122(4):193–197, August 18, 1978. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http:/

/links.jstor.org/sici?sici=0003-049X(19780818)122%
3A4%3C193%3ATFDP%3E2.0.CO%3B2-C; http://www.jstor.

org/stable/986530. This paper contains derivations of both Stigler’s Law and Benford’s Law, and receives strong criticism in [Rai85]. This paper contains an important historical note that is recorded in entry [Ben38].

Lu:2010:BLP

[LG10] Ocean Fan Lu and David E. Giles. Benford’s Law and psychological barriers in certain eBay auctions. Applied Economics Letters, 17(10):1005–1008, ???? 2010. CODEN ???? ISSN 1350-4851 (print), 1466-4291 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/13504850802631814.

Lee:2020:CFC

[LHJ20] Kang-Bok Lee, Sumin Han, and Yeasung Jeong. COVID-19, flattening the curve, and Benford’s law. Physica A: Statistical Mechanics and its Applications, 559:125090, December 2020. CODEN PHYADX. ISSN 0378-4371 (print), 1873-2119 (electronic).

Lin:2015:DES

[LHK^{+}15] Cheng-Yen Lin, Chung-Wen Huang, Chi-Bang Kuan, Shi-Yu Huang, and Jenq-Kuen Lee. The design and experiments of a SID-based power-aware simulator for embedded multicore systems. ACM Transactions on Design Automation of Electronic Systems (TODAES), 20(2):22:1–22:??, February 2015. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Li:1992:RTE

[Li92] W. Li. Random texts exhibit Zipf’s-law-like word frequency distribution. IEEE Transactions on Information Theory, 38 (6):1842–1845, ???? 1992. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Li:2011:ZLB

[Li11] Wentian Li. Zipf’s Law bibliography. Web site., August

2011. URL http://www.nslij-genetics.org/wli/zipf/.

The site holds 870 references (in untagged reference-list form) up to 11 August 2011.

Lim:2006:BRB

[Lim06] Nena Lim. Book review: The new forensics: investigating corporate fraud and the theft of intellectual property, J. Anastasi. John Wiley & Sons, New Jersey (2003). Digital Investigation, 3(4):245–246, 2006. ISSN 1742-2876 (print), 1873-

202X (electronic). URL http://www.sciencedirect.com/
science/article/pii/S174228760600106X.

Lines:1986:NYT

[Lin86] Malcolm E. Lines. A number for your thoughts: facts and speculations about numbers from Euclid to the latest computers. Adam Hilger Ltd., Bristol, UK, 1986. ISBN 0-85274-495-1 (paperback). vi + 214 pp. LCCN QA241 .L617 1986.

URL http://www.loc.gov/catdir/enhancements/fy0668/
86183277-d.html.

Linville:2008:PFN

[Lin08] Mark Linville. The problem of false negative results in the use of digit analysis. Journal of Applied Business Research, 24 (1):17–25, First Quarter 2008. ISSN 0892-7626 (print), 2157-

8834 (electronic). URL http://www.cluteinstitute.com/
ojs/index.php/JABR/article/view/1363.

Lindstrom:2020:VRC

[Lin20] Peter Lindstrom. Variable-radix coding of the reals. In Cornea et al. [CLT20], pages 111–116. ISBN 1-72817-120-2, 1-72817-121-0. ISSN 2576-2265. LCCN ???? URL https://

ieeexplore.ieee.org/servlet/opac?punumber=9146973.

Lipovetsky:2009:PLD

[Lip09] Stan Lipovetsky. Pareto 80/20 law: derivation via random partitioning. International Journal of Mathematical Education in Science and Technology, 40(2):271–277, ???? 2009. CODEN IJMEBM. ISSN 0020-739X (print), 1464-5211 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00207390802213609.

Liu:2008:MNS

[Liu08] Cheng-Shi Liu. Maximal non-symmetric entropy leads naturally to Zipf’s law. Fractals, 16(1):99–101, 2008. CODEN FRACEG. ISSN 0218-348X.

Liu:2015:ASN

[Liu15] Qun Liu. Analysis of a stochastic non-autonomous foodlimited Lotka–Volterra cooperative model. Applied Mathematics and Computation, 254(??):1–8, March 1, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0096300314018037.

Livio:2002:GRS

[Liv02] Mario Livio. The Golden Ratio: the Story of Phi, the World’s Most Astonishing Number. Broadway Books, New York, NY, USA, 2002. ISBN 0-7679-0815-5. viii + 294 pp. LCCN QA466 .L58 2002. US$24.95, CAN$37.95. Benford’s Law is treated on pages 231–237.

LimKaiJie:2012:CES

[LK12] Shawn Lim Kai Jie and Mohandas Kalaichelvan. A critical evaluation of the significance of round numbers in European equity markets in light of the predictions from Benford’s Law. International Research Journal of Finance and Economics, 95(??):196–210, August 28, 2012. ISSN 1450-2887.

URL http://mpra.ub.uni-muenchen.de/40960/;   http:
//mpra.ub.uni-muenchen.de/40960/1/IRJFE_95_15.pdf.

Luque:2008:FDF

[LL08] Bartolo Luque and Lucas Lacasa. The first digit frequencies of primes and Riemann zeta zeros tend to uniformity following a size-dependent generalized Benford’s law. ArXiv e-prints, November 2008. CODEN ???? ISSN ???? URL

http://adsabs.harvard.edu/abs/2008arXiv0811.3302L;
 http://arxiv.org/abs/0811.3302. See published paper

[LL09].

Luque:2009:FDF

[LL09] Bartolo Luque and Lucas Lacasa. The first digit frequencies of prime numbers and Riemann zeta zeros. Proceedings of the Royal Society A: Mathematical, Physical, & Engineering Sciences, 465(2107):2197–2216, July 8, 2009. CO-DEN PRLAAZ. ISSN 1364-5021 (print), 1471-2946 (elec-

tronic). URL http://rspa.royalsocietypublishing.org/
content/465/2107/2197.abstract.

LoBuglio:2015:WDT

[LLD15] David Lo Buglio, Vanessa Lardinois, and Livio De Luca. What do thirty-one columns say about a “theoretical” thirty-second? Journal on Computing and Cultural Heritage (JOCCH), 8(1): 6:1–6:??, February 2015. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Lowe:1999:FDL

[LMH99] Tomas Lowe, Sally Murphy, and Justin Hayward. The first digit law. Report ??, Department of Computing, Imperial College, London, UK, June 1999. URL http://users.isr.ist.

utl.pt/~jmrs/research/TopicsLinks/Benford/hayward.
pdf; http://www.doc.ic.ac.uk/~jjh97/suprema/main_page.
html.

Lolbert:2008:NEG

[Lol08] Tamás Lolbert. On the non-existence of a general Benford’s law. Mathematical Social Sciences, 55(2):103–106, March 2008. CODEN MSOSDD. ISSN 0165-4896 (print), 1879-

3118 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165489607000935.

Longfellow:1992:LER

[Lon92] Charles A. Longfellow, Jr. Letters to the Editor: Regarding “Lotka’s Law and the Kolmogorov–Smirnov Test: An Error in Calculation”. Journal of the American Society for Information Science, 43(7):518, August 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic). See [Lou92].

Leinicke:2006:FFN

[LOR06] Linda M. Leinicke, Joyce A. Ostrosky, and W. Max Rexroad. Fighting fraud by the numbers: Digital analysis primer. Fraud Magazine, ??(??):??, January/February 2006. ISSN 1553-

6645. URL http://www.fraud-magazine.com/article.
aspx?id=4294967620.

Losee:2001:TDB

[Los01] Robert M. Losee. Term dependence: a basis for Luhn and Zipf models. Journal of the American Society for Information Science and Technology, 52(12):1019–1025, 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lotka:1926:FDS

[Lot26] Alfred J. Lotka. The frequency distribution of scientific productivity. Journal of the Washington Academy of Sciences, 16 (12):317–323, June 19, 1926. CODEN JWASA3. ISSN 0043-

0439. URL http://www.jstor.org/stable/24529203. This

is the earliest known publication on the phenomenon known as Zipf’s Law, here applied to publication counts of chemists and physicists. Lotka’s Law (a term possibly first used in [Zip49]) says that the number of authors producing n publications is about 1/n^{2} of the number producing only one. This ‘law’ seems to have been misunderstood and misapplied in other fields; see [Pot81].

Loughner:1992:BCL

[Lou92] William Loughner. Brief communications: Lotka’s Law and the Kolmogorov–Smirnov test: an error in calculation. Journal of the American Society for Information Science, 43(2): 149–150, March 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic). See comment [Lon92].

Lowe:2000:BLF

[Low00a] R. Lowe. Benford’s Law and fraud detection. Chartered Accountants Journal of New Zealand, 79(10):32–36, ???? 2000. ISSN 1172-9929.

Lowe:2000:WBL

[Low00b] R. Lowe. When Benford’s Law is broken. Chartered Accountants Journal of New Zealand, 79(11):24–27, ???? 2000. ISSN 1172-9929.

Loynes:1973:SRP

[Loy73] R. M. Loynes. Some results in the probabilistic theory of asymptotic uniform distribution modulo 1. Zeitschrift für Wahrscheinlichkeitstheorie und verwandte Gebiete, 26 (1):33–41, ???? 1973. CODEN ZWVGAA. ISSN 0044-

3719. URL http://link.springer.com/article/10.1007/
BF00533958.

Llamocca:2015:DEP

[LP15] Daniel Llamocca and Marios Pattichis. Dynamic energy, performance, and accuracy optimization and management using automatically generated constraints for separable 2D FIR filtering for digital video processing. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 7(4):4:1–4:??, January 2015. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Lagarias:2005:BLF

[LS05] J. C. Lagarias and K. Soundararajan. Benford’s law for the 3x + 1 function. ArXiv Mathematics e-prints,

September 2005. CODEN ????  ISSN ????  URL

http://adsabs.harvard.edu/abs/2005math......9175L;
 http://arxiv.org/abs/math/0509175.

Lagarias:2006:BLF

[LS06] Jeffrey C. Lagarias and K. Soundararajan. Benford’s law for the 3x + 1 function. J. London Math. Soc. (2), 74(2):289–303, 2006. CODEN JLMSAK. ISSN 0024-6107 (print), 1469-7750 (electronic).

Lyons:2011:NSF

[LS11] M. Lyons and E. Carl-Ludwig Siegel. NON-Shor factorization via BEQS BEC: Watkins number-theory “pure”-mathematics ∪ with statistical-physics: Benford log-law inversion to ONLY BEQS digit d = 0 BEC!!! In APS March Meeting Abstracts, page K1142. American Physical Society, ????, March

2011. URL http://adsabs.harvard.edu/abs/2011APS..
MAR.K1142L.

Luca:2014:FDF

[LS14] Florian Luca and Pantelimon Stănică. On the first digits of the Fibonacci numbers and their Euler function. Uniform Distribution Theory, 9(1):21–25, 2014. ISSN 1336-913X (print), 2309-5377 (electronic). URL https://math.boku.

ac.at/udt/vol09/no1/03LucaStanica.pdf.

Leemis:2000:SDS

[LSE00] Lawrence M. Leemis, Bruce W. Schmeiser, and Diane L. Evans. Survival distributions satisfying Benford’s Law. The American Statistician, 54(4):236–241, November 2000. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.amstat.org/publications/tas/
Leemis.htm;   http://www.jstor.org/stable/2685773.

Lei:2000:EIN

[LSS00] Skylar Lei, Michael Smith, and Giancarlo Succi. Empirical investigation of a novel approach to check the integrity of software engineering measuring processes. In ACM [ACM00], page 773. ISBN 1-58113-206-9, 0-7803-6325-6 (casebound), 0-7803-6326-4 (microfiche). LCCN QA76 N271 2000.

Lu:2007:UFD

[Lu07] Fletcher Lu. Uncovering fraud in direct marketing data with a fraud auditing case builder. In Knowledge Discovery in Databases: PKDD 2007, pages 540–547. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc.,

2007. URL http://link.springer.com/chapter/10.1007/
978-3-540-74976-9_56.

Ley:1994:TPB

[LV94] Eduadro Ley and Hal R. Varian. Are there psychological barriers in the Dow–Jones index? Applied Financial Economics, 4(3):217–224, ???? 1994. CODEN ???? ISSN 0960-3107 (print), 1466-4305 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/758526902.

Lagarias:1992:PTS

[LW92] J. C. Lagarias and A. Weiss. The 3x + 1 problem: Two stochastic models. Annals of Applied Probability, 2(1):229– 261, February 1992. CODEN ???? ISSN 1050-5164. URL

http://projecteuclid.org/euclid.aoap/1177005779.

Lin:2014:CCE

[LW14] Fengyi Lin and Sheng-Fu Wu. Comparison of cosmetic earnings management for the developed markets and emerging markets: Some empirical evidence from the United States and Taiwan. Economic Modelling, 36(0):466–473, 2014. ISSN 0264-9993 (print), 1873-6122 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0264999313004288.

Lin:2014:RAA

[LWFW14] Fengyi Lin, Chung-Min Wu, Tzu-Yi Fang, and Jheng-Ci Wun. The relations among accounting conservatism, institutional investors and earnings manipulation. Economic Modelling, 37(0):164–174, 2014. ISSN 0264-9993 (print), 1873-

6122 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0264999313004471.

Liu:2010:CIE

[LWY10] Hongyan Liu, Xiaoyu Wang, and Yinghui Yang. Comments on “An integrated efficient solution for computing frequent and top-k elements in data streams”. ACM Transactions on Database Systems, 35(2):15:1–15:??, April 2010. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). See [MAE06].

Li:2005:ZLP

[LX05] Yujian Li and Chuangbai Xiao. Zipf’s law probably existing in protein sequences. J. Beijing Univ. Technol., 31(4):366–368, 2005. CODEN ???? ISSN ????

Li:2002:ZLI

[LY02] Wentian Li and Yaning Yang. Zipf’s Law in importance of genes for cancer classification using microarray data. Journal of Theoretical Biology, 219(4):539–551, December 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302931450.

Luo:2015:IAD

[LYH15] Da Luo, Rui Yang, and Jiwu Huang. Identification of AMR decompressed audio. Digital Signal Processing, 37 (0):85–91, 2015. ISSN 1051-2004 (print), 1095-4333 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1051200414003200.

Li:2009:PDD

[LYY09] Weihai Li, Yuan Yuan, and Nenghai Yu. Passive detection of doctored JPEG image via block artifact grid extraction. Signal Processing, 89(9):1821–1829, 2009. CO-DEN SPRODR. ISSN 0165-1684 (print), 1872-7557 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165168409001315.

Li:2012:DTR

[LZL^{+}12] Xiang Hua Li, Yu Qian Zhao, Miao Liao, Frank Y. Shih, and Yun Q. Shi. Detection of tampered region for JPEG images by using mode-based first digit features. EURASIP Journal on Advances in Signal Processing, 2012(1):190:1–190:??, August 2012. CODEN ???? ISSN 1687-6172 (print), 1687-6180

(electronic). URL http://link.springer.com/article/10.
1186/1687-6180-2012-190.

Liu:2011:BLC

[LZNR11] X. J. Liu, X. P. Zhang, D. D. Ni, and Z. Z. Ren. Benford’s law and cross-sections of A(n, α)B reactions. European Physical Journal A: Hadrons and Nuclei, 47:78, June 2011. CODEN EPJAFV. ISSN 1434-6001 (print), 1434-601X (electronic).

URL http://adsabs.harvard.edu/abs/2011EPJA...47...
78L.

Lu:2013:DZH

[LZZ13] Linyuan Lü, Zi-Ke Zhang, and Tao Zhou. Deviation of Zipf’s and Heaps’ Laws in human languages with limited dictionary sizes. Scientific Reports, 3(??):1–7, ???? 2013. CODEN SRCEC3. ISSN 2045-2322. URL http://www.nature.com/

srep/2013/130130/srep01082/full/srep01082.html.

Mir:2018:BLS

[MA18] Tariq Ahmad Mir and Marcel Ausloos. Benford’s law: a “sleeping beauty” sleeping in the dirty pages of logarithmic tables. Journal of the Association for Information Science and Technology, 69(3):349–358, March 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Macaulay:1922:PLG

[Mac22] F. R. Macaulay. Pareto’s law and the general problem of mathematical describing the frequency of distribution of income. In W. C. Mitchell, editor, Income in the United States: Its Amount and Distribution: 1909–1919, page ?? National Bureau of Economic Research, New York, NY, USA, 1922. LCCN ????

MacDougall:2014:AIC

[Mac14a] Margaret MacDougall. Assessing the integrity of clinical data: When is statistical evidence too good to be true? Topoi: An International Review of Philosophy, 33(2):323– 337, October 2014. ISSN 0167-7411 (print), 1572-8749

(electronic). URL http://link.springer.com/article/10.
1007/s11245-013-9216-5.

Mir:2014:BLP

[MAC14b] Tariq Ahmad Mir, Marcel Ausloos, and Roy Cerqueti. Benford’s law predicted digit distribution of aggregated income taxes: the surprising conformity of Italian cities and regions. European Physical Journal B: Condensed Matter and Complex Systems, 87(11), 2014. CODEN EPJBFY. ISSN 1434-6028 (print), 1434-6036 (electronic).

Martinson:2005:SBB

[MAdV05] Brian C. Martinson, Melissa S. Anderson, and Raymond de Vries. Scientists behaving badly. Nature, 435(7043):737– 738, June 9, 2005. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v435/n7043/pdf/435737a.pdf.

Metwally:2006:IES

[MAE06] Ahmed Metwally, Divyakant Agrawal, and Amr El Abbadi. An integrated efficient solution for computing frequent and top-k elements in data streams. ACM Transactions on Database Systems, 31(3):1095–1133, September 2006. CO-DEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). See comments in [LWY10].

Mandelbrot:1953:ITS

[Man53] Benoît B. Mandelbrot. An informational theory of the statistical structure of languages. In Willis Jackson, editor, Communication theory: papers read at a Symposium on “Applications of Communication Theory” held at the Institution of Electrical Engineers, London, September 22nd–26th 1952, pages 486–502. Butterworths, London, UK, 1953. LCCN Q350

.S92 1952. URL http://lccn.loc.gov/53004215;   http:
//www.nytimes.com/2010/10/17/us/17mandelbrot.html.

Mandelbrot:1959:NCS

[Man59] Benoît Mandelbrot. A note on a class of skew distribution functions: analysis and critique of a paper by H. A. Simon. Information and Control, 2(??):90–99, April 1959. CODEN IFCNA4. ISSN 0019-9958 (print), 1878-2981 (electronic). See [Sim55].

Mandelbrot:1960:PLL

[Man60] Benoît Mandelbrot. The Pareto–Levy law and the distribution of income. International Economic Review, 1(??): 79–106, ???? 1960. CODEN INERAE. ISSN 0020-6598 (print), 1468-2354 (electronic). URL http://www.ingenta.

com/journals/browse/bpl/iere.

Mandelbrot:1962:SNR

[Man62] Benoît Mandelbrot. Statistics of natural resources and the law of Pareto. Research Note NC-146, IBM, Yorktown Heights, NY, USA, June 29, 1962.

Manin:2008:ZLA

[Man08] Dmitrii Y. Manin. Zipf’s Law and avoidance of excessive synonymy. Cognitive Science, 32(7):1075–1098, 2008. CODEN COGSD5. ISSN 0364-0213 (print), 1551-6709 (electronic).

Marcus:2012:UFA

[Mar12] Adam Marcus. Update on Fujii: Anesthesia journal finds overwhelming statistical evidence of data fabrication. Web site

article., March 8, 2012. URL https://retractionwatch.
com/2012/03/08/update-on-fujii-anesthesia-journal-
finds-overwhelming-statistical-evidence-of-data-fabrication/

.

Maslov:2005:GTS

[Mas05a] V. P. Maslov. On a general theorem of set theory that leads to the Gibbs, Bose–Einstein, and Pareto distributions and to the Zipf–Mandelbrot law for the stock market. Matematicheskie Zametki, 78(6):870–877, 2005. CODEN ???? ISSN 0025-567X.

Maslov:2005:RZLb

[Mas05b] V. P. Maslov. Refinement of Zipf’s law for frequency dictionaries. Doklady Mathematics, 72(3):942–945, 2005. CODEN ???? ISSN 1064-5624 (print), 1531-8362 (electronic).

Maslov:2005:RZLa

[Mas05c] V. P. Maslov. Refining Zipf’s law for frequency dictionaries. Doklady Akademii Nauk SSSR, 405(5):591–594, 2005. CO-DEN DANKAS. ISSN 0869-5652.

Maslov:2005:ZML

[Mas05d] V. P. Maslov. The Zipf–Mandelbrot law: quantization and an application to the stock market. Russian Journal of Mathematical Physics, 12(4):483–488, 2005. CODEN RJMPEL. ISSN 1061-9208 (print), 1555-6638 (electronic).

Maslov:2006:BGA

[Mas06a] V. P. Maslov. Bose gas of anharmonic oscillators and a refinement of Zipf’s law. Teoreticheskaya i Matematicheskaya Fizika, 148(3):495–496, 2006. CODEN TMFZAL. ISSN 0564-

Maslov:2006:NAD

[Mas06b] V. P. Maslov. Negative asymptotic dimension, a new condensate, and their relation to the quantized Zipf law. Matematicheskie Zametki, 80(6):856–863, 2006. CODEN ???? ISSN 0025-567X.

Maslov:2006:RZM

[Mas06c] V. P. Maslov. A refinement of the Zipf–Mandelbrot law and lacunarity in an ideal gas. Teoreticheskaya i Matematicheskaya Fizika, 147(3):511–512, 2006. CODEN TMFZAL. ISSN 0564-6162.

Maslova:2006:ZLV

[Mas06d] T. V. Maslova. On the Zipf law and the V. P. Maslov formula for semiotic systems. Doklady Akademii Nauk SSSR, 411(6): 754–757, 2006. CODEN DANKAS. ISSN 0869-5652.

Maslova:2006:ZLM

[Mas06e] T. V. Maslova. Zipf’s law and Maslov’s formula as applied to semiotic systems. Doklady Mathematics, 74(3):917–920, 2006. CODEN ???? ISSN 1064-5624 (print), 1531-8362 (electronic).

Maslov:2007:PTZ

[Mas07] V. P. Maslov. Phase transitions of the zeroth kind and the quantization of the Zipf law. Teoreticheskaya i Matematicheskaya Fizika, 150(1):118–142, 2007. CODEN TMFZAL. ISSN 0564-6162.

Matthews:1999:PO

[Mat99] Robert Matthews. The power of one. New Scientist, 163 (2194):27–30, July 10, 1999. CODEN NWSCAL. ISSN 0262-4079 (print), 1364-8500 (electronic). URL http://www.

fortunecity.com/emachines/e11/86/one.html;   http:
//www.newscientist.com/article/mg16321944.600-the-
power-of-one.html.

Mayr:2008:EBE

[May08] Philipp Mayr. An evaluation of Bradfordizing effects. Collnet Journal of Scientometrics and Information Management, 2 (2):21–27, 2008. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Mantegna:1994:LFN

[MBG^{+}94] R. N. Mantegna, S. V. Buldyrev, A. L. Goldberger, S. Havlin, C. K. Peng, M. Simons, and H. E. Stanley. Linguistic features of noncoding DNA sequences. Physical Review Letters, 73(??): 3169–??, December 5, 1994. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145. URL http:

//link.aps.org/doi/10.1103/PhysRevLett.73.3169. See

comment [Vos96] and reply [MBG^{+}96].

Mantegna:1995:SAC

[MBG^{+}95] R. N. Mantegna, S. V. Buldyrev, A. L. Goldberger, S. Havlin, C.-K. Peng, M. Simons, and H. E. Stanley. Systematic analysis of coding and noncoding DNA sequences using methods of statistical linguistics. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 52(??): 2939–??, September 1, 1995. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-2376 (electronic). URL http://link.aps.

org/doi/10.1103/PhysRevE.52.2939.

Mantegna:1996:MAR

[MBG^{+}96] R. N. Mantegna, S. V. Buldyrev, A. L. Goldberger, S. Havlin, C.-K. Peng, M. Simons, and H. E. Stanley. Mantegna et al. reply. Physical Review Letters, 76:1979–1981, March 1996. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (elec-

tronic), 1092-0145. URL http://link.aps.org/doi/10.
1103/PhysRevLett.76.1979. See [MBG^{+}94, Vos96].

Malcai:1999:PLD

[MBS99] Ofer Malcai, Ofer Biham, and Sorin Solomon. Power-law distributions and Lévy-stable intermittent fluctuations in stochastic systems of many autocatalytic elements. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 60(2):1299–1303, August 1999. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-

2376 (electronic). URL http://link.aps.org/doi/10.
1103/PhysRevE.60.1299.

McAllister:2003:ARE

[McA03] James W. McAllister. Algorithmic randomness in empirical data. Studies in History and Philosophy of Science Part A, 34(3):633–646, September 2003. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0039368103000475. See replies [TGD05, McA05].

McAllister:2005:ACE

[McA05] James W. McAllister. Algorithmic compression of empirical data: reply to Twardy, Gardner, and Dowe. Studies in History and Philosophy of Science Part A, 36(2):403–410, June 2005. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368105000221. See [McA03, TGD05].

McGrayne:2011:TWH

[McG11] Sharon Bertsch McGrayne. The theory that would not die: how Bayes’ rule cracked the Enigma code, hunted down Russian submarines, and emerged triumphant from two centuries of controversy. Yale University Press, New Haven, CT, USA, 2011. ISBN 0-300-16969-8. xiii + 320 pp. LCCN QA279.5 .M415 2011.

McKee:1980:POR

[McK80] Arnold F. McKee. The Pareto optimum: Reconcilable with social thought? International Journal of Social Economics, 7 (7):366–375, ???? 1980. CODEN ISLEBC. ISSN 0306-8293 (print), 1758-6712 (electronic).

Mosimann:2002:TDE

[MDDK02] James Mosimann, John Dahlberg, Nancy Davidian, and John Krueger. Terminal digits and the examination of questioned data. Accountability in Research, 9(2):75–92, ???? 2002. CO-DEN ARQAEZ. ISSN 0898-9621 (print), 1545-5815 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/08989620212969.

McCowan:2005:AUZ

[MDJH05] B. McCowan, L. R. Doyle, J. M. Jenkins, and S. F. Hanser. The appropriate use of Zipf’s law in animal communication studies. Animal Behaviour, 69(1):F1–F7, January 20, 2005. CODEN ANBEA8. ISSN 0003-3472 (print), 1095-

8282 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S000334720400346X.

Moret:2006:NBL

[MdSPZ06] M. A. Moret, V. de Senna, M. G. Pereira, and G. F. Zebende. Newcomb–Benford law in astrophysical sources. International Journal of Modern Physics C [Physics and Computers], 17 (11):1597–1604, ???? 2006. CODEN IJMPEO. ISSN 0129-1831 (print), 1793-6586 (electronic). URL http://adsabs.

harvard.edu/abs/2006IJMPC..17.1597M.

Moret:2009:GSA

[MdSSZ09] M. A. Moret, V. de Senna, M. C. Santana, and G. F. Zebende. Geometric structural aspects of proteins and Newcomb–Benford Law. International Journal of Modern Physics C [Physics and Computers], 20(12):1981–1988, ???? 2009. CODEN IJMPEO. ISSN 0129-1831 (print), 1793-

6586 (electronic). URL http://adsabs.harvard.edu/abs/
2009IJMPC..20.1981M.

Mbona:2022:FSU

[ME22] Innocent Mbona and Jan H. P. Eloff. Feature selection using Benford’s law to support detection of malicious social media bots. Information Sciences, 582:369–381, January 2022. CODEN ISIJBC. ISSN 0020-0255 (print), 1872-6291

(electronic). URL https://dblp.org/db/journals/isci/
isci582.html#MbonaE22.

Mebane:2006:DAE

[Meb06a] Walter R. Mebane, Jr. Detecting attempted election theft: Vote counts, voting machines and Benford’s Law. In ????, editor, 2006 Annual Meeting of the Midwest Political Science Association, April 20–23, Palmer House, Chicago, pages i + 50. ????, ????, April 19, 2006. ISBN ???? LCCN ???? URL

http://www.umich.edu/~wmebane/mw06.pdf.

Mebane:2006:EFS

[Meb06b] Walter R. Mebane, Jr. Election forensics: The second-digit Benford’s Law test and recent American Presidential elections. In ????, editor, Election Fraud Conference, Salt Lake City, Utah, September 29–30, 2006, page ?? ????, ????, 2006. ISBN

???? LCCN ???? URL http://www-personal.umich.edu/~
wmebane/fraud06.pdf.

Mebane:2006:EFV

[Meb06c] Walter R. Mebane, Jr. Election forensics: Vote counts and Benford’s Law. In ????, editor, 2006 Summer Meeting of the Political Methodology Society, UC-Davis, July 20–22, 2006, pages ii + 48. ????, ????, July 18, 2006. ISBN ???? LCCN

???? URL http://www.umich.edu/~wmebane/pm06.pdf.

Mebane:2007:SD

[Meb07] Walter R. Mebane, Jr. Statistics for digits. In ????, editor, 2007 Summer Meeting of the Political Methodology Society, Pennsylvania State University, July 18–21, 2007, pages 1–21. ????, ????, July 2, 2007. ISBN ???? LCCN ???? URL http:

//www.umich.edu/~wmebane/pm07.pdf.

Mebane:2008:EFO

[Meb08a] Walter R. Mebane, Jr. Election forensics: Outlier and digit tests in America and Russia. In ????, editor, The American Electoral Process conference, Center for the Study of Democratic Politics, Princeton University, May 1–3, 2008, pages i + 26. ????, ????, 2008. ISBN ???? LCCN ???? URL http:/

/www.umich.edu/~wmebane/aep2008.pdf.

Mebane:2008:EFSa

[Meb08b] Walter R. Mebane, Jr. Election forensics: The second-digit Benford’s Law test and recent American Presidential elections. In R. Michael Alvarez, Thad E. Hall, and Susan D. Hyde, editors, The Art and Science of Studying Election Fraud: Detection, Prevention, and Consequences, pages ii + 30. Brookings Institution, Washington, DC, USA, 2008. ISBN ???? LCCN

???? URL http://www.umich.edu/~wmebane/fraud06.pdf.

Preprint dated 3 November 2006.

Mebane:2008:EFSb

[Meb08c] Walter R. Mebane, Jr. Election forensics: the second-digit Benford’s Law test and recent American presidential elections. In Alvarez et al. [AHH08], page ?? ISBN 0-8157-0138-1 (hardcover), 0-8157-0139-X (paperback). LCCN JK1994 .E43

2008. URL http://www.loc.gov/catdir/toc/ecip0810/
2008005719.html.

Mebane:2009:NPE

[Meb09] Walter R. Mebane, Jr. Note on the Presidential election in Iran. Report ????, University of Michigan, Ann Arbor, MI,

USA, June 2009. URL http://www.umich.edu/rvwmebane/
note29jun2009.pdf.

Mebane:2010:EFSa

[Meb10a] Walter R. Mebane, Jr. Election fraud or strategic voting? In ????, editor, Proceedings of the 2010 Annual Meeting of the Midwest Political Science Association, Chicago, IL, April 22–25, 2010, page ?? ????, ????, 2010.

Mebane:2010:EFSb

[Meb10b] Walter R. Mebane, Jr. Election fraud or strategic voting? Can second-digit tests tell the difference? In ????, editor, 2010 Summer Meeting of the Political Methodology Society, University of Iowa, July 22–24, 2010, pages 1–49. ????, ????,

2010. ISBN ???? URL http://www.umich.edu/~wmebane/
pm10.pdf.

Mebane:2010:FPE

[Meb10c] Walter R. Mebane, Jr. Fraud in the 2009 Presidential Election in Iran? Chance: a magazine for people interested in the analysis of data, 23(1):6–15, March 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280

(electronic). URL http://www.springerlink.com/content/
0337v415563q0753/abstract/.

Mebane:2010:MSD

[Meb10d] Walter R. Mebane, Jr. Memo on second-digit tests done on precinct counts for Democratic Senate primary in South Carolina, 2010. Report, University of Michigan, Ann Arbor,

MI, USA, June 12, 2010. URL http://www.umich.edu/~
wmebane/SCmemo12jun2010.pdf.

Mebane:2011:CBL

[Meb11] Walter R. Mebane, Jr. Comment on Benford’s Law and the detection of election fraud. Political Analysis, 19(3):269–272, Summer 18, 2011. CODEN ???? ISSN 1047-1987 (print),

1476-4989 (electronic). URL http://pan.oxfordjournals.
org/content/19/3/269.abstract. See [DMO11].

Mebane:2012:SDT

[Meb12] Walter R. Mebane, Jr. Second-digit tests for voters’ election strategies and election fraud. In ????, editor, Proceedings of the 2012 Annual Meeting of the Midwest Political Science Association, Chicago, IL, April 11–14, 2012, page ?? ????, ????, 2012.

Mebane:2015:CVC

[Meb15] Walter R. Mebane, Jr. Can vote counts’ digits and Benford’s Law diagnose elections? In Miller [Mil15a], pages 212– 222. ISBN 0-691-14761-2 (hardcover), 1-4008-6659-6 (e-book).

LCCN QA273.6 .B46 2015. URL http://press.princeton.
edu/titles/10527.html.

Mebane:20xx:EF

[Mebxx] Walter R. Mebane, Jr. Election Forensics. ????, ????, 20xx. ???? pp. Book in preparation.

Meintanis:2009:UAT

[Mei09] Simos G. Meintanis. A unified approach of testing for discrete and continuous Pareto laws. Statistical Papers, 50(3):569– 580, June 2009. CODEN STPAE4. ISSN 0932-5026 (print),

1613-9798 (electronic). URL http://link.springer.com/
article/10.1007/s00362-007-0103-2.

Mellor:2006:GKZ

[Mel06] Andy Mellor. Book review: G. K. Zipf, The Psychobiology of Language (1935) MIT Press, Cambridge, MA. System, 34(3):455–457, September 20, 2006. CODEN ???? ISSN 0346-251X (print), 1879-3282 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0346251X06000674.

Mumic:2021:MTD

[MF21] Nermina Mumic and Peter Filzmoser. A multivariate test for detecting fraud based on Benford’s law, with application to music streaming data. Stat. Methods Appl., 30(3):819–840, 2021. ISSN 1618-2510.

Mansury:2007:EZL

[MG07a] Yuri Mansury and László Gulyás. The emergence of Zipf’s Law in a system of cities: an agent-based simulation approach. Journal of Economic Dynamics and Control, 31(7):2438–2460, July 15, 2007. CODEN JEDCDH. ISSN 0165-1889 (print),

1879-1743 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165188906001527.

Morris:2007:MRT

[MG07b] Steven A. Morris and Michel L. Goldstein. Manifestation of research teams in journal literature: a growth model of papers, authors, collaboration, coauthorship, weak ties, and Lotka’s Law. Journal of the American Society for Information Science and Technology, 58(12):1764–1782, October 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

McCartney:2015:TSD

[MG15] Mark McCartney and David H. Glass. A three-state dynamical model for religious affiliation. Physica A: Statistical Mechanics and its Applications, 419(0):145–152, 2015. CO-DEN PHYADX. ISSN 0378-4371 (print), 1873-2119 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378437114008206.

Miraglia:2022:AIC

[MG22] Jorge E. Miraglia and Maria Silvia Gravielle. Atomic ionization, capture, and stopping cross sections by ion impact examined with the Benford law. Advances in Quantum Chemistry, 85:??, 2022. CODEN AQCHA9. ISSN 0065-3276.

URL https://www.sciencedirect.com/science/article/
pii/S0065327622000107.

Makse:1995:MUG

[MHS95] Hernán A. Makse, Shlomo Havlin, and H. Eugene Stanley. Modelling urban growth patterns. Nature, 377(6550):608– 612, October 19, 1995. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v377/n6550/abs/377608a0.html.

Miller:1957:SEI

[Mil57] George A. Miller. Some effects of intermittent silence. The American Journal of Psychology, 70(2):311–314, June 1957. CODEN AJPCAA. ISSN 0002-9556 (print), 1939-8298 (elec-

tronic). URL http://www.jstor.org/stable/1419346.

Miller:2004:STB

[Mil04] Steven J. Miller. Some thoughts on Benford’s law. Preprint, Williams College, Williamstown, MA 01267 USA,

2004. URL http://www.williams.edu/go/math/sjmiller/
public_html/BrownClasses/1/BenfordTreatise.pdf.

Miller:2015:BLT

[Mil15a] Steven J. Miller, editor. Benford’s Law: theory and applications. Princeton University Press, Princeton, NJ, USA, 2015. ISBN 0-691-14761-2 (hardcover), 1-4008-6659-6 (e-book). xxvi + 438 pp. LCCN QA273.6 .B46 2015. URL http://press.

princeton.edu/titles/10527.html.

Miller:2015:FAB

[Mil15b] Steven J. Miller. Fourier analysis and Benford’s Law. In Benford’s Law: theory and applications [Mil15a], pages 68– 105. ISBN 0-691-14761-2 (hardcover), 1-4008-6659-6 (e-book).

LCCN QA273.6 .B46 2015. URL http://press.princeton.
edu/titles/10527.html.

Miller:2015:HSOa

[Mil15c] Steven J. Miller. How a simple observation from the 1800s about patterns in big data sets can fight fraud. The Conversation, ??(??):??, December 10, 2015. URL https://

theconversation.com/how-a-simple-observation-from-
the-1800s-about-patterns-in-big-data-sets-can-fight-
fraud-42683.

Miller:2015:HSOb

[Mil15d] Steven J. Miller. How a simple observation from the 1800s about patterns in big data sets can fight fraud. Scientific Computing, ??(??):??, December 10, 2015. CODEN SCHRCU. ISSN 1930-5753 (print), 1930-6156 (electronic). URL http:/

/www.scientificcomputing.com/articles/2015/12/how-
simple-observation-1800s-about-patterns-big-data-sets-
can-fight-fraud.

Miller:2015:QIB

[Mil15e] Steven J. Miller. A quick introduction to Benford’s Law. In Benford’s Law: theory and applications [Mil15a], pages 3– 22. ISBN 0-691-14761-2 (hardcover), 1-4008-6659-6 (e-book).

LCCN QA273.6 .B46 2015. URL http://press.princeton.
edu/titles/10527.html.

Mir:2011:LLD

[Mir11] Tariq Ahmad Mir. Law of the leading digits and the ideological struggle for numbers. ArXiv e-prints, page 11, 2011. URL

http://arxiv.org/abs/1104.3948.

Mir:2012:LDD

[Mir12a] T. A. Mir. The leading digit distribution of the worldwide illicit financial flows. ArXiv e-prints, 2012. URL http://

arxiv.org/abs/1201.3432.

Mir:2012:LLD

[Mir12b] Tariq Ahmad Mir. The law of the leading digits and the world religions. Physica A: Statistical Mechanics and its Applications, 391(3):792–798, February 1, 2012. CO-DEN PHYADX. ISSN 0378-4371 (print), 1873-2119 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378437111007175.

Mir:2014:BLB

[Mir14] T. A. Mir. The Benford Law behavior of the religious activity data. Physica A: Statistical Mechanics and its Applications, 408(0):1–9, 2014. CODEN PHYADX. ISSN 0378-4371 (print),

1873-2119 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378437114002854.

Mir:2016:LDD

[Mir16a] T. A. Mir. The leading digit distribution of the worldwide illicit financial flows. Quality & Quantity, 50(1):271–281, January 2016. CODEN QQEJAV. ISSN 0033-5177 (print),

1573-7845 (electronic). URL http://link.springer.com/
article/10.1007/s11135-014-0147-z.

Mir:2016:CAC

[Mir16b] Tariq Ahmad Mir. Citations to articles citing Benford’s law: a Benford analysis. CoRR, abs/1602.01205(??):??, ???? 2016. CODEN ???? ISSN ???? URL http:

//arxiv.org/abs/1602.01205;   https://dblp.org/db/
journals/corr/corr1602.html#Mir16.

Mitzenmacher:2003:BHG

[Mit03] Michael Mitzenmacher. A brief history of generative models for power law and lognormal distributions. Internet Mathematics, 1(2):226–251, December 2003. CODEN ???? ISSN 1542-7951 (print), 1944-9488 (electronic). URL http:/

/projecteuclid.org/euclid.im/1089229510;   http://
www.eecs.harvard.edu/~michaelm/CS223/powerlaw.pdf;
 http://www.tandfonline.com/toc/uinm20/1/2.

Marulli:1979:BCB

[MK79] Luciana Marulli and Michael E. D. Koenig. Brief communications: Bradford distribution of data elements. Journal of the American Society for Information Science, 30(2):107–108, March 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Miller:2005:BLV

[MK05] Steven J. Miller and A. Kontorovich. Benford’s law, values of L-functions and the 3x + 1 problem. Acta Arithmetica, 120 (3):269–297, ???? 2005. CODEN AARIA9. ISSN 0065-1036 (print), 1730-6264 (electronic).

Mebane:2009:EFR

[MK09] Walter R. Mebane, Jr. and K. Kalinin. Electoral falsification in Russia: Complex diagnostics selections 2003–2004, 2007– 2008. Russian Electoral Review, 2(??):57–70, ???? 2009.

Mebane:2010:EFR

[MK10] Walter R. Mebane, Jr. and K. Kalinin. Electoral fraud in Russia: Vote counts analysis using second-digit mean tests. In ????, editor, Proceedings of the 2010 Annual Meeting of the Midwest Political Science Association, Chicago, IL, April 22–25, 2010, page ?? ????, ????, 2010.

Mebane:2013:SDI

[MK13] Walter R. Mebane, Jr. and T. Kent. Second digit implications of voters’ strategies and mobilizations in the United States during the 2000s. In ????, editor, Proceedings of the 2013 Annual Meeting of the Midwest Political Science Association, Chicago, IL, April 11–14, 2013, page ?? ????, ????, 2013.

McLaughlin:1984:DFI

[ML84] William I. McLaughlin and Sylvia A. Lundy. Digit functions of integer sequences. Fibonacci Quarterly, 22(2):105–115, May 1984. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/22-2/mclaughlin.pdf.

Macon:1950:DFD

[MM50] N. Macon and L. Moser. On the distribution of first digits of powers. Scripta Mathematica, 16(??):290–291, ???? 1950. CODEN ???? ISSN 0036-9713.

MacRoberts:1982:REL

[MM82] Michael H. MacRoberts and Barbara R. MacRoberts. A re-evaluation of Lotka’s Law of Scientific Productivity. Social Studies of Science, 12(3):443–450, August 1, 1982. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/030631282012003005.

Maia:1984:UBL

[MM84] M. J. F. Maia and M. D. Maia. On the unity of Bradford’s Law. Journal of Documentation, 40(3):206–216, ???? 1984. CODEN JDOCAS. ISSN 0022-0418 (print), 1758-7379 (electronic).

MacRoberts:1986:QMC

[MM86] Michael H. MacRoberts and Barbara R. MacRoberts. Quantitative measures of communication in science: a study of the formal level. Social Studies of Science, 16(1):151–172, February 1, 1986. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/
doi/pdf/10.1177/030631286016001008. See [Sti87] and re-

ply [MM87].

MacRoberts:1987:MFU

[MM87] Michael H. MacRoberts and Barbara R. MacRoberts. Measurement in the face of universal uncertainty: a reply to Stigler. Social Studies of Science, 17(2):334–336, May 1, 1987. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-

3659 (electronic). URL https://journals.sagepub.com/
doi/pdf/10.1177/030631287017002008. See [MM86].

Miller:1999:JFM

[MM99] Irwin Miller and Marylees Miller. John E. Freund’s mathematical statistics. Prentice-Hall, Upper Saddle River, NJ, USA, sixth edition, 1999. ISBN 0-13-123613-X. xii + 624 pp. LCCN QA276 .M4726 1999.

Miller:2004:JFM

[MM04] Irwin Miller and Marylees Miller. John E. Freund’s mathematical statistics with applications. Prentice-Hall, Upper Saddle River, NJ, USA, seventh edition, 2004. ISBN 0-13-142706-7. x + 614 pp. LCCN QA276 .M4726 2004.

Maslov:2006:ZLR

[MM06] V. P. Maslov and T. V. Maslova. On Zipf’s law and rank distributions in linguistics and semiotics. Matematicheskie Zametki, 80(5):718–732, 2006. CODEN ???? ISSN 0025-567X.

Miller:2014:JFM

[MM14] Irwin Miller and Marylees Miller. John E. Freund’s mathematical statistics with applications. Pearson, Boston, MA, USA, eighth edition, 2014. ISBN 0-321-80709-X (hardcover). xi + 529 pp. LCCN QA276 .M4726 2014.

Manack:2015:LDL

[MM15] Corey Manack and Steven J. Miller. Leading digit laws on linear Lie groups. Research in Number Theory, 1(1): 22:1–22:19, December 2015. CODEN ???? ISSN 2363-

9555. URL http://link.springer.com/article/10.1007/
s40993-015-0024-4.

Marchand:2021:BLC

[MM21] Chase Marchand and Dalton Maahs. Benford’s Law and COVID-19 data. Chance: a magazine for people interested in the analysis of data, 34(2):31–38, 2021. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Malacarne:2001:QED

[MML01] L. C. Malacarne, R. S. Mendes, and E. K. Lenzi. q-exponential distribution in urban agglomeration. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 65(??):017106, December 21, 2001. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-2376 (electronic). URL http:/

/link.aps.org/doi/10.1103/PhysRevE.65.017106.

Miller:2006:MCL

[MN06a] Steven J. Miller and M. J. Nigrini. The Modulo 1 Central Limit Theorem and Benford’s Law for products. International Journal of Algebra, 3(3):119–130, July 2006. CODEN ???? ISSN 1312-8868 (print), 1314-7595 (elec-

tronic). URL http://adsabs.harvard.edu/abs/2006math.
.....7686M;   http://arxiv.org/abs/math/0607686.

Miller:2006:OSB

[MN06b] Steven J. Miller and M. J. Nigrini. Order statistics and Benford’s Law. ArXiv Mathematics e-prints, January 2006. CO-

DEN ???? ISSN ???? URL http://adsabs.harvard.edu/
abs/2006math......1344M.

Miller:2008:MCL

[MN08a] Steven J. Miller and Mark J. Nigrini. The modulo 1 central limit theorem and Benford’s Law for products. International Journal of Algebra, 2(3):119–130, 2008. CO-DEN ???? ISSN 1312-8868 (print), 1314-7595 (electronic).

URL http://www.m-hikari.com/ija/ija-password-2008/
ija-password1-4-2008/millerIJA1-4-2008.pdf.

Miller:2008:OSB

[MN08b] Steven J. Miller and Mark J. Nigrini. Order statistics and Benford’s law. International Journal of Mathematics and Mathematical Sciences, pages 382948:1–382948:19, 2008. CO-DEN ???? ISSN 0161-1712 (print), 1687-0425 (electronic).

URL http://arxiv.org/abs/math/0601344;   https://
www.hindawi.com/journals/ijmms/2008/382948/. Article

ID 382948.

Mitov:2014:EAP

[MN14] Kosto V. Mitov and Saralees Nadarajah. Extremal and additive processes generated by Pareto distributed random vectors. ESAIM: Probability and Statistics, 18(??):667–??, ???? 2014. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Ma:2005:CBL

[MNW^{+}05] Y. G. Ma, J. B. Natowitz, R. Wada, K. Hagel, J. Wang, T. Keutgen, Z. Majka, M. Murray, L. Qin, P. Smith, R. Alfaro, J. Cibor, M. Cinausero, Y. El Masri, D. Fabris, E. Fioretto, A. Keksis, M. Lunardon, A. Makeev, N. Marie, E. Martin, A. Martinez-Davalos, A. Menchaca-Rocha, G. Nebbia, G. Prete, V. Rizzi, A. Ruangma, D. V. Shetty, G. Souliotis, P. Staszel, M. Veselsky, G. Viesti, E. M. Winchester, and S. J. Yennello. Critical behavior in light nuclear systems: Experimental aspects. Physical Review C (Nuclear Physics), 71(5): 054606, May 26, 2005. CODEN PRVCAN. ISSN 0556-2813 (print), 1089-490X, 1538-4497. URL http://link.aps.org/

doi/10.1103/PhysRevC.71.054606.

Mochty:2002:AMI

[Moc02] L. Mochty. Die Aufdeckung von Manipulationen im Rechnungswesen — Was leistet das Benford’s Law?. (German) [The disclosure of accounting manipulations — what does Benford’s Law mean?]. Die Wirtschaftsprüfung, 14(??):725–736, ???? 2002. CODEN ???? ISSN 0043-6313.

Mochty:2017:ABL

[Moc17] Ludwig Mochty. Die Achillesferse des Benford’s Law: KMUs. (German) [The Achilles heel of Benford’s Law: SMEs]. In B. Kaluza, K. Braun, H. Beschorner, and B. Rolfes, editors, Betriebswirtschaftliche Fragen zu Steuern, Finanzierung, Banken und Management, pages 39–55. Springer Fachmedien Wiesbaden, Wiesbaden, Germany, 2017.

Mocnik:2021:BLG

[Moc21] Franz-Benjamin Mocnik. Benford’s law and geographical information — the example of OpenStreetMap. International Journal of Geographical Information Science, 35(9): 1746–1772, ???? 2021. CODEN ???? ISSN 1365-8816 (print), 1365-8824 (electronic). URL https://dblp.org/db/

journals/gis/gis35.html#Mocnik21.

Mohanty:1979:MMQ

[Moh79] Siba N. Mohanty. Models and measurements for quality assessment of software. ACM Computing Surveys, 11(3):251– 275, September 1979. CODEN CMSVAN. ISSN 0010-4892. See also [Dun80, Moh80].

Mohanty:1980:SFW

[Moh80] Siba Mohanty. Surveyor’s forum: Working on interpretations. ACM Computing Surveys, 12(2):256, June 1980. CODEN CMSVAN. ISSN 0010-4892. See [Moh79, Dun80].

Moller:2009:MQA

[Möl09] M. Möller. Measuring the quality of auditing services with the help of Benford’s Law – an empirical analysis and discussion of this methodical approach. Report ??, Institute for Accounting and Control, University of Zürich, Zürich, Switzerland, April

24, 2009. 25 pp. URL http://ssrn.com/abstract=1529307.

Montemurro:2001:BZM

[Mon01] Marcelo A. Montemurro. Beyond the Zipf–Mandelbrot law in quantitative linguistics. Physica A: Statistical Mechanics and its Applications, 300(3–4):567–578, November 15, 2001. CODEN PHYADX. ISSN 0378-4371 (print), 1873-

2119 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378437101003557.

Moon:1936:SBI

[Moo36] Parry Moon. The Scientific Basis of Illuminating Engineering. McGraw-Hill, New York, NY, USA, 1936. xi + 608 pp. LCCN TH7703 .M65.

Morse:1981:IEB

[Mor81] Philip M. Morse. Implications of the exact Bradford distribution. Journal of the American Society for Information Science, 32(1):43–50, January 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Morters:2001:BGV

[Mör01] Peter Mörters. Benford’s Gesetz über die Verteilung der Ziffern. (German) [Benford’s law on the distribution of digits]. Habilitationsvorlesung, ????, Kaiserslauten, Germany, and Bath, UK, ???? 2001. 4 pp. URL http://people.bath.

ac.uk/maspm/benford.ps.

Morrow:2006:BLF

[Mor06] John Morrow. Benford’s Law, families of distributions and a test basis. Report, University of Wisconsin–Madison, Madison, WI, USA, ???? 2006. URL http://www.johnmorrow.

info/projects/benford/benfordMain.pdf. Revised draft of

9 October 2010.

Morrison:2010:MG

[Mor10] Kent E. Morrison. The multiplication game. Mathematics Magazine, 83(??):100–110, April 2010. CODEN MAMGA8. ISSN 0025-570X. URL http://www.calpoly.

edu/~kmorriso/Research/multgameMMfinal.pdf.

Manaris:2002:PTR

[MPM02] B. Manaris, T. Purewal, and C. McCormick. Progress towards recognizing and classifying beautiful music with computers: MIDI-encoded music and the Zipf-Mandelbrot Law. In IEEE, editor, Proceedings: IEEE SoutheastCon 2002: April 5–7, 2002, Embassy Suites Hotel, Columbia, South Carolina, USA, pages 52–57. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2002. ISBN 0-7803-7252-2. LCCN TK7801 .I56 2002. URL http://

ieeexplore.ieee.org/servlet/opac?punumber=7814.

Manaris:2006:IES

[MPPH06] Bill Manaris, Luca Pellicoro, George Pothering, and Harland Hodges. Investigating Esperanto’s statistical proportions relative to other languages using neural networks and Zipf’s law. In ????, editor, 2006 IASTED International Conference on Artificial Intelligence and Applications (AIA 2006), February 13–16, 2006, Innsbruck, Austria, pages 1–7. Acta Press, ????, 2006. ISBN ???? LCCN ???? URL http://www.cs.cofc.

edu/~manaris/uploads/Main/IASTED2006.pdf.

Malevergne:2011:TPA

[MPS11] Yannick Malevergne, Vladilen Pisarenko, and Didier Sornette. Testing the Pareto against the lognormal distributions with the uniformly most powerful unbiased test applied to the distribution of cities. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 83(3): 036111, March 22, 2011. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-2376 (electronic). URL http://link.aps.org/

doi/10.1103/PhysRevE.83.036111.

Moura:2006:ZLB

[MR06] Newton J. Moura, Jr. and Marcelo B. Ribeiro. Zipf law for Brazilian cities. Journal of Theoretical Biology, 219(4):441– 448, July 15, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378437105012422.

Manaris:2005:ZLM

[MRM^{+}05] Bill Manaris, Juan Romero, Penousal Machado, Dwight Krehbiel, Timothy Hirzel, Walter Pharr, and Robert B. Davis. Zipf’s law, music classification, and aesthetics. Computer Music Journal, 29(1):55–69, Spring 2005. CODEN CMUJDY. ISSN 0148-9267 (print), 1531-5169 (electronic).

URL http://muse.jhu.edu/journals/cmj/summary/v029/
29.1manaris.html;   http://muse.jhu.edu/journals/computer_
music_journal/v029/29.1manaris.pdf.

Manoochehrnia:2010:BLA

[MRR^{+}10] P. Manoochehrnia, F. Rachidi, M. Rubinstein, W. Schulz, and G. Diendorfer. Benford’s Law and its application to lightning data. IEEE Transactions on Electromagnetic Compatibility, 52(4):956–961, November 2010. CODEN IEMCAE. ISSN 0018-9375 (print), 1558-187X (electronic).

Martinez-Rodriguez:2011:MUE

[MRSCCS11] A. M. Martínez-Rodríguez, A. J. Sáez-Castillo, and A. Conde-Sánchez. Modelling using an extended Yule distribution. Computational Statistics & Data Analysis, 55(1):863–873, January 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-

7352 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167947310002872.

Montroll:1983:NOD

[MS83] Elliott W. Montroll and Michael F. Shlesinger. On 1/f noise and other distributions with long tails. Proceedings of the National Academy of Sciences of the United States of America, 79(10):3380–3383, May 15, 1983. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic). URL http://www.

pnas.org/content/79/10/3380.abstract.

Mitchell:2008:TST

[MS08] James G. Mitchell and Laurent Seuront. Towards a seascape topology II: Zipf analysis of one-dimensional patterns. Journal of Marine Systems, 69(3–4):328–338, February 2008. CO-DEN JMASE5. ISSN 0924-7963 (print), 1879-1573 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0924796307000632.

Mahdian:2009:UNI

[MS09] Babak Mahdian and Stanislav Saic. Using noise inconsistencies for blind image forensics. Image and Vision Computing, 27(10):1497–1503, 2009. CODEN IV-CODK. ISSN 0262-8856 (print), 1872-8138 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0262885609000146. Special Section: Computer Vision Methods for Ambient Intelligence.

Mahdian:2010:BBM

[MS10] Babak Mahdian and Stanislav Saic. A bibliography on blind methods for identifying image forgery. Signal Processing: Image Communication, 25(6):389–399, July 2010. CODEN SPICEF. ISSN 0923-5965 (print), 1879-2677 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0923596510000536.

Masse:2011:SWD

[MS11] Bruno Massé and Dominique Schneider. A survey on weighted densities and their connection with the first digit phenomenon. Rocky Mountain Journal of Mathematics, 41(5):1395–1415, ???? 2011. CODEN RMJMAE. ISSN 0035-7596 (print), 1945-

3795 (electronic). URL http://rmmc.asu.edu/abstracts/
rmj/vol41-5/surv-masspag1.pdf.

Masse:2012:RNS

[MS12] Bruno Massé and Dominique Schneider. Random number sequences and the first digit phenomenon. Electronic Journal of Probability, 17:86:1–86:17, 2012. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/1900.

Michalski:2013:DCF

[MS13] T. Michalski and G. Stoltz. Do countries falsify economic data strategically? Some evidence that they might. Review of Economics and Statistics, 95(2):591–616, May 2013. CODEN RECSA9. ISSN 0034-6535 (print), 1530-9142

(electronic). URL http://www.mitpressjournals.org/doi/
abs/10.1162/REST_a_00274.

Masse:2015:FGS

[MS15] Bruno Massé and Dominique Schneider. Fast growing sequences of numbers and the first digit phenomenon. International Journal of Number Theory, 11(3):705–719, May 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.worldscientific.com/doi/10.
1142/S1793042115500384.

Michel:2011:QAC

[MSA^{+}11] Jean-Baptiste Michel, Yuan Kui Shen, Aviva Presser Aiden, Adrian Veres, Matthew K. Gray, The Google Books Team, Joseph P. Pickett, Dale Hoiberg, Dan Clancy, Peter Norvig, Jon Orwant, Steven Pinker, Martin A. Nowak, and Erez Lieberman Aiden. Quantitative analysis of culture using millions of digitized books. Science, 331(6014):176–182, January 14, 2011. CODEN SCIEAS. ISSN 0036-8075 (print),

1095-9203 (electronic). URL http://www.sciencemag.org/
content/331/6014/176.abstract.

Maillart:2008:ETZ

[MSSvK08] T. Maillart, D. Sornette, S. Spaeth, and G. von Krogh. Empirical tests of Zipf’s Law mechanism in open source Linux distribution. Physical Review Letters, 101(21):218701, November 19, 2008. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145. URL http://link.aps.

org/doi/10.1103/PhysRevLett.101.218701.

Miller:2006:IMN

[MTB06] Steven J. Miller and Ramin Takloo-Bighash. An Invitation to Modern Number Theory. Princeton University Press, Princeton, NJ, USA, 2006. ISBN 0-691-12060-9 (hardcover). xx + 503 pp. LCCN QA241 .M5344 2006.

Milani:2013:AAB

[MTT13] S. Milani, M. Tagliasacchi, and S. Tubaro. Antiforensics attacks to Benford’s Law for the detection of double compressed images. In 2013 IEEE International Conference on Acoustics, Speech and Signal Processing, pages 3053–3057. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, May 2013. ISSN 1520-6149.

Mizuno:2008:PLE

[MTTT08] Takayuki Mizuno, Masahiro Toriyama, Takao Terano, and Misako Takayasu. Pareto law of the expenditure of a person in convenience stores. Physica A: Statistical Mechanics and its Applications, 387(15):3931–3935, 2008. CODEN PHYADX. ISSN 0378-4371 (print), 1873-2119 (electronic).

Muller:2003:AZP

[Mül03] Mathias Müller. Anwendungsmöglichkeiten der Ziffernanalyse in der Prüfungspraxis mit Schwerpunkt auf Benford’s Law. (German) [Applications of the digit analysis in the audit practice with a focus on Benford’s Law]. Diplomarbeit, Wirtschaftsuniversität Wien, Wien, Austria, ???? 2003. v + 75 pp. URL

http://permalink.obvsg.at/wuw/AC03895842.

Muller:2011:HAS

[Mül11] Hans Christian Müller. How an arcane statistical law could have prevented the Greek disaster. Economics Intelligence, ??(??):??, July 28, 2011. CODEN ???? ISSN ????

URL http://economicsintelligence.com/2011/07/28/
how-an-arcane-statistical-law-could-have-prevented-
the-greek-disaster/.

Murphy:1973:BCL

[Mur73a] Larry J. Murphy. Brief communications: “Lotka’s Law in the humanities?”. Journal of the American Society for Information Science, 24(6):461–462, November 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Murphy:1973:LLH

[Mur73b] Larry J. Murphy. Lotka’s Law in the humanities? Journal of the American Society for Information Science, 24(6):461–462, November/December 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Murtagh:2023:WBL

[Mur23] Jack Murtagh. What is Benford’s Law? Why this unexpected pattern of numbers is everywhere. Scientific American, ??(??):??, May 8, 2023. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL https://www.

scientificamerican.com/article/what-is-benfords-law-
why-this-unexpected-pattern-of-numbers-is-everywhere/

.

Manaris:2003:EMZ

[MVW^{+}03] Bill Manaris, Dallas Vaughan, Christopher Wagner, Juan Romero, and Robert B. Davis. Evolutionary music and the Zipf–Mandelbrot Law: Developing fitness functions for pleasant music. In Raidl [Rai03], pages 522–534. CO-DEN LNCSD9. ISBN 3-540-00976-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.618 .E899

2003. URL http://link.springer-ny.com/link/service/
series/0558/tocs/t2611.htm;   http://www.springerlink.
com/content/978-3-540-00976-4;   http://www.springerlink.
com/openurl.asp?genre=issue&issn=0302-9743&volume=
2611.

Molchanov:2004:BEL

[MW04] Stanislav Molchanov and Xian Wang. On the Benford’s empirical law. Random Oper. Stochastic Equations, 12(3):201–210, 2004. CODEN ???? ISSN 0926-6364 (print), 1569-397X (electronic).

Mosimann:1995:DFC

[MWE95] James E. Mosimann, Claire V. Wiseman, and Ruth E. Edelman. Data fabrication: Can people generate random digits? Accountability in Research, 4(1):31–55, ???? 1995. CO-DEN ARQAEZ. ISSN 0898-9621 (print), 1545-5815 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/08989629508573866.

Mercik:1999:SAI

[MWS99] Szymon Mercik, Karina Weron, and Zuzanna Siwy. Statistical analysis of ionic current fluctuations in membrane channels. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 60(6):7343–7348, December 1, 1999. CODEN PLEEE8. ISSN 1539-3755 (print),

1550-2376 (electronic). URL http://link.aps.org/doi/10.
1103/PhysRevE.60.7343.

Ma:2020:ERD

[MWX^{+}20] Chengcheng Ma, Baoyuan Wu, Shibiao Xu, Yanbo Fan, Yong Zhang, Xiaopeng Zhang, and Zhifeng Li. Effective and robust detection of adversarial examples via Benford–Fourier coefficients. CoRR, ??(??):??, ???? 2020. CODEN ???? ISSN

???? URL https://arxiv.org/abs/2005.05552 [;]   db/
journals/corr/corr2005.html#abs-2005-05552.

Marden:2014:APO

[MYP14] Jason R. Marden, H. Peyton Young, and Lucy Y. Pao. Achieving Pareto optimality through distributed learning. SIAM Journal on Control and Optimization, 52(5):2753–2770, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Mosteller:1967:DSR

[MYZ67] Frederick Mosteller, Cleo Youtz, and Douglas Zahn. The distribution of sums of rounded percentages. Demography, 4(2):850–858, June 1967. CODEN ???? ISSN 0070-3370 (print), 1533-7790 (electronic). URL http://muse.jhu.edu/

journals/dem/;   http://www.biomedsearch.com/nih/Distribution-
Sums-Rounded-Percentages/21318695.html; http://www.

jstor.org/stable/2060324. See [DF79] for further work.

Manrubia:1998:IMU

[MZ98a] Susanna C. Manrubia and Damián H. Zanette. Intermittency model for urban development. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 58(1):295–302, July 1998. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-2376 (electronic). URL http://link.aps.org/

doi/10.1103/PhysRevE.58.295.

Marsili:1998:IIL

[MZ98b] Matteo Marsili and Yi-Cheng Zhang. Interacting individuals leading to Zipf’s Law. Physical Review Letters, 80(??): 2741–??, March 23, 1998. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145. URL http:/

/link.aps.org/doi/10.1103/PhysRevLett.80.2741.

Manrubia:1999:SMP

[MZ99] Susanna C. Manrubia and Damián H. Zanette. Stochastic multiplicative processes with reset events. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 59(5):4945–4948, May 1999. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-2376

(electronic). URL http://link.aps.org/doi/10.1103/
PhysRevE.59.4945.

Mishra:2015:PGMa

[MZLH15a] Nikita Mishra, Huazhe Zhang, John D. Lafferty, and Henry Hoffmann. A probabilistic graphical model-based approach for minimizing energy under performance constraints. ACM SIGARCH Computer Architecture News, 43(1):267–281, March 2015. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Mishra:2015:PGMb

[MZLH15b] Nikita Mishra, Huazhe Zhang, John D. Lafferty, and Henry Hoffmann. A probabilistic graphical model-based approach for minimizing energy under performance constraints. ACM SIGPLAN Notices, 50(4):267–281, April 2015. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Manrubia:1999:TDS

[MZS99] Susanna C. Manrubia, Damián H. Zanette, and Ricard V. Solé. Transient dynamics and scaling phenomena in urban growth. Fractals, 7(1):1–8, March 1999. CODEN FRACEG.

ISSN 0218-348X. URL http://www.worldscientific.com/
doi/abs/10.1142/S0218348X99000025.

Nagasaka:1984:BL

[Nag84] Kenji Nagasaka. On Benford’s law. Annals of the Institute of Statistical Mathematics (Tokyo), 36(1):337–352, ???? 1984. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052

(electronic). URL http://link.springer.com/article/10.
1007/BF02481974.

Nagasaka:2005:BLD

[Nag05] Kenji Nagasaka. Benford’s law with different basis. In G. Barat, H. Daudé, M. Laurent, P. Liardet, S. Louboutin, and A. Thomas, editors, Journées Arithmétiques XXIV, Marseilles, July 4–8, 2005, Université de Provence and Université de la Méditerranée and CNRS Marseilles, France, pages 1–2. ????, ????, 2005. ISBN ???? LCCN ???? URL http://

atlas-conferences.com/c/a/r/e/47.htm.

Naldi:2003:CIZ

[Nal03] M. Naldi. Concentration indices and Zipf’s law. Economics Letters, 78(3):329–334, March 2003. CODEN ECLEDS. ISSN 0165-1765 (print), 1873-7374 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165176502002513.

NunesAmaral:1998:PLS

[NBH^{+}98] Luís A. Nunes Amaral, Sergey V. Buldyrev, Shlomo Havlin, Michael A. Salinger, and H. Eugene Stanley. Power law scaling for a system of interacting units with complex internal structure. Physical Review Letters, 80(7):1385–1388, February 16, 1998. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145. URL http://link.aps.

org/doi/10.1103/PhysRevLett.80.1385.

Newbold:2010:SBE

[NCT10] Paul Newbold, William L. Carlson, and Betty Thorne. Statistics for Business and Economics. Pearson, Boston, MA, USA, seventh edition, 2010. ISBN 0-13-507248-4. 986 + 16 pp. LCCN QA276.18 .N49 2010.

Newbold:2013:SBE

[NCT13] Paul Newbold, William L. (William Lee) Carlson, and Betty Thorne. Statistics for Business and Economics. Pearson Education, Harlow, Essex, UK, eighth edition, 2013. ISBN 0-273-76706-2. 792 pp. LCCN HF1017 .N48 2013.

Nisbet:2009:CFD

[NEM09] Robert Nisbet, John Elder, and Gary Miner. Chapter 17 — fraud detection. In Handbook of Statistical Analysis and Data Mining Applications, pages 347–

361. Academic Press, Boston, 2009. ISBN 0-12-374765-1.

URL http://www.sciencedirect.com/science/article/
pii/B9780123747655000176.

Newcomb:1881:NFU

[New81] Simon Newcomb. Note on the frequency of use of the different digits in natural numbers. American Journal of Mathematics, 4(1–4):39–40, 1881. CODEN AJMAAN. ISSN 0002-9327 (print), 1080-6377 (electronic). URL http://www.jstor.

org/stable/2369148.

Newman:2005:PLP

[New05] M. E. J. Newman. Power laws, Pareto distributions and Zipf’s law. Contemporary Physics, 46(5):323–351, September 2005. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Nabeshima:2004:ZLP

[NG04] Terutaka Nabeshima and Yukio-Pegio Gunji. Zipf’s law in phonograms and Weibull distribution in ideograms: comparison of English with Japanese. Biosystems (A6E), 73 (2):131–139, February 20, 2004. CODEN BSYMBO. ISSN 0303-2647 (print), 1872-8324 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0303264703002260.

Newby:2003:OSS

[NGJ03] Gregory B. Newby, Jane Greenberg, and Paul Jones. Open source software development and Lotka’s Law: Bibliometric patterns in programming. Journal of the American Society for Information Science and Technology, 54(2):169–178, January 15, 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic). See comments [Bur04].

Nicolaisen:2007:PPB

[NH07] Jeppe Nicolaisen and Birger Hjørland. Practical potentials of Bradford’s law: a critical examination of the received view. Journal of Documentation, 63(3):359–377, ???? 2007. CO-DEN JDOCAS. ISSN 0022-0418 (print), 1758-7379 (electronic).

Neubert:2016:IPP

[NHD16] T. Neubert, M. Hildebrandt, and J. Dittmann. Image preprocessing detection: Evaluation of Benford’s Law, spatial and frequency domain feature performance. In 2016 First International Workshop on Sensing, Processing and Learning for Intelligent Machines (SPLINE), pages 1–5. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, July 2016.

Nicholls:1987:BCE

[Nic87a] Paul Travis Nicholls. Brief communications: Estimation of Zipf parameters. Journal of the American Society for Information Science, 38(6):443–445, November 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Nicholls:1987:EZP

[Nic87b] Paul Travis Nicholls. Estimation of Zipf parameters. Journal of the American Society for Information Science, 38(6):443– 445, 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Nicolis:1988:PBZ

[Nic88] John S. Nicolis. On the parallel between Zipf’s law and 1/f processes in chaotic systems possessing coexisting attractors. In Synergetics, order and chaos (Madrid, 1987), pages 723– 734. World Sci. Publ., Teaneck, NJ, USA, 1988.

Nicholls:1989:RBM

[Nic89] Paul Travis Nicholls. Research: Bibliometric modeling processes and the empirical validity of Lotka’s Law. Journal of the American Society for Information Science, 40(6):379–385, November 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Nigrini:1992:DIE

[Nig92] Mark John Nigrini. The detection of income tax evasion through an analysis of digital distributions. Ph.D. dissertation, Department of Accounting/Business Law, University of Cincinnati, Cincinnati, OH, USA, August 28, 1992. xvi

+ 309 pp. URL http://search.proquest.com/docview/
304052348.

Nigrini:1993:CBL

[Nig93] Mark Nigrini. Can Benford’s Law be used in forensic accounting? The Balance Sheet, ??(??):7–8, June 1993. CODEN ???? ISSN ????

Nigrini:1994:UDF

[Nig94] Mark J. Nigrini. Using digital frequencies to detect fraud. The White Paper, ??(2):3–6, April/May 1994.

Nigrini:1996:DAR

[Nig96a] Mark J. Nigrini. Digital analysis and the reduction of auditor litigation risk. In M. Ettredge, editor, Proceedings of the 1996 Deloitte & Touche I University of Kansas Symposium on Auditing Problems, pages 69–81. University of Kansas, Lawrence, KS, USA, 1996.

Nigrini:1996:TCA

[Nig96b] Mark J. Nigrini. A taxpayer compliance application of Benford’s Law. Journal of the American Taxation Association, 18 (1):72–91, ???? 1996. CODEN ???? ISSN 0198-9073 (print), 1558-8017 (electronic).

Nigrini:1997:DAT

[Nig97] Mark J. Nigrini. Digital Analysis Tests and Statistics. The Nigrini Institute, Inc., Allen, TX, USA, ???? 1997.

URL http://www.nigrini.com/data_software/Program_
Details_2009.docx.

Nigrini:1998:DIB

[Nig98] Mark J. Nigrini. Dow illustrates Benford’s Law. New York Times, ??(??):F4, August 4, 1998. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://

search.proquest.com/hnpnewyorktimes/docview/109901848/
13C4AB0BB5661D397CE/1.

Nigrini:1999:AVD

[Nig99a] Mark J. Nigrini. Adding value with digital analysis. The Internal Auditor: Journal of the Institute of Internal Auditors, 56(1):21–23, ???? 1999. CODEN ITAUAB. ISSN 0020-5745.

Nigrini:1999:IGY

[Nig99b] Mark J. Nigrini. I’ve got your number: How a mathematical phenomenon can help CPAs uncover fraud and other irregularities. Journal of Accountancy, 187(2): 79–83, May 1999. CODEN JACYAD. ISSN 0021-8448 (print), 1945-0729 (electronic). URL http://www.

aicpa.org/pubs/jofa/may1999/nigrini.htm;   http://
www.journalofaccountancy.com/issues/1999/may/nigrini.
html.

Nigrini:1999:PPF

[Nig99c] Mark J. Nigrini. Peculiar patterns of first digits. IEEE Potentials, 18(2):24–27, April/May 1999. CODEN IEPTDF. ISSN 0278-6648 (print), 1558-1772 (electronic).

Nigrini:2000:CA

[Nig00a] Mark J. Nigrini. Continuous auditing. Preprint, Ernst & Young Center for Auditing Research and Advanced Technology, University of Kansas, Lawrence, KS, USA, August 30,

2000. 23 pp. URL https://aaahq.org/audit/midyear/
01midyear/papers/nigrini_continuous_audit.pdf.

Nigrini:2000:DAT

[Nig00b] Mark J. Nigrini. Digital analysis tests and statistics: using digit and number patterns and Benford’s law to detect errors, biases, fraud, irregularities, and processing inefficiencies. Nigrini Institute Inc., Allen, TX, USA, 2000. 242 pp. LCCN ????

Nigrini:2000:DAU

[Nig00c] Mark John Nigrini. Digital analysis using Benford’s law: tests and statistics for auditors. Global Audit Publications, Vancouver, BC, Canada, second edition, 2000. ISBN 189449709 (??invalid ISBN??). xvii + 278 pp. LCCN ????

Nigrini:2001:DAU

[Nig01] Mark Nigrini. Digital analysis using Benford’s law: Tests and statistics for auditors. EDPACS, 28(9):1–2, ???? 2001. CO-DEN EDPCDF. ISSN 0736-6981 (print), 1936-1009 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1201/1079/43266.28.9.20010301/30389.4.

Nigrini:2005:ACI

[Nig05] Mark Nigrini. An assessment of the change in the incidence of earnings management around the Enron–Andersen episode. Review of Accounting and Finance, 4(1):92–110, ???? 2005. ISSN 1475-7702 (print), 1758-7700 (elec-

tronic). URL http://www.emeraldinsight.com/journals.
htm?issn=1475-7702&volume=4&issue=1&articleid=1657234.

Nigrini:2011:BLA

[Nig11a] Mark J. Nigrini. Benford’s Law: Assessing conformity. In Forensic analytics: methods and techniques for forensic accounting investigations [Nig11e], pages 109–129. ISBN 0-470-89046-0, 1-118-08763-1 (e-book), 1-118-08766-6 (e-book), 1-118-08768-2 (e-book), 1-118-38679-5 (e-book). LCCN

HV6768 .N54 2011. URL http://catdir.loc.gov/catdir/
enhancements/fy1107/2011007210-d.html;   http://catdir.
loc.gov/catdir/enhancements/fy1107/2011007210-t.html;
 http://catdir.loc.gov/catdir/enhancements/fy1109/
2011007210-b.html.

Nigrini:2011:BLB

[Nig11b] Mark J. Nigrini. Benford’s Law: The basics. In Forensic analytics: methods and techniques for forensic accounting investigations [Nig11e], pages 85–108. ISBN 0-470-89046-0, 1-118-08763-1 (e-book), 1-118-08766-6 (e-book), 1-118-08768-2 (e-book), 1-118-38679-5 (e-book). LCCN

HV6768 .N54 2011. URL http://catdir.loc.gov/catdir/
enhancements/fy1107/2011007210-d.html;   http://catdir.
loc.gov/catdir/enhancements/fy1107/2011007210-t.html;
 http://catdir.loc.gov/catdir/enhancements/fy1109/
2011007210-b.html.

Nigrini:2011:BLN

[Nig11c] Mark J. Nigrini. Benford’s Law: The number duplication and last-two digits tests. In Forensic analytics: methods and techniques for forensic accounting investigations [Nig11e], pages 153–172. ISBN 0-470-89046-0, 1-118-08763-1 (e-book), 1-118-08766-6 (e-book), 1-118-08768-2 (e-book), 1-118-38679-5 (e-book). LCCN

HV6768 .N54 2011. URL http://catdir.loc.gov/catdir/
enhancements/fy1107/2011007210-d.html;   http://catdir.
loc.gov/catdir/enhancements/fy1107/2011007210-t.html;
 http://catdir.loc.gov/catdir/enhancements/fy1109/
2011007210-b.html.

Nigrini:2011:BLS

[Nig11d] Mark J. Nigrini. Benford’s Law: The second-order and summation tests. In Forensic analytics: methods and techniques for forensic accounting investigations [Nig11e], pages 130–152. ISBN 0-470-89046-0, 1-118-08763-1 (e-book), 1-118-08766-6 (e-book), 1-118-08768-2 (e-book), 1-118-38679-5 (e-book).

LCCN HV6768 .N54 2011. URL http://catdir.loc.gov/
catdir/enhancements/fy1107/2011007210-d.html;   http:
//catdir.loc.gov/catdir/enhancements/fy1107/2011007210-
t.html;   http://catdir.loc.gov/catdir/enhancements/
fy1109/2011007210-b.html.

Nigrini:2011:FAM

[Nig11e] Mark J. (Mark John) Nigrini. Forensic analytics: methods and techniques for forensic accounting investigations. Wiley corporate F and A. Wiley, New York, NY, USA, 2011. ISBN 0-470-89046-0, 1-118-08763-1 (e-book), 1-118-08766-6 (e-book), 1-118-08768-2 (e-book), 1-118-38679-5 (e-book). xvi + 463 pp.

LCCN HV6768 .N54 2011. URL http://catdir.loc.gov/
catdir/enhancements/fy1107/2011007210-d.html;   http:
//catdir.loc.gov/catdir/enhancements/fy1107/2011007210-
t.html;   http://catdir.loc.gov/catdir/enhancements/
fy1109/2011007210-b.html.

Nigrini:2012:BLA

[Nig12] Mark J. (Mark John) Nigrini. Benford’s Law: Applications for Forensic Accounting, Auditing, and Fraud Detection. Wiley, New York, NY, USA, 2012. ISBN 1-118-15285-9 (hardcover), 1-118-28226-4, 1-118-28284-1, 1-118-28686-3. xviii + 330 pp. LCCN HV6691 .N54 2012.

URL http://www.scribd.com/doc/86733139/Benford-s-
Law-Applications-for-Forensic-Accounting-Auditing-
and-Fraud-Detection.

Nigrini:2015:DFE

[Nig15] Mark Nigrini. Detecting fraud and errors using Benford’s Law. In Miller [Mil15a], pages 191–211. ISBN 0-691-14761-2 (hardcover), 1-4008-6659-6 (e-book). LCCN QA273.6 .B46 2015.

URL http://press.princeton.edu/titles/10527.html.

Nillsen:2010:RRD

[Nil10] Rodney Victor Nillsen. Randomness and recurrence in dynamical systems: a real analysis approach, volume 31 of The Carus mathematical monographs. Mathematical Association of America, Washington, DC, USA, 2010. ISBN 0-88385-043-

5. xviii + 357 pp. LCCN QA614.8 .N55 2010.

Nitsch:2005:ZZ

[Nit05] Volker Nitsch. Zipf zipped. Journal of Urban Economics, 57(1):86–100, January 15, 2005. CODEN JUECDW. ISSN 0094-1190 (print), 1095-9068 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0094119004000981.

Nadarajah:2007:CPG

[NK07] Saralees Nadarajah and Samuel Kotz. On the convolution of Pareto and gamma distributions. Computer Networks (Amsterdam, Netherlands: 1999), 51(12):3650–3654, August 22, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Nguyen:2003:DPL

[NKL03] H. T. Nguyen, V. Kreinovich, and L. Longpre. Dirty pages of logarithm tables, lifetime of the universe, and subjective (fuzzy) probabilities on finite and infinite intervals. IEEE Transactions on Fuzzy Systems, 1:67–73, 2003. CODEN IEFSEV. ISSN 1063-6706 (print), 1941-0034 (electronic). The 12th IEE International Conference on Fuzzy Systems. FUZZ’03.

Nguyen:2004:DPL

[NKL04] Hung T. Nguyen, Vladik Kreinovich, and Luc Longpré. Dirty pages of logarithm tables, lifetime of the universe, and (subjective) probabilities on finite and infinite intervals. Reliable Computing = Nadezhnye vychisleniia, 10(2):83–106, April 2004. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.springer.

com/article/10.1023/B%3AREOM.0000015848.19449.12/;
 http://www.springerlink.com/openurl.asp?genre=article&
issn=1385-3139&volume=10&issue=2&spage=83;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=10&issue=2&spage=83-106.

Noll:2015:FAO

[NKRS15] Tobias Nöll, Johannes Köhler, Gerd Reis, and Didier Stricker. Fully automatic, omnidirectional acquisition of geometry and appearance in the context of cultural heritage preservation. Journal on Computing and Cultural Heritage (JOCCH), 8 (1):2:1–2:??, February 2015. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Nagasaka:1990:BLL

[NKS90] Kenji Nagasaka, Shigeru Kanemitsu, and Jau-Shyong Shiue. Benford’s law: the logarithmic law of first digit. In Number theory, Vol. I (Budapest, 1987), volume 51 of Colloq. Math. Soc. János Bolyai, pages 361–391. North-Holland Publishing Co., Amsterdam, The Netherlands, 1990.

Niola:2012:RRA

[NKZ12] Vincenzo Niola, Michel Kadoch, and Alexander Zemliak, editors. Recent researches in automatic control and electronics: proceedings of the 14th International conference on automatic control, modelling and simulation (ACMOS ’12): proceedings of the 11th International conference on microelectronics, nanoelectronics, optoelectronics (MINO ’12): Saint Malo and Mont Saint-Michel, France, April 2–4, 2012. WSEAS, ???, Greece, 2012. ISBN 1-61804-080-4 (paperback). LCCN ????

Nigrini:1997:UBL

[NM97] Mark J. Nigrini and L. J. Mittermaier. The use of Benford’s Law as an aid in analytical procedures. Auditing: A Journal of Practice & Theory, 16(2):52–67, ???? 1997. CODEN ???? ISSN 0278-0380 (print), 1558-7991 (electronic). URL http:/

/aaapubs.org/loi/ajpt.

Nigrini:2006:BLA

[NM06] M. J. Nigrini and Steven J. Miller. Benford’s law applied to hydrology data — results and relevance to other geophysical data. Preprint, ????, ????, ???? 2006.

Nigrini:2007:BLA

[NM07a] Mark J. Nigrini and Steven J. Miller. Benford’s Law applied to hydrology data — results and relevance to other geophysical data. Mathematical Geology, 39(5):469–490, July 2007. CODEN MATGED. ISSN 0882-8121 (print), 1573-8868

(electronic). URL http://link.springer.com/article/10.
1007/s11004-007-9109-5;   http://www.springerlink.
com/content/m2274ng70435u3x6/.

Nye:2007:PEN

[NM07b] J. Nye and C. Moul. The political economy of numbers: On the application of Benford’s Law to international macroeconomic statistics. B. E. Journal of Macroeconomics, 7(1):??, ???? 2007. CODEN ???? ISSN 1935-1690. URL http://

www.bepress.com/bejm/vol7/iss1/art17.

Nigrini:2009:DDU

[NM09] Mark J. Nigrini and Steven J. Miller. Data diagnostics using second-order tests of Benford’s Law. Auditing: A Journal of Practice & Theory, 28(2):305–324, November 2009. CODEN ???? ISSN 0278-0380 (print), 1558-7991 (elec-

tronic). URL http://aaapubs.org/loi/ajpt;   http://
link.aip.org/link/AJPTXX/v28/i2/p305/s1.

Neves:2021:ABL

[NNF21] Gilberto A. Neves, Catarina S. Nunes, and Paula Odete Fernandes. Application of Benford’s law to the tourism demand: the case of the Island of Sal, Cape Verde. In Optimization, Learning Algorithms and Applications: First International Conference, OL2A 2021, Bragança, Portugal, July 19– 21, 2021, Revised Selected Papers, volume 1488 of Commun. Comput. Inf. Sci., pages 587–598. Springer, Cham, Switzerland, 2021. ISBN 3-030-91884-X; 3-030-91885-8. URL https:

//doi.org/10.1007/978-3-030-91885-9_43.

Nicolis:1989:CDM

[NNN89] G. Nicolis, C. Nicolis, and John S. Nicolis. Chaotic dynamics, Markov partitions, and Zipf’s law. Journal of Statistical Physics, 54(3–4):915–924, February 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01019781.

Norris:2013:NRA

[Nor13] Pippa Norris. The new research agenda studying electoral integrity. Electoral Studies, 32(4):563–575, 2013. ISSN 0261-3794 (print), 1873-6890 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0261379413001157.

Special Symposium: The new research agenda on electoral integrity.

Novikov:2010:APB

[Nov10] V. P. Novikov. Aspectual principle, Benford’s Law and Russell’s Paradox. ArXiv e-prints, December 2010. CODEN

???? ISSN ???? URL http://adsabs.harvard.edu/abs/
2010arXiv1012.5537N;   http://arxiv.org/abs/1012.5537.

Newman:2002:ME

[NP02] M. E. J. (Mark E. J.) Newman and Richard G. Palmer. Modeling extinction. Santa Fe Institute studies in the sciences of complexity. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2002. ISBN 0-19-515946-2 (paperback), 0-19-515945-4 (hardcover). xii + 102 pp. LCCN QH78 .N48 2002.

Nebel:2012:DHG

[NP12] J.-C. Nebel and S. Pezzulli. Distribution of human genes observes Zipf’s Law. Kingston University Research & Innovation Reports (KURIR), 8(??):1–9, ???? 2012. ISSN 1749-

5652. URL http://staffnet.kingston.ac.uk/~ku33185/
Papers/PDFformat/Kurir12.pdf.

Ni:2008:BLH

[NR08] D. Ni and Z. Ren. Benford’s law and half-lives of unstable nuclei. European Physical Journal A: Hadrons and Nuclei, 38:251–255, December 2008. CODEN EPJAFV. ISSN 1434-6001 (print), 1434-601X (electronic). URL http://adsabs.

harvard.edu/abs/2008EPJA...38..251N.

Nagasaka:1987:BLL

[NS87] Kenji Nagasaka and Jau-Shyong Shiue. Benford’s law for linear recurrence sequences. Tsukuba J. Math., 11(2):341–351, 1987. CODEN ???? ISSN 0387-4982.

Newman:1996:ASC

[NS96] M. E. J. Newman and Kim Sneppen. Avalanches, scaling, and coherent noise. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 54(6): 6226–6231, December 1996. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-2376 (electronic). URL http://link.aps.

org/doi/10.1103/PhysRevE.54.6226.

Neumann:2005:CAE

[NSGP05] László Neumann, Mateu Sbert, Bruce Gooch, and Werner Purgathofer, editors. Computational aesthetics 2005: Eurographics Workshop on Computational Aesthetics in Graphics, Visualization and Imaging, Girona, Spain, 18–20 May 2005, Computational aesthetics. Eurographics Association, Aire-la-Ville, Switzerland, 2005. ISBN 3-905673-27-4. ISSN 1816-0859.

Nicolis:1989:PBZ

[NT89] John S. Nicolis and Ichiro Tsuda. On the parallel between Zipf’s law and 1/f processes in chaotic systems possessing coexisting attractors. Progress of Theoretical Physics, 82(2): 254–274, 1989. CODEN PTPKAV. ISSN 0033-068X (print), 1347-4081 (electronic).

Nelson:2022:BLM

[NWNG22] Sybil Prince Nelson, Brian Wickman, Jack Null, and Eric Gazin. Benford’s law and music note frequencies. In Mathematics and computation in music, volume 13267 of Lecture Notes in Comput. Sci., pages 383–389. Springer, Cham, Switzerland, 2022. ISBN 978-30310-7-0-1-4-3; 978-30310-7-0-

1-5-0. URL https://doi.org/10.1007/978-3-031-07015-
0_34.

Ni:2009:BLD

[NWR09] Dong-Dong Ni, Lai Wei, and Zhong-Zhou Ren. Benford’s Law and β-decay half-lives. Communications in Theoretical Physics, 51(4):713–716, April 2009. CODEN CT-PHDI. ISSN 0253-6102 (print), 1572-9494 (electronic). URL

http://adsabs.harvard.edu/abs/2009CoTPh..51..713N;
 http://stacks.iop.org/0253-6102/51/i=4/a=25.

Ozer:2013:BLD

[OB13] G. Ozer and B. Babacan. Benford’s Law and digital analysis: Application on Turkish banking sector. Business and Economics Research Journal, 4(1):29–41, 2013. ISSN 1309-2448.

Orlov:1983:SSZ

[OC83] Yu. K. Orlov and R. Ya. Chitashvili. The statistical significance of Zipf’s distribution. Soobshch. Akad. Nauk Gruzin. SSR, 109(3):505–508, 1983. CODEN ???? ISSN 0132-1447.

Obradovic:2015:GFT

[OJM15] Marko Obradović, Milan Jovanović, and Bojana Milosević. Goodness-of-fit tests for Pareto distribution based on a characterization and their asymptotics. Statistics: A Journal of Theoretical and Applied Statistics, 49(5):1026–1041, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ogasawara:2006:ZLH

[OKO06] Osamu Ogasawara, Shoko Kawamoto, and Kousaku Okubo. Zipf’s law and human transcriptomes: an explanation with an evolutionary model. Comptes Rendus Biologies, 326(10–11):1097–1101, July 1, 2006. CODEN CR-BOCM. ISSN 1631-0691 (print), 1768-3238 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1631069103002294.

Oleksy:2010:DMB

[Ole10] Marius Oleksy. Data Mining und Benford’s Law als Controllinginstrumente. (German) [Data Mining and Benford’s Law as controlling instruments], volume 45 of Wismarer Schriften zu Management und Recht. Europäischer Hochschulverlag, Bremen, Germany, 2010. ISBN 3-86741-405-X. 181 (est.)

pp. LCCN ???? URL http://www.amazon.de/Data-Mining-
Benfords-Law-Controllinginstrumente/dp/386741405X.

Oleksy:2014:DMB

[Ole14] Marius Oleksy. Data Mining und Benford’s Law als Controllinginstrumente. (German) [Data mining and Benford’s Law as controlling instruments], volume 45 of Wismarer Schriften zu Management und Recht. EHV Academicpress, Bremen, Germany, 2014. ISBN 3-86741-923-X (paperback), 3-86741-928-0. 220 pp. LCCN ???? URL http://deposit.d-

nb.de/cgi-bin/dokserv?id=4627278\%26prov=M\%26dok\
_var=1\%26dok\_ext=htm.

Olver:2014:MIM

[Olv14] F. W. J. Olver. Mathematics that has intrigued me. Analysis and Applications (Singapore), 12(4):341–354, 2014. ISSN 0219-5305 (print), 1793-6861 (electronic). Lecture delivered at Asymptotics and Applied Analysis, Conference in Honor of Frank W. J. Olver’s 75th Birthday, January 10–14, 2000, San Diego State University, San Diego, California.

Osareh:2011:LLA

[OM11] Farideh Osareh and Esmaeel Mostafavi. Lotka’s Law and authorship distribution in computer science using Web of Science (WoS) during 1986–2009. Collnet Journal of Scientometrics and Information Management, 5(2):171–183, 2011. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Orita:2010:UBL

[OMNO10] Masaya Orita, Ayako Moritomo, Tatsuya Niimi, and Kazuki Ohno. Use of Benford’s Law in drug discovery data. Drug Discovery Today, 15(9–10):328–331, May 2010. CO-DEN DDTOFS. ISSN 1359-6446 (print), 1878-5832 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1359644610000796.

Oteo:2007:DPE

[OR07] J. A. Oteo and J. Ros. Double precision errors in the logistic map: Statistical study and dynamical interpretation. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 76:036214, September 2007. CODEN PLEEE8. ISSN 1539-3755 (print),

1550-2376 (electronic). URL http://link.aps.org/doi/10.
1103/PhysRevE.76.036214.

Oransky:2017:TCT

[Ora17] Ivar Oransky. Two in 100 clinical trials in eight major journals likely contain inaccurate data: Study. Web site article., June

5, 2017. URL https://retractionwatch.com/2017/06/05/
two-100-clinical-trials-eight-major-journals-likely-
contain-inaccurate-data-study/.

Orlov:1970:GZM

[Orl70] Ju. K. Orlov. A generalization of the Zipf–Mandelbrot law. Sakharthvelos SSR Mecnierebatha Akademiis Moambe, 57:37– 40, 1970. CODEN ???? ISSN ????

Orlov:1976:CBP

[Orl76] Ju. K. Orlov. The connection between the Pareto distribution and the generalized Zipf–Mandelbrot law. Sakharthvelos SSR Mecnierebatha Akademiis Moambe, 83(1):57–60, 1976. CO-DEN ???? ISSN ????

Okuyama:1999:ZLI

[OTT99] K. Okuyama, M. Takayasu, and H. Takayasu. Zipf’s law in income distribution of companies. Physica A: Statistical Mechanics and its Applications, 269(1):125–131, July 1, 1999. CODEN PHYADX. ISSN 0378-4371 (print), 1873-

2119 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378437199000862.

Oluic-Vukovic:1993:LEW

[OV93] Vesna Oluić-Vuković. Letters to the Editor: Why has Bradford’s Law been an elusive phenomenon so far? Journal of the American Society for Information Science, 44(3):182–183, April 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Oluic-Vukovic:1995:LED

[OV95] Vesna Oluić-Vuković. Letters to the Editor: On the dynamic behavior of Bradford’s Law. Journal of the American Society for Information Science, 46(10):798–799, December 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Oluic-Vukovic:1997:BDC

[OV97] Vesna Oluić-Vuković. Bradford’s distribution: From the classical bibliometric “law” to the more general stochastic models. Journal of the American Society for Information Science, 48 (9):833–842, September 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Odueke:2012:TFA

[OW12]  Adeola Odueke and George R. S. Weir.  Triage in

forensic accounting using Zipf’s Law. In George R. S. Weir and A. Al-Nemrat, editors, Issues in Cybercrime, Security and Digtal Forensics, pages 33–43. University of Strathclyde Publishing, Glasgow, UK, 2012. ISBN

????  LCCN ????  URL http://www.cis.strath.
ac.uk/cis/research/publications/papers/strath_cis_
publication_2590.pdf.

Pain:2008:BLC

[Pai08] Jean-Christophe Pain. Benford’s law and complex atomic spectra. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 77(1):012102:1– 012102:3, January 2008. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-2376 (electronic). URL http://adsabs.

harvard.edu/abs/2008PhRvE..77a2102P;   http://pre.
aps.org/abstract/PRE/v77/i1/e012102.

Pain:2013:RSA

[Pai13] Jean-Christophe Pain. Regularities and symmetries in atomic structure and spectra. High Energy Density Physics, 9(3): 392–401, September 2013. CODEN HEDPBW. ISSN 1574-1818 (print), 1878-0563 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1574181813000475.

Pao:1985:LLT

[Pao85] Miranda Lee Pao. Lotka’s law: a testing procedure. Information Processing and Management, 21(4):305–320, ???? 1985. CODEN IPMADK. ISSN 0306-4573 (print), 1873-5371 (electronic).

Pao:1986:EEL

[Pao86a] Miranda Lee Pao. An empirical examination of Lotka’s Law. Journal of the American Society for Information Science, 37 (1):26–33, January 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Pao:1986:REE

[Pao86b] Miranda Lee Pao. Research: An empirical examination of Lotka’s Law. Journal of the American Society for Information Science, 37(1):26–33, January 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Pareto:1896:CEP

[Par97] Vilfredo Pareto. Cours d’économie politique professé a l’université de Lausanne. (French) [Course on political economy given at the University of Lausanne]. F. Rouge, Lausanne, Switzerland, 1896–1897. ???? pp. LCCN HB173 .P22.

Partridge:1975:DDW

[Par75] D. Partridge. A dynamic database which automatically removes unwanted generalisation for the efficient analysis of language features that exhibit a disparate frequency distribution. The Computer Journal, 18(1):43–48, February 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_01/tiff/43.tif; http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_
01/tiff/44.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_01/tiff/45.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_18/Issue_01/
tiff/46.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_01/tiff/47.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_18/Issue_01/
tiff/48.tif.

Pavlov:1981:DFF

[Pav81] A. I. Pavlov. On the distribution of fractions and F. Benford’s law. Izvestiia Akademii Nauk. USSR Seriia Matematicheskaia (Moscow), 45(4):760–774, 927, 1981. CODEN IZAMAT. ISSN 0373-2436.

Pavlov:1982:DFP

[Pav82] A. I. Pavlov. On the distribution of fractional parts and Benford’s Law. Mathematics of the USSR — Izvestiya, 19(1):65– 77, ???? 1982. CODEN MUSIAE. ISSN 0025-5726 (print),

2169-5075 (electronic). URL http://stacks.iop.org/0025-
5726/19/i=1/a=A05.

Paech:2007:ZLN

[PBP07] Kerstin Paech, Wolfgang Bauer, and Scott Pratt. Zipf’s law in nuclear multifragmentation and percolation theory. Physical Review C (Nuclear Physics), 76(??):054603, November 15, 2007. CODEN PRVCAN. ISSN 0556-2813 (print),

1089-490X, 1538-4497. URL http://link.aps.org/doi/10.
1103/PhysRevC.76.054603.

Pedrotti:1994:ALU

[Ped94] A. Pedrotti. Analysis of a list-update strategy. Information Processing Letters, 52(3):115–121, November 11, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Peng:2010:ZLC

[Pen10] Guohua Peng. Zipf’s law for Chinese cities: Rolling sample regressions. Physica A: Statistical Mechanics and its Applications, 389(18):3804–3813, September 15, 2010. CO-DEN PHYADX. ISSN 0378-4371 (print), 1873-2119 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037843711000381X.

Perline:1982:EVM

[Per82] Richard K. Perline. An extreme value model of weakly harmonic (Pareto–Zipf type) laws. Thesis (Ph.D.), Department of Behavioral Sciences, The University of Chicago, Chicago, IL, USA, December 1982. 129 pp. URL http://search.

proquest.com/docview/251796701.

Peritz:1990:BDB

[Per90] Bluma C. Peritz. A Bradford distribution for bibliometrics. Scientometrics, 18(5–6):323–329, May 1990. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02020148.

Perline:1996:ZLC

[Per96] Richard Perline. Zipf’s law, the Central Limit Theorem, and the random division of the unit interval. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 54(1):220–223, July 1, 1996. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-2376

(electronic). URL http://link.aps.org/doi/10.1103/
PhysRevE.54.220.

Perline:2005:SWF

[Per05] Richard Perline. Strong, weak and false inverse power laws. Statistical Science, 20(1):68–88, ???? 2005. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

Perin:2007:CNA

[Pér07] Stephen Périn. The constructal nature of the air traffic system. In Constructal Theory of Social Dynamics, pages 119– 145. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. URL http://link.

springer.com/chapter/10.1007/978-0-387-47681-0_6.

Perc:2010:ZLL

[Per10] Matjaž Perc. Zipf’s law and log-normal distributions in measures of scientific output across fields and institutions:

40 years of Slovenia’s research as an example. Journal of Informetrics, 4(3):358–364, July 1, 2010. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157710000234.

Petersen:1966:RMT

[Pet66] Gordon Marshall Petersen. Regular Matrix Transformations. McGraw-Hill, New York, NY, USA, 1966. viii + 142 pp. LCCN QA263 .P415 1966.

Petruszewycz:1973:LLE

[Pet73] M. Petruszewycz. L’histoire de la loi d’Estoup–Zipf: documents. (French) [The history of the Estoup–Zipf law]. Math. Sci. Humaines, ??(44):41–56, 1973. CODEN ???? ISSN 0025-5815.

Peters:1981:EFB

[Pet81a] James V. Peters. An equivalent form of Benford’s Law. Fibonacci Quarterly, 19(1):74–76, February 1981. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/19-1/peters.pdf.

Peters:1981:SCC

[Pet81b] James V. Peters. Short communications: a combinatoric proof of Benford’s law. Aequationes Mathematicae, 23(1):122–123, 1981. CODEN AEMABN. ISSN 0001-9054 (print), 1420-8903

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02188018.

Peter:2003:ADE

[Pet03] Manfred Peter. The asymptotic distribution of elements in automatic sequences. Theoretical Computer Science, 301(1 3): 285–312, 2003. CODEN TCSCDI. ISSN 0304-3975 (print),

1879-2294 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030439750200587X.

Petek:2008:PNH

[Pet08] Marija Petek. Personal name headings in COBIB: Testing Lotka’s law. Scientometrics, 75(1):175–188, April 2008. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/s11192-007-1829-2.

Perez-Gonzalez:2007:GBL

[PGHA07] F. Pérez-González, G. L. Heileman, and C. T. Abdallah. A generalization of Benford’s Law and its application to images. In 2007 European Control Conference (ECC), pages 3613– 3619. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, July 2007.

Peng:2010:SFW

[PGLL10] Fei Peng, Re-Si Guo, Chang-Tsun Li, and Min Long. A semi-fragile watermarking algorithm for authenticating 2D CAD engineering graphics based on log-polar transformation. Computer-Aided Design, 42(12):1207–1216, 2010. CO-DEN CAIDA5. ISSN 0010-4485 (print), 1879-2685 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010448510001491.

Perez-Gonzalez:2015:ABL

[PGQA^{+}15] Fernando Pérez-González, Tu-Thach Quach, Chaouki T. Abdallah, Gregory L. Heileman, and Steven J. Miller. Application of Benford’s Law to images. In Miller [Mil15a], pages 338– 370. ISBN 0-691-14761-2 (hardcover), 1-4008-6659-6 (e-book).

LCCN QA273.6 .B46 2015. URL http://press.princeton.
edu/titles/10527.html.

Phatarfod:2013:SAB

[Pha13] Ravi Phatarfod. Some aspects of the Benford law of leading significant digits. The Mathematical Scientist, 38(2): 73–85, 2013. ISSN 0312-3685 (print), 1475-6080 (elec-

tronic). URL http://appliedprobability.org/content.
aspx?Group=tms&Page=tmsabstracts38_2#one.

Pinkham:1961:DFS

[Pin61] Roger S. Pinkham. On the distribution of first significant digits. Annals of Mathematical Statistics, 32(4):1223–1230, December 1961. CODEN AASTAD. ISSN 0003-4851. URL

http://www.jstor.org/stable/2237922.

Piva:2013:OIF

[Piv13] Alessandro Piva. An overview on image forensics. ISRN Signal Processing, 2013(??):496701, ???? 2013. ISSN 2090-5041 (print), 2090-505X (electronic). URL http://www.isrn.com/

journals/sp/2013/496701/.

Praunlich:1978:BDN

[PK78] Peter Praunlich and Michael Kroll. Bradford’s distribution: a new formulation. Journal of the American Society for Information Science, 29(2):51–55, March 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Posch:2006:ADP

[PK06] Peter N. Posch and Welf A. Kreiner. Analysing digits for portfolio formation and index tracking. Journal of Asset Management, 7(1):69–80, May 2006. CODEN ???? ISSN 1470-8272 (print), 1479-179X (electronic).

Pontigo:1986:QAB

[PL86] J. Pontigo and F. W. Lancaster. Qualitative aspects of the Bradford distribution. Scientometrics, 9(1–2):59–70, January 1986. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02016608.

Li:2004:DBL

[pLCjW04] Zhi peng Li, Lin Cong, and Hua jia Wang. Discussion on Benford’s Law and its application. ArXiv e-prints,

pages 1–13, 2004. CODEN ????  ISSN ????  URL

http://adsabs.harvard.edu/abs/2004math......8057L;
 http://arxiv.org/abs/math/0408057.

Piqueira:1999:ZLO

[PMdM^{+}99] J. R. C. Piqueira, L. H. A. Monteiro, T. M. C. de Magalhães, R. T. Ramos, R. B. Sassi, and E. G. Cruz. Zipf’s Law organizes a psychiatric ward. Journal of Theoretical Biology, 198(3):439–443, June 7, 1999. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519399909232.

Pocheau:1998:SRS

[Poc98] A. Pocheau. Scale ratios, statistical symmetries and intermittency. Europhysics Letters, 43(4):410–415, August 1998. CODEN EULEEJ. ISSN 0295-5075 (print), 1286-4854 (electronic).

Pocheau:2006:SDL

[Poc06] A. Pocheau. The significant digit law: a paradigm of statistical scale symmetries. European Physical Journal B: Condensed Matter and Complex Systems, 49(4):491–511, February 2006. CODEN EPJBFY. ISSN 1434-6028 (print), 1434-6036

(electronic). URL http://link.springer.com/article/10.
1140/epjb/e2006-00084-2.

Poincare:1912:CPL

[Poi12] H. Poincaré. Calcul des Probabilités: Leçons professées pendant le deuxiè me semestre 1893–1894. (French) [Calculation of Probabilities: Lectures from the second semester 1893– 1894]. Gauthier-Villars, Paris, France, 1912. ii + 274 pp. LCCN QA273 .P75.

Pope:1975:BLP

[Pop75] Andrew Pope. Bradford’s Law and the periodical literature of information science. Journal of the American Society for Information Science, 26(4):207–213, July 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Popescu:2002:LNZ

[Pop02] Ioan-Iovitz Popescu. On the Lavelette’s nonlinear Zipf’s law. Web preprint., February 2002. URL http://alpha2.infim.

ro/~ltpd/Zipf's_Law.html.

Posch:2004:BBH

[Pos04a] Peter N. Posch. Benford or not-Benford? How to test for the first digit law. Working paper., Department of Finance, University of Ulm, Ulm, Germany, 2004.

Posch:2004:SSD

[Pos04b] Peter N. Posch. A survey on sequences and distribution functions satisfying Benford’s law. Working paper., Department of Finance, University of Ulm, Ulm, Germany, Octo-

ber 25, 2004. URL http://www.posch.org/paper/posch_
benforddist.pdf.

Posch:2004:ZFB

[Pos04c] Peter N. Posch. Ziffernanalyse in der Fälschungsaufspürung. Benford’s Gesetz und Steuererklärungen in Theorie und Praxis. (German) [digit analysis in fake tracing. Benford’s Law and tax returns in theory and practice]. Arbeitspapier, Abteilung Finanzwirtschaft, University Ulm, Ulm, Germany, October 2004.

Posch:2005:ZTP

[Pos05] Peter N. Posch. Ziffernanalyse in Theorie und Praxis: Testverfahren zur Fälschungsaufspürung mit Benfords Gesetz. Berichte aus der Statistik. Shaker, Aachen, Germany, 2005. ISBN 3-8322-4492-1. iv + 87 + v pp. LCCN ????

EUR 29.80 (DE). URL http://www.gbv.de/dms/hbz/toc/
ht014717371.pdf.

Posch:2010:ZNB

[Pos10] Peter N. Posch. Ziffernanalyse mit dem Newcomb–Benford Gesetz in Theorie und Praxis. (German) [Digit analysis with the Newcomb–Benford law in theory and practice]. Verlag Europäische Wirtschaft, Munich, Germany, second edition, 2010. ISBN ???? 105 (est.) pp. LCCN ????

Potter:1981:LLR

[Pot81] William Gray Potter. Lotka’s Law revisited. Library Trends, 31(2):21–39, Summer 1981. CODEN LIBTA3. ISSN 0024-2594 (print), 1559-0682 (electronic). URL http:/

/citeseerx.ist.psu.edu/viewdoc/summary?doi=10.1.1.
172.9842.

Padmanabhan:2000:CAD

[PQ00] Venkata N. Padmanabhan and Lili Qiu. The content and access dynamics of a busy Web server (poster session). ACM SIGMETRICS Performance Evaluation Review, 28(1):122– 123, June 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Pollack:2022:DSB

[PR22] Paul Pollack and Akash Singha Roy. Dirichlet, Sierpiński, and Benford. Journal of Number Theory, 239(??):352–364, October 2022. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022314X22000099.

Prather:1988:CET

[Pra88] R. E. Prather. Comparison and extension of theories of Zipf and Halstead. The Computer Journal, 31(3):

248–252, June 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.

oup.co.uk/computer_journal/hdb/Volume_31/Issue_03/
tiff/248.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_03/tiff/249.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_31/Issue_03/
tiff/250.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_03/tiff/251.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_31/Issue_03/
tiff/252.tif.

Preece:1981:DFD

[Pre81] D. A. Preece. Distributions of final digits in data. Journal of the Royal Statistical Society. Series D (The Statistician), 30(1):31–60, March 1981. ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.org/stable/2987702.

Paredaens:2009:PTE

[PS09] Jan Paredaens and Jianwen Su, editors. Proceedings of the twenty-eighth ACM SIGMOD-SIGACT-SIGART symposium on Principles of database systems, PODS’09, Providence, Rhode Island, June 29–July 1, 2009. ACM Press, New York, NY 10036, USA, 2009. ISBN 1-60558-553-X. LCCN ????

Pan:2010:EJC

[PSM10] Da Pan, Lijing Shao, and Bo-Qiang Ma. Benford’s Law in statistical physics. Wolfram Online Demonstrations

Projects., 2010. URL http://demonstrations.wolfram.
com/BenfordsLawInStatisticalPhysics/.

Pericchi:2004:LNB

[PT04] Luis Rául Pericchi and David Torres. La Ley de Newcomb– Benford y sus aplicaciones al referendum revocatorio en Venezuela. (Spanish) [the Newcomb–Benford Law and its applications to the recall referendum in Venezuela]. In ????, editor, Third Universidad Simon Bolivar Seminar on Statistical Analyses of the Venezuelan Recall Referendum, pages 1–7. ????, ????, October 1, 2004. URL http://esdata.info/pdf/

pericchi-torres.pdf;   https://sites.google.com/a/
upr.edu/probability-and-statistics/home/techical-reports.

Reporte Técnico no-definitivo 2a. Presented on September 23,

Pericchi:2011:QAD

[PT11] Luis Pericchi and David Torres. Quick anomaly detection by the Newcomb–Benford law, with applications to electoral processes data from the USA, Puerto Rico and Venezuela. Statistical Science, 26(4):502–516, 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:/

/projecteuclid.org/euclid.ss/1330437932.

Pietronero:1998:UDN

[PTTV98] L. Pietronero, E. Tosatti, V. Tosatti, and A. Vespignani. The uneven distribution of numbers in nature. ArXiv e-prints,

page 9, August 27, 1998. URL http://arxiv.org/abs/cond-
mat/9808305.

Pietronero:2001:EUD

[PTTV01] L. Pietronero, E. Tosatti, V. Tosatti, and A. Vespignani. Explaining the uneven distribution of numbers in nature: the laws of Benford and Zipf. Physica A: Statistical Mechanics and its Applications, 293(1–2):297– 304, April 1, 2001. CODEN PHYADX. ISSN 0378-4371 (print), 1873-2119 (electronic). URL http://adsabs.

harvard.edu/abs/2001PhyA..293..297P;   http://www.
sciencedirect.com/science/article/pii/S0378437100006336.

Patil:1986:DFD

[PU86] S. A. Patil and V. R. R. Uppuluri. The distribution of first j digits. College Mathematics Journal, 17(3):240–243, May 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.jstor.org/stable/2686982;
http://www.tandfonline.com/doi/abs/10.1080/07468342.
1986.11972963.

Pustet:2004:ZHH

[Pus04] Regina Pustet. Zipf and his heirs. Language Sciences, 26(1): 1–25, January 1, 2004. CODEN ???? ISSN 0388-0001 (print),

1873-5746 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0388000103000184.

Quarantelli:2004:ISA

[QBP^{+}04] Mario Quarantelli, Karim Berkouk, Anna Prinster, Brigitte Landeau, Claus Svarer, Laszlo Balkay, Bruno Alfano, Arturo Brunetti, Jean-Claude Baron, and Marco Salvatore.

Integrated software for the analysis of brain PET/SPECT studies with partial-volume effect correction. Journal of Nuclear Medicine, 45(2):192–201, February 1, 2004. CO-DEN JNMEAQ. ISSN 0161-5505 (print), 1535-5667 (elec-

tronic). URL http://jnm.snmjournals.org/content/45/
2/192.abstract.

Quick:2003:BLD

[QW03] Reiner Quick and Matthias Wolz. Benford’s Law in deutschen Rechnungslegungsdaten. (German) [Benford’s Law in German accounting data]. Betriebswirtschaftliche Forschung und Praxis, 55(2):208–224, ???? 2003. CODEN ???? ISSN 0340-5370.

Qadir:2010:EJC

[QZH10] G. Qadir, X. Zhao, and A. T. S. Ho. Estimating JPEG2000 compression for image forensics using Benford’s Law. In Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, volume 7723 of Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series. SPIE Optical Engineering Press, Bellingham, WA, USA, April 2010. URL

http://adsabs.harvard.edu/abs/2010SPIE.7723E..14Q.

Qadir:2011:IFG

[QZHC11] G. Qadir, X. Zhao, A. T. Ho, and M. Casey. Image forensic of glare feature for improving image retrieval using Benford’s Law. In 2011 IEEE International Symposium of Circuits and Systems (ISCAS), pages 2661–2664. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, May 2011. ISSN 0271-4302.

Raimi:1969:PDF

[Rai69a] R. Raimi. The peculiar distribution of first significant digits. Scientific American, 221(6):109–120, December 1969. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Raimi:1969:DFS

[Rai69b] R. A. Raimi. On the distribution of first significant figures. American Mathematical Monthly, 76(4):342–348, April 1969. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Raimi:1976:FDP

[Rai76] Ralph A. Raimi. The first digit problem. American Mathematical Monthly, 83(7):521–538, August/September 1976. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/2319349.

Raimi:1985:FDP

[Rai85] Ralph A. Raimi. The first digit phenomenon again. Proceedings of the American Philosophical Society, 129(2):211–219, June 1985. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL http://www.jstor.org/stable/

986989. This paper contains strong criticism of a derivation of Benford’s Law [LG78].

Raidl:2003:AEC

[Rai03] Gunther Raidl, editor. Applications of evolutionary computing: EvoWorkshops 2003: EvoBIO, EvoCOP, EvoIASP, EvoMUSART, EvoROB, and EvoSTIM, Essex, UK, April 14–16, 2003: Proceedings, volume 2611 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. CO-DEN LNCSD9. ISBN 3-540-00976-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.618 .E899

2003. URL http://link.springer-ny.com/link/service/
series/0558/tocs/t2611.htm;   http://www.springerlink.
com/content/978-3-540-00976-4;   http://www.springerlink.
com/openurl.asp?genre=issue&issn=0302-9743&volume=
2611.

Ravikumar:2008:BND

[Rav08] Bala Ravikumar. The Benford–Newcomb distribution and unambiguous context-free languages. International Journal of Foundations of Computer Science (IJFCS), 19(3):717–727, June 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Raqab:2010:PPD

[RAV10] Mohammad Z. Raqab, A. Asgharzadeh, and R. Valiollahi. Prediction for Pareto distribution based on progressively Type-II censored samples. Computational Statistics & Data Analysis, 54(7):1732–1743, July 1, 2010. CO-DEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167947310000563.

Rubino:2015:ILB

[RBXM15] Irene Rubino, Claudia Barberis, Jetmir Xhembulla, and Giovanni Malnati. Integrating a location-based mobile game in the museum visit: Evaluating visitors’ behaviour and learning. Journal on Computing and Cultural Heritage (JOCCH), 8(3):15:1–15:??, May 2015. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Riccioni:2018:RPF

[RC18] Jessica Riccioni and Roy Cerqueti. Regular paths in financial markets: investigating the Benford’s law. Chaos, Solitons & Fractals, 107:186–194, 2018. CODEN CSFOEH. ISSN 0960-0779 (print), 1873-2887 (electronic).

Rybski:2023:ALZ

[RC23] Diego Rybski and Antonio Ciccone. Auerbach, Lotka, and Zipf: pioneers of power-law city-size distributions. Archive for History of Exact Sciences, 77(6):601–613, November 2023. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s00407-023-00314-0.

Rezayan:2010:SRG

[RDFM10] Hani Rezayan, Mahmoud Reza Delavar, Andrew Ulrich Frank, and A. Mansouri. Spatial rules that generate urban patterns: Emergence of the power law in the distribution of axial line length. International Journal of Applied Earth Observation and Geoinformation, 12(5):317–330, 2010. ISSN

0303-2434. URL http://www.sciencedirect.com/science/
article/pii/S0303243410000413.

Reed:2001:PZO

[Ree01] William J. Reed. The Pareto, Zipf and other power laws. Economics Letters, 74(1):15–19, December 20, 2001. CO-DEN ECLEDS. ISSN 0165-1765 (print), 1873-7374 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165176501005249.

Reed:2003:PLI

[Ree03] William J. Reed. The Pareto law of incomes—an explanation and an extension. Physica A: Statistical Mechanics and its Applications, 319(1–4):469–486, 2003. CODEN PHYADX. ISSN 0378-4371 (print), 1873-2119 (electronic).

Ree:2006:PLD

[Ree06] Suhan Ree. Power-law distributions from additive preferential redistributions. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 73(??): 026115, February 15, 2006. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-2376 (electronic). URL http://link.aps.

org/doi/10.1103/PhysRevE.73.026115.

Regazzini:1982:BFL

[Reg82] Eugenio Regazzini. The Benford–Furlan law as a statistical law. Statistica (Bologna), 42(3):351–370, 1982. CODEN STATDJ. ISSN 0390-590X (print), 1973-2201 (electronic).

Regan:2012:BLB

[Reg12] K. W. Regan. Benford’s Law and baseball. Web docu-

ment, July 29, 2012. URL http://rjlipton.wordpress.
com/2012/07/29/benfords-law-and-baseball/.

Rauch:2011:FFE

[RGBE11] Bernhard Rauch, Max Göttsche, Gernot Brähler, and Stefan Engel. Fact and fiction in EU-governmental economic data. German Economic Review, 12(3):243–255, August 2011. ISSN 1465-6485 (print), 1468-0475 (elec-

tronic). URL http://onlinelibrary.wiley.com/doi/10.
1111/j.1468-0475.2011.00542.x/abstract.

Rauch:2014:DVS

[RGBK14] Bernhard Rauch, Max Göttsche, Gernot Brähler, and Thomas Kronfeld. Deficit versus social statistics: Empirical evidence for the effectiveness of Benford’s Law. Applied Economics Letters, 21(3):147–151, ???? 2014. ISSN 1350-4851 (print),

1466-4291 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/13504851.2013.844319.

Rauch:2013:LME

[RGE13] Bernhard Rauch, M. Göttsche, and Florian El Mouaaouy. LI-BOR manipulation — empirical analysis of financial market benchmarks using Benford’s Law. Social Sciences Research Network, ??(??):26, December 5, 2013. URL http://papers.

ssrn.com/sol3/papers.cfm?abstract_id=2363895.

Rauch:2014:DPM

[RGL14] Bernhard Rauch, Max Göttsche, and Stephan Langenegger. Detecting problems in military expenditure data using digital analysis. Defence and Peace Economics, 25(2):97–111,

???? 2014. URL http://www.tandfonline.com/doi/full/
10.1080/10242694.2013.763438.

Reed:2002:GFG

[RH02] William J. Reed and Barry D. Hughes. From gene families and genera to incomes and Internet file sizes: Why power laws are so common in nature. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 66(6): 067103, December 2002. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-2376 (electronic). URL http://link.aps.org/

doi/10.1103/PhysRevE.66.067103.

Ristic:2008:GSP

[Ris08] Miroslav M. Ristić. A generalized semi-Pareto minification process. Statistical Papers, 49(2):343–351, April 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798

(electronic). URL http://link.springer.com/article/10.
1007/s00362-006-0017-4.

Radhakrishnan:1979:LLC

[RK79] T. Radhakrishnan and R. Kernizan. Lotka’s Law and computer science literature. Journal of the American Society for Information Science, 30(1):51–54, January 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Romano:2011:NLR

[RM11] Paul K. Romano and Harry McLaughlin. On non-linear recursive sequences and Benford’s Law. Fibonacci Quarterly, 49(2): 134–138, May 2011. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/49-2/romano.pdf.

Rane:2014:BLG

[RMB^{+}14] Ameya Deepak Rane, Utkarsh Mishra, Anindya Biswas, Aditi Sen(De), and Ujjwal Sen. Benford’s Law gives better scaling exponents in phase transitions of quantum XY models.

Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 90(2):022144:1–022144:8, August 29, 2014. CODEN PLEEE8. ISSN 1539-3755 (print),

1550-2376 (electronic). URL http://link.aps.org/doi/10.
1103/PhysRevE.90.022144.

Rousset:1998:CPV

[RME98] Olivier G. Rousset, Yilong Ma, and Alan C. Evans. Correction for partial volume effects in PET: Principle and validation. Journal of Nuclear Medicine, 39(5):904–911, May 1998. CODEN JNMEAQ. ISSN 0161-5505 (print), 1535-5667

(electronic). URL http://jnm.snmjournals.org/content/
39/5/904.full.pdf+html.

Robertson:1997:FEM

[Rob97] Jack C. Robertson. Fraud Examination for Managers and Auditors. Viesca Books, Austin, TX, USA, 1997 edition, 1997. ISBN 0-9656785-0-4. vii + 564 pp. LCCN HV8079.F7 R63 1997.

Robledo:2011:LZB

[Rob11] Alberto Robledo. Laws of Zipf and Benford, intermittency, and critical fluctuations. Chinese Science Bulletin, 56(34):3643–3648, 2011. ISSN 1001-6538 (print), 1861-9541

(electronic). URL http://link.springer.com/article/10.
1007/s11434-011-4827-y.

Rodriguez:2004:FSD

[Rod04] Ricardo J. Rodriguez. First significant digit patterns from mixtures of uniform distributions. The American Statistician, 58(1):64–71, February 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//oberon.ingentaselect.com/cgi-bin/linker?ini=asa&
reqidx=/cw/asa/00031305/v58n1/s13/p64.

Ross:2011:BLG

[Ros11] Kenneth A. Ross. Benford’s Law, a growth industry. American Mathematical Monthly, 118(7):571–583, August/September 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/10.4169;
 http://www.jstor.org/stable/pdfplus/10.4169/amer.
math.monthly.118.07.571.pdf;   http://www.maa.org/
history/presidents/ross.html.

Ross:2012:FDS

[Ros12] Kenneth A. Ross. First digits of squares and cubes. Mathematics Magazine, 85(1):37–43, February 2012. CODEN MAMGA8. ISSN 0025-570X. URL http://www.maa.org/

history/presidents/ross.html.

Rouault:1976:PAE

[Rou76a] Alain Rouault. Propriétés asymptotiques d’un n-échantillon d’une variable aléatoire dénombrable connues sous le nom de lois de Zipf. (French) [asymptotic properties of an n-sample of a denumerable random variable known under the name of Zipf’s law]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences. Séries A et B, 283(6):Aiv, A379– A380, 1976. CODEN CHASAP. ISSN 0151-0509.

Rouault:1976:SML

[Rou76b] Alain Rouault. Sources markoviennes et lois de Zipf. (French) [Markovian source and Zipf laws]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences. Séries A et B, 283(10):Aiii, A789–A790, 1976. CODEN CHASAP. ISSN 0151-0509.

Rouault:1978:LZS

[Rou78] Alain Rouault. Lois de Zipf et sources markoviennes. (French) [Zipf’s law and Markovian sources]. Ann. Inst. H. Poincaré Sect. B (N.S.), 14(2):169–188, 1978. CODEN AHPBAR. ISSN 0020-2347.

Rousseau:1992:RBR

[Rou92] Ronald Rousseau. Research: Breakdown of the robustness property of Lotka’s Law: the case of adjusted counts for multiauthorship attribution. Journal of the American Society for Information Science, 43(10):645–647, December 1992. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rousseau:2005:RFE

[Rou05] Ronald Rousseau. Robert Fairthorne and the empirical power laws. Journal of Documentation, 61(2):194–202, ???? 2005. CODEN JDOCAS. ISSN 0022-0418 (print), 1758-7379 (electronic).

Roukema:2009:BLA

[Rou09] Boudewijn F. Roukema. Benford’s Law anomalies in the 2009 Iranian presidential election. ArXiv e-prints,

June 16, 2009. CODEN ????  ISSN ????  URL

http://adsabs.harvard.edu/abs/2009arXiv0906.2789R;
 http://arxiv.org/abs/0906.2789.

Rousseau:2010:AFO

[Rou10] Christiane Rousseau. Apprendre à frauder ou à détecter les fraudes. (French) [Learning to defraud or to detect fraud]. Accromαth, 5:2–7, Été–automne 2010. ISSN 1911-0197. URL

http://accromath.uqam.ca/contents/pdf/Benford.pdf.

Roukema:2014:FDA

[Rou14] Boudewijn F. Roukema. A first-digit anomaly in the 2009 Iranian Presidential election. Journal of Applied Statistics, 41 (1):164–199, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Roukema:2015:CBL

[Rou15] Boudewijn F. Roukema. Complementing Benford’s Law for small N: a local bootstrap. In Miller [Mil15a], pages 223– 232. ISBN 0-691-14761-2 (hardcover), 1-4008-6659-6 (e-book).

LCCN QA273.6 .B46 2015. URL http://press.princeton.
edu/titles/10527.html.

Rowlands:2005:EAD

[Row05] Ian Rowlands. Emerald authorship data, Lotka’s law and research productivity. ASLIB Proceedings, 57(1):5–10, ???? 2005. CODEN ASLPAO. ISSN 0001-253X (print), 1758-3748 (electronic).

Rose:2003:TEF

[RR03] Anna M. Rose and Jacob M. Rose. Turn Excel into a financial sleuth: an easy-to-use digital analysis tool can red-flag irregularities. Journal of Accountancy, 196(2):58–??, August 2003. CODEN JACYAD. ISSN 0021-8448 (print), 1945-0729

(electronic). URL http://www.journalofaccountancy.com/
Issues/2003/Aug/TurnExcelIntoAFinancialSleuth.

Romero-Rochin:2009:DBL

[RR09] V. Romero-Rochin. A derivation of Benford’s Law … and a vindication of Newcomb. ArXiv e-prints, September 2009.

CODEN ???? ISSN ???? URL http://adsabs.harvard.
edu/abs/2009arXiv0909.3822R;   http://arxiv.org/abs/
0909.3822.

Rawlings:2004:EBP

[RRR04] Philip K. Rawlings, David Reguera, and Howard Reiss. Entropic basis of the Pareto law. Physica A: Statistical Mechanics and its Applications, 343(1–4):643–652, 2004. CODEN PHYADX. ISSN 0378-4371 (print), 1873-2119 (electronic).

Ravisankar:2011:DFS

[RRRB11] P. Ravisankar, V. Ravi, G. Raghava Rao, and I. Bose. Detection of financial statement fraud and feature selection using data mining techniques. Decision Support Systems, 50(2):491–500, January 2011. CODEN DSSYDK. ISSN 0167-9236 (print), 1873-5797 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167923610001879.

Ramanana-Rahary:2009:APR

[RRZR09] Suzy Ramanana-Rahary, Michel Zitt, and Ronald Rousseau. Aggregation properties of relative impact and other classical indicators: Convexity issues and the Yule–Simpson paradox. Scientometrics, 79(2):311–327, May 2009. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0420-4.

Ruankong:2019:CTB

[RS19] Pongpol Ruankong and Songkiat Sumetkijakan. Chains of truncated beta distributions and Benford’s law. Unif. Distrib. Theory, 14(2):27–32, 2019. ISSN 1336-913X.

Ramsden:1998:ZLB

[RV98] J. J. Ramsden and J. Vohradský. Zipf-like behavior in procaryotic protein expression. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 58(??):7777–??, December 1, 1998. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-2376 (electronic). URL http://link.

aps.org/doi/10.1103/PhysRevE.58.7777.

Rajabpour:2009:SRC

[RV09] M. A. Rajabpour and S. M. Vaez Allaei. Scaling relations for contour lines of rough surfaces. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 80(1):011115, July 13, 2009. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-2376 (electronic). URL http://link.aps.

org/doi/10.1103/PhysRevE.80.011115.

Raman:2015:CSC

[RVAN15] Raghu Raman, Smrithi Venkatasubramanian, Krishnashree Achuthan, and Prema Nedungadi. Computer science (CS) education in Indian schools: Situation analysis using Darmstadt model. ACM Transactions on Computing Education, 15 (2):7:1–7:??, May 2015. CODEN ???? ISSN 1946-6226.

Ribeiro:2015:MPE

[RZS^{+}15] Marco Tulio Ribeiro, Nivio Ziviani, Edleno Silva De Moura, Itamar Hata, Anisio Lacerda, and Adriano Veloso. Multiobjective Pareto-efficient approaches for recommender systems. ACM Transactions on Intelligent Systems and Technology (TIST), 5(4):53:1–53:??, January 2015. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Stanley:2000:SIU

[SAG^{+}00] H. E. Stanley, L. A. N. Amaral, P. Gopikrishnan, P. Ch. Ivanov, T. H. Keitt, and V. Plerou. Scale invariance and universality: organizing principles in complex systems. Physica A: Statistical Mechanics and its Applications, 281(1–4):60–68, June 15, 2000. CODEN PHYADX. ISSN 0378-4371 (print),

1873-2119 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378437100001953.

Salsburg:1997:DPB

[Sal97] David Salsburg. Digit preferences in the Bible. Chance: a magazine for people interested in the analysis of data, 10(4): 46–48, ???? 1997. CODEN CNDCE4. ISSN 0933-2480 (print),

1867-2280 (electronic). URL http://amstat.tandfonline.
com/doi/full/10.1080/09332480.1997.10542065.

Sanders:1987:PPUa

[San87a] Robert Sanders. The Pareto Principle: Its use and abuse. Journal of Consumer Marketing, 4(1):47–50, ???? 1987. CO-DEN ???? ISSN 0736-3761.

Sanders:1987:PPUb

[San87b] Robert Sanders. The Pareto Principle: Its use and abuse. Journal of Services Marketing, 1(2):37–40, ???? 1987. CO-DEN ???? ISSN 0887-6045.

Sanders:1988:PPU

[San88] Robert Sanders. The Pareto Principle: Its use and abuse. Journal of Business & Industrial Marketing, 3(2):37–40, ???? 1988. CODEN ???? ISSN 0885-8624.

Sanders:1992:PPU

[San92] Robert Sanders. The Pareto Principle: its use and abuse. Journal of Product & Brand Management, 1(2):37–40, ???? 1992. CODEN ???? ISSN 1061-0421.

Sandon:2002:DPC

[San02a] F. Sandon. Do populations conform to the law of anomalous numbers? Population, 57(4):755–761, ???? 2002.

URL http://www.persee.fr/doc/pop_1634-2941_2002_
num_57_4_18419.

Sandron:2002:DPC

[San02b] Frédéric Sandron. Do populations conform to the Law of Anomalous Numbers? Population, 57(4–5):755–761, ???? 2002. CODEN ???? ISSN 1169-1018. URL http:/

/www.cairn.info/article_p.php?ID_ARTICLE=POPE_204_

0753. Translation from French by S. R. Hayford.

Sandron:2002:PSE

[San02c] Frédéric Sandron. Les populations suivent-elles la loi des nombres anomaux? (French). [Do populations conform to the Law of Anomalous Numbers?]. Population (French edition), 57(4–5):761–768, ???? 2002. CODEN POPUAQ. ISSN 0032-4663 (print), 1957-7966 (electronic). URL http:

//www.persee.fr/web/revues/home/prescript/article/
pop_0032-4663_2002_num_57_4_16873.

Sarkar:1973:OSD

[Sar73] B. P. Sarkar. An observation on the significant digits of the binomial coefficients and factorials. Sankhyā (Indian Journal of Statistics), Series B. Methodological, 35(??):363–364, ???? 1973. CODEN SANBBV. ISSN 0581-5738.

Saville:2004:SWN

[Sav04] Adrian D. Saville. Sorry, wrong number. How accounting data are wrong and how the numbers can be fixed. Preprint, ????, ????, ???? 2004.

Saville:2006:UBL

[Sav06] Adrian D. Saville. Using Benford’s law to detect data error and fraud: An examination of companies listed on the Johannesburg Stock Exchange. South African Journal of Economic and Management Sciences, 9(3):341–354, ???? 2006. CODEN ???? ISSN 1015-8812 (print), 2222-3436 (electronic).

Sundheim:1993:FNT

[SB93] R. Sundheim and B. Busta. Fibonacci numbers tend to obey Benford’s Law: an extension of Wlodarski and Sentance. Working paper, St. Cloud State University, St. Cloud, MN, USA, ???? 1993.

Singh:2015:ABL

[SB15] N. Singh and R. Bansal. Analysis of Benford’s Law in digital image forensics. In Signal Processing and Communication (ICSC), 2015 International Conference on, pages 413–

418. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, March 2015.

Sellke:2001:CVT

[SBB01] Thomas Sellke, M. J. Bayarri, and James O. Berger. Calibration of p values for testing precise null hypotheses. The American Statistician, 55(1):62–71, February 2001. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.amstat.org/publications/tas/
Sellke.htm;   http://www.jstor.org/stable/2685531;
 http://www.tandfonline.com/doi/abs/10.1198/000313001300339950.

Stanley:1995:ZPS

[SBH^{+}95] Michael H. R. Stanley, Sergey V. Buldyrev, Shlomo Havlin, Rosario N. Mantegna, Michael A. Salinger, and H. Eugene Stanley. Zipf plots and the size distribution of firms. Economics Letters, 49(4):453–457, October 15, 1995. CO-DEN ECLEDS. ISSN 0165-1765 (print), 1873-7374 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016517659500696D.

Snyder:2001:SAO

[SCD01] Mark A. Snyder, James H. Curry, and Anne M. Dougherty. Stochastic aspects of one-dimensional discrete dynamical systems: Benford’s law. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 64(2):026222:1–026222:5, August 2001. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-2376 (electronic). URL

http://adsabs.harvard.edu/abs/2001PhRvE..64b6222S;
 http://pre.aps.org/abstract/PRE/v64/i2/e026222.

Swanson:2003:DPF

[SCE03] David Swanson, Moon Jung Cho, and John Eltinge. Detecting possibly fraudulent or error-prone survey data using Benford’s Law. In ????, editor, Proceedings of the Survey Research Section, pages 4172–4177. American Statistical Association, Washington, DC, USA, 2003. ISBN ????

LCCN ???? URL http://www.amstat.org/sections/srms/
proceedings/y2003/Files/JSM2003-000205.pdf..

Schatte:1973:VMG

[Sch73] Peter Schatte. Zur Verteilung der Mantisse in der Gleitkommadarstellung einer Zufallsgröße. (German) [distribution of the mantissa in the floating-point representation of a random variable]. Zeitschrift für Angewandte Mathematik und Mechanik, 53(??):553–565, ???? 1973. CODEN ZAMMAX. ISSN 0044-2267 (print), 1521-4001 (electronic).

Schorr:1974:LLL

[Sch74] Alan E. Schorr. Lotka’s Law and library science. RQ, 14(1): 32–33, Fall 1974. CODEN ???? ISSN 0033-7072.

Schorr:1975:LLM

[Sch75a] Alan E. Schorr. Lotka’s Law and map librarianship. Journal of the American Society for Information Science, 26(3):189–190, May/June 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Schorr:1975:LLH

[Sch75b] Alan E. Schorr. Lotka’s Law and the history of legal medicine. Research in Librarianship, 30(??):205–209, September 1975. CODEN ???? ISSN 0034-5245.

Schorr:1975:BCL

[Sch75c] Alan Edward Schorr. Brief communications: Lotka’s Law and Map Librarianship. Journal of the American Society for Information Science, 26(3):189–190, May 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Schatte:1981:RVL

[Sch81] Peter Schatte. On random variables with logarithmic mantissa distribution relative to several bases. Elektronische Informationsverarbeitung und Kybernetik (EIK), 17(??):293–295, ???? 1981. CODEN EIVKAX. ISSN 0013-5712.

Schatte:1983:SUD

[Sch83a] Peter Schatte. On H_{∞}-summability and the uniform distribution of sequences. Mathematische Nachrichten, 113(1): 237–243, ???? 1983. CODEN MTMNAQ. ISSN 0025-

584X. URL http://onlinelibrary.wiley.com/doi/10.
1002/mana.19831130122/abstract.

Schatte:1983:AUD

[Sch83b] Peter Schatte. On the asymptotic uniform distribution of sums reduced mod1. Mathematische Nachrichten, 115(1): 257–281, ???? 1983. CODEN MTMNAQ. ISSN 0025-

584X. URL http://onlinelibrary.wiley.com/doi/10.
1002/mana.19841150121/abstract.

Schatte:1984:AUD

[Sch84] Peter Schatte. On the asymptotic uniform distribution of sums reduced mod1. Mathematische Nachrichten, 115(1): 275–281, ???? 1984. CODEN MTMNAQ. ISSN 0025-

584X. URL http://onlinelibrary.wiley.com/doi/10.
1002/mana.19841150121/abstract.

Schatte:1986:ALD

[Sch86] Peter Schatte. On the asymptotic logarithmic distribution of the floating-point mantissas of sums. Mathematische Nachrichten, 127(1):7–20, ???? 1986. CODEN MTMNAQ.

ISSN 0025-584X. URL http://onlinelibrary.wiley.com/
doi/10.1002/mana.19861270102/abstract.

Schatte:1987:ABM

[Sch87a] Peter Schatte. On the asymptotic behaviour of the mantissa distribution of sums. Journal of Information Processing and Cybernetics: EIK, 23(??):353–360, ???? 1987. CODEN JI-CYE5. ISSN 0863-0593.

Schatte:1987:SEU

[Sch87b] Peter Schatte. Some estimates of the H_{∞}-uniform distribution. Monatshefte für Mathematik, 103(3):233–249, September 1987. CODEN MNMTA2. ISSN 0026-9255 (print), 1436-5081

(electronic). URL http://link.springer.com/article/10.
1007/BF01364342.

Schatte:1988:LIL

[Sch88a] Peter Schatte. On a law of the iterated logarithm for sums mod1 with application to Benford’s law. Probability Theory and Related Fields, 77(2):167–178, February 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF00334035.

Schatte:1988:MDC

[Sch88b] Peter Schatte. On mantissa distribution in computing and Benford’s Law. Journal of Information Processing and Cybernetics: EIK, 24(9):443–455, ???? 1988. CODEN JICYE5.

ISSN 0863-0593. URL https://dblp.org/db/journals/
eik/eik24.html#Schatte88.

Schatte:1988:ASC

[Sch88c] Peter Schatte. On the almost sure convergence of floatingpoint mantissas and Benford’s Law. Mathematische Nachrichten, 135(1):79–83, ???? 1988. CODEN MTMNAQ. ISSN 0025-584X.

Schatte:1988:UDC

[Sch88d] Peter Schatte. On the uniform distribution of certain sequences and Benford’s law. Mathematische Nachrichten, 136 (1):271–273, 1988. CODEN MTMNAQ. ISSN 0025-584X.

Schatte:1989:MUD

[Sch89] Peter Schatte. On measures of uniformly distributed sequences and Benford’s law. Monatshefte für Mathematik, 107(3):245–256, 1989. CODEN MNMTA2. ISSN 0026-9255 (print), 1436-5081 (electronic).

Schatte:1990:BLC

[Sch90] Peter Schatte. On Benford’s law for continued fractions. Mathematische Nachrichten, 148:137–144, 1990. CODEN MTM-NAQ. ISSN 0025-584X.

Schatte:1991:ULI

[Sch91a] P. Schatte. On a uniform law of the iterated logarithm for sums mod 1 and Benford’s law. Litovsk. Mat. Sb., 31(1):205– 217, 1991. CODEN ???? ISSN 0132-2818.

Schroeder:1991:FCP

[Sch91b] Manfred Schroeder. Fractals, Chaos, Power Laws. W. H. Freeman and Company, New York, NY, USA, 1991. ISBN 0-671-74217-5, 0-7167-2136-8, 0-7167-2357-3. xviii + 429 pp. LCCN QD921 .S3 1990.

Schatte:1998:BLV

[Sch98] P. Schatte. On Benford’s law to variable base. Statistics & Probability Letters, 37(4):391–397, March 30, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715297001429.

Schneider:2006:RAS

[Sch06] Reto U. Schneider. Das Rätsel der abgegriffenen Seiten: Vor über hundert Jahren stiess ein Astronom auf ein merkwürdiges statistisches Gesetz, das heute Steuerbetrüger entlarven soll. (German) [The puzzle of the worn pages: More than a hundred years ago an astronomer came across a strange statistical law that will now expose tax cheats]. NZZ Folio: Die Zeitschrift der Neuen Zürcher Zeitung, ??(1/06), ???? 2006. URL http://www-x.nzz.

ch/folio/curr/articles/schneider_2.html;   http://
www.nzzfolio.ch/www/d80bd71b-b264-4db4-afd0-277884b93470/
showarticle/4c9a5444-883e-4e6b-8a4c-6c31f32c02f3.aspx.

Schurger:2008:EBS

[Sch08] Klaus Schürger. Extensions of Black–Scholes processes and Benford’s Law. Stochastic Processes and Their Applications, 118(7):1219–1243, July 2008. CODEN STOPB7. ISSN 0304-4149 (print), 1879-209X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304414907001391.

Schraepler:2010:BLI

[Sch10] Jörg-Peter Schraepler. Benford’s Law as an instrument for fraud detection in surveys using the data of the Socio-Economic Panel (SOEP). SOEP paper 273, Socio-Economic Panel, Deutsches Institut für Wirtschaftsforschung, Berlin, Germany, March 2, 2010. ii + 56 pp. URL http://www.

diw.de/documents/publikationen/73/diw_01.c.349061.
de/diw_sp0273.pdf.

Schurger:2015:LPB

[Sch15] Klaus Schürger. Lévy processes and Benford’s Law. In Miller [Mil15a], pages 135–173. ISBN 0-691-14761-2 (hardcover), 1-4008-6659-6 (e-book). LCCN QA273.6 .B46 2015. URL http:

//press.princeton.edu/titles/10527.html.

Scollnik:2013:CTP

[Sco13] David P. M. Scollnik. Comments on two papers concerning estimation of the parameters of the Pareto distribution in the presence of outliers. Statistical Methodology, 13(??): 1–11, July 2013. CODEN ???? ISSN 1572-3127 (print),

1878-0954 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1572312712000822.

Samson:1978:STU

[SD78] W. B. Samson and R. H. Davis. Search times using hash tables for records with non-unique keys. The Computer Journal, 21(3):210–214, August 1978. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://www3.oup.co.uk/computer_journal/hdb/Volume_
21/Issue_03/tiff/210.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_21/Issue_03/tiff/211.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_
21/Issue_03/tiff/212.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_21/Issue_03/tiff/213.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_
21/Issue_03/tiff/214.tif.

Savanur:2014:LLA

[SDK14] Kiran Savanur, S. Gayathri Devi, and P. V. Konnur. Lotka’s Law and authorship distribution in the journal of ‘Columbia Law Review’. Collnet Journal of Scientometrics and Information Management, 8(1):193–208, 2014. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Seaman:2002:BLB

[Sea02a] R. S. Seaman. Benford’s law and and background field errors in data assimilation. Preprint, Bureau of Meteorology Research Centre, Melbourne, Australia, ???? 2002. URL http:

//amm.bom.gov.au/amoj/docs/2002/seaman.pdf.

Seaman:2002:RBL

[Sea02b] R. S. Seaman. The relevance of Benford’s Law to background field errors in data assimilation. Australian Meteorological Magazine, 51(1):25–33, ???? 2002. CODEN AMMGAS. ISSN

0004-9743. URL http://www.bom.gov.au/amm/docs/2002/
seaman.pdf;   http://www.bom.gov.au/amm/docs/2002/
seaman_hres.pdf.

Semboloni:2008:HCS

[Sem08] F. Semboloni. Hierarchy, cities size distribution and Zipf’s law. European Physical Journal B: Condensed Matter and Complex Systems, 63(3):295–301, 2008. CODEN EPJBFY. ISSN 1434-6028 (print), 1434-6036 (electronic).

Sentance:1973:FAB

[Sen73] W. A. Sentance. A further analysis of Benford’s Law. Fibonacci Quarterly, 11(5):490–494, December 1973. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/11-5/sentance.pdf.

Sigurd:2004:WLS

[SEOV04] Bengt Sigurd, Mats Eeg-Olofsson, and Joost Van Weijer. Word length, sentence length and frequency — Zipf revisited. Studia Linguistica, 58(1):37–52, 2004. ISSN 0039-3193 (print), 1467-9582 (electronic).

Scott:2001:BLE

[SF01] P. D. Scott and M. Fasli. Benford’s Law: An empirical investigation and a novel explanation. CSM Technical Report 349, Department of Computer Science, University Essex, Colchester, UK, October 26, 2001. URL http://www.essex.

ac.uk/csee/research/publications/technicalreports/
2001/CSM-349.pdf.

Shaki:2008:RSN

[SF08] Samuel Shaki and Martin H. Fischer. Reading space into numbers — a cross-linguistic comparison of the SNARC effect. Cognition, 108(2):590–599, August 2008. CO-DEN CGTNAU. ISSN 0010-0277 (print), 1873-7838 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010027708000978.

Shi:2011:ESP

[SF11] Yun-Quing Shi and DongDong Fu. Apparatus and method for a generalized Benford’s Law analysis of DCT and JPEG coefficients. US Patent 7940989., May 10, 2011. URL http:/

/www.google.com/patents/US7940989. US Patent Applica-

tion 11/772,636 filed 2 July 2007. US Patent issued 10 May 2011.

Sole:2000:SLH

[SG00] Ricard Vicente Solé and Brian Goodwin. Signs of Life: How Complexity Pervades Biology. Basic Books, New York, NY, USA, 2000. ISBN 0-465-01927-7, 0-465-01928-5 (paperback). xi + 322 pp. LCCN QH501 .S65 2000.

Sheng:2016:DSC

[SG16] Guorui Sheng and Tiegang Gao. Detection of seam-carving image based on Benford’s Law for forensic applications. International Journal of Digital Crime and Forensics, 8(1):51–61, ???? 2016. CODEN ???? ISSN 1941-6210 (print), 1941-6229

(electronic). URL https://dblp.org/db/journals/ijdcf/
ijdcf8.html#ShengG16.

Skousen:2004:AUP

[SGW04] Christopher J. Skousen, Liming Guan, and T. Sterling Wetzel. Anomalies and unusual patterns in reported earnings: Japanese managers round earnings. Journal of International Financial Management and Accounting, 15(3):212–234, October 2004. CODEN ???? ISSN 0954-1314 (print), 1467-646X (electronic).

Suh:2010:CAB

[SH10] Ikseon Suh and T. Christopher Headrick. A comparative analysis of the bootstrap versus traditional statistical procedures applied to digital analysis based on Benford’s Law. Journal of Forensic and Investigative Accounting, 2(2):144–175, July/December 2010. CODEN ???? ISSN 2165-3755. URL

http://epublications.marquette.edu/account_fac/31;

 http://www.bus.lsu.edu/accounting/faculty/lcrumbley/
jfia/Articles/FullText/2010v2n2a7.pdf;   http://www.
bus.lsu.edu/accounting/faculty/lcrumbley/jfia/Articles/
v2n2.htm.

Shanbhag:1989:BRBa

[Sha89] D. N. Shanbhag. Book review: Pareto Distributions, by B. C. Arnold. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(2):253–254, ???? 1989. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/2982920.

Shan:2005:GZD

[Sha05] Shi Shan. On the generalized Zipf distribution. Part I. Information Processing and Management, 41(6):1369–1386, December 1, 2005. CODEN IPMADK. ISSN 0306-4573 (print),

1873-5371 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030645730500018X.

Schupfer:2012:FWG

[SHC^{+}12] G. Schüpfer, J. Hein, M. Casutt, L. Steiner, and C. Konrad. Vom Finanz- zum Wissenschaftsbetrug: Methode, den Irrungen in der medizinischen Literatur beizukommen. (German) [From financial to scientific fraud: Method to overcome errors in the medical literature]. Der Anaesthesist, 61(6):537–542, June 2012. ISSN 0003-2417 (print), 1432-055X (electronic).

Shenton:1968:PDM

[She68] L. R. Shenton. Periodicity and density of modified Fibonacci sequences. Fibonacci Quarterly, 6(2):109–116, April 1968. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.

math.ca/Scanned/6-2/shenton.pdf.

Sheppard:1994:TNA

[She94] Lee A. Sheppard. Tax Notes audits the Clintons. Tax Notes Today, ??(??):18–23, April 4, 1994. URL http://www.

bankler.com/docs/Tax_Notes_Audits_the_Clintons.pdf.

Shelburne:2012:ED

[She12] Brian Shelburne. The ENIAC’s 1949 determination of π. IEEE Annals of the History of Computing, 99(PrePrints), 2012. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Shieh:1989:BZE

[Shi89] Yeung-Nan Shieh. The Beckmann–Zipf effect and plant location under uniform pricing. Economics Letters, 30(1):89–92, January 1, 1989. CODEN ECLEDS. ISSN 0165-1765 (print),

1873-7374 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165176589901626.

Shi:2009:FDL

[Shi09] Yun Q. Shi. First Digit Law and its application to digital forensics. Lecture Notes in Computer Science, 5450: 448–453, 2009. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/
content/pdf/10.1007/978-3-642-04438-0_37.

Slepkov:2013:BLT

[SID13] Aaron D. Slepkov, Kevin B. Ironside, and David DiBattista. Benford’s Law: Textbook exercises and multiple-choice testbanks. arXiv.org, ??(??):??, November 19, 2013. URL http:

//arxiv.org/abs/1311.4787v1.

Slepkov:2015:BLT

[SID15] Aaron D. Slepkov, Kevin B. Ironside, and David DiBattista. Benford’s Law: Textbook exercises and multiplechoice testbanks. PLoS One, 10(2):e0117972:1–e0117972:13, February 17, 2015. CODEN POLNCL. ISSN 1932-

6203. URL http://journals.plos.org/plosone/article?
id=10.1371/journal.pone.0117972.

Simon:1955:CSD

[Sim55] Herbert A. Simon. On a class of skew distribution functions. Biometrika, 42(3/4):425–440, December 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333389.

Simon:1998:ADC

[Sim98] Jonathan Simon. An analysis of the distribution of combinations chosen by UK National Lottery players. Journal of Risk and Uncertainty, 17(3):243–277, 1998. CODEN JRUNEN. ISSN 0895-5646 (print), 1573-0476 (electronic).

Simons:2007:SIP

[Sim07] John L. Simons. A simple (inductive) proof for the non-existence of 2-cycles of the 3x + 1 problem. Journal of Number Theory, 123(1):10–17, March 2007. CO-DEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06001223.

Singmaster:1974:RSB

[Sin74] D. Singmaster. Recurrent sequences and Benford’s law. Preprint submitted to Fibonacci Quarterly in 1974, but never published there., 1974.

Sinai:2003:SP

[Sin03] Ya. G. Sinai. Statistical (3x+1) problem. Communications on Pure and Applied Mathematics (New York), 56(7):1016–1028, July 2003. CODEN CPAMAT, CPMAMV. ISSN 0010-3640 (print), 1097-0312 (electronic). Dedicated to the memory of Jürgen K. Moser.

Sjoberg:2014:AAS

[Sjo14] Fredrik M. Sjoberg. Autocratic adaptation: The strategic use of transparency and the persistence of election fraud. Electoral Studies, 33(0):233–245, 2014. ISSN 0261-3794 (print),

1873-6890 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0261379413001273.

Serra:2015:POA

[SJP^{+}15] Edoardo Serra, Sushil Jajodia, Andrea Pugliese, Antonino Rullo, and V. S. Subrahmanian. Pareto-optimal adversarial defense of enterprise systems. ACM Transactions on Information and System Security, 17(3):11:1–11:??, March 2015. CODEN ATISBQ. ISSN 1094-9224 (print), 1557-7406 (electronic).

Solomon:1996:SSE

[SL96] Sorin Solomon and Moshe Levy. Spontaneous scaling emergence in generic stochastic systems. International Journal of Modern Physics C [Physics and Computers], 7(5):745–752, October 1996. CODEN IJMPEO. ISSN 0129-1831 (print),

1793-6586 (electronic). URL http://www.worldscientific.
com/doi/abs/10.1142/S0129183196000624.

Slijepcevic:1998:NID

[Sli98] Siniša Slijepčević. A note on initial digits of recurrence sequences. Fibonacci Quarterly, 36(4):305–308, August 1998.

CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.

math.ca/Scanned/36-4/slijepcevic.pdf.

Sheng:2016:DCA

[SLS^{+}16] Guorui Sheng, Tao Li, Qingtang Su, Beijing Chen, and Yi Tang. Detection of content-aware image resizing based on Benford’s law. Soft computing: newsletter of the Canadian Society for Fuzzy Information and Neural Systems, 21(19): 5693–5701, April 2016. ISSN 1432-7643 (print), 1433-7479 (electronic).

Sole:1996:ESO

[SM96] Ricard V. Solé and Susanna C. Manrubia. Extinction and self-organized criticality in a model of large-scale evolution. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 54(1):R42–R45, July 1996. CODEN PLEEE8. ISSN 1539-3755 (print),

1550-2376 (electronic). URL http://link.aps.org/doi/10.
1103/PhysRevE.54.R42.

Sanches:2006:IRU

[SM06a] João M. Sanches and Jorge S. Marques. Image reconstruction using the Benford Law. In 2006 IEEE International Conference on Image Processing, Atlanta, Georgia, 8–11 October 2006, pages 2029–2032. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2006. ISBN 1-4244-0480-0. ISSN 1522-4880.

Sugimoto:2006:AFP

[SM06b] Toyoaki Sugimoto and Naoto Miyoshi. On the asymptotics of fault probability in least-recently-used caching with Zipftype request distribution. Random Structures and Algorithms, 29(3):296–323, 2006. CODEN RSALFD. ISSN 1042-9832 (print), 1098-2418 (electronic).

Seuront:2008:TST

[SM08] Laurent Seuront and James G. Mitchell. Towards a seascape typology. I. Zipf versus Pareto laws. Journal of Marine Systems, 69(3–4):310–327, February 2008. CODEN JMASE5. ISSN 0924-7963 (print), 1879-1573 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0924796307000620.

Shao:2009:FDD

[SM09] Lijing Shao and Bo-Qiang Ma. First digit distribution of hadron full width. Modern Physics Letters A (MPLA), 24 (40):3275–3282, ???? 2009. CODEN MPLAEQ. ISSN 0217-7323 (print), 1793-6632 (electronic).

Shao:2010:EMD

[SM10a] Lijing Shao and Bo-Qiang Ma. Empirical mantissa distributions of pulsars. Astroparticle Physics, 33(4):255–262, May 2010. CODEN APHYEE. ISSN 0927-6505 (print), 1873-

2852 (electronic). URL http://arxiv.org/abs/1005.1702;
 http://www.sciencedirect.com/science/article/pii/
S0927650510000447.

Shao:2010:FDD

[SM10b] Lijing Shao and Bo-Qiang Ma. First digit distribution of hadron full width. ArXiv e-prints, April 19, 2010. URL http:

//arxiv.org/abs/1004.3077.

Shao:2010:FDLa

[SM10c] Lijing Shao and Bo-Qiang Ma. First-digit law in nonextensive statistics. ArXiv e-prints, 2010. URL http://arxiv.org/

abs/1010.2699.

Shao:2010:FDLb

[SM10d] Lijing Shao and Bo-Qiang Ma. First-digit law in nonextensive statistics. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 82(4):041110, October 2010. CODEN PLEEE8. ISSN 1539-3755 (print),

1550-2376 (electronic). URL http://link.aps.org/doi/10.
1103/PhysRevE.82.041110;   http://pre.aps.org/abstract/
PRE/v82/i4/e041110.

Shao:2010:SDL

[SM10e] Lijing Shao and Bo-Qiang Ma. The significant digit law in statistical physics. Physica A: Statistical Mechanics and its Applications, 389(16):3109–3116, May 2010. CO-DEN PHYADX. ISSN 0378-4371 (print), 1873-2119 (elec-

tronic). URL http://arxiv.org/abs/1005.0660;   http:
//inspirehep.net/record/855898; http://inspirehep.
net/record/873155;   http://www.sciencedirect.com/
science/article/pii/S0378437110003596.

Shikano:2011:WDD

[SM11] S. Shikano and V. Mack. When does 2nd digit Benford’s Lawtest signal an election fraud? Facts or misleading test results. Jahrbücher für Nationalökonomie und Statistik, 231(5–

6):719–732, ???? 2011. URL http://benfordonline.net/
fullreference/1015.

Smith:1997:EBL

[Smi97] Steven W. Smith. Explaining Benford’s Law. In The Scientist and Engineer’s Guide to Digital Signal Processing, chapter 34, pages 701–722. California Technical Publishing, San Diego, CA, USA, 1997. ISBN 0-9660176-3-3. LCCN TK5102.9 .S57

1997. URL http://www.dspguide.com/ch34.htm.

Smith:2013:BNS

[Smi13] Aaron Carl Smith. Benford–Newcomb subsequences for fraud detection. arXiv.org, ??(??):??, January 22, 2013. URL http:

//arxiv.org/abs/1301.6086v1.

Saichev:2010:TZL

[SMS10] Alexander Saichev, Yannick Malevergne, and Didier Sornette. Theory of Zipf’s law and beyond, volume 632 of Lecture Notes in Economics and Mathematical Systems. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 3-642-02945-0. xii + 171 pp.

Schatte:1991:NBL

[SN91] P. Schatte and K. Nagasaka. A note on Benford’s law for second order linear recurrences with periodical coefficients. Zeitschrift für Analysis und ihre Anwendungen, 10(2):251– 254, 1991. CODEN ???? ISSN 0232-2064 (print), 1661-4534 (electronic).

Soliman:2000:BPP

[Sol00] Ahmed A. Soliman. Bayes prediction in a Pareto lifetime model with random sample size. Journal of the Royal Statistical Society. Series D (The Statistician), 49(1):51–62, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/stable/2681255.

Soo:2005:ZLC

[Soo05] Kwok Tong Soo. Zipf’s Law for cities: a cross-country investigation. Regional Science and Urban Economics, 35(3):

239–263, May 1, 2005. CODEN RSUEDM. ISSN 0166-0462 (print), 1879-2308 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016604620400033X.

Soo:2013:ZGG

[Soo13] Kwok Tong Soo. Zipf, Gibrat and geography: Evidence from China, India and Brazil. Papers in Regional Science, page ??, 2013. ISSN 1056-8190 (print), 1435-5957 (electronic).

Sorell:2012:ZLV

[Sor12] C. Joseph Sorell. Zipf’s Law and vocabulary. In Carol A. Chapelle, editor, The Encyclopedia of Applied Linguistics. Wiley, New York, NY, USA, 2012. ISBN 1-4051-9843-5.

Speed:2009:YWP

[Spe09] T. Speed. You want proof? Bulletin — Institute of Mathematical Statistics, 38(??):11–??, ???? 2009. CODEN SM-

BCVA. ISSN 0146-3942. URL http://bulletin.imstat.
org/archive/.

Shebaro:2010:LTC

[SPGC10] Bilal Shebaro, Fernando Perez-Gonzalez, and Jedidiah R. Crandall. Leaving timing-channel fingerprints in hidden service log files. Digital Investigation, 7, Supplement(0): S104–S113, 2010. ISSN 1742-2876 (print), 1873-202X (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S174228761000040X. The Proceedings of the

Tenth Annual DFRWS Conference.

Sottili:2012:BLT

[SPGM12] Gianluca Sottili, Danilo M. Palladino, Biagio Giaccio, and Paolo Messina. Benford’s Law in time series analysis of seismic clusters. Mathematical Geosciences, 44(5):619–634, July 2012. CODEN ???? ISSN 1874-8953 (print), 1874-8961

(electronic). URL http://link.springer.com/article/10.
1007/s11004-012-9398-1.

Shukla:2017:BDE

[SPP17] Abhishek Shukla, Ankit Kumar Pandey, and Anirban Pathak. Benford’s distribution in extrasolar world: Do the exoplanets follow Benford’s distribution? Journal of Astrophysics and Astronomy, 38(1), March 2017. CODEN JASRD7. ISSN 0250-6335 (print), 0973-7758 (electronic).

Solomon:2001:SPZ

[SR01] Sorin Solomon and Peter Richmond. Stability of Pareto–Zipf law in non-stationary economies. In Economics with heterogeneous interacting agents (Marseille, 2000), volume 503 of Lecture Notes in Econom. and Math. Systems, pages 141–159. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001.

Srivastava:2013:FD

[Sri13] Tanya Kaushal Srivastava. The first digit 1. Resonance, 18 (12):1073–1085, December 2013. CODEN RESOFE. ISSN 0971-8044 (print), 0973-712X (electronic). URL http://

link.springer.com/article/10.1007/s12045-013-0135-
y.

Schuette:2006:TFZ

[SS06] Paul Schuette and Marcus C. Spruill. Tail fit and the Zipf– Pareto law. Extremes, 9(3–4):243–261, 2006. CODEN ???? ISSN 1386-1999.

Stosic:2009:CSM

[SS09] Borko D. Stosić and Tatijana Stosić. Comment on “Ising model on the scale-free network with a Cayley-tree-like structure”. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 79(April):048101, April 24, 2009. CODEN PLEEE8. ISSN 1539-3755 (print),

1550-2376 (electronic). URL http://link.aps.org/doi/10.
1103/PhysRevE.79.048101.

Singleton:2010:FAF

[SS10] Tommie Singleton and Aaron J. Singleton. Fraud auditing and forensic accounting. Wiley corporate F and A. Wiley, New York, NY, USA, fourth edition, 2010. ISBN 0-470-56413-X. xiv + 317 pp. LCCN HV8079.W47 B65

2010. URL http://catalogimages.wiley.com/images/
db/jimages/9780470564134.jpg;   http://catdir.loc.
gov/catdir/enhancements/fy1007/2010013504-d.html;
http://catdir.loc.gov/catdir/enhancements/fy1007/2010013504-
t.html;   http://catdir.loc.gov/catdir/enhancements/
fy1011/2010013504-b.html.

SenDe:2011:BLD

[SS11] Aditi Sen De and Ujjwal Sen. Benford’s law detects quantum phase transitions similarly as earthquakes. Europhysics Letters, 95(5):50008, September 2011. CODEN EULEEJ. ISSN 0295-5075 (print), 1286-4854 (electronic). URL http:/

/arxiv.org/abs/1103.5398.

Song:2012:QEM

[SS12] Jongwoo Song and Seongjoo Song. A quantile estimation for massive data with generalized Pareto distribution. Computational Statistics & Data Analysis, 56(1):143–150, January 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-

7352 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167947311002477.

Schaefer:2005:AIF

[SSMW05] Christin Schaefer, Jörg-Peter Schräpler, Klaus-Robert Müller, and Gert G. Wagner. Automatic identification of faked and fraudulent interviews in the German SOEP. Journal of Applied Social Science Studies, 125(??):183–193, 2005.

URL https://www.diw.de/documents/publikationen/73/
diw_01.c.42515.de/dp441.pdf.

Stamp:1914:NIP

[Sta14] J. C. Stamp. A new illustration of Pareto’s law. Journal of the Royal Statistical Society, 77(2):200–204, January 1914. ISSN

0952-8385. URL http://www.jstor.org/stable/2339802.

Stanley:1999:SUR

[Sta99] H. Eugene Stanley. Scaling, universality, and renormalization: Three pillars of modern critical phenomena. Reviews of Modern Physics, 71(2):S358–S366, March 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756.  URL http://link.aps.org/doi/10.1103/
RevModPhys.71.S358;   http://rmp.aps.org/abstract/
RMP/v71/i2/pS358_1.

Stadje:2005:APD

[Sta05] Wolfgang Stadje. Asymptotic properties of digit sequences of random numbers. Mathematische Nachrichten, 278(10):1209– 1229, ???? 2005. CODEN MTMNAQ. ISSN 0025-584X.

Stalcup:2010:BLH

[Sta10] Ken Stalcup. Benford’s Law: How a simple misconception can trip up a fraudster and how a savvy CFE can spot it. Fraud Magazine, ??(??):??, January/February 2010.

ISSN 1553-6645. URL http://www.fraud-magazine.com/
article.aspx?id=203.

Sambridge:2011:BLF

[STA11] Malcolm Sambridge, Hrvoje Tkalčić, and Pierre Arroucau. Benford’s law of first digits: from mathematical curiosity to change detector. Asia Pacific Mathematics Newsletter, 1(4): 1–6, October 2011. ISSN 2010-3484.

Stanton:2012:DFU

[Sta12] Gabriella Stanton. Detecting fraud: Utilizing new technology to advance the audit profession. Honors thesis, University of New Hampshire, Durham, NH 03824, USA, Spring 2012.

26 pp. URL http://scholars.unh.edu/cgi/viewcontent.
cgi?article=1017&context=honors.

Sterbenz:1974:FPC

[Ste74] Pat H. Sterbenz. Floating Point Computation. Prentice-Hall Series in Automatic Computation. Prentice-Hall, Upper Saddle River, NJ, USA, 1974. ISBN 0-13-322495-3. xiv + 316 pp. LCCN QA76.8.I12 S771 1974.

Stewart:1993:LAN

[Ste93] Ian Stewart. The law of anomalous numbers. Working paper, Mathematics Institute, University of Warwick, Warwick, UK, 1993. Published in Spektrum der Wissenschaft (April 1994).

Stigler:1945:DLD

[Sti45] George J. Stigler. The distribution of leading digits in statistical tables. Unpublished, but written about 1945–1946, and presented in an 1975 address at Haskell Hall, University of Chicago. Stigler’s distribution is more complex than that of Newcomb and Benford. Stigler has F_{d} = (d ln(d)−(d+1) ln(d+ 1) + (1 + (10/9) ln(10)))/9, which gives leading-digit frequencies of 0.2413, 0.1832, 0.1455, 0.1174, 0.0950, 0.0764, 0.0605, 0.0465, and 0.0342. See [LCJ10] for a comparison of the Benford and Stigler distributions, and their relations to Zipf and Pareto distributions. For derivations of Stigler’s distribution, see [LG78, Rai85]., 1945. URL http://www.nobelprize.

org/nobel_prizes/economics/laureates/1982/.

Stigler:1987:PMF

[Sti87] Stephen M. Stigler. Precise measurement in the face of error: a comment on MacRoberts and MacRoberts. Social Studies of Science, 17(2):332–334, May 1, 1987. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659

(electronic). URL https://journals.sagepub.com/doi/
pdf/10.1177/030631287017002007. See [MM86] and reply

[MM87].

Sambridge:2009:ABL

[STJ09] M. Sambridge, H. Tkalcic, and A. Jackson. On the applicability of Benford’s Law in the geosciences. AGU Fall Meeting Abstracts, pages A1756–??, December 2009. CODEN ???? ISSN

???? URL http://adsabs.harvard.edu/abs/2009AGUFM.
S33A1756S.

Sambridge:2010:BLN

[STJ10] M. Sambridge, Hrvoje Tkalčić, and A. Jackson. Benford’s law in the natural sciences. Geophysical Research Letters, 37(22): L22301–L22306, ???? 2010. CODEN GPRLAJ. ISSN 0094-8276 (print), 1944-8007 (electronic). URL http://adsabs.

harvard.edu/abs/2010GeoRL..3722301S;   http://www.
agu.org/pubs/crossref/2010/2010GL044830.shtml.

Stoerk:2016:SCB

[Sto16] Thomas Stoerk. Statistical corruption in Beijing’s air quality data has likely ended in 2012. Atmospheric Environment, 127(??):365–371, 2016. CODEN ATENBP. ISSN 1352-2310 (print), 1873-2844 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1352231015306336.

Strzalka:2010:SPB

[Str10] Dominik Strzałka. On some properties of Benford’s law. Journal of the Korean Mathematical Society, 47(5):1055–1075, 2010. CODEN JKMSDG. ISSN 0304-9914.

Strauch:2012:UP

[Str12] O. Strauch. Unsolved problems. Uniform Distribution Theory, ??(??):1–98, ???? 2012. ISSN 1336-913X. URL http:

//www.boku.ac.at/MATH/udt/unsolvedproblems.pdf. Un-

solved Problems Section on the home page of Uniform Distribution Theory.

Sudhier:2010:BLS

[Sud10] K. G. Sudhier. Bradford’s law of scattering revisited: A study based on the references in doctoral theses in the area of physics. Collnet Journal of Scientometrics and Information Management, 4(2):35–47, 2010. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Sun:1928:EFD

[Sun28] C. P. Sun. On the examination of final digits by experiments in artificial sampling. Biometrika, 20A(1/2):64–68, July 1928. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2331941.

Sury:1993:FPD

[Sur93] B. Sury. Fractional parts of log p and a digit function. Expositiones Mathematicae, 11(4):381–384, ???? 1993. ISSN 0723-0869 (print), 1878-0792 (electronic). URL ????.

Shengmin:2010:DSC

[SW10] Z. Shengmin and W. Wenchao. Does [sic] Chinese stock indices agree with Benford’s Law? In IEEE, editor, 2010 International Conference on Management and Service Science (MASS), 24–26 Aug. 2010, Wuhan, China, pages 1–3. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2010. ISBN 1-4244-5325-9, 1-4244-5326-7. LCCN ????

Szekely:1990:PKN

[Szé90] Gábor J. Székely. Paradoxa: klassische und neue Überraschungen aus Wahrscheinlichkeitsrechnung und mathematischer Statistik. (German) [Paradoxes: traditional and new surprises in probability and mathematical statistics]. Deutsch, Thun, Switzerland, 1990. ISBN 3-87144-850-8. 239 pp. LCCN ????

URL http://www.gbv.de/dms/hbz/toc/ht003363686.pdf;
 http://www.zentralblatt-math.org/zmath/en/search/

?an=0743.60004. Translation by Ulrike Leitner from the Hungarian original.

Szewczak:2010:LTR

[Sze10] Zbigniew S. Szewczak. A limit theorem for random sums modulo 1. Statistics & Probability Letters, 80(9–10):747–751, May 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715210000106.

Saldana-Zepeda:2010:GFT

[SZVHA10] Dayna P. Saldaña-Zepeda, Humberto Vaquera-Huerta, and Barry C. Arnold. A goodness of fit test for the Pareto distribution in the presence of type II censoring, based on the cumulative hazard function. Computational Statistics & Data Analysis, 54(4):833–842, April 1, 2010. CO-DEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167947309004071.

Tavangar:2012:SUC

[TA12] Mahdi Tavangar and Majid Asadi. Some unified characterization results on the generalized Pareto distributions based on generalized order statistics. Metrika, 75(7):997–1007, October 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X

(electronic). URL http://link.springer.com/article/10.
1007/s00184-011-0364-6.

Taguchi:1970:ZLC

[Tag70] Tokio Taguchi. On Zipf’s law. A characterization of distributions in linguistics and demography. Proceedings of the Institute of Statistical Mathematics, 17:83–90, 1970. CODEN ???? ISSN 0912-6112.

Tao:2009:BLZ

[Tao09] Terrence Tao. Benford’s law, Zipf’s law, and the Pareto distribution. Web document., 2009. URL http://terrytao.

wordpress.com/2009/07/03/benfords-law-zipfs-law-and-
the-pareto-distribution/.

Taylor:2005:TMT

[Tay05] Jonathan Taylor. Too many ties? An empirical analysis of the Venezuelan recall referendum counts. Technical report ??, Department of Statistics, Stanford University, Stanford, CA

94305, USA, 2005. URL http://esdata.info/pdf/Taylor-
Ties.pdf;   http://www.benfordonline.net/fullreference/
577.

Tapp:2000:UTD

[TB00] D. J. Tapp and D. B. Burg. Using technology to detect fraud. Pennsylvania CPA Journal, 71(4):20–23, ???? 2000. ISSN 0746-1062.

Troll:1998:ZLC

[TbG98] G. Troll and P. beim Graben. Zipf’s law is not a consequence of the Central Limit Theorem. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 57(3):1347–1355, February 1, 1998. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-2376 (electronic). URL http://

link.aps.org/doi/10.1103/PhysRevE.57.1347;   http:
//pre.aps.org/abstract/PRE/v57/i2/p1347_1.

Tolle:2000:DDS

[TBL00] Charles R. Tolle, Joanne L. Budzien, and Randall A. Laviolette. Do dynamical systems follow Benford’s law? Chaos (Woodbury, NY), 10(2):331–336, ???? 2000. CODEN CHAOEH. ISSN 1054-1500. URL http://adsabs.harvard.

edu/abs/2000Chaos..10..331T;   http://link.aip.org/
link/chaoeh/v10/i2/p331/s1;   http://link.aip.org/
link/doi/10.1063/1.166498.

Torres:2007:HDN

[TFGS07] J. Torres, S. Fernández, A. Gamero, and A. Sola. How do numbers begin? (the first digit law). European Journal of Physics, 28(3):L17–L25, May 2007. CODEN EJPHD4. ISSN 0143-0807 (print), 1361-6404 (electronic). URL http://iopscience.

iop.org/0143-0807/28/3/N04.

Twardy:2005:EDS

[TGD05] Charles Twardy, Steve Gardner, and David L. Dowe. Empirical data sets are algorithmically compressible: reply to McAllister? Studies in History and Philosophy of Science Part A, 36(2):391–402, June 2005. CODEN SH-PSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S003936810500021X. See [McA03, McA05].

Tavangar:2013:CGP

[TH13] Mahdi Tavangar and Marzieh Hashemi. On characterizations of the generalized Pareto distributions based on progressively censored order statistics. Statistical Papers, 54(2):381– 390, May 2013. CODEN STPAE4. ISSN 0932-5026 (print),

1613-9798 (electronic). URL http://link.springer.com/
article/10.1007/s00362-012-0434-5.

Thomas:1989:UPR

[Tho89] Jacob K. Thomas. Unusual patterns in reported earnings. The Accounting Review, 59(4):773–787, October 1989. CODEN ACRVAS. ISSN 0001-4826 (print), 1558-7967 (electronic).

URL http://www.jstor.org/stable/247861.

Tichy:1985:UDD

[Tic85] Robert F. Tichy. Uniform distribution and diophantine inequalities. Monatshefte für Mathematik, 99(2):147–152, June 1985. CODEN MNMTA2. ISSN 0026-9255 (print), 1436-5081

(electronic). URL http://link.springer.com/article/10.
1007/BF01304194.

Tichy:1986:GZU

[Tic86] Robert F. Tichy. Gleichverteilung und zahlentheoretische Ungleichungen II. (German) [Uniform distribution and numbertheoretic inequalities II]. Anzeiger der Österreichischen Akademie der Wissenschaften. Mathematisch-Naturwissenschaftliche Klasse, 122(1985)(??):95–99, ???? 1986. CODEN OSAWA8. ISSN 0065-535X. URL ????.

Tichy:1987:GS

[Tic87] Robert F. Tichy. Gleichverteilung zum Summierungsverfahren H_{∞}. (German) [Uniform distribution to the H_{∞} summation method]. Mathematische Nachrichten, 131(??):119–125, ???? 1987. CODEN MTMNAQ. ISSN 0025-584X. URL ????.

Tichy:1988:SRC

[Tic88] Robert F. Tichy. Statistische Resultate über computergerechte Darstellungen von Zahlen. (German) [statistical results on computer-oriented representations of numbers]. Anzeiger der Österreichischen Akademie der Wissenschaften. Mathematisch-Naturwissenschaftliche Klasse, 1987(??):1–8, ???? 1988. CODEN OSAWA8. ISSN 0065-535X. URL ??

??.

Tijms:2004:UPC

[Tij04] Henk C. Tijms. Understanding probability: chance rules in everyday life. Cambridge University Press, Cambridge, UK, 2004. ISBN 0-521-83329-9 (hardcover), 0-521-54036-4 (paperback). x + 380 pp. LCCN QA273 .T48 2004.

Tijms:2007:UPC

[Tij07] H. C. Tijms. Understanding probability: chance rules in everyday life. Cambridge University Press, Cambridge, UK, second edition, 2007. ISBN 0-511-34626-3 (e-book), 0-511-35003-1 (ebook), 0-521-70172-4 (paperback). x + 442 pp. LCCN QA273 .T48 2007.

Tijms:2012:UPC

[Tij12] Henk C. Tijms. Understanding probability: chance rules in everyday life. Cambridge University Press, Cambridge, UK, third edition, 2012. ISBN 1-107-65856-X. x + 562 pp. LCCN QA273.

Tripathi:2014:IEP

[TKP14] Yogesh Mani Tripathi, Somesh Kumar, and C. Petropoulos. Improved estimators for parameters of a Pareto distribution with a restricted scale. Statistical Methodology, 18(??):1–13, May 2014. CODEN ???? ISSN 1572-3127 (print), 1878-

0954 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1572312713000749.

Taylor:2002:STD

[TMS02] Rosemary N. Taylor, Damian J. McEntegart, and Eleanor C. Stillman. Statistical techniques to detect fraud and other data irregularities in clinical questionnaire data. Drug Information Journal, 36(1):115–125, January 2002. ISSN 0092-

8615. URL http://www.diahome.org/DIAHome/resources/
content.aspx?type=eopdf&file=%2Fproductfiles%2F8357%
2Fdiaj_11524.pdf????.

Tague:1987:MVZ

[TN87] Jean Tague and Paul Nicholls. The maximal value of a Zipf size variable: Sampling properties and relationship to other parameters. Information Processing and Management, 23(3):155–170, June 7, 1987. CODEN IP-MADK. ISSN 0306-4573 (print), 1873-5371 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030645738790001X.

Todter:2009:BLI

[Töd09] Karl-Heinz Tödter. Benford’s Law as an indicator of fraud in economics. German Economic Review, 10(3):339–351, August 2009. CODEN ???? ISSN 1465-6485 (print), 1468-0475 (electronic).

Too:1992:UDM

[Too92] Yeu-Hua Too. On the uniform distribution modulo one of some log-like sequences. Proceedings of the Japan Academy, Series A. Mathematical Sciences, 68(9):269–272, ???? 1992.

ISSN 0021-4280. URL http://projecteuclid.org/euclid.
pja/1195511634.

TorresNunez:2006:NBL

[Tor06] David A. Torres Núñez. Newcomb–Benford’s Law: Applications to electoral processes, bioinformatics, and the stock index. M.S. thesis, Center for Bioststatistics and Bioinformatics, University of Puerto Rico, Río Piedras Campus, PO Box 23355, San Juan, PR 00931-3355, May 2006.

Trubnikov:1991:SDZ

[TR91] B. A. Trubnikov and I. A. Rumynskiĭ. A simple derivation of the Zipf–Krylov law for words and the possibility of its “evolution” interpretation. Doklady Akademii Nauk SSSR, 321(2):270–275, 1991. CODEN DANKAS. ISSN 0002-3264.

Trachtenberg:1992:WFR

[Tra92] M. Trachtenberg. Why failure rates observe Zipf’s law in operational software. IEEE Transactions on Reliability, 41(3): 386–389, 1992. CODEN IEERAJ. ISSN 0018-9529 (print), 1558-1721 (electronic).

Taimori:2012:PTS

[TRB^{+}12] A. Taimori, F. Razzazi, A. Behrad, A. Ahmadi, and M. Babaie-Zadeh. A proper transform for satisfying Benford’s Law and its application to double JPEG image forensics. In 2012 IEEE International Symposium on Signal Processing and Information Technology (ISSPIT), pages 000240–000244. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, December 2012. ISSN

2162-7843. URL http://ieeexplore.ieee.org/document/
6621294/.

Tritch:1994:MAC

[TS94] Teresa Tritch and Mary L. Sprouse. Money audits the Clintons. Money Magazine, ??(??):84–98, April 1,

1994. URL http://money.cnn.com/magazines/moneymag/
moneymag_archive/1994/04/01/88775/index.htm.

Tsao:1974:DSD

[Tsa74] Nai Kuan Tsao. On the distributions of significant digits and roundoff errors. Communications of the ACM, 17(5):269–271, May 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Tsonis:1997:ZLS

[TST97] A. A. Tsonis, C. Schultz, and P. A. Tsonis. Zipf’s Law and the structure and evolution of languages. Complexity, 2(5):12–13, 1997. CODEN COMPFS. ISSN 1076-2787 (print), 1099-0526 (electronic).

Tsuji:1952:UDN

[Tsu52] M. Tsuji. On the uniform distribution of numbers mod1. Journal of the Mathematical Society of Japan, 4(3–4):313–322, ???? 1952. CODEN NISUBC. ISSN 0025-5645. URL http:/

/projecteuclid.org/euclid.jmsj/1261415780.

Tuldava:1985:FST

[Tul85] Yu. A. Tuldava. The frequency structure of a text and Zipf’s law. Tartu Riikl. Ül. Toimetised, ??(711):93–116, 1985. CO-DEN ???? ISSN ????

Turner:1982:DLS

[Tur82] Peter R. Turner. The distribution of leading significant digits. IMA Journal of Numerical Analysis, 2(4):407–412, October 1982. CODEN IJNADH. ISSN 0272-4979 (print), 1464-3642 (electronic).

Turner:1984:FRD

[Tur84] Peter R. Turner. Further revelations on l.s.d. IMA Journal of Numerical Analysis, 4(2):225–231, April 1984. CODEN IJNADH. ISSN 0272-4979 (print), 1464-3642 (electronic).

Turner:1987:DDI

[Tur87] Peter R. Turner. The distribution of l.s.d. and its implications for computer design. The Mathematical Gazette, 71(455):26– 31, March 1987. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). [l.s.d. = least significant digits]. The topic is variously known as Benford’s Law, the Law of Anomalous Numbers, and Zipf’s Law.

Tosic:2021:UBL

[TV21] Aleksandar Tošić and Jernej Vičič. Use of Benford’s law on academic publishing networks. Journal of Informetrics, 15(3):Article 101163, August 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157721000341.

Thorp:1971:PCD

[TW71] E. Thorp and R. Whitley. Poincaré’s conjecture and the distribution of digits in logarithm tables. Compositio Mathematica, 21(??):233–250, ???? 1971. CODEN CMPMAF. ISSN 0010-437X (print), 1570-5846 (electronic).

Tang:2011:ESP

[TW11] Lian Tang and Da Hui Wang. An explanation of the shift parameter ρ in the Zipf–Mandelbrot law. Beijing Shifan Daxue Xuebao, 47(1):97–100, 2011. CODEN BSDKDH. ISSN 0476-0301.

Uppuluri:1985:DFD

[UP85] V. R. R. Uppuluri and S. A. Patil. The distribution of the first j digits of beta related random variables. Communications in Statistics: Simulation and Computation, 14(2):467–472, 1985. CODEN CSSCDB. ISSN 0361-0918.

Urzua:2000:SET

[Urz00] Carlos M. Urzúa. A simple and efficient test for Zipf’s law. Economics Letters, 66(3):257–260, March 1, 2000. CO-DEN ECLEDS. ISSN 0165-1765 (print), 1873-7374 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165176599002153.

Urzua:2011:TZL

[Urz11] Carlos M. Urzúa. Testing for Zipf’s law: a common pitfall. Economics Letters, 112(3):254–255, September 1, 2011. CODEN ECLEDS. ISSN 0165-1765 (print), 1873-7374 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165176511002266.

Uzer:2010:MTC

[Uze10] Ali Uzer. Multiplicative type complex calculus as an alternative to the classical calculus. Computers and Mathematics with Applications, 60(10):2725–2737, November 2010. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0898122110006747.

Vandewalle:1999:ZAF

[VA99] N. Vandewalle and M. Ausloos. The n-Zipf analysis of financial data series and biased data series. Physica A: Statistical Mechanics and its Applications, 268(1–2):240–249, June 1, 1999. CODEN PHYADX. ISSN 0378-4371 (print), 1873-

2119 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037843719900031X.

Villasenor-Alva:2009:BGF

[VAGE09] José A. Villaseñor-Alva and Elizabeth González-Estrada. A bootstrap goodness of fit test for the generalized Pareto distribution. Computational Statistics & Data Analysis, 53(11):3835–3841, September 1, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167947309001406.

Vallee:2011:BMF

[Val11] Robert Vallée. Benoît Mandelbrot — further notes. Kybernetes, 40(9–10):??, ???? 2011. CODEN KBNTA3. ISSN 0368-492X (print), 1758-7883 (electronic). URL http://www.

emeraldinsight.com/journals.htm?issn=0368-492X&volume=
40&issue=9/10&articleid=17003220.

Valadier:2012:BPR

[Val12] Michel Valadier. The Benford phenomenon for random variables. Discussion of Feller’s way. ArXiv e-prints, page 23,

March 9, 2012. URL http://arxiv.org/abs/1203.2518.

VanCaneghem:2002:EMI

[Van02] T. Van Caneghem. Earnings management induced by cognitive reference points. The British Accounting Review, 34 (2):167–178, 2002. ISSN 0890-8389 (print), 1095-8347 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0890838902901903.

VanCaneghem:2015:NFSa

[Van15a] Tom Van Caneghem. NPO financial statement quality: An empirical analysis based on Benford’s Law. VOLUNTAS: International Journal of Voluntary and Nonprofit Organizations, 26(1):1–24, February 2015. ISSN 0957-8765 (print), 1573-7888

(electronic). URL http://link.springer.com/article/10.
1007/s11266-015-9629-4.

VanCaneghem:2015:NFSb

[Van15b] Tom Van Caneghem. NPO financial statement quality: An empirical analysis based on Benford’s Law. VOLUNTAS: International Journal of Voluntary and Nonprofit Organizations, 27(6):2685–2708, August 2015. ISSN 0957-8765 (print), 1573-7888 (electronic).

Varian:1972:LEB

[Var72] Hal R. Varian. Letter to the Editor: Benford’s Law. The American Statistician, 26(3):65–66, June 1972. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://links.jstor.org/sici?sici=0003-
1305%28197206%2926%3A3%3C62%3ALTTE%3E2.0.CO%3B2-Q.

Vardi:1999:PCS

[Var99] Ilan Vardi. Premiers chiffres significatifs et nombres algébriques. (French) [significant leading digits and algebraic numbers]. Comptes rendus de l’Académie des sciences. Série I, Mathématique, 328(9):749–754, May 1999. CODEN CASMEI. ISSN 0764-4442 (print), 1778-3577 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0764444299802651.

Volchenkov:2008:SUC

[VB08] D. Volchenkov and Ph. Blanchard. Scaling and universality in city space syntax: Between Zipf and Matthew. Physica A: Statistical Mechanics and its Applications, 387(10):2353–2364, April 1, 2008. CODEN PHYADX. ISSN 0378-4371 (print),

1873-2119 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378437107012630.

VandenBroeck:1990:LFB

[VK90] C. Van den Broeck and R. Kawai. Learning in feedforward Boolean networks. Physical Review A (Atomic, Molecular, and Optical Physics), 42(??):6210–??, November 15, 1990. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-

4446, 1538-4519. URL http://link.aps.org/doi/10.1103/
PhysRevA.42.6210.

Vlachy:1972:VFS

[Vla72] Jan Vlachý. Variable factors in scientific communities (observations on Lotka’s Law). Teorie a Metoda, 4(??):91–120, ???? 1972. CODEN ???? ISSN ????

Vlachy:1976:TFL

[Vla76] Jan Vlachý. Time factor in Lotka’s Law. Probleme de Informare si Documentare, 10(2):44–87, ???? 1976. CODEN PIDCA6. ISSN 0032-924X.

Vlachy:1979:FDS

[Vla79] Jan Vlachý. Frequency distribution of scientific performance: a bibliography of Lotka’s Law and related phenomena. Scientometrics, 1(1):107–130, September 1979. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/BF02016844;   http://www.springerlink.com/
content/0138-9130/.

vanLeijenhorst:2005:FDH

[vLvdW05] D. C. van Leijenhorst and Th. P. van der Weide. A formal derivation of Heaps’ law. Information Sciences, 170(2–4):263– 272, 2005. CODEN ISIJBC. ISSN 0020-0255 (print), 1872-6291 (electronic).

Vanichpun:2004:OCU

[VM04] Sarut Vanichpun and Armand M. Makowski. The output of a cache under the independent reference model: where did the locality of reference go? ACM SIGMETRICS Performance Evaluation Review, 32(1):295–306, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Vogt:2000:BGS

[Vog00] W. Vogt. Benford’s Gesetz: Steuer- und Budgetsündern auf der Spur — Zahlen lügen nichts. (German) [Benford’s Law: tax and budget sinners lying on the track — numbers do not lie]. Schweizer Versicherung, 9(??):27–29, ???? 2000. URL

????.

Volcic:2020:UDB

[Vol20] Aljoša Volčič. Uniform distribution, Benford’s law and scaleinvariance. Bollettino della Unione matematica italiana, 13 (4):539–543, 2020. CODEN BLUMAM. ISSN 1972-6724.

Voos:1974:BCL

[Voo74a] Henry Voos. Brief communications: Lotka and information science. Journal of the American Society for Information Science, 25(4):270–272, July 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Voos:1974:LIS

[Voo74b] Henry Voos. Lotka and information science. Journal of the American Society for Information Science, 25(4):270–272, July/August 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Voss:1996:CLF

[Vos96] Richard F. Voss. Comment on “Linguistic Features of Noncoding DNA Sequences”. Physical Review Letters, 76(11):1978, March 1996. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145. URL http://link.aps.

org/doi/10.1103/PhysRevLett.76.1978. See [MBG^{+}94]

and reply [MBG^{+}96].

Vuko:2008:UBL

[Vuk08] Tina Vuko. Using Benford’s Law to detect earnings management practice: the case of listed companies in Croatia. Global Business & Economics Anthology, 1(??):69–74, ???? 2008. CODEN ???? ISSN 1553-1333.

VandeWalle:2006:ZGK

[VW06] J. Van de Walle and K. Willems. Zipf, George Kingsley (1902–1950). Journal of Economic Dynamics and Control, 31(7):756–757, July 15, 2006. CODEN JED-CDH. ISSN 0165-1889 (print), 1879-1743 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/B008044854204726X.

Velleman:2022:ECD

[VW22] Paul Velleman and Howard Wainer. Exploring COVID data with Benford’s and Zipf’s laws. Chance: a magazine for people interested in the analysis of data, 35(2):11–15, 2022. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/09332480.2022.2066410.

vanZyl:2013:GPD

[vZ13] J. Martin van Zyl. The generalized Pareto distribution fitted to research outputs of countries. Scientometrics, 94(3): 1099–1109, March 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0798-2.

Wagon:2010:BLD

[Wag10] S. Wagon. Benford’s Law and data spread, 2010. URL http:/

/demonstrations.wolfram.com/BenfordsLawAndDataSpread.

Web site.

Wallace:2002:AQD

[Wal02] W. A. Wallace. Assessing the quality of data used for benchmarking and decision making. Journal of Government Financial Management, 51(3):16–22, ???? 2002. ISSN 1533-1385.

Wales:2003:IWI

[Wal03] Elspeth Wales. Is it worth investing in software to combat fraud? Computer Fraud & Security, 2003(5):6– 8, 2003. ISSN 1361-3723 (print), 1873-7056 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1361372303050097.

Wang:2011:BLR

[Wan11a] Jialan Wang. Benford’s law: a revised analysis. Blog post-

ing, October 24, 2011. URL http://econerdfood.blogspot.
com/2011/10/benfords-law-revised-analysis.html.

Wang:2011:BLD

[Wan11b] Jialan Wang. Benford’s Law and the decreasing reliability of accounting data for US firms. Studies in everyday life: Investigations and experiments on food, ethics, economics, and the environment. Blog posting, October 9,

2011. URL http://econerdfood.blogspot.com/2011/10/
benfords-law-and-decreasing-reliability.html.

Wang:2015:SPW

[Wan15] Xiaohuan Wang. Stability of planar waves in a Lotka– Volterra system. Applied Mathematics and Computation, 259(??):313–326, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0096300315002398.

Wapner:2012:UEC

[Wap12] Leonard M. Wapner. Unexpected expectations: the curiosities of a mathematical crystal ball. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2012. ISBN 1-56881-721-5 (hardback). xvi + 204 pp. LCCN QA95 .W34 2012.

Warren:2003:DLD

[War03a] Henry S. Warren. The distribution of leading digits. In Hacker’s delight [War03b], chapter 15.3, pages 264–267. ISBN 0-201-91465-4. LCCN QA76.6 .W375 2003. URL http://

www.awprofessional.com/bookstore/product.asp?isbn=
0201914654;   http://www.hackersdelight.org/; http:
//www.hackersdelight.org/hackerTOC.pdf;   http://www.
informit.com/content/images/chap3_0201914654/elementLinks/

0201914654.pdf. While this book does not specifically address computational aspects of floating-point arithmetic (apart from the nine-page Chapter 15), it has extensive coverage of, and clever algorithms for, integer arithmetic operations that are fundamental for implementing hardware floatingarithmetic and software multiple-precision arithmetic.

Warren:2003:HD

[War03b] Henry S. Warren. Hacker’s delight. Addison-Wesley, Reading, MA, USA, 2003. ISBN 0-201-91465-4. xiv + 306 pp. LCCN

QA76.6 .W375 2003. URL http://www.awprofessional.
com/bookstore/product.asp?isbn=0201914654;   http:
//www.hackersdelight.org/; http://www.hackersdelight.
org/hackerTOC.pdf;   http://www.informit.com/content/
images/chap3_0201914654/elementLinks/0201914654.pdf.

While this book does not specifically address computational aspects of floating-point arithmetic (apart from the nine-page Chapter 15), it has extensive coverage of, and clever algorithms for, integer arithmetic operations that are fundamental for implementing hardware floating-arithmetic and software multiple-precision arithmetic.

Washington:1981:BLF

[Was81] Lawrence C. Washington. Benford’s Law for Fibonacci and Lucas numbers. Fibonacci Quarterly, 19(2):175–177, April 1981. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/19-2/washington.pdf.

Wase:2021:BLB

[Was21] Viktor Wase. Benford’s Law in the Beale ciphers. Cryptologia, 45(3):282–286, 2021. CODEN CRYPE6. ISSN 0161-1194 (print), 1558-1586 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/01611194.2020.1821409.

Watanabe:1996:EZL

[Wat96a] Makoto S. Watanabe. Erratum: Zipf’s law in percolation. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 54(??):4483, October 1, 1996. CODEN PLEEE8. ISSN 1539-3755 (print),

1550-2376 (electronic). URL http://link.aps.org/doi/10.
1103/PhysRevE.54.4483. See [Wat96b].

Watanabe:1996:ZLP

[Wat96b] Makoto S. Watanabe. Zipf’s law in percolation. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 53(??):4187–4190, April 1, 1996. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-

2376 (electronic). URL http://link.aps.org/doi/10.
1103/PhysRevE.53.4187. See erratum [Wat96a].

Woods:2015:PDP

[WAT15] Louis Woods, Gustavo Alonso, and Jens Teubner. Parallelizing data processing on FPGAs with shifter lists. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 8(2):7:1–7:??, April 2015. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Wu:2003:IPD

[WC03] Shuo-Jye Wu and Chun-Tao Chang. Inference in the Pareto distribution based on progressive Type II censoring with random removals. Journal of Applied Statistics, 30(2):163–172, ???? 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wang:2009:UBL

[WCCK09] J. Wang, B. H. Cha, S. H. Cho, and C. C. J. Kuo. Understanding Benford’s Law and its vulnerability in image forensics. In IEEE, editor, 2009 IEEE International Conference on Multimedia and Expo, pages 1568–1571. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, June 2009. ISSN 1945-7871.

Wells:2007:QPV

[WCP^{+}07] K. Wells, J. Chiverton, M. Partridge, M. Barry, H. Kadhem, and B. Ott. Quantifying the partial volume effect in PET using Benford’s Law. IEEE Transactions on Nuclear Science, 54 (5):1616–1625, October 2007. CODEN IRNSAM. ISSN 0018-9499 (print), 1558-1578 (electronic). URL http://adsabs.

harvard.edu/abs/2007ITNS...54.1616W.

Wagner-Dobler:1995:DLL

[WDB95] Roland Wagner-Döbler and Jan Berg. The dependence of Lotka’s Law on the selection of time periods in the development of scientific areas and authors. Journal of Documentation, 51(1):28–43, ???? 1995. CODEN JDOCAS. ISSN 0022-0418 (print), 1758-7379 (electronic).

Wolf:1993:PSM

[WDY93] Joel L. Wolf, Daniel M. Dias, and Philip S. Yu. Parallel sort merge join algorithm for managing data skew. IEEE Transactions on Parallel and Distributed Systems, 4(1):70–86, January 1993. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Weaver:1963:LLT

[Wea63] Warren Weaver. Lady Luck: The Theory of Probability. Doubleday Anchor Series, New York, NY, USA, 1963. 392 pp. LCCN QA273 .W4. Benford’s Law is discussed on pp. 270–

Weaver:1982:LLT

[Wea82] Warren Weaver. Lady Luck: the theory of probability. Dover Publications, Inc., New York, NY, USA, 1982. ISBN 0-486-24342-7. 392 pp. LCCN QA273 .W4 1982. URL http://www.

loc.gov/catdir/description/dover031/82007396.html.

Webb:1975:DFD

[Web75] William Webb. Distribution of the first digits of Fibonacci numbers. Fibonacci Quarterly, 13(4):334–336, December 1975. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.

math.ca/Scanned/13-4/webb.pdf.

Weisstein:2011:BL

[Wei11] Eric W. Weisstein. Benford’s Law. MathWorld—A Wolfram

Web Resource, 2011. URL http://mathworld.wolfram.com/
BenfordsLaw.html.

Wells:2005:PNM

[Wel05] D. G. (David G.) Wells. Prime Numbers: the Most Mysterious Figures in Math. Wiley, New York, NY, USA, 2005. ISBN 0-471-46234-9 (hardcover), 0-471-71892-0 (ebook). xv + 272 pp. LCCN QA246 .W35 2005. URL http://

www.loc.gov/catdir/enhancements/fy0616/2004019974-
b.html;   http://www.loc.gov/catdir/enhancements/fy0616/
2004019974-d.html;   http://www.loc.gov/catdir/toc/
ecip0422/2004019974.html.

Weyl:1915:GZM

[Wey15] Hermann Weyl. Über die Gleichverteilung von Zahlen mod. Eins. (German) [On the uniform distribution of numbers mod. one]. Mathematische Annalen, 77(3):313–352, ???? 1915. CO-DEN MAANA3. ISSN 0025-5831 (print), 1432-1807 (electronic).

Wilk:1968:PPM

[WG68] M. B. Wilk and R. Gnanadesikan. Probability plotting methods for the analysis of data. Biometrika, 55(1):1–17, March 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334448.

Woodroofe:1975:ZL

[WH75] Michael Woodroofe and Bruce Hill. On Zipf’s law. Journal of Applied Probability, 12(3):425–434, September 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212857.

Whitney:1972:MNI

[Whi72] R. E. Whitney. Mathematical notes: Initial digits for the sequence of primes. American Mathematical Monthly, 79(2): 150–152, February 1972. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

White:1982:FSS

[Whi82] Warren H. White. On the form of steady-state solutions to the coagulation equations. Journal of Colloid and Interface Science, 87(1):204–208, May 1982. CO-DEN JCISA5. ISSN 0021-9797 (print), 1095-7103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0021979782903824.

Whittaker:1983:SID

[Whi83] James V. Whittaker. On scale-invariant distributions. SIAM Journal on Applied Mathematics, 43(2):257–267, April 1983. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wainer:2013:VRS

[WHN13] Howard Wainer, Polina Harik, and John Neter. Visual revelations: Stigler’s Law of Eponymy and Marey’s train schedule: Did Serjev do it before Ibry, and what about Jules Petiet? Chance: a magazine for people interested in the analysis of data, 26(1):53–56, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Walthoe:1999:LNO

[WHP99] Jon Walthoe, Robert Hunt, and Mike Pearson. Looking out for number one. Plus Magazine, ??(??):1–9, September 1999.

URL http://plus.maths.org/issue9/features/benford/

2pdf/index.html/op.pdf. Published by the University of Cambridge Millennium Mathematics Project.

Wilkinson:1972:ABL

[Wil72] Elizabeth A. Wilkinson. The ambiguity of Bradford’s Law. Journal of Documentation, 28(2):122–130, ???? 1972. CO-DEN JDOCAS. ISSN 0022-0418 (print), 1758-7379 (electronic).

Wilkinson:2006:RPC

[Wil06] Leland Wilkinson. Revising the Pareto chart. The American Statistician, 60(4):332–334, November 2006. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wirsching:1998:DSG

[Wir98] Günther J. Wirsching. The dynamical system generated by the 3n + 1 function, volume 1681 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 3-540-63970-5 (softcover). ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 158 pp. LCCN

QA3 .L28 no. 1681. URL http://www.loc.gov/catdir/
enhancements/fy0818/97051329-d.html;   http://www.
loc.gov/catdir/enhancements/fy0818/97051329-t.html.

Wlodarski:1971:FLN

[Wlo71] J. Wlodarski. Fibonacci and Lucas numbers tend to obey Benford’s Law. Fibonacci Quarterly, 9(1):87–88, February 1971. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.

math.ca/Scanned/9-1/wlodarski2.pdf.

Wjcik:2013:HFI

[Wój13a] Michał Ryszard Wójcik. How fast increasing powers of a continuous random variable converge to Benford’s Law. Statistics & Probability Letters, 83(12):2688–2692, 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213002927.

Wojcik:2013:HFI

[Wój13b] Michal Ryszard Wójcik. How fast increasing powers of a continuous random variable converge to Benford’s Law. Statistics & Probability Letters, 83(12):2688–2692, December 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213002927.

Wojcik:2014:CBL

[Wój14] Michał Ryszard Wójcik. A characterization of Benford’s Law through generalized scale-invariance. Mathematical Social Sciences, 71(??):1–5, September 2014. ISSN 0165-4896 (print),

1879-3118 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165489614000262.

Wong:1969:GDS

[Won69] James Teng Wong. On the Generalization of the Distribution of the Significant Digits Under Computation. Ph.D. thesis, Department of Mathematics, Oregon State University, Corvallis,

OR, USA, 1969. 52 pp. URL http://search.proquest.com/
docview/302476508?accountid=14677.

Worthen:1975:ABL

[Wor75] Dennis B. Worthen. The application of Bradford’s Law to monographs. Journal of Documentation, 31(1):19–25, ???? 1975. CODEN JDOCAS. ISSN 0022-0418 (print), 1758-7379 (electronic).

Wu:2003:PDC

[WR03] Wei Biao Wu and Chinya V. Ravishankar. The performance of difference coding for sets and relational tables. Journal of the ACM, 50(5):665–693, September 2003. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Whyman:2015:ICC

[WSB15] G. Whyman, E. Shulzinger, and Ed. Bormashenko. Intuitive considerations clarifying the origin and applicability of the Benford Law. arXiv.org, ??(??):12, October 25, 2015. URL

http://arxiv.org/abs/1510.07220v2.

Whyman:2016:ICC

[WSB16] G. Whyman, E. Shulzinger, and Ed. Bormashenko. Intuitive considerations clarifying the origin and applicability of the Benford Law. Results in Physics, 6(??):3–6, 2016. ISSN

2211-3797. URL http://www.sciencedirect.com/science/
article/pii/S2211379715000728.

Watrin:2008:BLI

[WSU08] C. Watrin, R. Struffert, and R. Ullmann. Benford’s Law: An instrument for selecting tax audit targets? Review of Managerial Science, 2(3):219–237, November 2008. CO-DEN ???? ISSN 1863-6683 (print), 1863-6691 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11846-008-0019-9.

Winter:2011:DFU

[WSY11] Christian Winter, Markus Schneider, and York Yannikos. Detecting fraud using modified Benford analysis. In Gilbert Peterson and Sujeet Shenoi, editors, Advances in Digital Forensics VII: 7th IFIP WG 11.9 International Conference on Digital Forensics, Orlando, FL, USA, January 31– February 2, 2011, Revised Selected Papers. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 3-642-24211-1 (paperback), 3-642-24212-X (e-book). LCCN HV8079.C65 I357 2011. US$. URL http://link.

springer.com/chapter/10.1007/978-3-642-24212-0_10/.

Wu:2008:IEP

[Wu08] Shu-Fei Wu. Interval estimation for a Pareto distribution based on a doubly type II censored sample. Computational Statistics & Data Analysis, 52(7):3779–3788, March 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167947308000030.

Wolf:1993:PHJ

[WYTD93] Joel L. Wolf, Philip S. Yu, John Turek, and Daniel M. Dias. A parallel hash join algorithm for managing data skew. IEEE Transactions on Parallel and Distributed Systems, 4(12):1355– 1371, December 1993. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2012:LPW

[WZS12] Dongbo Wang, Danhao Zhu, and Xinning Su. Lotka phenomenon in the words’ syntactic distribution complexity. Scientometrics, 90(2):483–498, February 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0546-z.

Xie:2014:GTK

[XLW14] Min Xie, Laks V. S. Lakshmanan, and Peter T. Wood. Generating top-k packages via preference elicitation. Proceedings of the VLDB Endowment, 7(14):1941–1952, October 2014. CO-DEN ???? ISSN 2150-8097.

Xekalaki:1989:APY

[XP89] E. Xekalaki and J. Panaretos. On the association of the Pareto and the Yule distribution. Theory of Probability and its Applications, 33(1):191–195, ???? 1989. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Xie:2009:FAS

[XS09] Tao Xie and Yao Sun. A file assignment strategy independent of workload characteristic assumptions. ACM Transactions on Storage, 5(3):10:1–10:??, November 2009. CODEN ???? ISSN 1553-3077 (print), 1553-3093 (electronic).

Xu:2011:PCG

[XWLD11] B. Xu, J. Wang, G. Liu, and Y. Dai. Photorealistic computer graphics forensics based on leading digit law. Journal of Electronics (China), 28(1):95–100, ???? 2011. URL http://link.

springer.com/article/10.1007/s11767-011-0474-3.

Yavuz:1974:ZLE

[Yav74] Davras Yavuz. Zipf’s law and entropy. IEEE Transactions on Information Theory, 20(5):650, 1974. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Yang:2015:ERP

[YD15] Ziyun Yang and Xiaobo Dong. Earnings roundup in private and public bank holding companies. Advances in Accounting, 31(1):96–99, 2015. ISSN 0882-6110. URL http://www.

sciencedirect.com/science/article/pii/S088261101500005X.

Yeh:1990:ODT

[Yeh90] Hsiaw-Chan Yeh. One discrete time series model for fat-tailed integer random variables: Zipf process. Bulletin of the Institute of Mathematics. Academia Sinica, 18(1):19–33, 1990. CODEN BIMSDG. ISSN 0304-9825.

Yeh:2002:SMZ

[Yeh02] Hsiaw-Chan Yeh. Six multivariate Zipf distributions and their related properties. Statistics & Probability Letters, 56 (2):131–141, January 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715201001493.

Yeh:2007:SMD

[Yeh07] Hsiaw-Chan Yeh. Some multivariate discrete time series models for dependent multivariate Zipf counts. Bulletin of the Institute of Mathematics. Academia Sinica. New Series, 2(1): 29–53, 2007. CODEN ???? ISSN 0304-9825.

Ylvisaker:1977:TR

[Ylv77] Donald Ylvisaker. Test resistance. Journal of the American Statistical Association, 72(359):551–556, September 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2286216.

Yang:2014:RAE

[YLX14] Z. K. Yang, D. M. Lin, and M. Z. Xu. The re-applicability explore of Lotka’s Law in patent documents. Collnet Journal of Scientometrics and Information Management, 8(1):183–191, 2014. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

York:2000:ATB

[Yor00] D. York. Auditing technique — Benford’s law. Accountancy: The journal of the Institute of Chartered Accountants in England and Wales, 126(1283):126–??, ???? 2000. CODEN ACTYAD. ISSN 0001-4664.

Young:2013:ATL

[You13] D. S. Young. Approximate tolerance limits for Zipf– Mandelbrot distributions. Physica A: Statistical Mechanics and its Applications, 392(7):1702–1711, April 1, 2013. CO-DEN PHYADX. ISSN 0378-4371 (print), 1873-2119 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378437112010497.

Yan:2013:CNE

[YSC13] Su Yan, W. Scott Spangler, and Ying Chen. Chemical name extraction based on automatic training data generation and rich feature set. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 10(5):1218–1233, September 2013. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Yuan:1984:LOC

[Yua84] Y. Yuan. On the least Q-order of convergence of variable metric algorithms. IMA Journal of Numerical Analysis, 4 (2):233–239, April 1984. CODEN IJNADH. ISSN 0272-4979 (print), 1464-3642 (electronic).

Yule:1925:MTE

[Yul25] G. Udny Yule. A mathematical theory of evolution, based on the conclusions of Dr. J. C. Willis, F.R.S. Philosophical transactions of the Royal Society of London Series B, Biological sciences, 213(??):21–87, May 14, 1925. CODEN PTRBAE. ISSN 0962-8436 (print), 1471-2970 (electronic). URL http://rstb.

royalsocietypublishing.org/content/213/402-410/21.

Yule:1927:RS

[Yul27] G. Udny Yule. On reading a scale. Journal of the Royal Statistical Society, 90(3):570–587, ???? 1927. ISSN 0952-8385.

URL http://www.jstor.org/stable/2341205. See related

later work [Pre81].

Yang:2016:WSZ

[YZ16] Yue Yang and Jianwen Zhu. Write skew and Zipf distribution: Evidence and implications. ACM Transactions on Storage, 12 (4):21:1–21:??, August 2016. CODEN ???? ISSN 1553-3077 (print), 1553-3093 (electronic).

Zornig:1995:URZ

[ZA95] Peter Zörnig and Gabriel Altmann. Unified representation of Zipf distributions. Computational Statistics & Data Analysis, 19(4):461–473, April 1, 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0167947394000098.

Zeller:1970:TLG

[ZB70] Karl Zeller and Wolfgang Beekmann. Theorie der Limitierungsverfahren. (German) [Theory of limiting processes], volume 15 of Ergebnisse der Mathematik und ihrer Grenzgebiete. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1970. xii + 314 pp. LCCN QA295 .Z4 1970.

Zhang:2009:ZDT

[ZCW09] Jianhua Zhang, Qinghua Chen, and Yougui Wang. Zipf distribution in top Chinese firms and an economic explanation. Physica A: Statistical Mechanics and its Applications, 388(10):2020–2024, May 15, 2009. CODEN PHYADX. ISSN 0378-4371 (print), 1873-2119 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378437109000806.

Zgela:2011:ATC

[ŽD11] Mario Žgela and Jasminka Dobša. Analysis of top 500 Central and East European companies net income using Benford’s Law. Journal of Informational and Organizational Sciences, 35(2):215–228, December 2011. ISSN 1846-3312 (print), 1846-9418 (electronic). URL http://hrcak.

srce.hr/file/111658;   http://hrcak.srce.hr/index.
php?show=clanak&id_clanak_jezik=111658. Prosinac 2011

conference.

Zeng:2012:TMB

[ZDCW12] Jianping Zeng, Jiangjiao Duan, Wenjun Cao, and Chengrong Wu. Topics modeling based on selective Zipf distribution. Expert Systems with Applications, 39(7):6541–6546, June 1, 2012. CODEN ESAPEH. ISSN 0957-4174 (print), 1873-

6793 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0957417411017222.

Zeller:1958:TLG

[Zel58] Karl Zeller. Theorie der Limitierungsverfahren. (German) [Theory of limiting processes], volume 15 of Ergebnisse der Mathematik und ihrer Grenzgebiete. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1958. 242 pp. LCCN QA295 .Z4.

Zenil:2011:RTC

[Zen11] Hector Zenil, editor. Randomness through computation: some answers, more questions. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2011. ISBN 981-4327-74-3. xviii + 419 pp. LCCN QA274 .R363 2011.

Zhao:2015:UCA

[ZF15] Peng Zhao and Engui Fan. A unified construction for the algebro-geometric quasiperiodic solutions of the Lotka– Volterra and relativistic Lotka–Volterra hierarchy. Journal of Mathematical Physics, 56(4):043501, April 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhao:2009:IFU

[ZHS09] Xi Zhao, A. T. S. Ho, and Y. Q. Shi. Image forensics using generalised Benford’s Law for accurate detection of unknown JPEG compression in watermarked images. In 2009 16th International Conference on Digital Signal Processing, pages 1–8. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, July 2009. ISSN 1546-1874.

Zhao:2010:IFU

[ZHS10] Xi Zhao, Anthony Tung Shuen Ho, and Yun Q. Shi. Image forensics using generalised Benford’s Law for improving image authentication detection rates in semi-fragile watermarking. International Journal of Digital Crime and Forensics, 2(2):1–20, ???? 2010. CODEN ???? ISSN 1941-6210 (print), 1941-6229 (electronic). URL https://dblp.org/db/

journals/ijdcf/ijdcf2.html#ZhaoHS10.

Zipf:1929:RFD

[Zip29] George Kingsley Zipf. Relative frequency as a determinant of phonetic change. Harvard Studies in Classical Philology, 40(??):1–95, ???? 1929. CODEN ???? ISSN 0073-0688.

URL http://www.jstor.org/journals/00730688.html;
http://www.jstor.org/stable/310585.

Zipf:1932:SSP

[Zip32] George Kingsley Zipf. Selected studies of the principle of relative frequency in language. Harvard University Press, Cambridge, MA, USA, 1932. 51 pp. LCCN P123 .Z5. This book is the origin of Zipf’s Law, but the phenomenon was apparently first reported two decades earlier in [Aue13] for city populations, in [Est16] for word frequencies, and in [Lot26] for researcher publication counts.

Zipf:1935:PBL

[Zip35] George Kingsley Zipf. The psycho-biology of language: an introduction to dynamic philology. Houghton-Mifflin, Boston, MA, USA, 1935. ix + 336 pp. LCCN P105 .Z5.

Zipf:1949:HBP

[Zip49] George Kingsley Zipf. Human behaviour and the principle of least effort. Hafner Pub. Co., New York, NY, USA, 1949. xi + 573 pp. LCCN H91 .Z5 1949a.

Zipf:1965:PBL

[Zip65] George Kingsley Zipf. The psycho-biology of language: an introduction to dynamic philology. MIT Press, Cambridge, MA, USA, 1965. xv + 336 pp. LCCN P105 .Z5 1965.

Zahran:2014:EAS

[ZIWU14] Sammy Zahran, Terrence Iverson, Stephan Weiler, and Anthony Underwood. Evidence that the accuracy of self-reported lead emissions data improved: A puzzle and discussion. Journal of Risk and Uncertainty, 49(3):235–257, 2014. CODEN JRUNEN. ISSN 0895-5646 (print), 1573-0476 (electronic).

Krakar:2009:ABL

[ŽK09] Mario Žgela and Zdravko Krakar. Application of Benford’s Law in payment system auditing. Journal of Informational and Organizational Sciences, 33(1):39–51, ???? 2009. CODEN ???? ISSN 1846-3312 (print), 1846-

9418 (electronic). URL http://bib.irb.hr/datoteka/
457819.04_Application_of_Benford_s_Law_in_Payment_
Systems_Auditing1.pdf;   http://jios.foi.hr/index.
php/jios/article/view/108/66;   http://jios.foi.hr/
index.php/jios/issue/view/13.

Zanette:1997:RIU

[ZM97] Damián H. Zanette and Susanna C. Manrubia. Role of intermittency in urban development: A model of large-scale city formation. Physical Review Letters, 79(3):523–526, July 3, 1997. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145. URL http://link.aps.org/

doi/10.1103/PhysRevLett.79.523.

Zanette:2001:VTC

[ZM01] Damián H. Zanette and Susanna C. Manrubia. Vertical transmission of culture and the distribution of family names. Physica A: Statistical Mechanics and its Applications, 295(1–2):1– 8, June 1, 2001. CODEN PHYADX. ISSN 0378-4371 (print),

1873-2119 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378437101000462.

Ziliak:2008:CSS

[ZM08] Stephen Thomas Ziliak and Deirdre N. McCloskey. The Cult of Statistical Significance: How the Standard Error Costs Us Jobs, Justice, and Lives. Economics, cognition, and society. University of Michigan Press, Ann Arbor, MI, USA, 2008. ISBN 0-472-07007-X (cloth), 0-472-05007-9 (paperback). xxiii + 321 pp. LCCN HB137 .Z55 2008. URL http://www.loc.

gov/catdir/enhancements/fy0809/2007035401-d.html.

Zomaya:2014:POC

[Zom14] Albert Y. Zomaya. Pareto-optimal cloud bursting. IEEE Transactions on Parallel and Distributed Systems, 25(10): 2670–2682, October 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://www.computer.

org/csdl/trans/td/2014/10/06587242-abs.html.

Zorzi:2011:BL

[Zor11] Alberto Zorzi. 95.22 Benford’s law and π. The Mathematical Gazette, 95(533):264–266, July 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Zornig:2015:ZLR

[Zör15] Peter Zörnig. Zipf’s law for randomly generated frequencies: explicit tests for the goodness-of-fit. Journal of Statistical Computation and Simulation, 85(11):2202–2213, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhang:2011:ETO

[ZS11] Qunzhi Zhang and Didier Sornette. Empirical test of the origin of Zipf’s law in growing social networks. Physica A: Statistical Mechanics and its Applications, 390(23–24): 4124–4130, November 1, 2011. CODEN PHYADX. ISSN 0378-4371 (print), 1873-2119 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378437111005103.

Zea:2012:BEP

[ZSB^{+}12] Luz Milena Zea, Rodrigo B. Silva, Marcelo Bourguignon, Andrea M. Santos, and Gauss M. Cordeiro. The beta exponentiated Pareto distribution with application to bladder cancer susceptibility. International Journal of Statistics and Probability, 1(2):8–??, ???? 2012. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/16301.

Zunini:2019:BL

[Zun19] Mar´ ıa Caputi Zunini. Benford’s law. Publ. Mat. Urug., 17: 285–290, 2019. ISSN 0797-1443.